A Complete Bibliography of Publications in the Journal of Fish Biology (2020–2029)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

23 April 2025 Version 1.26

Title word cross-reference

#TimesUp [750].

11 [209]. 3 [225]. ^{+} [578, 1076]. ^{0} [394]. ^{13} [1233, 306, 115, 454]. ^{14} [1313]. ^{15} [450, 1233, 306, 115, 454]. ^{2} [498]. ^{2+} [379]. _{1} [27, 1107]. _{2} [159, 981, 1034, 1107, 546]. _{2α} [292]. _{50} [319]. _{max} [1451]. α [578]. β [48, 1468, 797]. : [78]. δ [450, 1233, 306, 115, 454]. γ [1076]. κ [913]. μ [1155]. × [1284, 1097, 560, 459, 649]. z [225].

-1 [48]. -1/3 [48]. -axis [225]. -glucan [1468]. -ketotestosterone [209].

0 [233].

1 [1503]. 1/3 [48]. 1/6 [48]. 1/6-glucan [48]. 11-ketotestosterone [797]. 13-year [1211]. 1880 [241].

2* [1036]. 20th [99]. 242 [721]. 2744-kg [1091]. 28/$33 [571]. 2D [1457].

36^{o} [841]. 3D [1457].

4 [571]. 42^{o} [841]. 434 [721].

5 [93]. 52^{o} [696]. 56^{o} [696]. 5m [93].

6 [465]. 6-glucan [48].

8 [526, 1240, 767].

978 [465, 526, 233, 93, 571]. 978-0-691-20598-4 [571]. 978-1-119-12726-0 [233]. 978-1-4214-3252-6 [465]. 978-1-78918-069-5 [93]. 978-1-78972-770-8 [526].

A. [1390]. abandonment [1371]. Abant [1181]. abanticus [1181]. abdominal [510]. abilities [706, 1255]. ability [1350, 1147, 1510]. abiotic [1439]. abnormalities [1082]. abode [858]. Abrolhos [1498]. absence [1462, 1363]. Abudefduf [1446]. abundance [891, 1431, 643, 206, 1297, 903, 46, 1375]. abundances [1200, 311]. Acanthistius [398]. Acanthopagrus [1214, 1012, 986, 711]. Acanthuriformes [1500, 1330]. Acanthurus [1125]. access [732, 1147]. Acclimation [592, 730, 1451, 231]. account [1435]. Accuracy [34, 274, 88, 797]. accurate [348]. Acestrorhynchidae [1163]. Acestrorhynchus [1163]. Acheilognathus [865, 965, 704]. Acid [306, 980, 1171, 321, 1263, 1306, 1219, 492, 1477, 646, 110, 177, 1149, 1060]. acid-binding [1149]. acidification [1292]. acidilactici [1468]. acids [430]. Acıgöl [191]. Acipenser [101, 730, 732, 908, 145, 1474, 1080, 1390, 315, 215, 222, 550]. acipenseriform [864, 52]. Acoustic [506, 800, 1511, 1037, 305, 1136, 396, 521, 425, 290]. acquisition [53]. Acropomatiformes [1051]. Across [1405, 862, 243, 1270, 196, 940, 80, 1354, 358, 814, 414, 1464, 1360, 1230, 110, 1249, 903, 937, 904]. Across- [904]. act [1387]. Actinopterygii [1500, 42, 281, 22, 1106, 1134]. action [1127]. activated [993, 967, 844]. Active [1440, 53, 289, 960]. Activin [346]. activities [933, 462]. activity [1312, 1247, 82, 1110, 539, 1520, 210, 315, 333, 443, 911, 666, 784, 228, 1504, 546, 1502, 550]. aculeatus [1151, 790, 337]. acusticus [1330]. acute [1517, 648]. acyl [177]. Acyrtus [693]. adaptation [600]. adaptations [281, 314]. Adaptive [1293, 865, 274, 1409, 183, 1419]. adaptors [913]. additional [964]. addresses [558]. adipose [1018]. adjacent [212, 453]. administration [228]. admixed [161]. admixture [1341]. Adriatic [42, 779, 1033]. Adult [163, 678, 189, 404, 618, 444, 431, 856, 995, 274, 73, 3, 305, 1245, 689, 985, 333, 905, 1246, 938, 38, 1441]. adult-age [618]. adults [1214]. Advances [1509, 533, 1402]. Advancing [337]. advantageous [262]. advantages [491]. Aegean [200, 552, 1279, 139, 1046]. aeglefinus [892]. aerial [709]. Aerobic [790, 1249, 53, 190, 1058]. Aeromonas [967]. aestuaria [1009]. Aetobatus [843, 384, 845, 804]. afer [134]. aff. [729]. affect [101, 537, 1003, 1395, 73, 908, 540, 192, 646, 1446, 666, 1182, 1473]. affected [1427]. affecting [429, 1346, 895]. affects [237, 1422, 530, 516, 968, 75, 1153, 346]. affinis [6, 1488, 1449]. affinities [916]. Africa [120, 1096, 852, 126, 130, 91, 631, 595, 185, 602, 125, 1321]. African [833, 118, 355, 328, 1325, 131, 132, 987, 1361, 133, 1129, 53, 1487, 559, 863, 1029, 1418]. africanum [1362]. After [949, 439, 1374, 464, 1465, 1096, 1306, 553, 1517, 698, 1212, 1449]. against [578, 394]. Age [668, 727, 299, 630, 656, 1407, 644, 1376, 848, 660, 984, 869, 41, 1196, 488, 1408, 639, 730, 618, 407, 792, 227, 490, 203, 582, 1516, 683, 1360, 1498, 180, 1258, 1472]. age-0 [730]. Age-based [848, 984]. age-determination [227]. age-specific [180]. ageing [839]. Agelas [1394]. Ageneiosus [4]. agents [1155]. ages [1313]. aggregating [1442]. aggregation [678, 1035]. aggregations [740, 458, 398, 287]. aggression [800, 185]. aggressive [871, 499]. aglomerular [873]. Agonistic [618, 883, 9]. agricultural [570]. aid [267, 936]. aided [85]. Air [1317, 107, 667]. Air-breathing [1317, 107, 667]. akajei [563]. akamai [493]. Akarotaxis [1103]. alalunga [768]. alarm [189, 1511, 134]. albacares [1340]. albacore [768]. albicaudus [302]. albinism [160, 1257, 1260]. albolineatus [751]. albonubes [853]. albopunctatus [24]. albotaeniatus [1258]. Albula [87, 719, 456, 1358]. alburnoides [1397]. Alburnus [318]. albus [1180, 52, 353]. aldehyde [1269]. alewife [675]. alewives [659]. alexandrini [1091, 29]. Alfaro [1332]. alfredi [89, 962, 1129, 1111]. Algeria [1211]. alginolyticus [1099]. alginolyticus-induced [1099]. alizarin [464, 1467]. alleynei [764]. allis [386, 1465]. allocation [1263, 841]. Allometric [1105, 1256]. Allopatric [631, 277]. allopatry [144]. Allothunnus [299]. allow [250]. allozyme [1036]. almaco [259]. along [845, 591, 603, 1129, 1089, 841]. Alopias [1002, 1379]. Alopiidae [1002]. Alosa [941, 386, 1465, 659, 675]. alpine [437, 1320]. alpinus [589, 494, 1045, 74, 1178, 1476, 889]. altavela [1213]. alter [1034, 561, 75]. alteration [1052]. alterations [637, 837]. altered [164]. alternating [813]. Alternative [1339, 319]. alters [1411]. altivelis [1502, 1502]. Amatitlania [1098]. Amazon [753, 789, 257, 1515, 151, 504, 291, 452, 25, 1083, 1493, 391, 279, 220]. Amazonas [327, 1319, 1081]. Amazonian [1052, 104, 405, 60, 485]. amazonicus [26]. Ambassis [1380]. amberjack [830]. ambient [448]. ambigua [244]. ambiguities [1430]. Amblyceps [1424]. Amblycipitidae [1424, 950]. Amblyopinae [1032]. Amblyraja [1506]. amblyrhynchos [1450]. amboinensis [607]. Ambon [607]. Ameiurus [282]. America [90, 685]. American [603, 1260, 110, 1242, 314]. americanus [154]. Amia [1026, 1343]. amid [280]. amidst [1194]. Amiidae [1343]. Amiiformes [1343]. Amino [321]. Ammocrypta [758]. Ammodytes [281]. Ammodytidae [281]. Ammoglanis [327]. ammonia [856, 689]. among [806, 502, 1045, 382, 74, 1385, 670, 15, 185, 1183, 909, 161]. amongst [50]. amount [1151]. amphibious [235]. amphidromous [549, 49, 187, 1502]. Amphidromy [33, 1327]. Amphiliidae [562]. Amphilius [562]. Amphiprion [28]. Amphistichus [103]. amplification [1305]. ampullae [390]. Amur [1240]. amyloliquefaciens [1240]. Anabantidae [120, 401]. Anabantiform [107]. Anabantiformes [120]. anabantoid [798]. Anadromous [683, 1465, 1045, 659, 379, 525, 545, 635, 786, 889, 675, 844]. Anadromy [155]. anaesthesia [911]. Analyses [1126, 1270, 620, 1390, 1184, 1373]. Analysis [784, 696, 1345, 832, 1171, 281, 277, 141, 720, 153, 1151, 993, 138, 1440, 1378, 946, 1325, 1348, 957, 37, 83, 428, 508, 703, 1077, 1457, 1144, 695, 1275, 958, 306, 769, 1185, 65, 221, 18, 154, 1205, 1035, 544, 1375, 1074, 1174, 1112, 1172, 19, 728, 705, 1217, 1009]. Analytical [891]. Anarhichas [772, 1482, 749]. Anatolia [952, 734]. Anatomy [390, 1247]. ancestral [887]. Ancherythroculter [838]. Anchoviella [954]. anchovies [21, 805]. anchovy [1046, 722, 1408]. ancient [1227, 1021]. Ancistrus [1461]. Andaman [1016]. Andean [340, 1031]. anderssoni [41]. Andes [350]. andrica [672]. Andrology [1277]. Andrusov [872]. Anelasma [1235]. Anemonefish [28, 754, 304]. anemonefishes [1307]. anemones [754]. Angel [436, 435]. angelfish [1172]. angelshark [963, 1257]. anglerfish [831]. anglers [1183]. angling [973, 210]. Anguilla [657, 1187, 832, 1233, 416, 1234, 130, 766, 317, 365, 63, 974, 91, 791, 1250, 1037, 820, 1109, 570, 1077, 1062, 164, 425, 1236, 646, 777, 1406, 313, 1059, 999, 55]. anguillicaudatus [141, 877, 12, 10, 784]. Anguilliformes [326, 1273, 890, 824, 1188, 1435, 512, 881, 1466]. anguineus [188]. angular [779]. Anisakidae [844]. Anisotremus [894]. Anne [30]. Annotated [743, 76]. Annual [339, 17, 618, 268, 197, 680, 329, 906, 3, 1014]. annually [1432]. annulatus [1010]. annulus [620]. anomaly [992]. anoplus [631]. anosteocytic [474]. Anostomidae [142, 1493]. ansorgii [595]. antagonism [1157]. antagonist [1157]. Antarctic [862, 1103, 861, 737, 400, 538, 926]. Antarctica [1460]. antarcticus [538]. antennas [598]. anthias [1092]. anthocyanin [476]. Anthropocene [936]. Anthropogenic [1209, 1293, 218, 1420, 687, 1194, 828]. anti [1312, 1245, 1099, 1172]. anti-enzymatic [1312]. anti-lipopolysaccharide [1099]. anti-Mullerian [1172]. anti-predator [1245]. Antibacterial [1520]. anticoagulants [608]. antioxidant [870, 1510]. antioxidative [923]. antisera [578]. anvils [334]. apapa [1242]. apex [1325]. Aphanius [191]. Aplocheilichthys [584]. Aplochiton [156]. Apparently [113]. appearance [1212]. appetite [1219]. appetite-regulating [1219]. Application [396, 52]. applying [138, 782]. appraisal [862]. approach [1022, 1315, 1268, 597, 510, 1068, 1202, 994, 1381, 1242]. approaches [517, 387, 1328, 407, 409, 319]. approaching [1109]. appropriate [592, 454]. Apteronotidae [452]. aquaculture [479, 1412, 718, 233]. aquaponics [922, 1504]. Aquaporin [809]. aquarium [1041, 1287, 1282]. aquatic [933, 742, 1325, 667, 1027]. aqueous [1230]. Arabia [681]. Arabian [1188, 268, 591, 1228]. arabicum [268]. arabicus [1214]. Araguaia [736]. Arapaima [193]. Araújo [954]. Archipelago [695, 284, 564, 1356, 1336, 623, 1489, 1251, 1518]. architect [309]. Architectural [469]. architecture [470]. archival [384, 1169]. Arctic [589, 494, 1045, 74, 1178, 426, 155, 1476, 1304, 573, 1280, 889, 1235, 1355, 892]. arctifrons [1017]. Arctozenus [404]. area [256, 709, 785, 200, 1232, 280, 1063, 686, 928, 1057, 777, 522, 1489, 636, 1502]. areas [969, 387, 761, 1267, 131, 558, 187, 453, 311, 1190, 55]. areolatus [727]. argenteus [1126, 608, 582, 1222]. Argentina [1022, 668, 702, 239]. Argentine [785, 387]. argentinensis [22]. argentiventris [503]. Argulus [884, 363]. Argyrops [509, 69]. Argyrosomus [1015]. Arhynchobatidae [964]. Ariidae [1023, 1197, 602]. Ariommatidae [1335]. Ariopsis [575, 221]. Ariosoma [890, 824]. armored [1319]. armour [1143]. armoured [555, 667, 60, 629, 279]. Arms [311]. Arnold [465]. arrangement [1204]. artificial [368, 271, 731, 553, 999]. artificially [799]. artisanal [1400]. Ascanius [1499]. Ascending [333]. Asia [621]. Asian [60, 1189]. asp [273]. aspect [338]. Aspects [349, 1386]. Aspidoras [279]. aspius [273]. Aspredinidae [493]. assays [1311]. assemblage [1146, 754, 1272, 782]. assemblages [1354, 484, 643, 628, 671, 11, 238]. assess [387, 274, 430, 1255]. assessed [919, 1507, 1184, 769]. Assessing [517, 695, 1439, 66, 1241]. Assessment [566, 8, 364, 1125, 1318, 268, 812, 157, 4, 1273, 543, 1474, 419, 409, 945, 410]. assessments [628]. assimilation [760]. associate [850, 1511]. associated [1118, 1488, 20, 641, 1264, 582, 1334, 215, 790, 896]. Association [576, 798, 1458, 1050, 1335, 1074, 300]. associations [1096, 755, 948, 401, 782, 1321]. Assortative [185]. assumptions [1174]. asterias [371, 66]. Astragaloside [923]. Astroblepidae [1031]. Astroblepus [1031]. astromycterus [855]. astyanax [1114, 402, 707, 1114, 250]. Asymmetric [751, 851, 618]. asymmetry [1267, 399, 230, 1153, 1279, 1224, 1104]. Atabapo [327]. atelodus [792]. Atelomycterus [356, 1394]. ater [1038]. Atherinella [393]. Atherinopsidae [22]. Atlantic [261, 360, 763, 284, 1057, 1014, 378, 836, 1248, 1227, 515, 625, 871, 113, 1484, 930, 269, 696, 1118, 310, 772, 102, 1458, 90, 620, 1086, 1244, 1354, 912, 1318, 778, 904, 981, 1431, 1506, 404, 935, 1108, 1161, 479, 992, 483, 799, 1368, 1091, 357, 656, 329, 968, 979, 1191, 37, 83, 274, 605, 385, 398, 997, 44, 301, 1136, 428, 186, 568, 1517, 953, 1025, 1392, 1144, 596, 146, 1158, 1175, 647, 1316, 695, 1469, 67, 1432, 212, 540, 315, 306, 1353, 492, 1360, 749, 1301, 283, 453, 9, 1429, 633, 198]. Atlantic [181, 1489, 488, 938, 441, 1313, 1357, 38, 363, 1036, 724, 1441, 1381, 1032, 821, 1330, 432, 349, 1428]. atlanticus [1244, 656]. Atlantoraja [668]. Atoll [812, 1166, 414]. Atractoscion [792]. Atractosteus [585]. atresia [617]. atrioventricular [246]. atripinnis [502]. atromaculatus [516]. attached [463]. attachment [361, 665, 585, 1343, 945]. attendance [218]. attenuates [6]. attraction [402, 700]. attractive [700]. Auchenipteridae [4, 25]. Auchenoglanididae [1203]. audax [322]. auditory [521]. Aulopiformes [404]. aurata [324, 258, 533, 399, 194, 733, 462, 482, 140]. auratus [331, 641, 645, 1050, 797, 489]. Australasian [641]. Australia [951, 727, 320, 1019, 845, 697, 1053, 286, 46]. Australian [89, 684, 845, 792, 1140, 746, 347]. Austroglanididae [863]. Austrolebias [17]. autofluorescence [617]. autofluorescence-based [617]. automatic [1368]. autophagy [915]. Autumn [1199, 865, 1507, 926]. autumn-spawning [865]. availability [208, 916, 926]. avoid [1367]. avoidance [566, 438, 521, 84]. Awaous [1327, 634, 1421]. Awareness [35]. away [348]. axis [1219, 79, 225]. ayu [1502]. azaghal [279]. Azores [1484, 249, 564, 1297].

B [913, 1076, 1114]. Bacillus [1240]. Back [1287]. background [492]. bacteria [48]. bacterial [141, 720, 1266, 958]. bad [517, 262]. baerii [101, 1080, 215, 222]. baffled [731]. Bagridae [421]. Bahamas [1161, 903]. Bahia [1056]. bainsii [401]. bait [463, 262]. Baja [433]. Balbiani [868]. Balistes [488, 1054]. Balistoides [335]. Balitoridae [1084]. ballan [1499, 816, 1306]. Baltic [444, 1044, 44, 1358, 972, 777]. Baltimore [465]. banana [634]. band [3, 1202, 1379]. band-pair [3]. banded [1030, 1394, 584]. Bangalore [1011]. Bangladesh [743]. banjo [493]. Bank [261]. Baraka [244]. barb [604, 1112]. barbatulus [704]. barbel [139]. barbels [518]. Barbodes [661, 604]. Barbosa [954]. Barbus [139, 518, 262]. barcodes [1134]. barcoding [952, 1140, 424, 1068, 947]. Barents [1001, 1440]. barnacle [1235]. barracuda [1484, 210, 1484]. barracudina [404]. barred [971]. Barrier [697, 286, 243, 819, 1240, 1141]. barriers [1135]. basal [1298]. based [842, 569, 696, 440, 354, 245, 100, 935, 1038, 1201, 752, 617, 848, 1005, 1159, 409, 984, 619, 1015, 773, 1150, 947, 225, 964, 547, 713, 407, 410]. Basin [42, 1068, 555, 142, 354, 1315, 1225, 99, 753, 651, 151, 104, 504, 139, 789, 736, 710, 1231, 576, 255, 914, 1494, 1081, 1271, 1324, 122, 1106, 452, 513, 679, 672, 345, 25, 1493, 1056, 1229, 1397, 16, 279, 601, 485, 1319, 950, 314, 352, 24, 220, 40, 774, 121, 548, 257]. basins [1081]. basis [636]. basking [1008, 1014]. Bass [1256, 1481, 918, 1329, 1130, 1294, 766, 447, 1034, 910, 841, 747, 173, 1076]. Bathydraconidae [1103]. Bathymyrinae [890]. Bathyraja [239, 964]. Bathysphyraenops [1051]. batoid [1196]. batoids [449]. Batticaloa [1146]. Bay [580, 241, 1420, 676, 962, 1329, 1267, 1377, 1435, 512, 881, 633, 743, 1104]. be [754, 33, 846]. beach [1507]. beaches [1472]. beaching [335]. beam [567]. Bean [241, 899]. become [700]. beds [916]. before [57, 1465, 1517]. beginnings [508]. behavior [1317, 1450, 1332, 1478, 1482, 1457, 1494, 1316, 1469, 1386, 556, 1497, 1397]. behavioral [1476]. behaviors [1437]. behaviour [607, 625, 871, 1079, 516, 486, 1178, 1004, 554, 1148, 968, 175, 131, 793, 438, 973, 996, 820, 1109, 1110, 514, 1025, 614, 1113, 353, 51, 1275, 109, 294, 282, 207, 9, 828, 1014, 1035, 313, 663, 325, 335, 768, 999, 883, 1162, 457, 1259, 499]. Behavioural [292, 1293, 742, 465, 133, 632, 744, 813, 148, 153, 1034, 924, 592, 1124, 1008, 183, 173, 1182]. behaviours [566, 376]. behind [1008]. Beibu [219]. Beijerinck [1284]. bejeus [1030]. belizanus [1278]. bellied [230, 1083]. Bellotti [1038]. belly [552]. Belonesox [1278]. beluga [209]. benchmarking [397]. benefits [184]. Bengal [512, 881, 1377, 1435, 743]. bengalensis [884]. bentfin [1518]. benthic [760, 551, 1043, 289]. Benthosema [1048]. Berg [1396]. bergylta [1499, 816, 1306]. berthalutzae [1079]. Best [170]. beta [241, 899]. Better [879, 888, 1517, 319]. Between [120, 148, 1298, 318, 1458, 1095, 382, 1179, 759, 1127, 628, 14, 399, 459, 948, 803, 1281, 860, 545, 393, 1471, 1046, 576, 1445, 1245, 695, 88, 632, 1334, 176, 115, 232, 1335, 129, 1168, 770, 105, 238, 1278, 909, 663, 363, 851, 1502]. biannually [1196]. bias [498, 1217]. biased [278, 1125]. biases [1073]. Bichir [526]. bicolor [1187, 365]. Bidirectional [1386, 1173]. Biega [76]. biennially [360]. bifasciatus [870, 1189]. big [897]. big-eye [897]. bigeye [714, 1002]. bighead [71, 1342]. Bight [1353, 1364]. bigscale [1391]. Bilateral [1267, 1224, 1104, 739]. bile [733, 110]. bilineata [1124]. bimaculatus [247, 1114]. Bimini [903]. binding [1359, 1149]. binotatus [661, 604]. bio [1055]. bio-synthesizing [1055]. biochemical [387, 884, 997, 505, 1474, 511, 719]. biochemistry [1284, 1148]. biocides [1412]. Biodiversity [465, 827, 284, 60, 35]. bioenergetic [825]. biofilm [760]. biofluorescence [1000, 831]. biogeographic [350, 256, 1324]. Biogeographical [261]. biogeography [1081, 693]. biologger [759]. biologger-derived [759]. Biological [1425, 926, 912, 436, 565, 1002, 226]. biologists [35]. Biology [93, 602, 118, 237, 530, 843, 622, 426, 599, 1302, 605, 1471, 1391, 615, 1197, 1123, 31, 1119, 575, 1163, 1010, 869, 182, 451, 349, 1395, 118]. biomarker [1082]. Biomass [417, 256, 643, 1363]. Biomass-dependent [417]. biometry [1085]. biomineralization [1373]. biomonitor [637]. biophysical [949]. Biosecurity [783, 1183]. Biosphere [156]. biosynthesis [177]. biparental [1341]. birostris [89, 1513, 1111]. Birth [1075, 54]. Biscay [633]. bispinosa [697]. bite [804, 1489]. bitterling [865, 965, 704, 1334, 858, 161]. black [1052, 800, 599, 1012, 1425, 917, 282, 640, 841, 544, 45, 711, 285, 1046, 673]. black-chinned [800]. blackchin [1398]. blackfin [502]. blackmouth [298]. Blackspotted [1367]. blacktip [372, 163, 1378, 422, 921, 636]. bleak [318]. Bleeker [1450, 1410, 1490]. bleekeri [69]. Blenniidae [90, 795]. Blenniiformes [627]. blennioides [14]. blenny [90, 795, 362, 1119, 9, 228]. blind [402]. Bloch [1204, 1359, 1483, 692, 1501, 1226, 1032, 837]. blochii [1138]. blood [1244, 856, 1148, 764, 997, 379, 170, 1266, 689, 1218, 1477]. blood-feeding [764]. blood/interstitial [379]. blooded [400]. bloodworms [839]. blue [297, 1116, 223, 195, 157, 554, 1433, 207]. blue-streak [297]. bluefin [1227, 310, 1318, 633]. bluegill [486]. blueline [1313]. bluntnose [115]. Bockmann [1496]. Boddart [177]. boddartia [177]. Bodianus [1258]. bodied [112, 82, 818, 1363]. bodies [328, 370]. body [536, 21, 868, 1434, 252, 904, 642, 618, 865, 1135, 37, 885, 1046, 928, 553, 1445, 987, 146, 510, 442, 1011, 1388, 34, 1024, 1168, 938, 1049, 1344]. bolder [95]. Boldness [1295]. Boleophthalmus [1089, 177]. bomb [1313]. bonaci [917]. bonariensis [1072]. bone [483, 476, 478, 181]. bonefish [87, 719, 456]. Bonnaterre [696, 1400]. bonnet [1040]. bonnethead [729, 357, 178]. bony [1091, 882]. Book [465, 526, 233, 166, 93, 571, 225]. boost [467]. Boraras [1385]. borari [1229]. boreal [1354, 1110, 11, 1308]. Boreogadus [548, 1207]. borneensis [1197]. boscii [1033]. both [683, 701]. botiid [1410]. bottom [1259]. bottom-feeding [1259]. Boulenger [987, 1271, 562, 1092]. boulengeri [1092]. bound [663]. bounds [903]. bourgeois [1119]. boweni [577]. bowfin [1026, 1343]. bowfishing [1026]. box [112]. brachycephalus [1491]. Brachychalcinus [104]. brachygnathus [1408]. Brachymystax [1412]. brachypomus [459, 1083]. brachyurus [778]. brackish [425, 584, 55]. brackishwater [245]. Brahmaputra [588, 70]. brain [361, 1156, 968, 455, 1219, 109, 1496, 909, 275, 636]. brainer [901]. Braithwaite [30]. Bransfield [926]. brasiliensis [658, 393, 211, 432]. Brauer [41]. braueri [41]. Brazil [695, 917, 843, 288, 555, 1006, 142, 1515, 1086, 907, 1225, 358, 98, 651, 104, 729, 789, 736, 1231, 1324, 827, 991, 513, 284, 679, 672, 1168, 7, 796, 1260, 522, 1083, 1190, 1056, 629, 1224, 441, 16, 279, 855, 874, 314, 352, 241, 40, 774]. Brazilian [220, 208, 256, 740, 664, 1079, 992, 393, 291, 58, 1069, 211, 432]. bream [1214, 194, 462, 1012, 1437, 897, 482]. breathing [1317, 107, 667]. breed [1418]. breeder [1263]. breeding [956, 678, 642, 339, 337, 1061]. brevirostris [1277, 1075]. brevirostrum [145]. brevispina [950]. Brevoort [333]. Bringing [976, 1094]. Bristol [1483]. Britain [1183]. British [364]. Broad [1027]. Broad-scale [1027]. Broadbill [596]. broadnose [458, 811, 272]. Brood [165]. brooding [956, 1160, 1197]. broods [1160]. Broodstock [981, 245, 582, 489]. brook [100, 925, 1451, 231, 851]. brotula [664]. Brown [1178, 574, 589, 1382, 1349, 1444, 755, 888, 525, 545, 840, 305, 817, 1087, 573, 36, 886, 1074, 851, 1199]. brucus [1400]. Brunswick [1131]. BRUVS [58]. Brycon [26]. Bryconamericus [1396]. Bryconops [257]. Bryx [1479]. buchholzi [328]. bugensis [872]. building [429, 309, 1113]. bulk [454]. Bull [463, 998, 644, 568, 1336, 1133]. bullhead [68, 282]. bullseye [1010]. Bump [29]. Bump-head [29]. burbot [815]. Burgess [616]. Burmeister [685]. Burrowing [313, 1032]. burst [1488, 1303]. bursts [329]. bury [708]. bushynose [1461]. butterfly [1213, 47, 1123]. butterflyfish [328, 823]. butts [1005]. bycatch [976, 749].

C [1233, 78, 306, 115, 454, 1313]. Ca [379]. cabezon [1174]. cachara [649]. Cádiz [1472]. cadmium [648]. caeruleopunctatus [951]. caerulescens [669]. Caesionidae [1356]. cage [1368]. cages [1329]. calbasu [982]. calcified [811, 469, 1492]. calcium [1230]. Caledonia [848]. calico [103]. California [415, 688, 368, 229, 336, 508, 1078, 287, 433, 1150]. call [1511, 1364]. callainos [176]. Callichthyidae [279, 601]. Calliclinus [362]. callidus [344]. Callionymidae [594]. callionymids [594]. calva [1026]. Camargue [174]. Cambeva [907, 40, 774]. came [1223]. Can [408, 1218, 846, 494, 498, 1511, 754, 33, 1160, 1387, 898]. Canada [1137, 453, 364, 758, 1131]. Canadian [1431, 155, 889, 1235, 1355]. canadum [1375]. canal [390]. canals [418, 388]. candiru [736]. candirus [764]. canicula [537, 246, 1520, 776, 192]. Can’t [1444]. canteri [221]. canthaxanthin [1218]. capabilities [521]. capacity [1135, 923, 790]. Cape [120, 1096, 134, 156]. capensis [120, 271, 83]. Capoeta [1028]. capriscus [488]. captive [245, 339, 1004, 613, 1362]. captivity [652, 906, 230, 1075, 612]. capture [1431, 272, 226, 559]. capture-induced [226]. capture-recapture [272]. captured [843]. Carangidae [39]. Caranx [1388, 660]. Carassius [331, 645, 797, 489, 1308]. carbohydrates [742]. carbon [1270]. carbunculus [1005]. carcass [282]. carcasses [1141, 819]. Carcharhiniformes [622, 246, 356]. Carcharhinus [372, 360, 1450, 778, 163, 1378, 463, 520, 998, 644, 875, 568, 422, 921, 1336, 1133, 636]. carcharias [369, 1333]. Carcharodon [369, 1333]. Cardiac [911, 790]. cardinal [564]. cardinalfish [1490]. cardinalis [219]. care [1462]. Caribbean [1400, 729, 875, 303, 616, 1057]. carnivorous [658, 839]. carolinus [1468]. carotenoid [494]. carp [1065, 980, 331, 224, 738, 1055, 48, 793, 703, 71, 923, 521, 561, 147, 1503, 1176, 394, 45, 725, 960, 1427, 1342, 767, 1308]. carpetshark [268]. carpio [725, 1065, 980, 738, 48, 521, 1503, 1176, 546]. cartilage [469, 470]. Carvalho [954]. cascudinhos [358]. case [435, 415, 229, 983, 533, 1521, 552, 1201, 1411, 1264, 1208, 298, 363, 1223, 987]. cases [356, 964]. Caspian [359, 1339]. castelnaui [668]. catabolic [706]. catadromous [549]. catappa [1100]. catch [696, 949, 635, 1254, 1002, 1517, 1371]. catch-and-release [1517]. catchment [1382, 1431]. catchments [416, 423, 317, 817]. categories [274]. categorization [494]. catfish [235, 341, 247, 555, 260, 493, 907, 1225, 993, 651, 1312, 1148, 327, 710, 1231, 967, 389, 419, 405, 562, 1031, 1461, 958, 575, 667, 60, 679, 25, 221, 863, 602, 77, 165, 279, 601, 1319, 1424, 950]. Catfishes [421, 1023, 4, 15, 1496, 629]. Catla [1055]. Catoprion [240]. Catostomus [1326, 508]. catshark [1520, 298, 1321]. catsharks [537, 192, 356, 1394, 559]. caudal [170, 77]. caught [1274, 230, 1425, 815, 1174]. Caulolatilus [1313]. causation [985]. Cause [864]. Cave [569, 402]. Cave-dwelling [569]. cavefish [1414]. cayennensis [954]. ccdc3 [1427]. CD8 [578]. cdca9 [877]. Celebes [63]. celebesensis [63]. celiac [705]. cell [1065, 242, 669, 455, 1490, 784, 77, 45, 1013]. cell-specific [669]. cells [12, 1266, 1122, 394]. cellular [65, 1347]. cemiculus [1398]. census [1117]. Central [90, 914, 1078, 621, 125, 931, 963, 761, 276, 155, 736, 929, 1254, 206, 1336, 629, 46, 391, 239, 1033, 1356, 1415, 626]. central-southeast [206].

Centrarchidae [1106]. centrina [779, 970]. Centrolophus [1425]. Centropomidae [605]. Centropomus [605, 1278, 312]. Centropristis [841]. Centropyge [697, 1172]. Centroscymnus [67]. century [99]. Cépède [745]. cepedianus [264, 458, 811, 272]. cephalic [388]. Cephalocassis [1197]. cephalopod [232]. Cephaloscyllium [622]. Ceratioidei [1448]. Ceratoscopelus [1491]. cerebellar [968]. cerevisiae [258]. cernuus [631, 397]. certified [1276]. cesium [1445]. Cetopsorhamdia [753]. Cetorhinus [1014]. cf. [223, 1331, 1456]. Chaenodraco [926]. Chaenopsidae [627]. Chaenopsis [627]. Chaetodontidae [823]. Chaetostoma [485]. Chagos [1518]. chain [246, 177]. chalk [647]. challenge [141, 1232, 967, 78, 958]. challenged [1285]. challenges [1290, 1455, 147, 1418]. chamber [172]. Champsocephalus [1263]. Chang [950]. Chang-Jiang [950]. change [772, 507, 643, 971, 1386, 1173, 251, 749, 1262, 1370, 18, 978, 821, 1347]. Changes [965, 739, 1201, 1317, 912, 586, 1420, 998, 1299, 393, 563, 414, 1517, 1468, 966, 1429, 837]. changing [836, 295, 657, 1329, 43]. Channa [1204, 1156, 1359, 1351, 837]. Channel [1483, 1333, 993, 1495]. Channidae [1351]. Chapada [1324]. Char [166, 1045, 889]. characid [495]. Characidae [1396, 151, 104, 504, 97, 395, 1324, 707, 672, 1229, 16, 1114]. Characidium [350, 391, 314, 352]. Characiformes [350, 1006, 142, 355, 354, 151, 104, 504, 789, 97, 1271, 1324, 240, 1163, 672, 7, 1083, 1493, 257, 1229, 16, 391, 1396, 26, 352, 1114]. Character [1434, 27]. Characterisation [12, 20, 177, 26, 19]. Characteristics [1395, 727, 1097, 912, 382, 429, 365, 63, 974, 1044, 592, 1159, 211]. Characterization [578, 222, 1114, 324, 720, 763, 946, 455, 147, 370, 65, 544, 1375, 719]. characters [591, 391]. charged [718]. charr [589, 494, 74, 1178, 426, 988, 1476, 1304, 573, 1182]. chase [348]. Chasing [348, 172]. checking [1394]. checklist [569, 76, 743, 254]. Cheilinus [670]. Cheirodontinae [354]. Chelidoperca [624]. chemical [1447, 292, 164, 911, 994, 499, 1373]. chemically [700]. chemistry [288, 1244, 227, 860]. chemoreception [235, 341, 260]. Chen [233]. chest [460]. Chile [156]. Chilean [458, 490, 380]. chilensis [524]. chilkensis [1435]. Chiloscyllium [630, 268]. chimaera [249, 814, 249]. Chimaeras [571, 293]. Chimaeridae [249]. China [1491, 1466, 219, 1403, 5, 1410, 1407, 13, 853, 255, 1415, 1402, 1411, 914, 1414, 1406, 950, 1408, 728]. Chinese [720, 1474, 915, 19]. chinned [800]. Chinook [1141, 1127, 323, 78, 819]. Chionodraco [737]. Chirostoma [276]. Chlamydoselachus [188]. Chlorella [1284]. chloride [837]. chloride-induced [837]. Chocó [350]. Choerodon [334]. choice [865, 72, 164, 176, 691, 75, 1206]. choices [1311]. cholesterol [809]. chondrichthyan [469]. chondrichthyans [1514, 64]. Chondrichthyes [481, 264, 1277, 1086, 565, 188, 246, 1075, 198, 171]. choose [619]. chromis [208]. chromophores [937]. Chromosomal [27, 1107, 1082]. chromosome [1107]. Chronic [1252, 1058, 173]. chronologies [895]. chronology [203]. Chrosomus [230]. Chrysiptera [1511]. Chrysophrys [641, 1050]. chrysopterus [28]. chuatsi [1221, 915, 713]. chub [516, 1480]. chum [521, 822]. chupare [198]. Cichla [1130, 447]. Cichlasoma [392]. cichlid [1098, 392, 1041, 72, 303, 1011, 176, 666, 165, 125, 316]. Cichlidae [1041, 303, 1042, 332, 316, 1475]. cichlids [715, 937]. Cichliformes [303]. cigarfish [336]. Ciona [1285]. Circles [1014, 1008]. circular [320, 838]. circumtropical [259]. cirratum [442, 1090]. Cirrhimuraena [512]. cirrosa [557]. citations [1395]. Citharichthys [1017]. Citizen [955, 899, 96, 919]. clade [1447, 42]. Clarias [77]. class [895]. classification [930]. Claudins [81]. clavata [813, 957, 1423]. cleaned [297]. cleaner [297, 479, 1000]. cleaning [1397]. clearance [1192, 856, 689, 662]. Climate [1357, 237, 530, 772, 539, 971, 749, 1262, 978]. climate-driven [237, 530]. climatic [256]. Climbing [1370]. Clinostomus [1162]. cloacal [1353]. Cloning [1099, 1427, 141, 946, 1149]. close [589]. closed [200]. closely [176, 604]. closure [510]. closures [912]. Cloud [853]. clown [366]. clue [636]. Clupea [1044, 186, 972]. clupeaformis [895, 179]. clupeid [786]. Clupeidae [386, 1009]. Clupeiform [868, 1070]. Clupeiformes [1501, 954, 833, 1242]. clupeoides [833]. Clyde [1025]. co [318, 639, 1123, 1184, 147, 64, 1145, 1487, 782, 1321, 159, 1034]. co-occurrence [1487, 782]. co-occurring [318, 639, 1123, 1184, 64, 1145, 1321]. co-receptor [147]. coach [1514]. Coast [90, 64, 208, 256, 843, 845, 591, 824, 1188, 1435, 1254, 1053, 776, 1089, 616, 1069, 1061, 46, 1515]. Coastal [522, 549, 89, 1382, 963, 949, 805, 120, 458, 761, 358, 814, 996, 1272, 1392, 158, 822, 966, 602, 1224, 441, 1357, 1384]. coastline [1129]. coasts [547]. cobble [179]. cobia [1375]. Cobitidae [27, 723]. Cobitis [27, 1124, 723]. cocama [1464]. cochliodon [220]. Cochlodinium [556]. cod [981, 444, 329, 548, 453, 994, 1207]. coelolepis [67]. coexistence [250]. coexisting [951, 156]. cognitive [1311]. Coho [156, 1153, 364]. cohorts [1495]. Coilia [1408, 844]. coioides [507, 251]. cold [730, 455, 1264, 846, 1262, 859]. cold-sensitive [455]. cold-shock [1264]. collaborative [1111]. collagen [960]. collapse [1465, 539]. collected [591, 570, 991]. collections [517, 35]. collective [701]. Collichthys [1267, 752, 1104, 1217]. colliei [242]. Colombia [716, 106, 151, 97, 485, 1319]. Colombian [350, 616]. colonial [827]. colonic [1338]. Colonization [276, 1442, 1323]. color [1480]. Colorado [1210]. colorful [1324]. colour [907, 1218, 1206, 821]. colouration [440, 795, 1438, 356]. Columbia [364]. comber [1349]. combined [536, 252]. come [407]. comes [513, 573, 909]. coming [169]. comments [569, 867, 1103, 1324, 67, 704, 388, 1051, 1421, 1056, 16, 485, 314]. Commercial [8, 449, 1159, 1297]. commercially [657, 111, 39, 1000, 286, 1083, 1498, 180, 1243]. Common [212, 1065, 842, 765, 318, 516, 665, 738, 1412, 48, 558, 762, 521, 207, 1503, 1176, 1278, 1379, 1381, 725, 312]. communication [800, 499]. communities [216, 412, 1145]. Community [1168, 1509, 416, 317, 414, 676].

Community-scale [1168]. company [1194]. Comparative [1086, 1393, 799, 39, 1305, 1471, 1390, 286, 7, 797, 1278, 180, 1217, 316, 1318, 103, 1177, 510, 994]. compare [832, 217]. compared [1047, 1024]. Comparing [450, 1186, 1154, 1321]. Comparison [550, 731, 763, 973, 594, 1362, 1457, 521, 474, 1153, 903]. comparisons [178]. compensating [606]. compensation [1496]. competition [500, 73, 615, 1278, 851]. complementary [1171]. complete [652, 300]. complex [361, 1443, 756, 789, 1273, 853, 631, 634, 595, 661, 905]. Complexity [654, 671, 325]. compliant [409]. complicates [1382]. components [361, 949, 960]. Composition [563, 1003, 1377, 1263, 1254, 739, 1387, 1388, 474, 291, 492, 646, 66, 666, 784, 994, 1355, 1060, 1373]. comprehensive [1349, 373, 1268, 719]. compressed [1340]. computed [324]. Computer [85]. Computer-aided [85]. concentration [1499, 379, 1445]. concept [494]. conceptual [1488]. concerned [1417]. concolor [1273]. condition [1458, 730, 387, 904, 430, 442, 1024, 1106, 619]. conditioning [1255]. conditions [295, 1192, 1329, 586, 448, 1314, 1110, 894, 1105, 8, 598, 938, 662]. cone [567, 937]. confirmation [927]. confirmed [293, 652, 1366, 762]. conflicts [1364]. conger [57, 1188, 1228, 57]. Congo [123, 76, 576, 124, 1068, 1203, 125, 121, 1271, 122]. congrid [890, 824, 881]. Congridae [281, 890, 824, 1188, 881]. connected [580]. connectivity [50, 1444, 1272, 154, 380, 579]. connects [111]. consecutive [1160]. consequence [808]. consequences [836, 515, 1293, 43]. Conservation [655, 758, 1470, 1096, 1403, 1063, 1402, 1411, 1324, 259, 1111, 457]. conservatism [583]. Conserved [1361]. considerations [50, 66]. consistency [148, 1253]. consistent [619]. consistently [1372]. Conspecific [325, 1447, 1115, 1511, 134, 109]. conspecifics [850, 1367, 898]. Constance [1436]. constrained [56]. constraints [995, 1245]. consumption [536, 112, 252, 759, 393, 282]. contactless [1247]. Contas [142]. Contemporary [1291, 779]. content [696, 1345, 1171, 277, 809, 957, 860, 81, 1175, 695, 1184, 1477, 646, 784, 1100]. contents [193]. context [148, 532, 103, 539, 43]. continental [130, 776, 1017]. Continuous [525]. contraction [400]. contradictory [113]. contrast [1155]. Contrasting [1385, 1062, 9, 92, 304, 363]. contribute [84]. contributed [635]. Contribution [937, 1151]. contributions [1059]. Control [233, 213, 403, 746]. controlled [732, 1314, 1207]. convergence [987]. convergent [757]. conversion [1127]. cooki [366]. cookiecutter [1489]. cool [647]. cooling [86]. cope [354]. Copp [1452]. copper [778]. copulation [1031]. coral [607, 694, 278, 111, 995, 849, 670, 1208, 1145, 445, 898, 286, 869, 238, 335, 1363, 697, 286]. coral-reef [849]. coregonid [464]. Coregonus [887, 939, 1358, 895, 179]. coriiceps [737]. cornerstone [693]. Correction [1192, 1141]. correlated [865]. correlates [1419]. Corrib [423]. Corrigendum [167, 307]. corruscans [294, 1170, 649]. Corsica [806]. cortisol [1298]. Corydoras [667, 601]. Corynopomini [395]. Coryogalops [1473]. Coryphaena [1270, 212]. Coryphaenoides [1449]. Corythoichthys [1134]. Costa [1133, 1496]. cottid [726]. Cottidae [726]. Cottus [68, 11]. Coulson [93]. counterparts [1024]. counting [1409]. Counts [278]. coupling [803]. Courtship [1332, 1014]. covered [889]. cownose [845, 1260, 46, 432]. Coyote [1133]. cozy [858]. crassioris [1424]. create [320]. creating [75]. creek [516, 723]. Crenuchidae [350, 391, 314, 352]. Crepori [789]. crested [942, 802]. critical [1422, 962, 1179, 1167, 1378, 462, 385, 1428]. critically [437, 740, 832, 1233, 1398, 280, 779, 1268, 970, 1407, 1257, 682, 1211, 191, 460, 963, 366]. croaker [810, 1201, 752, 745, 1217]. Croatia [42]. cross [148, 341, 260]. cross-context [148]. crosses [751]. Crosstalk [1095]. crowding [821]. crucial [65]. crucian [331, 394, 1308]. cruising [1303]. crustacean [1132]. cryopreservation [77]. crypsis [133]. Cryptic [1021, 518, 661, 577, 120, 1315, 729, 1048, 150, 75, 1134, 1344]. crypto [1043]. crypto-benthic [1043]. CT [567, 1155, 1451]. Ctenopharyngodon [923, 960, 1427]. Cubiceps [336]. cuckoo [165]. cues [1447, 788, 321, 680, 1511, 525, 1392, 134, 164]. cuffyi [389]. culling [1497]. cultratus [1332]. Culture [93, 1105]. cultured [425, 209, 815]. cultures [1065]. culverts [438]. Curcumin [1219]. curema [288]. current [813, 98, 1302]. currents [1505]. curves [1371]. cusk [524]. cusk-eel [524]. cutaneous [797]. Cutthroat [31]. Cuvier [805, 1351, 210, 971, 1384, 650, 903]. CXCR3a [946]. CXCR3b [946]. cyanellus [896]. cycle [360, 1070, 268, 357, 906, 1441]. Cycleptus [195]. cycles [337, 1418]. cyclicity [1246]. Cyclopterus [1000]. cylinder [793, 838]. cylindrospermopsin [1192, 662]. Cynoglossus [720, 706]. cyp11a [544]. cyp11b [544]. Cyphomyrus [122]. cyprinid [669, 129, 838, 701]. Cyprinidae [1028, 139, 631, 558, 576, 518, 1414, 124, 128, 661, 604, 581]. Cypriniformes [13, 1414, 1410, 853, 128, 1084, 581, 723, 734]. Cyprinodon [556]. cyprinodontiform [340]. Cyprinodontiformes [17, 502, 440, 1332, 603, 143, 1224]. cyprinoid [491]. Cyprinus [1065, 980, 738, 48, 521, 1503, 1176, 546, 725]. cyrtogaster [257]. Cystidicolidae [1456]. Cyto [637]. Cyto-histopathological [637]. Cytological [281]. cytoplasmic [868]. cytotype [1114].

D [474, 571, 225, 793]. D-cylinder [793]. Dabry [914]. dace [230, 1162]. Daily [1498, 1191]. Dalatias [565]. Dalatidae [565]. Dam [980, 846]. dammed [1052]. damsel [1486]. damselfish [1251, 607, 28, 1511, 898, 1454, 1446, 1462]. Dando [571]. danio [751, 148, 850, 189, 523, 751, 1299, 213, 472, 1252, 473, 1152, 346]. darkness [162]. darter [1443, 14, 1101, 818, 380, 758, 314]. darters [355]. Dascyllus [28, 898, 1454, 1462]. Dasyatidae [1164, 1514, 283, 1069]. data [350, 435, 354, 100, 919, 812, 1268, 861, 274, 762, 1002, 1464, 244, 855, 782, 254]. data-limited [812]. data-poor [244]. database [284]. Date [620, 1274]. day [696, 828]. dead [1199]. dead-end [1199]. deani [599]. death [57]. decades [44, 1212]. deceitful [691]. decibels [1194]. decision [1381]. Decline [772, 435, 1365, 846]. decoupled [1460]. decreases [990]. Deep [150, 595, 931, 1132, 872, 261, 1354, 162, 803, 1188, 1228, 552, 539, 188, 831, 1208, 991, 1053, 1005, 616, 1216, 23, 1042, 180]. deep-sea [261, 1354, 162, 803, 188, 831, 1208, 991, 23]. deep-water [1132, 1188, 539, 1053, 1005, 1042, 180]. deeper [1400, 427]. deeply [513]. deeply-impacted [513]. deepwater [577, 893, 1223]. defence [733, 870]. defined [555]. deformities [930, 651, 479, 480, 472]. degradation [858]. degraded [786]. delimitation [952, 358, 240]. Delineation [921]. delphinidin [476]. Delta [1351, 174]. demanding [376]. demersal [1354, 1377, 1168]. demographic [130, 276, 697, 1199]. demographics [60]. demographies [1378]. Demography [917, 547, 286, 1446, 180]. dense [1372]. densities [1117]. density [256, 1422, 200, 197, 448, 1158, 472, 1423, 1087, 723, 325, 961]. dental [481, 171]. Denticeps [833]. Denticipitidae [833]. Denticle [833]. denticles [920]. denticulatus [1112, 564, 1112]. dentine [481, 171]. dependence [1426, 642, 1220]. dependency [536, 1508, 252, 300]. dependent [1374, 1156, 500, 417, 659, 1191, 1445, 282, 453, 221, 1072]. deposited [757]. deposition [3, 1379]. depressed [883]. deprivation [645, 511]. depth [444, 1448]. depth-specific [444]. depths [320, 771]. derived [1097, 759, 1376, 211, 1297, 77, 45, 1013]. dermis [688]. derzhavini [1438]. desaturase [177]. described [341, 260, 119]. describing [1027]. Description [42, 440, 588, 1273, 123, 641, 356, 1421, 624, 1430, 867, 5, 1210, 1438, 373, 1213, 764, 326, 631, 565, 255, 914, 518, 240, 562, 122, 1449, 1084, 257, 1224, 125, 964, 316]. Descriptions [1410, 853, 688, 756, 1048, 1071, 1463, 873]. descriptors [1186]. design [131, 1311, 746, 1413]. designation [1396]. despite [1420, 777]. destruction [513]. details [1380, 613]. detect [1232, 898, 457]. detected [27]. detecting [891]. Detection [969, 1101, 413, 416, 317, 1305, 831]. determinants [362]. Determination [980, 1351, 639, 498, 1191, 227, 203, 592, 715, 1072]. determined [214, 277, 1207]. determines [270]. Determining [454, 186]. devastating [749]. Development [694, 1001, 487, 1500, 247, 112, 868, 245, 483, 63, 87, 1407, 877, 908, 149, 1362, 739, 670, 1011, 1275, 584, 745, 511, 1105, 707, 722, 815, 796, 181, 311, 263, 1246, 346, 179, 397, 892, 69, 1222, 1172, 1102]. Developmental [224, 791, 481, 468, 277, 373, 37, 1451, 1476, 815, 171]. deviations [928]. devices [1442]. devil [214, 919, 1247, 1518]. Diademichthyinae [756]. diadromous [1348, 105]. diagnosed [354]. diagnosis [664, 440]. diameter [346]. Diaphus [1491]. DiceCT [1155]. Dicentrarchus [1481, 1256, 1329, 766, 1034]. Dichloroacetate [296]. did [341, 260]. Diel [554, 854, 59]. Diet [1398, 449, 291, 66, 623, 23, 768, 1510, 696, 223, 277, 1440, 362, 1377, 957, 1445, 1460, 1123, 405, 430, 1219, 1185, 724, 1355]. Dietary [217, 1284, 1171, 918, 258, 1325, 483, 476, 47, 1281, 1306, 703, 645, 1175, 15, 1388, 492, 181]. diets [214, 1353, 1015, 1278]. difference [1179, 707, 1060]. Differences [14, 399, 545, 115, 925, 1476, 841]. different [295, 101, 1214, 1374, 1098, 1156, 949, 387, 1178, 643, 1267, 799, 1306, 592, 521, 771, 503, 598, 1059, 489, 127, 1222, 19, 1104]. Differential [1374, 982]. differentially [1285, 275]. Differentiating [138]. differentiation [17, 502, 371, 358, 1506, 1385, 476, 591, 631, 105]. difficulty [68]. Digestion [760]. Digestive [763, 439, 1374, 658, 1312, 1471, 699, 1380, 26, 649]. digestive/metabolic [439]. digital [138]. dihydrotestosterone [1359]. dimension [225]. dimensional [1109, 140, 1373]. dimensions [1457]. dimidiatus [297]. dimorphic [1156, 382, 452, 1056]. dimorphism [557, 1197, 124]. Dinematichthyidae [664]. Dining [1194]. dinoflagellate [556]. diploid [480, 27, 925]. Diplostomidae [1473]. Dipturus [652, 223, 1268, 682]. Direct [1214, 659, 1303, 813]. Directive [409]. discards [1019]. Discerning [47]. discerns [1348]. discharge [549]. discordant [1405]. discovered [42]. discovery [331, 773, 119, 758, 1223]. discreteness [810]. discriminate [1046, 1344]. Discrimination [428, 88, 116]. discussion [97, 1081, 1496]. discussions [543]. disease [918, 700]. diseased [728]. Disentangling [1044]. disorders [1260]. disparity [1005]. dispersal [1270, 1405, 1239, 241]. Dispersion [108, 899]. displacement [1182]. display [1178]. displays [683]. dissolved [1130, 738, 1499, 1342]. Dissostichus [862, 861]. distance [990, 861]. distances [289]. Distichodontidae [355, 1271]. distinct [14, 47, 674, 1242]. distinctively [1461]. distinguishing [161]. distress [1486]. Distribution [673, 1310, 1111, 401, 1309, 89, 331, 963, 684, 371, 196, 1265, 852, 1354, 1288, 98, 919, 1268, 1302, 27, 83, 44, 1507, 340, 248, 206, 156, 1062, 1089, 828, 1061, 487, 222, 587, 1502, 254, 1425]. distributional [1502]. distributions [827, 1027, 1321]. disturbances [1293]. diurnal [849, 438]. dive [931]. Divergence [190, 1021, 265, 595, 726]. Divergent [1481, 568, 491]. diverse [1447, 1328, 754, 615]. diversification [704, 144, 1419]. Diversity [852, 1159, 350, 655, 121, 787, 355, 1403, 74, 988, 643, 1146, 4, 1180, 1048, 1402, 987, 518, 1043, 259, 1487, 1087, 654, 1122, 702, 1442, 1134, 1176, 1203, 441, 1375, 874, 1459, 456, 254]. Diving [1185]. dmrt1 [209]. DNA [1082, 517, 969, 952, 677, 1398, 1232, 795, 416, 216, 317, 417, 407, 418, 1140, 411, 414, 568, 703, 71, 1361, 704, 424, 1068, 1353, 1134, 1101, 947, 1036, 896, 410, 1502, 547, 713]. DNA-based [407, 410]. do [529, 850, 680, 908, 205, 217, 934, 1473, 225]. Doce [345]. documentation [519]. documenting [415, 229]. Does [501, 665, 766, 698, 69, 800, 1034, 540, 1487, 75, 1209, 1157, 648]. dog [497]. dogfish [302, 67, 776]. dolomieu [910]. dolphinfish [1270, 212]. domain [1359]. domain-like [1359]. Dominance [806, 1154]. dominant [392]. dominated [1272]. Domino [898]. Domperidone [6]. Don [93]. Donsol [461]. dopaminergic [109]. dorsal [832]. dose [997]. dourado [658]. down [976, 1448]. downs [534, 1301]. downstream [1141, 1465, 1440, 365, 63, 1250, 819, 1182]. downstream-migrating [365]. downward [55]. DR [123, 76, 122]. draco [1516]. drag [10]. dragonfish [1103]. dragons [835]. drainage [1006, 120, 588, 907, 70, 1461, 777, 1181]. drainages [14, 1464, 1224, 152]. draining [76]. drawer [944]. DRC [576, 1068]. Dreissena [872]. driftfishes [1335]. driftwood [651, 25]. drive [671]. Driven [137, 237, 530]. driver [1436, 54]. Drivers [484, 849, 503, 98, 287, 1008]. drives [230, 380]. drum [557, 1283]. drunk [795]. dry [86]. Duméril [584]. dumerili [830]. dummy [305]. dunckeri [1479]. duo [749]. duration [1284, 626, 1498]. Durbin [151]. duriense [1135]. during [361, 1323, 868, 1070, 268, 586, 483, 462, 906, 908, 1346, 385, 379, 1281, 706, 248, 739, 1025, 1476, 1011, 744, 333, 187, 1236, 511, 334, 911, 181, 1441, 768, 999, 1347, 1428, 1102]. dussumieri [1089, 77]. Dussumieria [773]. dwarf [1021, 1389, 446]. dwelling [569, 281]. dynamic [1290, 677, 1127, 910]. Dynamics [906, 1070, 271, 1318, 597, 1346, 206, 412, 64, 801, 363, 1441, 1384, 1347, 219, 701]. dynorphin [79]. Dysomma [1466]. Dysommina [1466].

eagle [843, 384, 845, 804]. Early [929, 888, 796, 1083, 1493, 69, 439, 247, 868, 1070, 242, 99, 597, 483, 656, 87, 1274, 791, 908, 1281, 794, 1025, 54, 511, 815, 700, 181, 1278]. early-juvenile [1278]. early-life-stage [656]. Early-season [888]. ears [879, 994]. earth [1405, 279]. east [1484, 845, 1129, 1061, 46, 892, 126, 83]. Eastern [76, 576, 433, 1023, 557, 142, 200, 904, 106, 779, 1512, 558, 124, 991, 1320, 1069, 679, 503, 796, 633, 1014, 693, 627, 378, 1355, 855, 874, 758, 349, 1085, 241, 113, 1270, 650, 484, 134, 801]. easternmost [212]. Eating [696]. Ebert [571]. Echinorhinus [1400]. Echinorhynchus [459]. ecolabel [1276]. Ecological [1382, 1345, 836, 916, 1291, 583, 1293, 1167, 449, 365, 193, 409, 1049, 782, 316]. Ecology [118, 437, 1251, 1132, 963, 942, 802, 223, 1171, 1086, 977, 1296, 1439, 426, 599, 865, 1137, 1491, 939, 1253, 1358, 422, 1423, 1184, 207, 1069, 1133, 1168, 640, 1017, 182, 602, 889, 723, 936, 127, 1131, 465]. ecomorphology [224]. economic [427]. ecophenotype [1488]. Ecoregion [1096]. ecoregions [1239]. ecosystem [891, 735, 935, 362, 1377, 718]. ecosystem-based [935]. ecosystems [933, 534, 602]. ecotourism [369]. ectotherm [1198]. Ectreposebastes [803]. Ecuador [655, 1031]. Ecuadorian [493]. edge [1309]. eDNA [295, 891, 413, 408, 412, 409, 1027]. Eds [233, 93]. Edward [126, 127]. Eel [93, 295, 657, 1187, 57, 832, 1233, 524, 809, 416, 1234, 766, 317, 974, 91, 791, 1273, 1037, 820, 890, 824, 1188, 1228, 1435, 1226, 570, 1077, 1043, 1062, 881, 164, 291, 1236, 646, 777, 1406, 313, 1211, 1059, 999, 1466, 55, 961, 1269]. eel-like [791]. eelpout [1482]. eels [1323, 365, 63, 791, 1250, 1180, 1109, 570, 425, 110, 1467, 883, 961, 93]. Effect [1187, 738, 258, 1135, 48, 997, 922, 608, 1113, 353, 228, 915, 1240, 1214, 536, 252, 1359, 665, 524, 197, 1445, 1175, 592, 1262, 722, 1220, 222, 1504, 489, 538, 225, 330]. effective [418]. Effects [1204, 531, 1233, 918, 741, 1130, 236, 1191, 1064, 953, 528, 53, 1388, 1477, 226, 870, 1413, 313, 1222, 701, 1475, 532, 607, 730, 1244, 912, 560, 137, 732, 593, 659, 1044, 1314, 149, 885, 1110, 1124, 306, 54, 511, 1497, 1503, 209, 1209, 1280, 978, 114, 1292, 818, 179, 304, 173, 960, 767]. efficiency [731, 84]. effort [830, 1111]. Egg [840, 1141, 757, 652, 1003, 524, 1521, 1004, 1044, 2, 149, 1476, 1175, 1208, 1520, 1275, 356, 1067, 1216, 682, 1462, 819, 964, 1223]. egg-guarding [1462]. egg-laying [1004, 2, 682]. Eggs [262, 464, 1073, 1265, 524, 321, 1001, 431, 864, 18, 1174, 711]. Egypt [606]. eicosapentaenoic [1306]. Eigenmann [1224, 380, 1022]. Eigenmannia [388]. eight [1393]. ejaculate [75]. ekmekciae [1028, 1389]. elaborate [320]. Elasmobranch [435, 159, 1115, 298, 1241]. Elasmobranchii [268, 1164, 1514, 590]. elasmobranchs [450, 468, 716, 1232, 1348, 1254, 420]. Elazig [980]. electric [1391, 53, 291, 211, 119, 587, 116]. electric-rays [587]. electroanaesthesia [911]. electrofishing [216]. Electrona [1391]. electronic [1316, 945]. Electrophorus [291]. electrosensory [53, 390]. electrostunning [911]. element [980]. Elemental [1348, 893, 394]. elements [826]. Elevated [549, 1034, 323, 981, 173]. elevates [1058]. elongase [177]. elongatus [810, 195, 1162]. Elovl5 [177]. elucidate [1247]. elvers [164]. emanating [321]. embryo [1514, 965]. Embryonic [189, 722, 537, 247, 586, 87, 192, 584, 115, 179]. embryonic-larval-transition [586]. embryos [450, 930, 523, 560, 86, 520, 1412, 52, 965, 1451, 188, 1334, 964]. Emergence [642]. Emerging [1364]. emission [417]. emperors [812]. empty [675]. Emptying [944]. Encephalon [316]. encounter [232]. encountering [525]. encrasicolus [1046]. end [1199]. endangered [59, 437, 740, 96, 668, 17, 142, 832, 1233, 271, 1398, 280, 779, 1268, 927, 970, 712, 1407, 669, 906, 1257, 52, 1140, 1474, 1138, 134, 156, 682, 215, 1061, 1035, 77, 1211, 723, 380, 1162, 191, 460, 219, 963, 366]. endemic [765, 1251, 17, 1079, 120, 988, 792, 669, 1494, 518, 129, 7, 559, 401, 876, 380, 1013, 1258, 250]. endemicity [1021]. endocrine [268]. endocrinology [360, 593, 499]. endogenous [706]. endorheic [217]. endoskeleton [905]. endotherm [1198]. endurance [145, 323]. enemy [680]. energetics [902, 486, 1241, 1473]. Energy [1158, 1011, 215, 1214, 1374, 343, 659, 1263, 706, 1175, 442, 353, 870, 841]. Engin [1028]. England [647, 1236, 783]. English [68]. Engraulidae [954, 21, 1501, 1046, 1408]. Engraulis [805, 1046, 722]. enhanced [735, 416, 317, 1306]. enhancement [1272]. enhances [549, 523, 1219]. Enhancing [674]. enigmatic [1198, 1038, 1029, 855]. enlarged [97, 388]. enlargement [109]. enrichment [520, 613, 1437, 612]. entering [938]. Enteromius [631, 124, 128]. Entomacrodus [1367]. Entosphenus [56, 253, 1027]. entry [51]. environment [1045, 146, 938, 786, 1072]. Environmental [416, 362, 317, 407, 1200, 418, 132, 411, 1437, 287, 666, 1419, 969, 1354, 1232, 1439, 216, 417, 1505, 525, 1305, 576, 414, 561, 966, 1282, 598, 1230, 1101, 938, 325, 191, 1502]. environments [742, 584, 16, 179]. enzymatic [1312]. enzyme [1055, 1036, 546]. enzymes [1214, 1374, 1312]. eos [230]. Epactionotus [358]. epaulette [445, 1473]. EPC [394]. eperlanus [149]. ephemeral [329]. ephemeridarum [1456]. epibranchial [224]. epidermal [1328, 315]. epidermis [739]. Epigonidae [564]. Epigonus [564]. Epinephelidae [1053, 917]. Epinephelus [727, 1097, 507, 1144, 251]. Episodic [747]. Epitheliocystis [687]. Eptatretus [1161, 599]. Equatorial [695, 146, 9, 1489, 1057]. Equilibrium [447]. era [1026]. eregoodoo [214, 825]. erinacea [3]. Eriocheir [1099]. erk1 [65]. Erratum [201, 942]. Erythrinidae [789]. erythrocyte [608]. escape [80]. Escaped [463]. escapees [718]. Escaping [1436]. esmarkii [1482]. Esox [580, 1305, 1131]. especially [1417]. Essequibo [710]. established [872]. Establishment [455, 77, 1013, 1065, 1139]. estherae [176]. estimate [369, 1446, 319]. estimates [348, 1382, 310, 759, 811, 1376, 34, 1303]. Estimating [1323, 1431, 272, 1371, 112, 1360, 1010]. estimation [1151, 1202]. estradiol [797]. estradiol-17 [797]. estuaries [1505, 393, 300, 1418]. estuarine [1374, 810, 56, 602, 786, 676, 1009]. estuarine-dependent [1374]. estuary [218, 610, 992, 847, 1471, 15, 422, 1390, 1133]. eteline [577, 893]. Etelis [577, 1005]. ethanol [1233]. Etheostoma [14]. ethical [204, 543]. ethylmaleimide [361]. Etmopteridae [160, 162]. Etmopterus [162, 616]. EU [409]. Euchiloglanis [1350]. Eucommia [960]. Eucyclogobius [415, 229]. Euphrates [1389, 152]. EUR [571]. Eurasian [553, 548]. Europe [571, 657, 939, 1282, 1381, 1189]. European [1481, 348, 536, 1256, 57, 1345, 832, 1233, 805, 413, 1329, 252, 741, 100, 416, 1312, 1294, 317, 939, 1250, 1034, 1037, 265, 149, 1046, 1109, 1077, 68, 1186, 164, 1236, 646, 409, 777, 313, 1211]. eurythermal [566]. Eusphyra [1138]. eutrophic [1345]. evacuation [885]. evaluate [703]. Evaluating [732]. Evaluation [1065, 1077, 945, 55, 360, 692, 1305, 114, 1100]. evansii [1103]. Even [1238]. event [1482, 1517]. events [747]. Evidence [243, 1045, 368, 962, 22, 1333, 610, 1271, 1068, 682, 633, 1139, 1399, 876, 662, 1116, 569, 806, 757, 778, 642, 1506, 594, 446, 302, 568, 1365, 395, 68, 1464, 1089, 575, 129, 1170, 1061, 1203, 747, 432, 1192]. evidence-based [569]. Evidences [17]. evolution [757, 21, 1383, 693]. Evolutionary [468, 987, 836, 569, 502, 795, 31, 873, 491, 1049, 1242]. Evynnis [219]. ex [584]. examination [1286]. Examining [173, 882]. example [1467]. Exceptional [786]. Exclusive [372]. Exercise [473, 1317, 296]. Exercise-induced [473]. exhaustion [172]. exhaustive [348]. exhibit [1157]. exiguus [1173]. exilis [43]. existence [568]. Exo [705]. Exo-celiac [705]. exogenous [489]. Exomegas [685, 829]. exotic [731, 1411]. expand [493]. Expanding [1328, 210]. expansion [208, 113, 684, 1366, 971, 747]. expansive [1364]. expensive [909]. experienced [278]. experiment [1207]. experimental [1058]. experimentally [884]. experimentation [75]. experiments [1431, 149, 883]. explain [1487]. explains [1298, 928]. Exploitation [93, 436, 244]. exploited [668, 243, 39, 178, 786, 1243]. exploiting [1400]. exploits [708]. exploration [1295]. Exploratory [1009]. explore [154]. Exploring [1349, 1359, 935, 759, 1195, 227, 315, 1087]. export [775]. Exposure [516, 159, 981, 553, 911, 648]. express [1266]. expressed [275]. Expression [877, 1172, 361, 1156, 730, 141, 720, 1359, 1055, 809, 993, 946, 669, 81, 967, 923, 752, 1080, 427, 1219, 958, 492, 1099, 797, 65, 609, 1283, 222, 173, 546, 19, 1427, 1102, 713, 1217]. expressional [147]. extant [420, 882]. Extending [1260]. extension [710, 1138, 1129, 67, 830, 432]. external [945]. extinct [1128, 469, 1121, 58]. extinction [668, 852, 1212, 244]. extra [542]. Eye [804, 897, 994]. eyed [351, 177]. eyeless [493]. Eyes [994].

F [233, 292, 27, 1313]. F1 [560]. faced [1290]. facilitate [507]. facilitated [241]. facilitates [1151, 447]. factor [361, 904, 1285, 20]. Factors [1250, 626, 895, 1363, 1439, 1346, 1505, 248, 1099, 290, 1492]. facultative [514, 667]. Fads2 [177]. falcatus [50]. fall [289, 1276]. fallai [299]. fam [390]. familiar [850]. Familiarity [800]. families [863]. Family [41, 1458, 160, 336, 1247, 1356, 967, 97, 752, 958, 1394, 492, 882, 559, 947, 1239, 19, 1473]. family-related [1458]. Fang [1413]. fare [1166]. farmed [560, 1127, 533, 479, 1000, 608, 883, 961, 1428]. farming [1329]. farms [459, 1012]. fasciata [1030]. fasciatus [1495, 1437, 1394]. Fast [80, 1426, 1417, 1112, 314]. fast- [1112]. fast-flowing [314]. Fast-start [80, 1426]. fasting [222, 701]. fat [605]. fathead [742]. Fatty [1263, 1149, 980, 1171, 430, 1219, 492, 1477, 646, 177, 1060]. Fauna [716, 126, 1505, 76]. favourable [8]. Fc [1076]. feature [1023]. features [887, 791, 469, 699, 10, 250]. fecundity [360, 1512, 357, 37, 777]. fed [820, 1015]. feed [1127, 1504, 839]. Feeding [1098, 1171, 1491, 1148, 193, 15, 1469, 1334, 646, 676, 1284, 1374, 785, 741, 218, 1178, 426, 1440, 444, 554, 107, 329, 970, 764, 1200, 1425, 820, 1306, 514, 556, 638, 207, 9, 1472, 221, 487, 675, 228, 346, 768, 999, 1259, 1207]. feel [1195]. felis [575]. Female [392, 176, 1057, 931, 359, 1447, 292, 1097, 560, 268, 157, 1004, 1263, 72, 698, 1175, 1107, 75, 312]. female-sign [1107]. females [450, 367, 431, 1147, 672, 700, 1462]. feminization [610]. fenthion [1475]. Fergana [621]. ferganaensis [621]. Fernando [695, 991]. ferox [423, 1366]. fertilisation [981]. fertilize [1160]. fertilizing [603]. fetal [199]. fibre [1015]. fibroblast [960]. fibrosis [844]. fidelity [195, 136, 762, 364]. Field [200, 891, 112, 606, 1005, 114, 1303, 571]. field-based [1005]. fields [813]. fighting [1147]. file [944]. Filho [459]. Filling [603]. fin [832, 1018, 672, 77, 45]. Finding [1454]. fine [849]. fine-scale [849]. finetooth [360]. finfish [1377]. fingerlings [1317]. finned [28]. First [1116, 1484, 652, 1079, 310, 384, 264, 90, 650, 249, 564, 519, 1331, 778, 1438, 160, 1521, 1312, 1004, 1518, 1257, 1000, 340, 1415, 831, 68, 1089, 294, 1336, 1335, 1092, 93, 433, 263, 1397, 1235, 1189, 1259, 341, 589, 757, 867, 42, 260, 940, 1366, 1210, 1213, 1137, 302, 1107, 1031, 830, 298, 934, 513, 1276, 319, 487, 1032, 1223, 64, 1394]. Fish [557, 677, 1395, 1146, 1505, 284, 702, 828, 410, 118, 300, 330, 1052, 655, 532, 1298, 1204, 256, 341, 529, 1192, 517, 1198, 872, 1098, 318, 113, 261, 1290, 17, 969, 1345, 1465, 159, 260, 1403, 1073, 1265, 392, 1070, 501, 153, 218, 1354, 137, 126, 70, 1221, 416, 983, 1439, 484, 1328, 681, 566, 479, 643, 86, 731, 763, 1135, 628, 1091, 216, 459, 766, 317, 671, 1191, 407, 849, 1195, 754, 1407, 1200, 1478, 175, 448, 906, 418, 737, 613, 1311, 1000, 1507, 438, 1455, 1425, 558, 76, 396, 1038, 72, 414, 739]. fish [170, 1197, 1272, 1387, 514, 1402, 1457, 205, 614, 1264, 697, 606, 262, 133, 1371, 592, 695, 521, 617, 210, 412, 461, 1047, 510, 53, 11, 775, 1011, 1078, 51, 289, 1186, 34, 176, 1159, 109, 846, 934, 654, 497, 612, 1067, 1166, 66, 409, 75, 1404, 1152, 184, 443, 771, 859, 1168, 870, 770, 7, 143, 1209, 945, 1015, 598, 869, 311, 1083, 1493, 18, 1249, 1442, 666, 798, 873, 238, 1409, 401, 1467, 35, 1058, 676, 909, 947, 335, 837, 1504, 874, 383, 1013, 1418, 467, 783, 1502, 782, 254, 728, 701, 711, 662, 27]. fish-based [409]. fish-community [414]. fish-health [1328]. fish-plant [1504]. fished [727]. fisheries [529, 696, 1019, 50, 750, 218, 280, 1400, 812, 123, 408, 1026, 543, 205, 262, 528, 879, 8, 1297, 893, 244]. fishery [949, 912, 919, 1177, 1254, 1358, 966, 1276, 495, 180]. fishes [956, 655, 549, 1248, 531, 639, 1470, 243, 1447, 1291, 1096, 902, 694, 112, 1293, 852, 868, 278, 940, 1434, 677, 977, 904, 1167, 533, 236, 826, 107, 270, 551, 49, 400, 132, 1356, 1391, 1155, 82, 446, 928, 703, 1494, 146, 1158, 469, 1107, 528, 535, 991, 217, 1184, 985, 746, 1145, 1169, 1320, 687, 84, 1335, 1040, 1262, 1492, 638, 726, 129, 1496, 1230, 882, 1442, 1224, 1357, 495, 1292, 119, 1419, 116, 290, 312, 1339, 499, 118]. fishing [1132, 772, 1346, 539, 674, 702, 1243, 1363]. fishway [1413]. fission [701]. fitness [537, 192, 1024, 777]. Five [606, 1248, 44, 594, 637, 903, 239]. fixation [517]. flannelmouth [1326]. flapper [652, 1268, 682]. flat [739, 445]. flatfish [1374, 43]. flatfishes [196, 1253]. flathead [1459]. flats [50]. flavipinnis [1242]. Flexibility [825, 49]. flexible [33]. floating [1329]. flooding [1432]. floodplain [1387, 8]. floodplains [1078]. Flora [716]. Florida [64, 1468, 1363]. flossada [223]. flounder [1108, 1445, 578, 1017]. flow [1135, 793, 1110, 548, 1487, 838]. flow-modified [1487]. flowing [1265, 381, 314]. flows [343, 732]. fluctuating [399]. fluctuating-asymmetry [399]. fluctuation [1267]. Fluctuations [448, 1191]. fluid [321, 379, 75]. fluorescence [1492, 982]. fluorescent [475]. fluviatilis [348, 536, 1345, 252, 741, 1003, 100, 73, 553]. focus [1393, 613]. focused [764]. foetuses [1213]. Fold [1096]. foliaceus [363]. follicle [242]. follicular [617]. follow [835]. following [1511, 490, 1272, 147, 804]. fontinalis [925, 1451, 231, 851]. Food [1017, 218, 444, 393, 511, 410]. forage [11]. Foraging [270, 500, 1352, 596, 84, 1489, 457]. forensic [2]. forever [1198]. forgotten [444]. Form [1330, 791, 683]. formalin [517, 34]. formalin-preserved [34]. formation [620, 868, 328]. formosa [624]. forms [398]. forskalii [987]. Forsskal [1388]. forsteri [88]. fossil [1296, 420]. fossilis [413]. Foul [497]. found [264, 742]. four [1048, 1463, 15, 562, 1279, 180, 1033]. four-spotted [1033]. fourcroi [745]. foxl2 [209]. fragility [608]. fragmented [808]. Framework [409, 902, 1296, 543]. frameworks [782]. France [1465]. francisci [1150]. Francisco [7, 1390, 513]. Fraser [364]. freckled [88].

Fredrick [377]. free [1265, 1034, 1026, 804]. free-flowing [1265]. free-living [1026]. free-ranging [804]. free-swimming [1034]. Freezing [75]. French [812]. frequencies [453]. Frequency [515, 1304, 972]. Frequent [334]. fresh [1140, 584]. Freshening [538]. Freshwater [51, 1230, 118, 655, 1481, 658, 1447, 1291, 1403, 1073, 852, 940, 153, 1328, 1127, 216, 927, 1407, 175, 131, 545, 996, 570, 1402, 534, 745, 187, 1236, 497, 858, 1152, 226, 1429, 770, 1209, 1493, 1061, 401, 1419, 397, 675]. friction [10]. frisii [359, 431]. frog [460]. Frogfishes [465]. Front [862]. frontiers [1349]. fry [755, 854]. fuelling [825]. fulleri [601]. fulvescens [730, 732, 550]. fulviadraco [967]. fulvidraco [958]. function [388, 718, 470, 183, 1330, 1240]. Functional [1487, 177, 1296, 643, 1267, 1420, 1387, 702, 544, 767]. functioning [735]. fundamental [208]. Fundulus [1508, 80, 1426]. fungal [728]. Further [806, 964, 1061, 1203]. furzeri [325, 1206, 839]. fuscata [504]. fuscoguttatus [1097]. fusilier [1356]. fusion [701]. future [655, 1198, 612, 409, 1292].

G [233]. G. [383]. Gabon [440, 128]. Gadiformes [1449]. Gadus [444, 329, 453]. Galápagos [264, 1234, 1340, 929]. Galaxias [608, 582, 105, 1459]. galaxiid [437, 1179, 133]. galeatus [942, 802, 651]. Galeocerdo [650, 903]. Galeus [298]. Gambusia [6, 1488, 746]. gamete [981]. gametogenesis [237, 530]. ganglion [1122, 1490]. gaps [1073, 603]. gar [585]. GARP [1201]. garra [1016]. gas [738, 1342]. Gasterosteus [1151, 790, 337]. gastric [1164, 885, 1380]. gastrointestinal [107]. gattorugine [1119]. gelatinous [1335]. gender [750]. Gene [361, 1283, 713, 641, 967, 923, 752, 548, 958, 251, 797, 609, 222, 19]. genera [1071, 1466]. General [688]. generalised [677]. generalist [939]. generates [120, 1049]. generating [116]. generation [1126]. generic [122]. genes [331, 524, 1055, 993, 1285, 1305, 1077, 209, 1149, 275, 1102, 713, 1217]. Genetic [502, 714, 1097, 162, 130, 83, 1180, 265, 419, 259, 704, 726, 1170, 432, 456, 941, 17, 1341, 243, 90, 358, 1385, 285, 86, 1021, 799, 1273, 641, 39, 136, 52, 558, 508, 762, 68, 595, 1106, 1087, 453, 503, 105, 1176, 758]. genetics [1443, 468, 729]. Genidens [262]. geniguttatus [362]. genital [1224]. Genome [1151, 993, 967, 752, 958, 1375, 19, 720, 186, 568, 1074]. Genome-wide [1151, 993, 967, 752, 958, 1375, 19, 720, 1074]. Genomic [1125, 862]. genotoxicity [837]. genotype [871]. genotyping [161]. genus [350, 664, 952, 805, 355, 440, 907, 1225, 795, 1161, 995, 1410, 736, 1231, 326, 824, 1188, 1228, 1435, 1226, 421, 340, 685, 255, 1071, 1415, 1351, 823, 150, 512, 356, 1394, 370, 388, 144, 283, 1051, 829, 1496, 143, 1029, 1421, 1134, 1203, 1224, 693, 391, 239, 734, 485, 125, 108]. Genypterus [524]. Geographic [1309, 670, 89, 366, 503]. Geographical [792, 1385, 1360, 198]. geometric [1434, 1330, 225]. Geophagini [316]. Geophagus [1041, 316]. George [1460]. Gerais [16]. Germ [669]. Gerreidae [1434]. gestational [563]. get [297, 542]. Getting [96, 712]. GH [1219]. GH-GHR-IGF [1219]. ghost [452]. GHR [1219]. ghrelinergic [645]. giant [1513, 1234, 1091, 927, 608, 582, 1062, 1388, 232, 1061]. giants [882]. gibbosus [1106]. Gigantactis [1448]. Gilchristella [1009]. Gill [928, 1350, 667, 636, 404, 561, 685, 699, 381, 585, 829]. gilled [264]. gills [1013]. gilthead [324, 258, 399, 194, 733, 462, 482, 1149]. Ginglymostoma [442, 1090]. Girard [880]. Gironde [1465]. gladius [596, 1376, 905, 579, 547]. gland [427, 1380]. glanis [1312]. Glass [999, 1323, 766, 570, 388, 1467]. glauca [1116, 157]. Glaucostegus [1398]. glimpse [1132, 1137]. Global [1073, 1429, 532, 531, 533, 236, 1302, 528, 1417, 1280, 1239]. Glossogobius [551, 344]. glucagon [725, 1427]. glucan [48, 1468]. glucose [1503, 725]. glutamine [1477]. glutathione [1510]. Glycan [427]. Glyptonsternine [1071]. Glyptosternum [705]. Gmelin [608, 582]. Gnathopogon [669]. go [835]. goatfish [869]. gobies [277, 995, 33, 446, 1043, 634, 1370]. Gobiesocidae [519, 756, 383]. gobiid [1189]. Gobiidae [1032, 551, 1038, 1383, 344, 1089, 1386, 177, 1473]. Gobiiformes [1421, 1032]. Gobius [1038]. goby [500, 1352, 1200, 1281, 600, 344, 1386, 187, 1173, 1016, 177, 1032, 1189, 330, 1295, 1339, 1473]. goggle [177]. goggle-eyed [177]. golden [1480, 140]. goldfish [1422, 645, 797, 489]. goliath [1144, 405]. Gomes [954]. gomesi [459]. gonad [1500, 1263]. Gonadal [707, 1347, 974, 670, 1173, 251, 1475]. Gonadosomatic [1010]. gonadotropin [209]. Gonorhynchiformes [1068]. good [435, 517, 498, 1387]. Goode [241, 899]. goodea [502, 502]. Goodeidaea [502]. gorbuscha [1001, 361, 1440, 158, 769, 1139, 1399]. Gordon [1452]. Gorgasia [281]. Gorgeous [1386, 1370]. Gorges [1459]. Gorgona [367, 433]. Gouania [519, 383]. gourami [1030]. gradient [1230]. gradients [80, 1354]. grading [1187]. Gram [1266]. Gram-negative [1266]. Granada [221]. grandis [1508]. grandispinis [951]. grandoculis [897]. graphic [334]. graphicus [334]. grass [923, 960, 1427, 767]. gravid [1213]. gray [1450]. Great [697, 1484, 163, 979, 1352, 286, 815, 1183]. greater [591, 1482, 1516, 830]. Green [580, 431, 1390, 896]. Greenland [74, 158, 1280, 1235, 1355]. greenling [1122]. greenside [14]. Gregory [1251]. grenadiers [23]. grey [661, 488, 140, 1239]. griseofasciatus [1053]. griseus [264, 115]. gross [316, 1496]. ground [1132, 785, 444, 1201]. grounds [682, 1243]. Group [486, 358, 429, 551, 987, 46, 715, 546, 701, 220]. group-living [701]. grouper [1097, 507, 398, 1144, 1053, 286, 251, 917]. groupers [984]. groups [1098, 618, 1034]. grow [888, 1417]. growing [1112]. Growth [735, 811, 188, 776, 769, 211, 1284, 101, 537, 247, 1187, 668, 269, 639, 727, 299, 730, 1097, 620, 1349, 918, 218, 382, 904, 642, 630, 586, 197, 14, 656, 659, 1274, 1407, 792, 644, 490, 1126, 641, 3, 1037, 794, 1314, 545, 1362, 1437, 1351, 1075, 540, 192, 1219, 1388, 1024, 54, 1360, 1477, 660, 1105, 895, 1163, 1429, 722, 815, 1202, 945, 869, 41, 1196, 488, 1498, 487, 378, 1381, 1504, 883, 961, 1222, 1258, 915, 1510, 960, 1408]. Growth-enhanced [735]. grunt [894]. GSH [1269]. Guadalupe [369]. Guadiana [1397]. Guangxi [13]. Guaratuba [241]. guarding [1462]. gudgeon [318, 1177, 1194]. Guiana [389]. Guide [571]. guides [1206]. guilds [638, 676]. guilinensis [13].

Guinea [1288]. guitarfish [740, 1277, 1398, 367, 1075, 433]. guitarfishes [1086, 433]. Guizhou [1414]. Güldenstädt [1028]. Gulf [219, 688, 269, 1508, 1495, 1318, 280, 268, 1137, 656, 591, 955, 422, 921, 178, 287, 984, 1472, 1017, 1205, 241]. gunnari [1263]. Günther [355, 778, 340, 303, 228]. guppies [376, 403]. guppy [98, 54]. gurnard [674]. gut [1345, 1312, 48, 695, 1184]. gut-content [1345]. guttatus [1430, 692]. guttulatus [1384]. Guyana [710]. Gymnocanthus [108]. Gymnocephalus [397]. Gymnocypris [1013]. Gymnoscopelus [41]. Gymnotidae [291]. Gymnotiformes [388, 452]. Gymnura [1213, 47, 1123]. gynogenetic [1151].

H [1452, 233]. Habitat [1248, 1096, 692, 755, 671, 82, 287, 842, 765, 916, 963, 785, 384, 368, 271, 761, 200, 1333, 270, 155, 73, 175, 814, 1062, 1432, 910, 461, 425, 898, 1050, 646, 858, 771, 322, 1340, 1406, 401, 1150, 1049, 676, 38, 1206, 1372, 1069]. habitats [709, 1436, 50, 1378, 803, 414, 921, 1249]. habiting [505]. habits [47, 584, 1472, 676]. habituation [666]. haddock [892]. Haematological [505, 582, 997, 608, 837]. haematological-biochemical [997]. haemulid [971]. haffara [509]. hagfish [1161, 599]. Haiti [143]. Haizhou [1267, 1104]. hake [785, 387]. halibut [373, 1137, 1316, 1301, 1280, 1355]. hamatus [737]. Hamilton [884, 419, 473, 1380, 982]. hammer [976]. hammerhead [59, 931, 761, 280, 1063, 979, 206, 199]. hammerheads [163]. Handbook [526]. handfish [1470]. handling [1252]. Haplochromis [332]. haplotilapiine [715]. haplotype [453]. hard [415, 229, 1376]. hardhead [575]. harengus [972, 1044, 186]. harlequin [1347]. Harmful [556]. harness [463]. Harpagifer [538]. harvested [657]. hassi [281]. Hatch [1274]. hatchery [515, 912, 983, 525, 1024, 1153, 184, 1519]. hatchery-origin [1519]. hatchery-reared [515, 912, 983, 525, 1024]. hatching [1004, 680]. hatchlings [372]. haven [740]. having [901]. Hawai’i [1327]. Hawai’ian [49, 33, 1125, 1258]. hawkfish [88]. haystack [782]. head [1041, 1310, 144, 578, 29, 767]. headed [1491]. headfirst [567]. headwater [16]. healing [159, 804]. health [1328, 1474, 539]. healthy [48]. hearing [1255]. Heart [385, 1428, 759, 462, 573]. Heat [1476, 721, 923, 870]. heatwave [1294]. heaviest [1091, 882]. heavily [178]. heavy [1082, 246, 637]. Heemstra [1092]. height [965]. Heishui [1413]. Hellenic [565]. helleri [1218]. help [542]. helps [203]. Hematological [1468, 1388]. Hemiclepsis [1334]. Hemigrammus [97]. Hemigymnus [670]. hemioliopterus [1148]. Hemiscylliidae [268]. Hemiscyllium [445]. Hemitrygon [563]. Henle [1277, 1004, 1075]. henlei [886]. hepatic [492]. hepatocyte [960]. hepatocytes [923]. hepatus [1349]. heptapterid [319]. Heptapteridae [1022, 569, 1315, 753]. hepuensis [1099]. her [698]. herbertaxelrodi [1152]. herbivorous [514, 1145]. here [182]. Herichthys [144]. heritability [1126]. Heritage [1519]. hermaphroditic [727]. herring [269, 1044, 929, 186, 833, 972, 786, 487, 1009]. Heterochrony [1383]. heteroclitus [80, 1426]. Heteroconger [281]. Heterodontus [678, 942, 802, 1150]. heterogametic [186]. Heterogeneity [1373, 102, 1049]. heterogeneous [701]. heterosis [491]. heterospecific [1511, 134, 798]. heterospecifics [133]. Heterotis [1317]. Hexagrammidae [1122]. Hexagrammos [1122]. Hexanchidae [264]. Hexanchiformes [264]. Hexanchus [264, 115]. Hidden [121, 395, 355, 1088, 1203]. Hierarchical [37, 771]. High [1422, 1118, 548, 561, 893, 1205, 819, 1243, 1473, 727, 159, 120, 498, 1117, 1420, 88, 777, 1280, 275, 1141, 340]. high-CO [159]. High-density [1422]. High-resolution [893, 1205]. High-throughput [1205]. high-tide [275]. higher [483, 1516, 181, 961]. highest [1405]. Highlands [910]. highly [17, 1361, 1353, 660, 401]. hilsa [245, 847]. Himalayan [505, 1246]. Himalayas [1405, 1415]. Himantura [1514]. hindcast [149]. Hippocampus [880, 271, 992, 712, 1384]. hippoglossoides [1280, 1355]. Hippoglossus [373, 1316]. Hippopotamyrus [595]. Hipposcarus [848]. hippurus [1270, 212]. histochemical [699, 26]. Histological [1380, 194, 974, 446, 482, 1286]. histology [510, 1388]. Histomorphological [268]. histopathological [637, 837]. Historical [98, 732, 1254, 1002, 1081, 936]. historically [555]. history [1309, 1227, 569, 57, 502, 727, 1509, 271, 382, 74, 795, 988, 812, 130, 276, 597, 995, 929, 1281, 33, 697, 1043, 1324, 344, 848, 897, 984, 92, 211, 1054, 1406, 1292, 723, 244, 69, 1258, 456, 1308]. hit [749]. hobby [1041]. hoevenii [561]. hogfish [1258]. holbrooki [746]. Holding [808]. holistic [1268]. Holland [703]. hollandi [703]. Holocephalan [481, 171]. Holocephali [249]. holotype [685]. holubi [1372]. Homatula [914]. home [898, 573, 663]. home-bound [663]. homeostatic [379]. homeward [361, 989]. homing [525, 822]. homogenates [733]. homogeneity [506]. homogeneous [701]. Honduran [303]. Honghu [1408]. Hongshui [1407]. Honmorokoo [669]. hook [949]. hooklets [672]. hope [1212]. Hopkins [465]. Hoplias [789]. Hoplostethus [947]. Hora [255, 1016]. horkelii [740]. hormone [1284, 641, 251, 1172]. hormones [157, 489]. Horn [156, 942, 802, 1150]. horse [83]. host [343, 865, 948, 754, 965, 1387, 1334, 1335, 363, 1473]. hotel [1394]. hotspot [655, 971, 60]. hotspots [1423]. hound [371, 762, 66, 886, 92]. house [1238]. housekeeping [1305, 1077]. housing [1307]. Hoven [561]. Howellidae [1051]. hsinglungshanensis [581]. Huallaga [391]. Huang [914]. Huang-He [914]. Hubbs [618]. hubbsi [785, 387]. Human [241, 933, 651, 1364, 936]. Human-facilitated [241]. humboldtianum [276]. humeralis [894]. humphead [554]. humpnose [897]. hunt [858]. hunting [335]. husbandry [1307, 1252]. Huso [209]. hybrid [1284, 1097, 459, 27, 491, 1100, 649]. hybridisation [129]. Hybridization [941, 1097]. hybrids [480, 491]. hydriforme [864]. hydroacoustic [1507]. Hydrocynus [1325, 131]. hydrogen [523]. Hydrolagus [242]. Hydrological [380, 8]. hydrolysate [1388]. hydrophila [967]. hydropower [625, 1109]. hydropower-regulated [625]. hydrostatic [275]. Hylopanchax [440]. Hypanus [1079, 692, 283]. hypermineralized [481, 171]. hyperoxia [1058]. Hyphessobrycon [151, 1152, 16]. hypomelanosis [1090]. Hypophthalmichthys [224, 738, 793, 71]. hypophthalmus [922]. Hypopthalmichthys [1342]. Hypoptopomatinae [1056]. hypoptopomatine [679]. Hyporthodus [1053]. Hypostominae [220]. Hypostomus [555, 513, 629, 24, 220]. hypothalamic [230]. hypotheses [782]. hypothesis [665, 1365, 1357]. hypoxia [537, 566, 860, 53, 192, 43]. hypoxia-tolerant [43]. hypoxic [1221].

I3K413 [913]. I3KL45 [913]. Iberian [1472]. Icariin [478]. ice [86, 889]. ice-covered [889]. icefish [1263, 926]. Iceland [1399]. Icelandic [1425]. ichthyofauna [1420]. ichthyofaunal [1288, 126]. Ictalurus [993]. idella [923, 1427]. idellus [960]. Identification [280, 20, 186, 1038, 1080, 1503, 1409, 546, 1023, 1210, 993, 551, 47, 52, 85, 272, 967, 752, 1005, 958, 1205, 19, 725, 728, 711, 1009, 720]. identifications [517]. identified [1482, 1014]. identifies [267, 1378, 283]. identify [396, 782]. identities [882]. identity [1030, 255, 257]. IGF [1219]. ignobilis [1388, 660]. Iguacu [907, 774]. Iguanodectidae [257]. Iguape [99]. Ihering [555]. III [546]. IL-8 [767]. ilisha [245, 847]. Illuminating [1054]. images [138]. imaging [475, 274, 1409]. Imbabura [1031]. immaculatus [345]. immigration [549]. Immune [733, 918, 1095, 48, 78, 844]. immunity [737, 553, 978, 491, 1510, 1217]. immunoactivities [394]. immunological [431]. impact [1382, 933, 137, 651, 1499, 1311, 746, 497, 858, 1282, 938, 325]. impacted [513]. Impacts [772, 1291, 1439, 533, 1294, 846]. impair [648]. Impaired [777, 159]. impairs [981, 556]. Imparfinis [1022, 1315]. imperilled [1096]. implantation [91]. implanted [1037, 290]. implants [209]. implementing [1232]. implication [92, 180]. Implications [131, 1026, 510, 261, 502, 1345, 1470, 620, 1021, 1514, 605, 438, 1474, 1320, 238, 495, 179, 69, 758, 312, 1373, 1413]. importance [435, 1444, 329, 558, 427, 944]. important [916, 1132, 111, 123, 525, 461, 286, 1498, 180]. impounded [895]. improve [614]. Improved [52, 1444, 923]. improves [1437, 323, 394, 1510]. improving [1183]. imus [803]. In-river [996]. inappropriate [376]. incapacity [177]. including [1411, 991, 140, 964]. income [1263]. Inconsistent [204]. increase [532, 766, 438, 635, 1015]. Increased [1480, 864, 830, 1429]. increases [102, 990, 1368]. increasing [1333]. incremental [172]. incubation [652, 840, 323, 1476]. Incursions [747, 1481]. Independent [526]. index [1360, 619, 1010]. Indexing [1474]. India [1430, 588, 70, 824, 1188, 1435, 1254, 421, 419, 512, 881, 1011, 1016, 60, 1061, 1421, 773, 1424]. Indian [644, 1356, 322, 640, 138, 505, 890, 1226, 147, 1336, 638, 1051, 1134]. indica [640]. indicate [242, 1263, 150, 231]. indication [1107]. indications [1239]. indicator [315, 443, 702, 1428]. indicators [207, 619]. indices [553, 1297, 489]. indicum [890]. Indirect [607]. Individual [429, 102, 267, 1314, 1160, 1245, 88, 938]. individuals [1119, 161]. Indo [577, 974, 1126, 39, 670, 971]. Indonesian [630]. induce [189]. Induced [87, 1079, 6, 476, 793, 477, 133, 54, 478, 473, 226, 1099, 837, 844]. induces [1299, 109]. industrial [702]. inequities [750]. inertia [296]. infected [864, 1047]. Infection [1456, 720, 997, 1106]. infections [1266]. infer [727, 1296]. inference [1202]. inferences [1186, 1489]. inferred [1270, 1171, 1434, 729, 861, 957, 1304, 704, 855]. infestation [1473]. inflammatory [767]. inflicted [804]. inflowing [425]. Influence [247, 1055, 992, 816, 1272, 765, 607, 788, 501, 1135, 680, 1200, 794, 1365, 146, 647, 770, 311, 723, 191]. influenced [1505]. influences [21, 1420, 840]. influencing [1250, 626, 1492, 290, 1363]. influential [1346]. influx [490]. inform [396]. Information [1, 32, 61, 94, 117, 135, 168, 202, 234, 266, 308, 342, 375, 406, 434, 466, 496, 527, 541, 572, 611, 653, 690, 717, 748, 780, 807, 834, 857, 878, 900, 932, 943, 975, 1007, 1039, 1066, 1093, 1120, 1142, 1165, 1193, 1215, 1237, 1261, 1289, 1300, 1337, 1369, 1401, 1416, 1453, 1485, 1136, 1425, 1119]. informing [1159]. ingens [880]. ingestion [567]. inhabiting [402, 1011, 55]. Inherent [431, 408]. inheritance [27]. inhibiting [476]. Inhibitory [213, 1285, 403]. injuries [804]. inland [1264, 370]. Inle [1021]. Innate [553, 134]. inner [1483, 879, 346]. Innervation [1018]. innovations [543]. insectivore [632]. inseminating [328]. inshore [654]. insight [1353, 1150, 1059]. Insights [813, 1345, 1254, 1379, 767, 1309, 384, 795, 1312, 83, 422, 863, 1134, 1054, 495, 827, 1088]. instability [191]. insular [303, 1469, 1057]. insulin [1157, 725, 1427]. intake [444]. integrated [1171, 741, 1037, 818]. integrating [193]. Integrative [555, 355, 283, 1163, 773, 1022, 562, 1068]. integrity [86]. integumentary [281]. intense [280]. intensify [1218]. intensities [438]. intensity [247]. intensive [1329]. Intentional [335]. intentionally [463]. intentions [498]. inter [500, 290]. inter- [500, 290]. interacting [369, 1080]. Interaction [1284, 1342, 696, 948, 968, 1334, 1152]. Interactions [238, 507, 618, 1377, 449, 695, 1088, 936]. interactive [140]. Interannual [206]. Interdecadal [1420]. interferon [1266]. interferon-I [1266]. interior [364]. interleukins [1266]. intermedius [652, 1268, 682]. internal [1247, 1394]. internalized [1198]. International [93]. Interpopulation [1404, 1182]. interpreted [1316]. intersex [263]. interspecific [480, 641]. interstitial [379, 1361, 383]. intertidal [277, 638, 228]. intervention [1096]. intestinal [918, 1047, 1510, 1240]. intestine [1312]. intimacy [162]. Intra [329, 968, 641]. intra- [641]. Intra-annual [329]. intra-specific [968]. intramuscular [960]. intraretinal [959]. Intraspecific [403, 535, 956, 493, 500, 1464, 1397, 290]. introduced [799, 508, 190, 1154, 1131]. introductions [71]. invaded [132]. invaders [60]. invalid [829]. invariances [985]. Invasion [1519, 1436, 1382, 1341, 1006, 447, 65]. invasive [208, 235, 341, 872, 1382, 318, 260, 1291, 1328, 500, 1001, 651, 799, 1302, 1177, 68, 801, 769, 1262, 1497, 351, 1209, 851, 460]. inverted [1448]. Investigating [218, 817, 1025, 1489]. investigation [1520, 1223]. involved [913, 715, 1076, 1269]. iodine [1155]. Ion [645, 81]. Ion-poor [645].

Irciniidae [1394]. Ireland [1366, 423, 301]. iridescence [1143]. Iridescent [1147]. Irish [963, 1392, 1212]. irrasa [964]. irrigation [1184, 1338]. irwini [380]. ISBN [465, 526, 233, 93, 571]. ishikawae [683]. Island [644, 1460, 1248, 650, 303, 9, 1035, 369]. Islands [635, 1423, 1016, 1505, 991, 58, 426, 1234, 1333, 929]. isodon [360]. isoforms [246]. isolated [707, 728]. isolation [120, 968]. isothermal [1305]. Isotope [18, 832, 1270, 277, 1440, 1325, 1390, 1184, 306, 454, 1185, 770, 221]. isotopes [214, 957, 2, 695, 363]. Isotopic [1052, 1227, 450, 520, 1207, 1047]. Issue [1, 32, 61, 94, 117, 135, 168, 202, 234, 266, 308, 342, 375, 406, 434, 466, 496, 527, 541, 572, 611, 653, 690, 717, 748, 780, 807, 834, 857, 878, 900, 932, 943, 975, 1007, 1039, 1066, 1093, 1120, 1142, 1165, 1193, 1215, 1237, 1261, 1289, 1300, 1337, 1369, 1401, 1416, 1453, 1485, 1194, 118]. Isthmin [1503]. Isthmin-1 [1503]. Istiompax [640]. Istiophorus [866]. istlanum [392]. Isurus [1202, 1303]. itajara [1144]. Italian [518, 1124]. Iterative [345, 1315]. itself [411]. Ituglanis [1496]. IUCN [219]. IV [923]. Ivaí [40]. iyonis [1275].

J [465]. jack [417, 1274, 490, 259, 660]. Jackson [678]. jacksoni [679]. jamaicensis [1196]. Japan [554, 570, 1519]. Japanese [295, 1274, 791, 794, 820, 570, 1445, 1062, 425, 578, 1406, 1059, 723, 55]. japonica [791, 820, 570, 1062, 425, 1406, 1059, 55]. japonicus [417, 1274]. Jari [257]. javelin [971]. jawed [468]. jawfish [1275]. Jeddah [681]. Jenyns [398]. Jenynsia [340]. Jequitinhonha [16]. jewelled [1147, 1143]. Jiang [950]. jigsaw [24]. jnk1 [65]. John [465, 233, 1131]. jordani [1407]. Joshua [526]. Journal [1395, 118]. journey [279]. junction [246]. junior [1028, 590]. jurisdictional [903]. Juruena [314]. Just [1450, 754]. Juvenile [1465, 985, 1372, 813, 148, 549, 607, 871, 372, 688, 918, 1130, 1244, 983, 1440, 1378, 1333, 847, 593, 659, 968, 1274, 791, 1034, 1037, 1287, 840, 817, 145, 739, 323, 521, 1423, 1432, 894, 1388, 356, 1360, 1133, 1202, 1015, 666, 1150, 1278, 1379, 1357, 1182]. Juvenile-to-adult [985]. juveniles [1214, 439, 1374, 417, 594, 1344, 1174].

kaakan [971]. Kahuzi [76]. Kahuzi-Biega [76]. kaibarai [723]. Kajikia [322]. Kamchatka [988]. Kanehira [704]. karyotypes [1361]. kasmiana [1334]. Keap1 [923]. Keban [980]. keep [1450]. Keeping [589, 1282]. kelberi [447]. kenojei [1004]. Kerala [60]. Kerguelen [426, 964]. keta [521, 822]. ketotestosterone [209, 797]. key [654, 433, 627]. kg [1091]. kidney [887, 1310, 578, 873, 767]. killifish [1508, 583, 618, 197, 680, 1361, 584, 1278, 325, 1206, 191, 839]. kills [1264]. kinase [993, 967, 1080]. kinds [784]. kinematics [875]. King [1460]. Kingdom [941, 1366]. Kirankaya [1028]. kishinouyei [1350]. kisspeptins [489]. kisutch [156, 1153, 364]. kitefin [565]. kitt [597]. Klamath [508]. km [303]. kner [257]. Kneriidae [1068]. knife [1332]. knifefish [388, 452]. knight [1143]. Knodus [1229]. knowledge [1509, 436, 193, 674]. known [1041, 1089, 616, 874]. koelzi [103]. kokopu [608, 582]. Kong [1004]. Korean [965]. Kottelat [1028, 588]. Kouilou [125]. Köycegiz [1279]. Krefftichthys [41]. Krka [42]. krkae [1020]. Kronichthys [441]. Kryptolebias [603]. kurper [120]. kutum [359, 431]. Kyphosus [1480].

L [258, 319, 1381, 725, 1428]. L. [1065, 1481, 657, 980, 494, 557, 1178, 429, 285, 423, 357, 379, 1037, 922, 817, 996, 635, 1110, 1025, 987, 647, 68, 1087, 178, 1503, 777, 1036, 378, 397, 1074, 999, 546, 1100]. labelling [1140]. Labeo [884, 987, 1266, 147, 65, 982]. Labeobarbus [576]. Labeotropheus [185]. labiatus [505]. Laboratory [1150, 1192, 766, 662]. Laboratory-based [1150]. Labrador [747]. labrax [1481, 1256, 1329, 766, 1034]. Labridae [686, 334]. Labroides [297]. Labrus [1499, 816, 1306]. Laccadive [1356]. Lacépède [513]. lack [1361, 606, 298]. Lacunae [866]. lacustrine [924]. ladder [1370]. laevis [286]. Lago [1031]. Lagoon [1146, 1279, 415, 1006, 1323, 229, 425, 602, 1384]. lagowskii [1240]. lake [872, 1345, 292, 730, 732, 1021, 155, 1507, 273, 817, 895, 18, 179, 173, 550, 1436, 126, 1021, 396, 176, 185, 143, 784, 1181, 1042, 127, 1013, 937, 1408]. lake-produced [18]. lake-wide [173]. lakes [11, 895, 1308, 1352, 815]. Lambert [440]. lamellae [1056]. Lamna [696, 920]. Lamniformes [1002]. Lampetra [1003, 100, 56]. lampeye [584]. Lamprey [56, 891, 1003, 100, 429, 301, 81, 685, 829, 747, 253, 1027]. landings [919]. landlocked [22, 105, 18]. Landscape [248]. lane [1309]. Lanka [1146]. lanternfish [1018, 1491, 1391, 1048, 41]. lanternshark [160, 552]. lanternsharks [162]. Lapillus [1023]. Large [754, 909, 625, 872, 112, 1431, 1001, 1440, 216, 817, 82, 419, 889, 663, 1363]. large-bodied [112, 82, 1363]. largemouth [918, 173, 1076]. largest [916, 657, 461]. Larimichthys [1201]. larvae [101, 247, 867, 891, 310, 1073, 1265, 387, 1103, 1210, 983, 1438, 431, 1412, 594, 739, 353, 1067, 1283, 1205, 487, 346, 719, 1174, 649]. Larval [1103, 1043, 405, 745, 148, 1125, 586, 362, 1137, 87, 149, 204, 1048, 556, 1105, 585, 1343, 722, 815, 7, 311, 1498, 253, 550, 1102]. laser [28]. last [57]. late [791, 1514, 1136, 617]. late-phase [1136]. late-term [1514]. later [1451, 323]. later-life [323]. lateral [388, 390]. lateral-line [388]. lateralization [1034]. Lates [190]. lathami [724]. laticlavius [1500]. latipes [474, 476, 478]. latipinna [1064]. latipinnis [1326]. latitudinal [814, 697, 178]. lato [1020]. Laurentian [1495]. lavaretus [939]. Lawrence [377, 1137, 1495]. laying [1004, 2, 682]. lead [1067, 1429]. leadership [1152]. leaf [1100]. learning [396, 457]. lectotype [1396]. led [1476, 1015]. leeches [1334]. left [1166]. Leiarius [699]. Leiognathidae [743]. Lemon [597]. Length [904, 274, 146, 940, 628, 816, 1123, 1010, 938, 319, 253]. lens [12, 1338]. lenses [994]. lentic [1011]. leopard [1150]. leopardus [286]. Lepadichthys [756]. Lepeophtheirus [363]. Lepidorhombus [1033]. Lepisosteidae [585]. Lepisosteiformes [585]. Lepomis [486, 1106, 896]. leptin [1156]. Leptobarbus [561]. Leptobotia [1410]. leptocephali [820]. leptocephalus [791, 719]. lessae [47]. lesser [776]. lesser-spotted [776]. let [785]. lethal [832, 157, 1455, 170, 231]. lethally [770]. Lethrinops [1042]. Lethrinus [812]. leucas [463, 998, 644, 568, 1336, 1133]. Leuciscidae [787, 42, 1397, 734, 1181]. Leuciscus [273]. Leucism [1286]. leucocytes [578]. leucomaenis [1182]. Leucoraja [3]. leukocytes [767]. level [1284, 1298, 1272, 66]. levels [980, 1219, 209, 1269, 1473]. lewini [59, 931, 761, 280, 1063, 199]. lgp2 [147]. Li [1412]. Lib [677]. library [677]. lice [363]. licha [565]. lidwilli [446]. Life [271, 988, 995, 344, 897, 612, 453, 723, 244, 1418, 1258, 1309, 1227, 57, 727, 1509, 382, 74, 812, 597, 593, 656, 62, 929, 1281, 794, 33, 1136, 1451, 323, 697, 1043, 848, 442, 1338, 984, 92, 1429, 1472, 211, 110, 1054, 1406, 313, 1441, 69, 456, 314, 1308]. Life-history [988, 995, 244, 1227, 727, 1509, 597, 1281, 697, 1043, 92, 211, 1406, 456, 1308]. Life-stage-dependent [453]. lifespan [909]. lifespans [1026]. lifetime [1024]. light [517, 247, 438, 552, 553, 340, 289, 1492, 978]. lightly [786]. lights [1204]. like [720, 1359, 242, 791, 1180, 1266, 1486]. limbata [208]. limbatus [163, 422, 921, 636]. Limia [143]. limit [156]. Limitation [381]. limitations [218, 1058]. Limited [1338, 1497, 1232, 812, 184]. limits [1195, 885, 231, 1357]. Lindberg [1438]. line [455, 388, 77, 45, 1013, 390]. lineage [568]. lineages [1385, 1089]. linear [1371]. lineatus [760, 756]. lines [395]. Linkage [1502, 715, 378]. linked [448, 472, 1085]. Linking [871, 154, 994]. Links [1281]. Linnaeus [916, 1082, 1256, 1349, 779, 399, 1257, 206, 776, 1472, 1088, 1375, 1114]. Linneus [547]. Liobagrus [950]. lion [781]. Lionfish [169, 801, 781, 1497, 182]. lionfishes [1117]. lipid [646, 960]. lipids [430]. lipopolysaccharide [1285, 1266, 1099]. lipopolysaccharide-challenged [1285]. lipoprotein [1422]. List [374, 219]. Listrura [874]. literature [1073, 1144, 1067]. little [3]. littoral [806, 828]. livebearer [1332, 1302, 1224]. livebearing [143]. lived [180]. liver [1221, 258, 733, 1388, 454, 1380, 844, 705]. lives [712]. Living [298, 980, 1026, 616, 174, 701]. Liza [140, 439]. lizard [1084]. lizardfish [1393, 591]. loach [413, 13, 877, 12, 1389, 914, 1124, 10, 621, 723]. loaches [1405, 1084, 784]. loads [1236]. Lobotes [1483]. local [123, 193, 1212, 190, 1152, 154]. localities [287]. locality [255]. locally [58]. located [1502]. location [423]. Loch [397]. locomotion [1148]. locomotor [1098, 462]. locus [1036]. Lomond [397]. London [526]. Long [912, 861, 305, 1264, 208, 214, 216, 274, 509, 1169, 898, 114, 177, 180, 1384, 1119]. long-chain [177]. Long-distance [861]. long-lived [180]. long-range [274]. long-snouted [1384]. Long-term [912, 305, 1264, 208, 214, 216, 509, 1169, 114, 1119]. longevity [639, 786, 909]. longfin [336]. longhorned [214]. longiceps [848]. longitudinal [380]. longjaw [1383]. longnose [692, 812, 848]. Longsnout [992]. Looking [415, 229, 1040]. loop [1305]. loop-mediated [1305]. lordosis [473]. Lorenzini [390]. Loricaria [389]. loricariid [220]. Loricariidae [1225, 555, 358, 389, 1081, 1463, 1464, 1461, 513, 679, 1056, 629, 441, 485, 1319, 24]. loricariin [389]. Loricariinae [1225, 1081, 1463]. Loss [404, 478]. louder [351]. Lough [423, 423, 301]. Low [438, 1236, 1284, 891, 997, 561, 894, 674, 683, 1087, 790, 1015, 1149]. low-abundance [891]. low-density [1087]. low-dose [997]. low-level [1284]. low-oxygen [894]. low-value [674]. Lowe [1002]. Lower [1288, 1130, 988, 327, 257, 1229, 1131, 220, 580]. lowland [82]. LSG2 [1240]. LSG2-8 [1240]. Ltd [233]. Lualaba [122, 1203]. luapulae [1271]. Lubricogobius [1173]. lucidus [1267, 752, 1104, 1217]. lucioperca [586]. lucius [580, 1305, 1351]. Lufira [1068, 123, 122]. Luhoho [576]. lumaireul [1020]. luminescent [162]. lump [604]. lumpfish [1000]. lumpus [1000]. lunar [847]. lunare [297]. lupus [772, 749]. lutjanid [180]. Lutjanidae [577, 1210, 1005, 771, 1344]. Lutjanus [1309, 1210, 511, 1105, 503]. Lutz [1079]. Lycodes [1482]. lymphocytes [1076]. Lyopsetta [43]. Lythrypnus [1386, 1370].

M [571]. M. [176]. Macabre [1454]. Machado [459]. machine [396, 1121, 457]. mackerel [417, 1274, 83, 490, 1263, 794, 623, 821]. macracanthus [590]. Macrocephenchelys [1228]. macrochirus [486]. macrophage [1285]. macrostomus [685]. macrotidal [847]. macrourid [639]. Macrouridae [1449, 23]. Maculabatis [244]. maculatum [705]. maculatus [1006, 1385, 1512, 506, 105]. Madagascar [130]. Madeira [352, 354, 1231, 1081, 1324, 623, 601]. maderensis [806]. magnetic [813]. Mahanadi [1421]. mailed [1225]. Maine [269]. maintain [1230, 604]. maintained [187, 790]. maintenance [1065]. major [1312, 1464, 147, 1446, 1243]. Makaira [1433]. make [543]. mako [1303]. Malaŵi [185, 176]. malabaricus [789]. Malapteruridae [370]. Malapterurus [370]. Malawi [775, 1042]. Malay [39]. Malaysia [372, 1385, 604]. Malaysian [1123]. Maldivian [1336]. Male [1447, 850, 72, 956, 292, 1097, 560, 1151, 642, 974, 186, 1107, 1147, 1119, 109, 1143, 1462]. male-sign [1107]. males [392, 1115, 185, 700]. malic [1036]. malnutrition [1436]. Malpelo [716]. Malvan [1254]. Mamanguape [679]. mammalian [1157]. manage [408]. Management [93, 267, 50, 935, 812, 605, 396, 1026, 259, 1159, 879, 60, 226, 238, 1364, 180]. managing [312]. mandarin [1221]. mandelai [631]. manganese [860]. mangrove [1404, 221]. mangrove-dependent [221]. mangroves [603]. manipulation [567]. Manta [919, 89, 1515, 962, 1513, 1247, 1129, 1111]. many [574]. MAPK [993, 967]. MAPKs [65]. MAPKs-jnk1 [65]. mapping [1373]. Mar [1384]. marble [1154]. marbled [1189]. march [1262]. Marcusenius [121, 53]. marginatus [351]. Marine [264, 74, 697, 1340, 1498, 836, 1374, 256, 589, 515, 709, 969, 1470, 694, 750, 1045, 368, 281, 620, 940, 22, 200, 912, 1232, 1178, 533, 1480, 763, 1294, 1348, 270, 686, 155, 1136, 739, 1025, 1304, 51, 687, 1236, 626, 286, 184, 1057, 1168, 522, 1190, 1297, 938, 1381, 1060]. marine-protected [256, 1190]. Marini [785]. marinus [429, 81, 747]. mark [1431, 1489]. markers [1341, 729, 136, 896, 1020]. Market [927, 1159]. markets [1140]. marking [464, 1467]. markings [88]. marlin [1433, 322, 640]. Marmara [970, 957, 1046]. marmaraensis [1322]. marmorata [1234, 974, 1062]. marmoratus [699, 1154, 1174]. marsilii [1020]. Mask [423]. masked [1122]. masou [683, 1456, 333]. masquinongy [1131]. mass [464, 34, 1467, 253]. mass-marking [464]. masses [757]. massive [398]. masu [683, 1456, 333]. mate [72, 176, 691]. maternal [1044, 54, 115]. mating [1079, 1004, 698, 294, 1057]. matrinxã [26]. matrotrophy [1514]. matter [616, 221]. matters [872]. Maturation [312, 1458, 245, 906, 1175, 1381, 1222]. mature [3, 251]. maturity [668, 951, 102, 1495, 940, 1512, 381, 776, 660, 934, 797, 1010, 1196, 319, 397, 1408]. maturity/maximum [381]. Maurolicus [796]. mawsoni [861]. MaxEnt [1201]. maxima [1167]. maximum [348, 1508, 940, 336, 381, 1516, 172]. maximus [285, 739, 1014]. may [150]. Mayaheros [303]. Mayan [303]. Maze [1311]. M’diq [1329]. me [830]. meadows [111]. meagre [1015]. meal [885]. mean [529, 205]. measurement [274, 1303]. measurements [498, 253]. measures [1150]. Measuring [172]. mechanical [474, 771]. mechanism [665]. mechanisms [733, 767, 1373]. mechanosensory [390]. medaka [476, 474, 109, 478]. media [927, 606, 366]. mediated [659, 669, 1305, 471, 1076]. medirastre [929]. medirostris [1390]. Mediterranean [806, 801, 806, 557, 714, 1323, 519, 200, 1318, 983, 1213, 169, 1002, 1038, 606, 212, 1376, 781, 1497, 174, 1088, 182, 1211, 768, 383, 1085, 571]. medium [1505]. medium-sized [1505]. meet [126]. mega [8]. Megaleporinus [142]. Megalops [1244, 656]. Megatrygon [85]. megrim [1033]. meiotic [1151]. Mekong [243, 1197, 1351]. Melamphaes [1391]. Melamphaidae [1391]. melampygus [660]. melanin [1438]. Melanogrammus [892]. melanophores [1204]. melanopterus [372, 1378]. melanostomus [500, 1200, 330, 1295]. melanotheron [800]. melas [282]. melastomus [298]. melatonin [1055, 1040]. melting [1205]. membras [972]. Mendelian [27]. Menezes [1163]. Menor [1384]. mentella [1495]. menu [169]. Mercuric [837]. meristic [591]. Merluccius [785, 387]. meso [847]. meso-macrotidal [847]. Mesoamerica [583]. mesopelagic [904, 146, 1158, 29]. mesophotic [320]. mesopotamicus [459]. mesopredatory [849, 15]. mesotrophic [1507]. Meta [677]. Meta-Fish-Lib [677]. metabarcoding [677, 1398, 1232, 416, 216, 317, 418, 703, 1353, 1338, 409]. metabarcoding-based [409]. Metabolic [1245, 78, 1214, 439, 348, 310, 498, 501, 1130, 816, 928, 582, 1150, 1303, 296, 550]. metabolism [706, 540, 53, 1503, 870, 960]. metabolomes [1328]. metacercarial [1473]. metagenetics [47]. metal [637]. metals [1082]. metamorphosis [791, 253]. metapopulation [453]. methionine [1015]. Method [1508, 1022, 1513, 1371, 1169, 1058]. methodological [66]. methodology [1167, 1124]. methods [1318, 47, 1177, 409, 945, 550]. methylation [609]. methyltransferase [713]. Metriaclima [176].

Mexican [276, 392, 761, 367, 402, 1202]. mexicanus [402]. Mexico [569, 288, 1318, 280, 1063, 656, 422, 921, 178, 984, 1205]. micro [842, 1433]. micro-habitat [842]. micro-tooth [1433]. microanatomy [887]. microbiome [1455]. microbiota [1510]. microcephalus [1235]. microchemistry [359, 692, 203, 338, 364, 1373]. microchromosome [1114]. microlepis [1493]. Micromyzon [493]. Microphysogobio [581]. microplastic [1236]. microps [85, 1313]. Micropterus [918, 910, 173, 1076, 1510]. Microrasbora [1021]. microsatellite [1020]. microsatellites [1375, 547]. microscopy [1492]. microsphincters [1164]. Microstomus [597]. microstructure [1174]. Mid [284, 1353, 1489]. Mid-Atlantic [284, 1353, 1489]. Middle [279, 123, 914, 1421]. middling [929]. midwater [803]. Mif1 [1285]. Mif2 [1285]. mIgM [1076]. migrants [888, 573]. migrating [1440, 365, 1392]. Migration [1326, 273, 1392, 425, 361, 836, 589, 625, 359, 1323, 1465, 778, 1285, 989, 847, 63, 379, 1110, 953, 1025, 1304, 822, 683, 333, 187, 338, 999, 854]. migratorius [887]. Migratory [49, 1141, 1109, 1110, 1169, 51, 1353, 1338, 443, 311, 18, 663, 819]. Mikinkidi [1068]. Mikrogeophagus [72]. miles [801, 781, 1497, 182]. milk [563]. Miller [618]. Millerichthys [618]. Min [1411]. Minas [16]. mineralization [483, 181]. mineralogy [1153]. miniature [440, 1385, 327, 710, 1231]. miniaturisation [21]. minimises [746]. minjiriya [370]. minnow [787, 853, 124, 746, 1240]. minnows [742, 556]. minor [1482]. minute [965, 672]. minutus [1281]. miR [721]. miR-242 [721]. miR-434 [721]. mirage [574]. Miragoane [143]. Miranda [774]. miRNA [492]. mirror [980]. Misgurnus [413, 141, 877, 12, 10, 784]. misleading [1140]. missing [481, 171]. Mitochondrial [1356, 568, 1176, 100, 71, 704, 547]. mitogen [993, 967]. mitogen-activated [993, 967]. Mitogenomic [1506]. Mitotic [1082]. mixed [84, 883]. mixed-species [84]. mizolepis [1391]. mm [1037]. mMEP [1036]. mMEP-2* [1036]. mobility [102, 477]. Mobula [89, 214, 962, 1513, 1518, 825, 1129, 1111]. mobular [1518]. Mobulidae [1247]. Mochokidae [1288, 123, 863]. modakandai [773]. Mode [21]. model [274, 476, 477, 88, 1238, 478, 619, 839]. Modeling [1314, 1321]. modelling [885, 825]. models [1201]. modern [1227, 469]. modification [828, 705]. modified [580, 566, 1487]. modify [800, 735]. Modifying [179]. modulates [1219]. modulating [248]. Modulation [492, 1374]. modulatory [147]. Moenkhausia [1324, 672]. mojarra [392]. mokarran [163, 979]. Mola [1091, 29]. Molecular [523, 141, 1210, 946, 826, 594, 240, 634, 1464, 147, 370, 129, 65, 441, 544, 725, 728, 350, 42, 354, 4, 154, 1149, 855, 1020, 604]. Molidae [882]. molitrix [224, 738, 793, 71]. molleri [162]. mollusc [127]. Mollusca [1335]. molly [1064]. Molva [539]. Monitoring [216, 93, 208, 1328, 407, 418, 1305, 1457, 58]. monochromatic [1204]. monogamous [72]. monophyly [1393, 594]. Monopterus [1180]. Monotaxis [897]. monstrosa [249]. moon [297]. moorii [121]. moray [1273]. mordax [1365]. morhua [444, 329, 453]. Mori [581]. Mormyridae [121, 122, 595]. Moroccan [1329, 1128, 1121]. Morone [747]. morph [1480]. Morphochromatic [1500]. morphologic [1068]. Morphological [439, 810, 1256, 787, 387, 651, 365, 974, 699, 511, 920, 174, 10, 347, 26, 350, 862, 354, 493, 358, 1210, 1385, 86, 14, 1273, 1281, 594, 1476, 419, 595, 638, 129, 441, 16, 1182, 314]. morphologically [1103]. Morphology [658, 1460, 1494, 607, 440, 918, 795, 998, 1038, 67, 1433, 905, 388, 667, 707, 1496, 947, 127, 1223, 316, 250, 1475]. Morphometric [565, 1349, 1299, 1464, 1009]. morphometrics [1434, 1330, 225]. morphometry [667, 1344]. morphospace [362]. morphotypes [1045]. morphs [939]. mortality [1371, 353, 1408]. mosaic [1190]. mosquitofish [6, 1488]. mossambica [130]. mossambicus [1284, 1011, 1219, 79]. most [96, 8]. mothers [542]. motif [1361]. motility [215, 648]. Motion [1152]. motivation [1444]. motor [1494, 403]. motoro [708]. mottled [1234, 1062]. mountain [1405, 562, 1320, 853]. mouth [956, 990, 1135, 1197, 676]. mouth-brooding [956, 1197]. mouthless [1514]. Movement [1270, 686, 1432, 910, 322, 640, 1340, 59, 678, 942, 802, 977, 1478, 131, 965, 82, 1154, 598, 154, 889, 380, 1295, 701]. movements [195, 554, 1348, 861, 1287, 596, 1129, 55]. moving [1238]. Mozambique [962, 1089]. mRNA [730, 1055, 1219, 222]. mtDNA [729, 1365]. much [1481]. mucin [331]. mucosa [1164, 26]. mucosal [797]. mucous [1328, 1474, 10]. mucus [832, 737, 739, 315]. mud [1108]. mudflat [638]. mudskipper [1089]. mudsucker [1383]. Mugil [439, 288]. mugilid [1279]. Mugilidae [439, 140, 1239]. Mugiliformes [140]. Müller [1277, 1004, 1075]. Mullerian [1172]. mullet [288, 140, 1239]. Mulloidichthys [869]. multi [1230, 1381]. multi-population [1381]. multi-trait [1230]. multidisciplinary [597]. Multilocus [1242]. multimodel [1202]. Multiple [1160, 1078, 1116, 1431, 302, 886, 1209, 876, 395]. Multiple-brooding [1160]. multiple-predator [1209]. multiscale [470]. mummichog [80, 1426]. mummification [199]. Muraenidae [326, 1273]. murphyi [490]. muscle [1198, 832, 1164, 1263, 430, 1015, 346, 915, 1112, 1207]. muscles [462]. muscular [281]. Museomics [1128]. muskellunge [1131]. mussel [1334]. mussel-associated [1334]. mussels [872, 965]. Mustelus [371, 762, 66, 886]. Mustura [588]. muzzled [90]. Myanmar [1021, 927]. Mycteroperca [917]. Myctophidae [1491, 1391, 1048, 41]. myd88 [141]. Myers [122]. mykiss [1341, 560, 486, 755, 593, 721, 632, 1477, 1154, 1220, 1504, 296]. Myliobatiformes [843, 283, 198]. Myliobatoidei [1069]. Mylopharyngodon [45]. myocardium [246]. Myosin [246]. Myoxocephalinae [726]. mysid [835]. mystax [1501]. mysteries [1121]. Mystery [320, 1216, 1008]. Myxinidae [1161].

N [361, 450, 1233, 306, 115, 454, 841]. N-ethylmaleimide-sensitive [361]. nākea [1327]. nahlongthai [70]. naked [298]. nalua [1380]. namaycush [292, 155]. Naming [1041]. Nandujiang [1176]. Nannocharax [355, 1271].

Nanpanjiang [1414]. Narcine [211, 587]. nares [1056]. narinari [843, 384, 804]. narrow [401]. narrow-range [401]. Narrower [1047]. narrows [981]. nasal [183]. nase [1135]. nasus [696, 920, 844]. Natal [562]. natalensis [562]. natgeorum [327]. National [59, 293, 76, 1498]. Native [71, 90, 852, 1117, 731, 799, 1352, 49, 1200, 132, 82, 703, 156, 217, 424, 746, 1487, 1320, 1154, 1399, 1278, 161, 851]. Natural [251, 1519, 243, 580, 586, 1272, 262, 419, 88, 1153]. naturalized [1341]. naturally [799]. navigation [1392]. Neagh [301]. near [586, 1213, 673, 540, 903]. near-natural [586]. near-term [1213]. nearshore [256, 50, 1460]. necrosis [20]. need [225]. needle [782]. negative [1266]. negatively [647]. neglecta [845, 46]. negligible [306]. Negro [1225]. neighbouring [111]. Nemacheilidae [952, 588, 13, 1389, 914, 621, 1322]. Nematalosa [786]. Nematoda [1456]. nematoptera [1490]. Nemipterus [1085]. Nemo [1454, 304]. Neogobius [500, 1200, 330, 1295]. Neolissochilus [558]. neonate [650]. neonates [1489]. Neosebastes [382]. neotropical [1439, 1148, 1387, 655, 1265, 197, 680, 447, 443, 311, 319, 220, 250]. nephrons [1310]. nerka [924]. nest [429, 309, 1113]. Nesting [1482, 372]. nests [320]. Net [746, 1329]. Network [526, 1035]. neural [924, 455, 901]. neuronal [361]. neurons [109]. neuropeptides [1219]. neutral [453]. newberryi [415, 229]. newly [119, 335]. next [1166]. NF [913]. NF- [913]. Niche [916, 1052, 208, 583, 845, 74, 132, 1047, 11, 239]. niches [318, 1404]. niger [1425]. Nigerian [370]. night [1368, 553]. nights [1059]. nigricans [1433]. nigrita [1288]. nigrocauda [838]. nigrofasciatus [17]. Nile [800, 1082, 946, 913, 20, 514, 190, 691, 700, 609, 637, 1475]. Nilo [1288]. Nilo-Sudan [1288]. niloticus [1317, 1284, 800, 1192, 1082, 946, 20, 514, 455, 190, 700, 609, 637, 662, 1475]. nine [1113]. nine-spined [1113]. niphonius [794]. nitidus [228]. nitrogen [1270]. nitrosative [1269]. NJ [571]. No [1179, 948, 1005, 148, 256, 694, 1480, 901, 1338, 1269]. no-brainer [901]. no-take [256, 694, 1480]. NO/GSH [1269]. nobilis [71, 1342]. nocturnal [333]. NOD [720, 1266]. NOD-like [720, 1266]. noise [1194, 1209]. Nomeidae [336]. nominal [106, 1501]. Non [832, 157, 359, 90, 1328, 1352, 27, 1348, 365, 1200, 132, 1455, 170, 133, 156, 1487, 51, 1320, 944, 770, 1399, 1278, 161]. non-invasive [1328]. Non-lethal [832, 157, 1455, 170]. non-lethally [770]. non-marine [1348]. non-Mendelian [27]. non-migrating [365]. non-migratory [51]. non-native [90, 1352, 1200, 132, 156, 1487, 1320, 1399, 1278, 161]. non-predatory [133]. non-reproductive [359]. non-significance [944]. noninvasive [891]. Nonlethally [1241]. nonnative [217]. nonspecific [1510]. norm [1458]. normal [705]. Norman [257]. Noronha [695, 991]. Norris [1022]. North [1161, 64, 1023, 1484, 727, 200, 106, 248, 596, 921, 1376, 991, 1069, 679, 633, 29, 1061, 892, 969, 772, 904, 1506, 404, 935, 597, 1091, 979, 552, 917, 433, 1014, 724, 378, 432, 579, 1243]. north-east [1484, 1061, 892]. north-eastern [1023, 200, 106, 991, 1069, 679, 633]. north-west [596, 29]. north-western [727, 921, 1376]. northcentral [656]. Northeast [841, 70, 1168]. northeastern [440, 886, 1379, 757]. Northerly [1366]. Northern [301, 1382, 668, 580, 1213, 1512, 1135, 597, 449, 939, 1231, 230, 743, 239, 152, 1519, 1333, 1282, 453]. northernmost [1502]. northward [747]. northwest [1354, 357, 360, 1393, 1346, 1048, 1357]. Norway [626]. Norwegian [487]. Notacanthidae [1226]. Notacanthiformes [1226]. Notacanthus [1226]. Notes [639, 1030, 565, 1002]. Nothobranchius [1361, 325, 1206, 839]. Notorynchus [264, 458, 811, 272]. Notothenia [737]. Notothenioidea [862]. Notothenioidei [1103]. notothenoids [1460]. Notropis [787, 788]. nov [577, 890, 128, 881, 773]. Novel [477, 1334, 161, 665, 1368, 1136, 183]. Nrf2 [923]. nubilus [787]. nucleotide [1359, 1087]. nucleotide-binding [1359]. nucleotides [918]. nudibranchs [1335]. nudiceps [1103]. Nui [830]. number [369, 278, 628, 1311]. Numerical [643]. Nunavik [1304]. nurse [442, 1090]. nursehound [1520, 1088]. nurseries [1208]. Nursery [761, 579, 785, 387, 280, 1063, 692, 1521, 921, 1489, 1372]. nutrient [1127, 1477]. nutrition [1427]. nutritional [387, 706, 1175, 1149]. NW [806].

O [474]. obesus [714]. obituary [1452]. obscures [987]. Observation [820, 1521, 1004, 1031, 1492]. observational [175]. Observations [931, 458, 586, 1090, 942, 802, 360, 803, 1119, 263, 1061, 1448, 1450]. observer [278, 304]. obstacles [1110]. obtained [431, 1279]. occupancy [50, 56]. occupying [381]. Occurrence [1483, 1479, 249, 1288, 98, 919, 1234, 1480, 970, 1487, 158, 1235, 69, 1033, 1239, 782]. occurring [318, 639, 1123, 1184, 64, 1145, 1321]. Ocean [537, 192, 120, 882, 38, 1292, 936, 43, 360, 484, 1108, 992, 763, 357, 644, 605, 1356, 1226, 446, 568, 206, 1158, 1469, 67, 1336, 9, 1051, 1057, 322, 640, 1489, 41, 29, 1134, 1014, 1379, 724, 378, 1032, 579, 349]. oceanic [89, 650, 58, 1111]. ocellate [708]. ocellatum [445]. ocellatus [845, 1283, 161]. ocellicauda [1343]. ochriamkini [867]. oconnorae [986]. octogrammus [1122]. ocurrence [1114]. Odisha [1421, 1424]. Odontaspis [1366]. Odontesthes [22, 1072]. odontodes [468]. Odontostilbe [354]. odours [924]. odyssey [1125]. oesophagus [346]. off [668, 628, 1245, 991, 616, 682, 1069, 796, 433, 1321]. offs [393]. offshore [533, 814, 303, 1035]. offspring [545]. ogunensis [987]. oHypanus [154]. Okamejei [1004]. Olaf [377]. old [712]. oleracea [922]. Olfactory [924, 361, 1056]. oligomerization [1359]. oligotrophic [872]. olivaceus [812, 1445, 578]. Oman [591]. omics [1121]. omnivorous [1439]. Omobranchus [90]. Ompok [247]. omul [887]. on-site [1305]. Oncorhynchus [361, 1141, 1341, 560, 486, 1127, 1001, 1440, 755, 593, 1314, 924, 323, 721, 78, 156, 521, 632, 158, 822, 683, 1456, 333, 1477, 769, 1154, 1153, 1220, 1139, 1399, 1504, 364, 819, 296]. one [1470, 1265, 814, 446, 302, 378]. ones [1047, 319]. only [256]. onset [1458, 153, 642]. Ontario [396, 1295]. ontogenesis [706]. Ontogenetic [998, 957, 724, 439, 223, 393, 594, 1449, 1404, 1241]. ontogenic [221]. Ontogeny [983, 982, 649, 1393, 670, 7, 1083, 1493, 1249]. Oocyte [1318, 868, 1070, 373, 892]. oocytes [575]. Oogonial [1070]. ’O’opu [1327]. opal [249]. opalescens [249]. open [165]. open-water [165]. Operant [1255]. operational [619]. operations [1329]. Ophichthidae [1435, 512]. Ophichthus [1435]. Ophidiiformes [664]. opinion [836]. opioid [79]. Opisthonema [929]. Opistognathus [1275]. Oplegnathus [1437]. Opportunistic [1515, 1174]. Opportunities [1455]. opportunity [1049]. Opsanus [241, 899]. opsin [1102]. opsins [937]. optimal [1510]. Optimizing [1467]. optimum [885, 540]. Oral [183, 920, 127]. oral-mollusc [127]. Orange [28, 507, 431, 251]. Orange-finned [28]. orange-spotted [507, 251]. orangethroat [1443]. oraniensis [631]. orbignyanus [1374, 1108]. orbitalis [196]. order [196]. Oreochromis [1284, 800, 1192, 1082, 946, 20, 514, 455, 424, 1011, 1219, 700, 609, 637, 79, 1100, 662, 1475]. Oreonectes [13]. organ [361, 224, 1055, 81, 585, 1343, 10]. organ-specific [81]. organic [221]. organism [839]. organisms [1238]. organs [1496]. oriental [141]. orientation [1448]. origin [1021, 68, 1153, 873, 579, 1519]. original [1449]. origins [242]. Orinoco [1315, 485]. Orkney [682]. ornament [72]. Ornamental [775, 1098, 614, 783]. Orthonopias [757]. Oryzias [476, 478]. Osmeridae [959]. Osmerus [149, 1365]. osmoregulatory [922]. osmotic [1422, 538]. Osphronemidae [1030]. ossicles [833]. ossification [905, 982]. Osteichthyes [1434, 1350]. osteoclasts [476]. osteocytic [474]. osteogenesis [523]. Osteoglossiformes [328, 615, 122]. Osteology [764, 440, 140]. Osteology-focused [764]. osteopenia [477]. osteoporosis [476, 478]. Otocinclus [1464]. Otolith [203, 1304, 1516, 1153, 1023, 288, 359, 557, 1349, 399, 227, 47, 860, 1046, 428, 1186, 895, 36, 1446, 338, 893, 1344, 364, 1330, 330, 1174, 1085, 1373]. Otolith-inferred [1304]. Otoliths [1026, 515, 519, 1267, 1425, 972, 1054, 994, 1104]. our [1195, 1514, 808]. outer [330]. outlook [1292]. outmigrants [1199]. outmigration [580, 525, 840]. Ovalentaria [125]. ovarian [360, 242, 321, 63, 906, 877, 1011, 75, 1246, 1222]. ovaries [617]. ovary [524, 209, 79]. ovata [711]. ovate [711]. ovatus [1472]. overall [1431, 1015]. overfishing [1159]. overlap [15, 217]. Overlapping [318]. overlooked [338]. overview [1239]. Overwintering [625, 889, 1201, 575]. ovifrons [554]. oviparous [630]. ovoviviparous [115]. ovulatory [524]. Oxford [233]. oxidative [524, 553, 837, 915]. Oxudercidae [634, 1421]. Oxudercinae [1089]. oxy [386]. oxy-thermic [386]. Oxycheilinus [670]. Oxygen [462, 381, 1317, 536, 1508, 112, 498, 252, 1130, 759, 1499, 894, 1058, 915, 172]. Oxynoemacheilus [952, 1322]. Oxynotus [779, 970]. oxyrhinchus [315]. oxyrinchus [315, 1202, 1303]. oxytetracycline [1492]. oyster [1012]. Ozark [787, 910].

P [93]. P. [286]. Pablo [1031]. pace [1429]. Pachitea [1461]. Pachypops [745]. Pacific [484, 579, 880, 931, 577, 1023, 757, 1187, 1270, 650, 1393, 1331, 761, 367, 321, 373, 599, 1480, 1512, 799, 365, 974, 1126, 1346, 446, 1048, 206, 670, 848, 971, 511, 1105, 503, 886, 1202, 433, 29, 1276, 1379, 947, 693, 627, 253, 1243, 1027]. pacifica [365, 1187]. pacificus [1331]. pacu [1083]. paddlefish [263]. Pagrinae [69]. Pagrus [916, 428]. pain [1195]. pair [100, 3, 1271, 185, 1379, 116]. pairing [1511, 1155]. pairs [800]. pajama [1490]. palate [1338]. paleatus [667]. pallaryi [1128, 1121]. Pallas [1122, 330]. pallid [52, 353]. Palmyra [414]. palometa [1006]. PAMP [147]. PAMPs [65]. Pampus [1126, 1222]. pancreas [1380]. Pandaka [446]. pandus [382]. panel [1087]. Pangasianodon [922]. pantaneiro [1163]. Pantodon [328]. Pantodontidae [328]. papers [1395]. papilla [97]. Parablennius [1119]. Paracanthopoma [736, 764]. Paracirrhites [88]. paradoxus [336]. Paraguay [1081]. parahybae [906]. Paraíba [843]. Parakneria [1068]. Paralabrax [894]. Paralepididae [404]. Paralichthys [1374, 1108, 1445, 578]. parallelus [605]. parameters [1470, 431, 997, 302, 272, 1474, 608, 1477, 178]. Paraná [555, 651, 40, 24, 774]. Parapercis [1347]. parasema [1511]. parasite [343, 1312, 948, 1387, 1047, 1145, 1456, 1106, 363]. Parasites [1088, 1334, 207]. parasitic [1235, 363]. parasitism [165, 844]. parasitized [884]. Parauchenoglanis [1203]. Paravandellia [764]. pardalis [60]. Pardo [1056]. Parecis [1324]. parental [27, 1059]. Pareuchiloglanis [5, 255, 1071]. Park [76, 59, 293, 1498]. Parks [697, 286]. Parotocinclus [679, 1056]. parr [102]. parrotfish [554, 848, 1363]. part [1298, 124]. Parthenogenesis [1115]. partial [763, 1257, 338, 335, 300]. partially [1357]. partition [1459]. partitioning [15, 1069, 239]. partner [698]. parts [1156, 799, 1376]. parturition [1079, 226]. parva [318, 681, 764, 1177]. parvus [987]. Pasco [1461]. pass [1444]. passage [731, 1250, 744]. passive [741, 1037, 818]. past [1239]. patachonicus [398]. Patagonia [1382, 702, 239]. Patagonian [989, 398, 248, 724]. patch [838]. paternal [1197]. paternity [1116, 302, 886, 876]. pathogenic [147, 65]. pathogens [728]. pathway [913, 923, 147, 1076]. pathways [1341, 1078, 215]. Patinga [459]. Patos [1006]. Pattern [558, 359, 907, 1055, 847, 27, 1464, 378, 587, 1172, 854]. Patterns [638, 59, 678, 1465, 1097, 1349, 741, 761, 444, 1325, 686, 1405, 132, 273, 248, 1392, 1264, 1304, 210, 453, 815, 1406, 401, 495, 1381, 69, 1217]. Paul [1251, 284, 1251]. Pausing [543]. pCO [981]. PCR [1305, 1077]. peacock [1130, 447]. pearl [751, 5, 71, 1406, 1176]. pearlscale [1172]. pearlside [796]. Peciformes [624]. pectinata [339]. Pediococcus [1468]. Pehuenche [460]. pejerrey [22, 1072]. Peladillas [156]. Pelagic [803, 1198, 785, 175, 448, 1425, 825, 695, 1186, 232, 1335, 1442, 1498, 349]. Pellona [1242]. pellucida [758]. Pelteobagrus [967, 958]. Pelvicachromis [666]. pencil [564, 327, 710, 1496]. pencil-catfish [710]. peninsula [433, 1472]. Peninsular [1385, 604]. people [1094]. peptide [79]. peptide2 [228]. Perca [348, 536, 1345, 252, 741, 73, 553]. perch [348, 536, 1345, 252, 741, 73, 553, 190, 915]. perchlet [624]. Perciformes [194, 1030, 482, 577, 867, 1484, 564, 1210, 39, 1351, 823, 1053, 745, 743, 1344, 69, 1475]. Percilia [380]. Percina [1101]. perennial [1011]. perezi [875]. perfasciata [954]. performance [1284, 1118, 902, 918, 80, 1426, 983, 1499, 1126, 1311, 1306, 1351, 540, 1477, 846, 75, 1015, 1413, 347, 818, 1504, 1342, 1473]. period [563, 698, 592, 768]. periodicity [1202]. periods [706, 1057]. permanent [1432]. permit [50]. Pernambuco [843]. perryi [616]. Persian [268, 36]. persistent [574]. persists [400]. Personality [500]. Personality-dependent [500]. perspective [468, 1095, 1017, 783]. perspectives [1291, 226, 316, 985]. Perth [786]. Peru [880, 151, 504, 1461, 495, 601, 511, 1105, 503]. Peruvian [449, 894, 722, 391]. Peter [1251, 284]. Peters [130, 1011]. Petromyzon [429, 81, 685, 747]. petronius [332]. pH [561, 1049]. phagocytosis [1076]. phalericus [673]. Phalloceros [1224]. pharyngalis [332]. pharynx [1285, 346]. phase [1156, 974, 1136]. phase-dependent [1156]. phases [847, 1173]. Phenolic [1100]. phenological [1502]. phenology [788, 817, 1162]. Phenotypic [105, 471, 561, 1230]. Philippine [694]. Philippines [297, 1273, 1111]. philopatry [154]. phoenix [1112]. phosphorus [483, 181]. Phosphorylated [705]. photo [85, 272]. photo-identification [85]. Photoconvertible [475]. photogenis [788]. photographed [29]. photographs [1038]. photoperiod [247, 1055, 816]. photoperiodism [1012]. photoperiods [1222]. Photoreceptors [959]. Phoxinus [42, 1181, 1020]. Phractocephalus [1148]. Phylliroidae [1335]. Phyllopteryx [842, 765]. Phylogenetic [583, 855, 1393, 1128, 704, 370, 1343, 726, 863, 1029, 728]. phylogenetically [1328]. phylogenetics [574]. phylogeny [826, 1356]. phylogeographic [1405, 150]. Phylogeography [100, 729, 603]. Physical [507, 613, 612, 926]. Physiological [1329, 1280, 559, 153, 365, 63, 1245, 534, 231, 1386, 174, 1058]. Physiology [499, 532, 529, 107, 1437, 205, 846]. Piaractus [459, 1083]. piceus [45]. Pickett [526]. picocyanobacteria [820]. piece [836, 24]. Pietsch [465]. Piferrer [233]. pigment [784, 937]. pigmentation [494, 1464, 1074]. pigments [959]. pigra [383]. pike [580, 1278]. pikeblenny [627]. pikeperch [586]. pilchardus [1186]. Pimelodidae [649, 906, 1170]. pineal [1055]. pink [361, 1001, 1440, 799, 158, 769, 1139, 1399, 854, 1519]. pinnimaculatus [128]. pipefish [1134, 1479]. pipeline [677]. piranha [240, 506]. piranhas [351]. Pisces [558, 959, 421, 925, 1048, 914, 291, 1122, 1009]. Piscirickettsia [997]. piscivore [273, 632, 164]. piscivores [759]. piscivory [153]. PIT [1124, 598, 114]. PIT-tagging [1124]. PIT-tags [598]. pituitary [79]. pituitary-ovary [79]. place [120]. placement [1029]. placental [450]. plagiostomus [1246]. plague [746]. Plains [195]. planeri [100]. planktivorous [924]. Plant [1015, 1109, 1504]. Plant-based [1015]. plasma [1298, 1422, 442]. plastic [137, 402]. Plasticity [472, 481, 230, 561, 1173, 171]. plateau [276, 340, 1013]. plates [481, 171]. Platycephalidae [951]. Platycephalus [951]. platypterus [866]. play [497]. player [916]. plea [1371]. Plecoglossus [1502]. Plectropomus [286]. Pleuronectidae [43]. Pleuronectiformes [196, 665, 1033]. plight [718]. plunderfish [538]. plunketi [590]. Poecilia [1064, 403, 54]. Poeciliidae [1332, 143, 1224]. poecilura [1123]. Poey [954, 520, 605, 917]. pointer [28]. Polar [862, 548, 1207]. politus [1470]. Poll [440, 1068]. pollutant [516]. pollution [1082, 687, 846, 637]. poly [78]. polyactis [1201]. polyandry [162]. polykrikoides [556]. polylepis [927, 1391, 1061]. polymorphism [1082, 1087, 637, 1036]. Polynemidae [826]. Polynesia [812]. Polyodon [263]. polyphekadion [1097]. polyploid [908]. Polypodium [864]. polyunsaturated [177]. Pomacentridae [208]. Pomacentrus [607]. Pomadasys [971]. Pomatoschistus [1281]. pomfret [1222]. pompano [1468, 1472]. Pomphorhynchus [1047]. pond [245]. pond-based [245]. ponds [1177, 1184]. ponyfishes [1471, 743]. pool [82]. poor [645, 539, 244]. poorer [1118]. poorly [555, 811]. pop [384, 1169]. pop-up [384, 1169]. Poptella [504]. Population [208, 1470, 1108, 285, 44, 1423, 1320, 1106, 1384, 219, 941, 657, 288, 1256, 1341, 243, 727, 1465, 299, 1270, 267, 730, 271, 22, 200, 1506, 197, 729, 83, 840, 272, 206, 955, 1186, 616, 351, 1170, 154, 1211, 723, 1381, 1020, 758, 579]. population-specific [840]. populations [569, 502, 14, 1352, 407, 265, 996, 1362, 71, 539, 1351, 419, 381, 190, 626, 707, 488, 637, 663, 347, 1459, 937, 547]. porbeagle [696, 920]. Porcupine [261]. porgy [916, 428, 219, 711]. Poroderma [1362]. porosus [520]. Port [678]. portable [112, 598]. Portugal [1256, 623, 1397]. Portuguese [67, 776]. portusjacksoni [678]. position [1128, 623, 855]. Possible [1138, 567, 449, 54, 1239]. post [589, 625, 524, 983, 1025, 1392, 1050, 378]. post-larvae [983]. post-ovulatory [524]. post-settlement [1050]. post-smolt [378]. post-smolts [589, 1025, 1392]. post-spawned [625]. postglacial [1419]. posts [366]. potamodromous [273, 817]. Potamoglanis [710]. Potamotrygon [708]. potassium [922]. Potential [684, 740, 515, 694, 650, 1312, 1302, 428, 721, 1360, 1429, 1462, 1139, 936, 1243]. power [1094, 1490]. Poyang [784]. pp [465, 526, 233, 93, 571]. practical [839]. practices [170]. prahli [367, 433]. Pranburi [1471]. pre [439, 827]. pre-colonial [827]. pre-juveniles [439]. preadaptation [1496]. preclara [281]. Predation [73, 84, 460, 872, 372, 766, 1511, 301, 1482, 396, 1365, 798, 1308]. predator [1052, 607, 1488, 368, 1325, 680, 270, 811, 1245, 1404, 184, 1209, 221, 1035, 1131]. predator-associated [1488]. Predators [1050, 1377, 1200, 396, 410]. predatory [1248, 133, 695, 66]. predict [956, 494]. Predictability [1488]. predicted [963]. Predicting [938, 253, 290, 842]. predictions [289]. predictive [274]. predictor [1517]. predictors [1387, 347]. predominance [1420]. prefer [698]. Preference [187, 392, 1003, 1150, 183, 275]. preferences [814, 570, 514, 1406]. preferentially [850]. preferred [536, 252, 1206]. pregnant [450, 367]. preleptocephali [1059]. Preliminary [802, 1059, 1273, 1520, 942]. Premature [54]. prenatal [188]. preparatory [1306]. Presence [367, 1022, 415, 714, 229, 1115, 97, 955, 109, 304, 296]. present [301]. presented [733]. preservation [1233, 8]. preserved [34]. preserves [86]. Press [465, 571]. pressure [772, 275]. pressures [836]. pretreatment [454]. prevents [476]. previously [1022, 1501]. Previtellogenic [209]. prey [842, 765, 567, 270, 329, 924, 289, 1088, 1278, 410, 926]. Price [233, 93, 571, 526]. Primary [884, 1065]. Princeton [571]. Prionace [1116, 157]. Prionodraco [1103]. Prionurus [1500]. prior [306]. prioritization [1411]. priority [1063, 1297]. Pristigasteridae [1242]. Pristiophoridae [390]. Pristis [339]. prized [1276]. pro [767]. pro-inflammatory [767]. Probabilistic [1458]. probe [1205]. problematic [551]. Procatopodidae [440]. procedures [454]. Proceedings [93]. process [905]. processes [242, 926]. processing [763]. Prochilodus [760]. produce [351, 337, 110]. produced [18]. Production [580]. productivity [727]. Professor [377]. profile [439, 884, 1055, 737, 1477, 649]. profiles [1171, 877, 582, 251, 110, 19, 1102]. profiling [967, 752, 1080, 713]. Profundulidae [583]. progeny [27]. Programmed [1136]. progress [543]. projected [1429]. Prolabeops [1029]. prolactin [228]. prolactin-releasing [228]. proliferating [12]. proliferation [1070]. prolonged [791]. Promoter [609]. promoting [346]. Proof [494]. properties [1422, 474, 771]. proportional [18]. proposal [106]. prospects [655]. prostaglandin [292]. protandrous [312]. protease [315]. proteases [649]. protected [256, 709, 969, 1232, 686, 131, 558, 1057, 522, 1190]. protective [1312]. protein [361, 993, 967, 1080, 1220, 637, 1149, 705]. protein-sparing [1220]. proteinases [763]. proteins [475]. proteomic [737]. protocol [462, 816]. protocols [348, 1446]. protogynous [727, 694, 507, 251, 1347]. protrusion [183]. provide [927, 1464, 1061, 1134, 1150]. provides [795, 416, 317, 83, 1353, 636]. Province [914, 125]. provinces [1288, 126]. provisioned [1185]. proxies [1296]. proximate [985]. proxy [487]. przewalskii [1013]. Psalidodon [870]. PSATs [1169]. Pseudobarbus [134]. Pseudobatos [740, 1086, 367, 433]. Pseudochondrostoma [1135]. Pseudocorynopoma [395]. pseudocryptic [345]. Pseudogilbia [664]. pseudoharengus [659, 675]. Pseudohomaloptera [1084]. Pseudoplatystoma [294, 1170, 649]. Pseudorasbora [318, 1177]. Pseudotolithus [810]. Psilorhynchidae [70]. Psilorhynchus [70]. Psychrolutidae [1438]. Pterois [801, 781, 1497, 182]. pterotum [1048]. Pterygoplichthys [60]. Ptychidio [1407]. puberty [642]. public [1287]. Published [526, 233, 93, 571]. Publishing [526, 93, 944]. puffer [1010]. Pufferfish [309, 320, 1024]. pugnose [1479]. pulchella [1347]. pulchellus [1386, 1370]. pulse [116]. pulse-generating [116]. pumped [416, 317]. pumpkinseed [1106]. punctata [1204, 1156, 1359]. punctatum [630]. punctatus [90, 665, 993, 837]. Pungitius [1113]. pupfishes [183]. putative [1324, 1089]. puyen [105]. Puyo [954]. pygmy [214]. pyjama [1362]. pyriproxyfen [1475].

Qinghai [1013]. quagga [872]. quality [981, 524, 1294, 1504]. quality-related [524]. Quantification [464, 413, 147, 598]. quantify [550]. Quantifying [1019, 1253, 338, 225, 1444]. Quantitative [1457, 1077]. quantity [1294].

r [498]. rabbitfish [249]. Rachel [465]. Rachycentron [1375]. rack [75]. radiata [1506]. radiation [1049, 183, 383]. radio [91]. radiocarbon [1313]. Radiological [324]. Radulinopsis [1438]. Rafinesque [283]. Rainbow [1365, 1341, 560, 486, 755, 856, 593, 885, 721, 689, 632, 1477, 1154, 1220, 1157, 1504, 296, 460, 172]. rainbow/steelhead [593]. rainfall [311]. Raja [813, 957, 1423]. Rajidae [451]. Rajiformes [451, 964]. rakers [404]. Ramakrishna [884]. ramirezi [72]. ramp [731]. randalli [1085]. Range [710, 971, 493, 1117, 1366, 358, 799, 979, 274, 814, 1138, 1129, 67, 366, 212, 210, 830, 1089, 1087, 198, 401, 747, 432]. range-expanding [210]. range-restricted [358]. ranging [804]. Rangiroa [812]. Rankl [476, 478]. Rankl-induced [476, 478]. Ranzani [520, 1091]. Rapa [830]. Rapid [86, 930, 628, 418, 966]. rare [415, 229, 1480, 779, 1231, 85, 1425, 565, 991, 1190, 1033]. rarest [1470]. rasbora [1021]. rasborine [1385]. Rasptail [451]. rassf4 [609]. rate [348, 310, 1508, 498, 990, 1130, 759, 462, 816, 794, 840, 149, 928, 846, 626, 1498, 1379, 1303, 550, 1428]. rates [536, 252, 659, 385, 996, 697, 1371, 938, 495, 290]. rather [1437]. ratio [940, 1127]. ration [659, 1314]. ration-dependent [659]. ratios [531, 236, 306, 770]. ray [813, 843, 384, 214, 324, 1513, 919, 1213, 1512, 47, 825, 1423, 1212, 1111, 804, 211, 46, 432]. Rays [571, 89, 293, 1515, 962, 845, 1247, 1518, 1123, 1129, 542, 1260, 522, 587]. razor [1500]. rDNAs [27]. Re [360, 914, 1202, 862, 1212, 1084]. re-appearance [1212]. re-appraisal [862]. Re-description [914, 1084]. Re-estimation [1202]. Re-evaluation [360]. reach [934]. reaches [1494]. reaction [1458]. reactive [915]. real [1077]. real-time [1077]. Realistic [1003]. realized [1126]. really [901]. reared [515, 912, 983, 399, 525, 820, 1306, 1024]. rearing [101, 472, 883, 1174]. rearranged [1361]. rearrangements [1107]. reassignment [122]. rebuild [539]. recapture [1431, 861, 762, 272]. recently [872]. receptor [361, 1156, 1359, 20, 1180, 1080, 147, 1157, 1076]. receptor-interacting [1080]. receptor-mediated [1076]. receptors [720, 1266, 209]. recessus [196]. reciprocal [751]. recognition [1447, 395, 176]. recognized [1119]. Recognizing [680]. recommendations [1155]. reconciles [795]. Reconstructing [2]. reconstructions [140]. record [931, 757, 1484, 1006, 1079, 90, 650, 249, 564, 1296, 160, 336, 1213, 1257, 1000, 340, 303, 1415, 1089, 294, 420, 1336, 1335, 1092, 433, 198, 1397, 1448, 1032, 1189, 1259]. recorded [99, 1235, 335]. records [208, 293, 264, 493, 1331, 1366, 779, 1518, 1356, 1254, 1002, 67, 991, 830, 284, 1057, 1260, 1190]. recovering [368]. recovery [385, 1245, 911]. recreational [262, 1183]. Recruiting [898]. recruitment [439, 1125, 732, 1365, 647, 1015, 1399]. recruits [913]. recurrent [1400]. Red [681, 1266, 394, 1284, 916, 1198, 464, 1470, 1041, 524, 1021, 1148, 230, 428, 563, 683, 511, 1105, 351, 503, 1083, 1467, 1283, 1344, 1100, 219, 1159, 1092, 986, 456]. red-bellied [1083]. red-eyed [351]. red-spotted [683]. red-tail [1148]. Redefinition [1071]. Redescription [954, 764, 581, 774]. redescriptions [1501, 383]. redfin [134]. redfish [1171, 1495]. rediscovery [1449]. redoxcline [673]. redside [1162]. reduce [1159, 846]. Reduced [1024]. reduces [800, 968, 478]. reductase [1269]. Reduction [1495, 1015, 10]. Redundancy [1074]. reef [256, 1248, 89, 607, 709, 372, 50, 694, 1450, 962, 278, 484, 1378, 628, 995, 554, 849, 875, 697, 1129, 1145, 445, 1111, 1057, 869, 238, 335, 286]. reefs [806, 664, 111, 670, 1208, 1363]. refeeding [222, 915]. Referees [374]. reference [517, 677, 505, 723]. references [1077]. reflection [543]. refuge [163, 184]. refuges [1367]. Regan [1103, 590]. regard [828]. regarded [1501]. regime [190]. regimes [1317]. Region [120, 350, 484, 968, 505, 822, 291, 453, 676, 1211, 1025]. Regional [697, 886, 841, 1198, 778, 1264]. regionally [56]. regions [1434, 186, 1469]. regius [1015]. regression [1371]. Regulate [445, 926]. regulated [625, 1285, 82]. regulating [1219, 1269]. Regulation [400, 379, 204, 1149]. reidi [992]. reinhardt [391]. Reinhardtius [1280, 1355]. related [1458, 1329, 524, 1438, 877, 1012, 923, 298, 176, 190, 491, 604, 1217]. relation [1055, 63, 37, 600, 687, 667, 470, 275]. Relationship [1445, 450, 382, 459, 860, 928, 1123, 704, 909, 550, 1373]. relationships [343, 557, 583, 1393, 904, 759, 1127, 1438, 270, 146, 381, 370, 1154, 726, 1168, 363, 69]. relative [903]. release [949, 1368, 1517, 110]. released [1368, 1287]. releases [846]. releasing [228]. relevance [1167]. Reliable [413]. reliance [1064]. remaining [657]. remake [1454]. remarkable [907, 124]. remarks [954, 106, 1438]. remodeling [1422]. remote [1251, 1248, 1382, 650, 628, 9, 495]. removal [1345]. Repeat [145]. repeatability [1488, 1508]. repeats [331]. replacement [495, 839]. Repomucenus [594]. Report [199]. reproducing [360, 22]. reproduction [6, 732, 1064, 1407, 825, 1272, 1258]. Reproductive [437, 843, 1156, 622, 599, 865, 357, 63, 605, 103, 1391, 446, 1123, 1316, 801, 575, 1010, 451, 397, 1162, 1131, 237, 740, 530, 757, 359, 727, 360, 694, 788, 1070, 1095, 382, 268, 1247, 157, 995, 447, 906, 1346, 615, 302, 1197, 1351, 262, 461, 427, 1275, 674, 178, 1163, 707, 443, 777, 869, 337, 1196, 723, 587, 489, 349, 1243, 191, 1339]. Republic [125]. required [1338]. requirement [483, 181]. rerio [474, 148, 850, 189, 523, 751, 1299, 213, 472, 1252, 473, 1152, 346]. Rescue [1513]. Research [1402, 1307, 613, 204, 1455, 612, 944, 467]. Reserve [156, 1480, 264, 1340]. reserves [1374, 694, 659]. Reservoir [980, 625, 447, 311, 347]. reshapings [723]. residence [626]. Residency [59, 678, 1352, 425]. resident [445]. residual [328]. resilience [702]. resistance [918]. resistant [560]. resistin [222]. resolution [274, 1409, 893, 1205]. resolving [68, 1490]. resorption [476]. Resource [64, 1232, 1407, 15, 1069, 1278]. resource-limited [1232]. Resource-use [64]. resources [966]. respiration [667]. respiratory [1350, 636]. Respirometer [112]. respirometry [172]. Response [1317, 28, 494, 189, 141, 720, 918, 884, 993, 402, 922, 793, 1451, 721, 1266, 744, 958, 65, 559, 1249, 821, 538, 844, 1269]. Responses [966, 838, 813, 949, 292, 1293, 1329, 218, 1221, 742, 107, 48, 81, 924, 1245, 78, 134, 1432, 1388]. responsive [730]. responsiveness [725]. resting [1450]. restocking [184]. restoration [238]. restricted [358, 477]. restriction [896, 1504]. restriction-site [896]. restructuring [251]. resulting [758]. results [348, 944, 319, 172]. Resurrection [1501, 829]. retained [1227]. retention [1141, 1003, 1127, 1037, 305, 1169, 945, 818, 819]. reticulata [403, 54]. reticulated [1454]. reticulatum [649]. reticulatus [1454, 1462]. retina [1055, 12]. Retinal [1490, 1102, 1122]. Retracing [359]. retrieve [319]. Retrieving [1190]. Retropinna [347]. retrospective [620]. return [996, 626, 938]. Reunion [644]. revalidation [631, 518]. reveal [567, 493, 643, 1136, 366, 1119, 58, 363]. revealed [288, 384, 1398, 1315, 1440, 814, 136, 71, 1043, 895, 1008]. revealing [1216, 254]. reveals [1022, 555, 355, 1458, 1117, 1128, 99, 309, 1356, 1140, 955, 1516, 129, 345, 221, 18, 1035, 279, 296, 1242, 24, 1217]. reversal [1033, 1072]. Review [880, 756, 1307, 465, 1002, 526, 233, 166, 93, 1084, 571, 125, 225, 1403, 436, 326, 562, 34, 859, 882, 391]. Revillagigedo [59, 293]. revised [758]. Revision [1271, 249, 328, 1463, 240, 581, 24]. rewarded [1370]. Rhône [174]. Rhabdoblennius [228]. Rhabdosargus [509, 1372]. Rhadinoloricaria [1225]. Rhamdia [569]. rheophilic [452, 1319, 314]. Rhincodon [297, 688, 461, 287, 1259]. Rhinobatidae [1086, 1075]. Rhinoptera [845, 46, 432]. Rhizoprionodon [520]. Rhodeus [161]. rhodolith [916]. rhombeus [704]. rhombic [41]. Rhynchobatus [366]. Rhynchoconger [1188, 881]. Rhynchocypris [1240]. rhythm [1098]. rhythms [1148]. Rhytiodus [1493]. Ribeira [99]. Ribeiro [774]. ribereños [495]. Rica [1133]. rich [1239]. Richardson [1343]. richardsoni [951]. Ricker [1371]. Ridge [284, 991]. rig [463]. rimiculus [508]. Rineloricaria [1107, 1319]. ringens [722]. Rio [40, 1324, 555, 142, 354, 907, 1225, 513, 679, 672, 345, 257, 1229, 16, 279, 601, 314, 352, 24, 220, 40, 774, 97, 391, 1056]. rise [1320, 1276]. risk [668, 852, 766, 1511, 73, 1159, 109, 798, 244, 1308]. risk-taking [109]. risks [119]. Risso [673, 1033, 404, 1391]. River [42, 5, 651, 104, 789, 1407, 327, 736, 710, 1231, 123, 1471, 301, 508, 576, 634, 1184, 7, 1421, 1406, 1203, 152, 1322, 1319, 1027, 1131, 1052, 549, 589, 625, 1382, 891, 1265, 990, 1003, 100, 416, 1439, 1001, 1440, 989, 317, 274, 1505, 505, 132, 273, 996, 76, 1110, 1062, 1464, 217, 344, 1487, 1106, 513, 8, 174, 18, 708, 889, 347, 874, 1519, 1502, 250, 243, 1315, 99, 988, 753, 1197, 71, 255, 1411, 1494, 1081, 1414, 124, 1461, 294, 1101, 1176, 1397, 1413, 876, 364]. river- [18]. riverine [243, 1444, 82, 1062, 412, 1236, 290]. rivers [549, 358, 799, 216, 551, 1505, 248, 1415, 1392, 1320, 822, 425, 808, 409, 441, 961]. Riverscape [1443, 195, 910]. rivoliana [259]. rivularis [707]. Rivulidae [17, 603]. RLR [147]. RLR-signalling [147]. Roa [823]. roach [1345]. road [341, 260, 438]. roads [1067]. robusta [1285]. robustus [618]. Rocha [954]. rock [1437, 894]. rockcod [727]. rockfish [544]. rockfishes [1160]. rockskippers [1367]. rocky [256, 806, 709, 1166]. rocky-reef [709]. rod [635]. rogersi [1512]. rohita [884, 1266, 147, 65]. Role [847, 449, 957, 803, 721, 606, 144, 410, 767, 1308]. roseus [947]. Ross [1521]. rostral [1433, 905]. rostralis [866]. rostriformis [872]. Rostroraja [451]. rostrum [866]. rosy [161]. rough [779, 724]. roughening [1413]. roughy [947]. round [500, 1512, 1352, 1200, 330, 1295, 1009]. routes [113, 276, 1239]. rubescens [1021]. rubripes [1024, 1102]. rubropictus [1396]. ruff [1425]. ruffe [397]. Running [675, 1027]. Rüppell [979]. Russell [1085]. ruthenus [908]. Rutilus [359, 1345, 431].

S [233, 464, 696, 1467, 546]. S-56^{o} [696]. S. [1495]. S961 [1157]. Sábalo [760]. sac [450]. Saccharomyces [258]. Sacura [1092]. safe [740]. sagitta [1279]. sagittal [330, 1085, 1330]. Sagua [370]. saida [548, 1207]. sailfin [1064, 60, 225]. sailfish [866]. Saint [284, 1251, 284, 1131]. salar [1381, 1428, 625, 871, 930, 1118, 102, 1458, 620, 912, 479, 483, 1368, 968, 37, 274, 385, 997, 44, 973, 1136, 1517, 953, 1025, 647, 1432, 540, 306, 492, 1360, 1429, 181, 938, 38, 363, 1036, 378, 1441]. sale [1140]. saline [187]. salinities [1214]. Salinity [570, 1422, 1244, 80, 993, 593, 600, 1113]. Salminus [658]. Salmo [589, 625, 1382, 871, 930, 1118, 949, 102, 1458, 620, 990, 912, 1178, 1128, 1444, 755, 479, 423, 483, 989, 1368, 968, 37, 274, 385, 997, 44, 379, 888, 525, 545, 840, 973, 817, 996, 1136, 635, 1110, 1517, 953, 1025, 647, 1432, 540, 306, 1087, 492, 1360, 1154, 1121, 1429, 181, 938, 38, 663, 363, 1036, 378, 1441, 1381, 1074, 851, 1199, 1428]. salmoides [918, 173, 1076, 1510]. Salmon [819, 361, 836, 1141, 515, 625, 871, 930, 1118, 102, 1458, 620, 912, 735, 1431, 321, 1127, 1001, 1440, 479, 483, 799, 1368, 968, 37, 274, 385, 997, 44, 301, 1136, 924, 1245, 1517, 953, 1025, 1392, 323, 78, 1175, 156, 647, 521, 540, 158, 822, 683, 306, 1456, 333, 492, 1360, 769, 718, 1153, 1429, 181, 938, 1139, 1399, 38, 363, 1036, 378, 1441, 1381, 364, 854, 1519, 1428, 1432]. Salmonema [1456]. salmonid [480, 248, 381, 683]. Salmonidae [988, 152]. Salmoniformes [152]. salmonis [997, 363]. salt [645]. Salvelinus [589, 494, 292, 1045, 74, 1178, 155, 925, 1451, 1476, 231, 889, 851, 1182]. same [402, 172]. sample [295]. sampled [1141, 1045, 973, 442, 770, 819]. samples [1520, 905]. sampling [832, 1177, 1455, 170, 1027]. San [1031, 1390]. sanctipauli [1251]. Sanctuary [716]. sand [709, 281, 1366, 1281, 1394, 758]. sand-dwelling [281]. Sandelia [120, 401]. Sander [586, 1365, 911]. sandpaper [964]. sandsmelt [1347]. sandvicensis [1125]. sandwicensis [1480]. sanfranciscana [954]. Sanriku [822]. São [513, 7]. Sardina [1186]. sardine [1186]. Sargasso [1479]. Sarotherodon [800]. satellite [384, 1361, 1169, 1166]. Saudi [681]. Saurida [591]. saury [1346]. Sauvage [922]. savanna [1265]. sawfish [339]. sawshark [390]. sawsharks [567]. saxatilis [747]. scad [724]. scale [849, 769, 36, 1168, 244, 1027]. scales [620, 379, 1147, 306, 1087, 1360, 771, 503]. Scaling [1426, 1256]. scalloped [59, 931, 761, 1063, 199]. scaly [980]. scans [567]. Scaphirhynchus [52, 353]. scapularis [894]. Scaring [304]. Scartichthys [795]. Scarus [554]. Scavenging [282]. Schindler [681]. Schindleria [681]. Schindleriidae [681]. Schizothoracine [1013]. Schizothorax [505, 1415, 1246, 1413]. schlegelii [1012, 544, 711]. schmardae [198]. Schneider [692, 1501, 1032]. schooling [175]. Schultz [1029]. sciaenid [792]. Sciaenidae [810, 745, 1330]. Sciaenops [1283]. Science [93, 96, 750, 1509, 919, 1402, 1144, 955, 1094, 1364, 899]. scientific [1073, 830]. Scientometric [1144]. scleral [833]. Scomber [821]. Scomberomorus [1430, 794]. scombrus [821]. scope [190, 790, 1058]. Scopelogadus [1391]. Scophthalmidae [665, 1033]. Scophthalmus [285, 739]. Scorpaena [806]. Scorpaenichthys [1174]. Scorpaenidae [1117, 1310]. scorpaenids [806]. Scorpaeniformes [382, 1122, 1438]. scorpionfish [803]. Scotia [41]. Scotland [1423, 68, 1139]. Scottish [799]. sculpin [757]. sculpins [11, 108]. Scyliorhinidae [622, 356, 559]. Scyliorhinus [537, 246, 1520, 776, 192, 1088]. Scymnodon [590]. Se [394]. Sea [1256, 949, 990, 697, 1279, 41, 663, 1214, 1481, 1023, 891, 261, 372, 1329, 1354, 429, 162, 1294, 194, 989, 766, 462, 803, 379, 1034, 888, 1012, 996, 81, 1110, 188, 831, 1208, 991, 683, 626, 769, 221, 602, 1014, 841, 482, 23, 378, 747, 806, 359, 969, 963, 557, 772, 519, 200, 1400, 681, 1001, 1440, 1521, 285, 444, 779, 1213, 597, 799, 1491, 970, 957, 1274, 44, 591, 1188, 1228, 552, 1002, 1358, 1201, 1392, 212, 673, 1376, 1212, 801, 1159, 286, 1092, 781, 986, 777, 1088, 182, 1479, 1033, 768, 1466, 383, 456, 1085]. Sea [1339, 219]. sea-migration [683]. sea-winter [378]. seabass [894]. seabream [324, 258, 399, 733, 986, 1149, 69]. seabreams [509]. seadragon [842, 765]. seadragons [842]. seafood [1140]. seagrass [111]. seahorse [880, 1509, 271, 200, 992, 712, 698, 1384]. seahorses [610]. Seamount [1208]. seamounts [991]. Searching [842]. Seas [1046]. seascape [671]. seascapes [654, 1372]. Season [770, 1346, 888, 1351, 442, 841]. Seasonal [50, 277, 195, 925, 412, 371, 1265, 444, 1512, 1044, 2, 955, 287, 1246, 676]. seasonality [1442]. seaward [361, 683]. Sebastes [1171, 1495, 1160, 1310, 544]. second [1126]. secondary [391]. secretion [10]. Secretive [874]. secretory [10]. secrets [1216]. sectioning [1516]. sediment [708, 1342]. sediments [891]. see [415, 229, 830]. Seeking [227]. seerfish [1430]. segregation [443]. seining [1507]. select [1372]. selection [836, 615, 510, 619, 18]. selective [746]. selenium [394]. self [1125, 603]. self-fertilizing [603]. self-recruitment [1125]. semi [200, 786, 838]. semi-anadromous [786]. semi-circular [838]. semi-closed [200]. semifasciata [1150]. semilaevis [720, 706]. seminatural [1072]. semoni [347]. Semotilus [516]. sense [1496]. sensing [289]. sensitive [361, 455]. sensitivity [959, 1150]. senso [1020]. sensory [1496]. separate [1434]. separation [698]. sequence [331]. sequencing [517, 896, 713]. sequential [525]. sequentially [295]. sergeant [1446]. series [1308]. serine [1080]. serine/threonine [1080]. serine/threonine-protein [1080]. Seriola [259, 830]. Serranidae [624]. Serranus [1349]. serrasalmid [495]. Serrasalmidae [1006, 240, 1083]. Serrasalmus [1006, 351, 506]. serum [1284, 1055, 258, 505, 1474]. set [1020]. Setarchidae [803]. settlement [1125, 1050, 719, 108]. seven [264, 1330]. seven-gilled [264]. sevengill [458, 811, 272]. several [1065]. Sex [956, 371, 691, 1008, 896, 1085, 1060, 1072, 531, 507, 157, 236, 1299, 1191, 186, 582, 1107, 442, 1386, 1173, 251, 1370, 715, 489, 1347, 1217, 233]. sex-change [1386]. sex-determining [186]. Sex-linked [1085]. Sex-specific [896, 1299]. sexmaculata [1084]. Sexual [557, 102, 1197, 1458, 1512, 615, 670, 124, 1005, 666, 391]. sexually [1156, 382, 3, 452, 1056]. Shaanxi [914]. shad [386, 1465, 847]. shads [138]. shallow [1345, 163, 1507, 762, 425, 828]. shape [1023, 557, 21, 904, 138, 399, 1046, 428, 1245, 987, 146, 1464, 1186, 144, 1153, 36, 1168, 330]. shapes [1478]. shaping [638]. shared [1334]. shark [931, 96, 1116, 709, 942, 802, 688, 696, 622, 1019, 458, 650, 1331, 1296, 761, 280, 778, 1366, 1378, 630, 1063, 1333, 779, 463, 729, 970, 644, 811, 1287, 875, 762, 565, 1002, 568, 188, 1138, 199, 921, 430, 744, 445, 115, 178, 232, 616, 1336, 1008, 1338, 58, 920, 207, 804, 1057, 886, 1088, 1190, 1489, 1276, 1364, 1090, 1150, 903, 1379, 1235, 1303, 1223, 1259, 636, 1243]. shark-inflicted [804]. Sharks [293, 1394, 571, 435, 1132, 297, 684, 372, 678, 369, 360, 267, 264, 1450, 163, 436, 157, 449, 998, 814, 1362, 272, 955, 422, 827, 442, 1255, 1353, 454, 207, 1133, 522, 1014, 1286, 457, 349]. sharpen [782]. shed [517]. Shedding [552]. sheds [340]. sheepshead [556]. Sheffield [93]. shelf [1017, 724, 917, 841]. shellers [127]. shelling [183]. shelter [501]. shelters [754]. Shen [233]. Shetland [635, 1423]. shi [557]. Shield [389]. shift [223, 957, 1445, 1432]. shifting [1436]. shifts [277, 251, 1404, 221, 724, 1241]. shimofuri [1189]. Shiner [788]. shining [1143]. Shiretoko [1519]. shoaling [84, 666]. shoals [701]. shock [1264, 859, 870]. shocking [119]. shocks [1476]. shoemaker [111]. shore [589]. Short [214, 379, 230, 741, 995, 1491, 78, 822, 909]. Short- [214]. short-headed [1491]. Short-term [379, 230, 741, 78]. shortfin [1187, 1303]. shortnose [1277, 145, 1075]. shows [148, 1451, 1280]. shrimp [835]. shuttlebox [566]. Siberian [101, 1080, 215, 222]. sicklefin [1518]. sickness [691]. sicydiine [1016]. Sicyopterus [1016]. Sierra [569]. Siganus [111]. sightings [1515]. sign [651, 1107]. signal [414, 1147, 65]. signalling [767, 147]. signatures [1227, 893]. signifer [965]. significance [944, 1343]. significant [912, 856, 689]. significantly [1505]. Silas [255]. silent [899]. silondia [419]. Silonia [419]. siluri [1312]. Silurifomes [370]. Siluriformes [1350, 1022, 569, 1023, 555, 493, 1315, 907, 1225, 1288, 106, 99, 753, 4, 327, 736, 710, 1231, 906, 123, 1197, 389, 1107, 562, 1464, 1031, 1461, 513, 679, 345, 1496, 7, 25, 1170, 1203, 279, 601, 485, 1319, 1424, 855, 874, 950, 24, 649, 40, 774].

Silurotaenia [1312]. Silurus [1312]. Silva [954]. silver [224, 738, 974, 1250, 793, 1109, 71, 1222, 788]. silver-phase [974]. silvering [646]. silverside [1191, 393]. silversides [225]. Similar [127, 1103]. similarities [16]. simple [331]. simulated [1294]. since [1449]. Sindh [505]. sinensis [255, 1474, 1113]. single [1087]. Siniperca [1221, 915, 713]. Sinocyclocheilus [1414]. sires [1160]. Sisoridae [5, 1350, 1071]. sisoroid [1424]. sister [704]. Site [136, 762, 372, 678, 962, 732, 856, 1305, 689, 18, 896, 364, 1519]. site-selection [18]. sites [1361, 822]. situ [584, 1335, 1448]. six [264, 853, 945, 23, 1060]. six-gilled [264]. sixgill [115]. Size [872, 1512, 536, 1187, 557, 269, 639, 1458, 1495, 252, 618, 336, 865, 1135, 270, 659, 968, 37, 274, 44, 1311, 864, 885, 928, 1445, 381, 1047, 164, 109, 1360, 616, 282, 619, 1049, 909, 46, 330, 636]. size-at-age [1360]. size-at-maturity [1495]. size-based [619]. size-dependency [536, 252]. size-dependent [1445, 282]. size-mediated [659]. size-specific [872, 44]. sized [123, 1505]. sizes [1104]. skate [668, 652, 223, 1506, 1521, 1004, 1268, 1212, 427, 682, 451, 239, 964]. skates [3, 1208]. Skeletal [479, 475, 471, 472, 982, 467]. skeleton [472]. skilled [309]. skills [1283]. Skin [331, 258, 856, 737, 965, 739, 689, 298, 1260, 10, 821, 1217]. skin-related [298]. skipped [892]. sleeper [1331]. Sleeping [1040]. slender [299, 43]. slot [1413]. slow [1112]. slow-growing [1112]. Slower [961]. slowing [846]. sluice [1250]. Small [289, 537, 686, 123, 1505, 508, 1201, 1520, 192, 1169, 1186, 1166, 9, 818, 244, 854, 1308]. small-bodied [818]. small-scale [244]. small-sized [123]. small-spotted [537, 1520, 192]. smallest [446, 616]. smallmouth [910]. smallscale [508]. smalltooth [1366, 339]. smells [691]. smelt [149, 1365, 347]. smelts [959]. Smith [508]. smithi [881]. smolt [1431, 938, 378]. smoltification [379]. smolts [589, 1368, 301, 973, 953, 1025, 1392, 78]. Smooth [1164, 371, 47, 762, 206, 66, 1185, 886, 92]. smooth-hound [371, 762, 66, 886, 92]. snack [1008]. snail [1101]. snake [1435, 1027]. snakehead [1204, 1156, 1359]. snapper [1309, 577, 1210, 641, 1005, 1050, 511, 1105, 503, 1498]. snappers [893, 1344, 180]. sneaker [1119]. snook [605, 1278, 312]. snouted [1384]. snow [505, 1246]. snubnose [757]. snyderi [51]. Social [968, 366, 1098, 927, 606, 1370]. sockeye [924, 1245]. soft [1155]. Soldatov [1438]. sole [720, 597, 706, 43, 711]. Solea [711]. soluble [361]. somata [109]. somatic [1106]. Some [542, 106, 33, 376, 1496]. Somniosidae [590]. Somniosus [1331, 1235]. sonar [274, 1409]. Sons [233]. soon [169]. sordidus [1473]. soroides [558]. Sounds [1486, 351]. source [988]. sources [221]. South [120, 90, 644, 827, 1313, 950, 219, 1248, 359, 696, 1108, 428, 824, 1188, 124, 67, 796, 1472, 724, 874, 241, 852, 763, 1491, 1410, 91, 631, 685, 603, 133, 1466, 1418, 1242, 314, 1321]. south-eastern [124, 796, 874, 241]. south-west [824, 1188, 1472, 724, 763]. south-western [1248, 696, 1108, 428, 67]. southeast [951, 1435, 206, 1086]. southeastern [1411, 1260, 1190, 903]. Southern [41, 208, 256, 1006, 355, 368, 962, 907, 1318, 280, 358, 605, 853, 570, 255, 1201, 647, 128, 776, 595, 178, 984, 522, 154, 1224, 1211, 1381, 1256, 195, 139, 559, 1295]. Southernmost [156, 683, 198]. Southward [1129]. southwest [1236, 778, 398]. Southwestern [283, 1086, 74, 624, 1330, 734]. Sox [752, 19]. sp [577, 1098, 1171, 1041, 56, 193, 890, 128, 881, 1047, 773, 1100]. sp.n [764]. Space [765, 102, 979, 849, 851]. span [995]. Spanish [794]. sparid [69, 150]. Sparidae [194, 428, 509, 986, 482, 69]. sparids [1372]. sparing [1220]. Sparus [324, 258, 533, 399, 194, 733, 462, 482]. spathula [263]. Spatial [1265, 1378, 1507, 443, 1049, 587, 684, 1087, 1490, 254]. spatio [1253]. spatio-temporal [1253]. Spatiotemporal [256, 1346, 1355, 1502]. spawn [949, 212]. spawned [625, 1059]. spawners [1318]. Spawning [1251, 951, 1012, 1110, 584, 237, 530, 1326, 1003, 732, 865, 423, 87, 1044, 398, 822, 781, 1301, 633, 841, 487, 675, 165, 1206, 892, 364, 275, 768]. special [723, 118]. specialist [939]. speciation [661]. Species [358, 175, 1430, 350, 1022, 1298, 931, 96, 709, 121, 113, 261, 969, 664, 952, 142, 42, 440, 120, 1509, 1070, 519, 278, 1393, 588, 354, 493, 1041, 153, 1103, 1315, 907, 1179, 1225, 137, 904, 5, 100, 70, 106, 99, 983, 919, 1328, 681, 1438, 753, 1161, 479, 731, 756, 151, 104, 1137, 27, 789, 574, 970, 1410, 91, 4, 1501, 754, 83, 1231, 13, 326, 1273, 123, 39, 853, 631, 1356, 438, 1140, 594, 890, 824, 1228, 1435, 959, 1226, 302, 1048, 255, 1411, 600, 603, 914, 987, 539, 389, 1081, 1463]. species [518, 606, 592, 395, 124, 1271, 1324, 405, 240, 823, 827, 562, 88, 150, 68, 217, 704, 512, 424, 510, 1053, 1487, 122, 595, 1461, 674, 661, 176, 1310, 1159, 84, 660, 286, 388, 452, 513, 283, 284, 986, 1016, 1449, 1152, 9, 1051, 1084, 226, 443, 672, 345, 1279, 870, 92, 770, 7, 25, 623, 311, 221, 882, 873, 1421, 621, 1056, 257, 1229, 1498, 491, 1203, 629, 604, 1224, 441, 581, 119, 693, 627, 16, 391, 279, 601, 304, 239, 152, 1322, 1181, 1042, 275, 485, 1319, 1330, 874, 1466, 127, 383, 125]. species [1223, 300, 1239, 1418, 915, 390, 950, 1027, 839, 352, 782, 219, 254, 701, 711, 1363, 220, 1072, 250, 40, 774, 1321, 872]. Species- [872]. species-record [931]. species-rich [1239]. Species-specific [175]. Specific [737, 872, 1431, 195, 1127, 444, 1299, 968, 669, 175, 44, 794, 840, 81, 896, 180]. specificity [217]. specimens [404, 964]. speckled [1130]. spectral [959]. spectrum [1500]. speculation [137]. speed [1488, 1323, 462, 385, 333, 347, 1428]. speeds [1179, 1303]. Sperata [421]. Sperm [615, 600, 1294, 1064, 1160, 75, 648]. Spermatogenesis [328]. spermatozoon [215]. Sphaeramia [1490]. Sphoeroides [1010]. Sphyraena [1484, 210]. Sphyraenidae [1484]. Sphyrna [59, 931, 761, 280, 163, 1063, 729, 357, 979, 206, 199, 178]. Spice [62, 75]. spilauchen [584]. Spinacia [922]. spinax [162]. spined [1436, 1151, 1113, 1124, 790]. spinefoot [111]. spinetail [1518]. Spinibarbus [703, 1112]. spinifer [509]. Spinipterus [25]. spiny [1213, 1226, 302, 25, 538, 926, 460]. spiny-chest [460]. spinyhead [752, 1217]. Spiral [207]. sPLA [546]. splenic [400]. split [75, 604]. split-ejaculate [75]. splitfin [1147, 1143]. Spondyliosoma [150]. sponge [1394]. sponges [1394]. spot [28, 1114]. spotted [1430, 1204, 537, 1156, 507, 1359, 1512, 1482, 1520, 776, 192, 683, 294, 251, 1084, 604, 1033, 1182]. spp [91, 1160, 820, 576, 11, 1205, 456]. sprattus [673, 673]. spring [1044, 487, 394]. spring-spawning [1044]. Springer [616]. springs [647]. sprint [1179]. spy [1492]. squalicola [1235]. Squalidae [302]. Squaliformes [162, 302]. Squalius [1397]. Squalus [302, 1243]. Squatina [963, 1257]. Squatinidae [436]. Sri [1146]. St [1251, 1495, 1137, 1251]. stability [1420, 1230]. Stable [221, 363, 832, 1270, 214, 277, 1440, 1325, 957, 2, 1445, 695, 1390, 1184, 454, 1185, 770]. Stable-isotope [221]. Stage [794, 57, 157, 656, 442, 646, 453, 110, 313, 487, 397, 719]. Stage-specific [794]. stages [1500, 439, 148, 1103, 373, 1512, 593, 791, 1451, 700]. Staining [1446, 982]. stakeholder [783]. stamineus [1327]. standard [498, 1492]. stanza [904]. starring [1454]. starry [371, 66]. start [80, 1426, 296]. Starter [839]. starvation [915, 725]. State [1056, 40, 774, 956, 1427]. States [903]. station [218]. status [101, 1403, 1407, 1030, 1456, 1519]. stawiarski [774]. stay [182]. steelhead [593]. Stegastes [1251]. stehmanni [796]. Steindachner [1397, 1189]. Steindachneridion [906]. stellaris [1520, 1088]. stem [12]. stenohaline [51]. stenolepis [373]. stenotherms [1280]. stenotopic [1385]. stereological [1318]. sterlet [908]. Sternopygidae [388]. steroid [157]. steroids [1217]. Stethaprioninae [16]. Stethaprionini [504]. Stevardiinae [1396]. Stichaeidae [867]. Stichaeus [867]. stickleback [1436, 1151, 948, 1113, 790, 337]. sticklebacks [174]. still [698]. stimulation [65]. stimulations [1266, 147]. stimuli [292, 521]. sting [1262]. stingray [1079, 692, 1512, 85, 283, 226, 1196, 154, 708, 876]. stingrays [1164, 136, 563, 1185, 1069]. Stock [591, 812, 39, 1201, 453, 36, 1009]. stocked [396, 961]. stocking [635]. stocks [1046, 428, 1297]. stomach [696, 1171, 277, 957, 193]. stone [1405]. stored [1064, 1160]. stoutii [599]. straight [1135]. straight-mouth [1135]. Straightforward [822]. strains [1065]. Strait [926]. stranded [1513]. strategies [1178, 995, 2, 1489]. strategy [757, 382, 447, 103, 33, 15, 445, 674, 587, 191]. Stratoni [200]. straying [990]. streak [297]. stream [735, 558, 647, 1432, 718, 1017]. streams [1118, 1352, 598, 1162, 854]. strength [895]. strengthen [1209]. Strengths [509]. stress [101, 532, 494, 386, 6, 1329, 884, 524, 1221, 993, 922, 553, 925, 721, 923, 1264, 561, 315, 54, 1238, 1252, 859, 1099, 559, 1249, 837, 43, 821, 915, 1428, 1269]. stress-induced [6, 54]. stretch [1421]. striata [841]. striatus [1367]. strike [289]. striped [1461, 322, 723, 747]. Stromateiformes [1335]. Stromateoidei [336]. strong [1495, 948]. Strophidon [326]. structural [828, 873]. structurally [1372]. structure [941, 288, 688, 1341, 639, 243, 1018, 200, 416, 1108, 285, 194, 317, 866, 83, 490, 39, 265, 558, 762, 1350, 955, 150, 259, 1186, 474, 1106, 616, 1122, 36, 771, 503, 1170, 1176, 482, 1380, 758]. structures [242, 298, 1492]. structuring [676, 1308]. studies [862, 200, 606, 454, 1074, 1020]. study [415, 468, 1349, 1055, 99, 229, 533, 194, 766, 175, 1177, 1201, 1411, 987, 1264, 1043, 88, 1163, 920, 797, 1246, 482, 363]. stunts [54]. sturgeon [101, 730, 732, 52, 145, 1474, 1080, 1390, 353, 315, 215, 222, 19, 550]. Sturisoma [1081, 1463]. Styracura [198]. Sub [231]. Sub-lethal [231]. subarctic [1354, 939, 663]. subclade [106]. subfamily [726]. suboptimal [908]. subpopulations [402]. subsequent [385]. subsistence [123, 495]. subspecies [161]. substances [189]. substitute [481, 171]. Substrate [164, 1206, 1003, 708, 1413, 179]. Substrate-size [164]. substrates [442]. subterranean [1496]. subtropical [684, 422]. success [781, 487]. successful [525, 1399]. such [96]. sucker [1326, 195, 508]. suckermouth [629]. suckermouth-armoured [629]. suckleyi [1243]. Sudan [1288]. suggest [1026, 1057]. suggested [825]. suggests [803, 508, 177]. suitability [287, 1010]. suitable [1087]. Sulcus [1330]. Sumatra [1084]. summer [1118, 248, 1338]. Sundaland [661]. Sundarbans [638]. Sundays [1184]. sunfish [1091, 29, 896]. sunfishes [882]. superciliosus [1002]. supergene [453]. supersaturation [738, 1342]. supplement [1510]. Supplementary [1449]. supplementation [1284, 258, 922, 1388, 1015]. supplemented [1468]. support [812, 395, 1078]. supported [350, 1357]. supporting [1439, 1365]. supports [358, 1381, 432]. supposedly [874]. suppresses [79]. Suppression [376, 6]. surf [103, 1472]. surface [589, 378, 636]. Surfing [989]. surfperch [103]. surgeonfish [1500, 1125]. surgery [953]. Surgical [91, 510]. surinamensis [1483]. surubim [294]. survey [709, 720, 1328, 216, 617, 1159, 1211, 1375]. surveys [927, 814, 1507]. survivability [1024]. Survival [1037, 537, 247, 515, 516, 1244, 912, 738, 197, 1368, 973, 1517, 192, 1124, 598, 1283, 818, 883, 290]. survive [888]. susceptibility [925, 45]. suspended [1342]. sustainability [509]. sustainable [1276]. sustainably [727]. sustained [215, 43]. Susurluk [1322]. sutor [111]. Svalbard [548]. sveni [316]. swabs [1353]. Swainson [1463, 1081]. Swamp [1269, 1180]. swell [622]. swim [385]. swimbladder [809]. Swimming [875, 793, 1450, 516, 1179, 486, 983, 1499, 1135, 462, 1034, 145, 323, 540, 75, 1413, 1448, 347, 818, 1504, 1303, 1428, 172, 1342, 1473]. swordfish [596, 1376, 579, 547]. swordtail [1218]. sympatric [1248, 277, 1086, 1377, 1253, 509, 1062, 185, 92, 1344, 108]. sympatry [1178, 144]. Symposium [93]. synagris [1309]. Synaphobranchidae [1466]. synbiotic [1468]. Synechococcus [820]. Syngnathiformes [1134]. Synodontidae [1393, 591]. Synodontis [1288, 123]. synonym [1028, 590]. synonymization [954]. synonyms [1501]. synonymy [332]. synthesis [1144, 1040]. synthesizing [1055]. system [658, 1465, 245, 126, 566, 132, 273, 1110, 645, 1107, 1487, 1101, 161, 390, 649, 250]. Systematics [805, 634]. systems [1095, 76].

T. [83, 345]. tackling [750]. tactics [567, 1339]. taeniatus [156, 666]. taeniolatus [842, 765]. Tag [818, 861, 1037, 762, 953, 945]. tag-recapture [861, 762]. tagging [766, 1124, 114, 818]. tags [384, 741, 91, 1037, 1136, 396, 1316, 1169, 1166, 945, 598]. tail [1148]. Taiwan [703, 624]. take [256, 694, 1480]. Takifugu [51, 1024, 1102]. taking [109]. tanasi [1101]. Tanichthyidae [853]. Tanichthys [853]. Tapajós [672, 220, 1041, 1229]. tapetail [1408]. Taranetz [867]. tarapacana [1518]. target [278]. Targeted [1117, 660]. targets [1409]. tarpon [1244, 656]. task [376]. Taste [514]. taurine [1015]. taxa [909]. Taxonomic [862, 1463, 581, 24, 1430, 1023, 1030, 1435, 987, 240, 562, 1020]. Taxonomical [99]. taxonomy [569, 555, 867, 355, 551, 31, 68, 283, 345, 773, 604, 383]. TCR [578]. technique [982]. teeth [481, 468, 1296, 404, 194, 171, 482]. telemetry [396, 425]. teleost [237, 530, 1019, 501, 928, 617, 978, 1249, 397]. teleostean [1095]. Teleostei [810, 121, 952, 90, 281, 519, 1434, 588, 1041, 1028, 5, 795, 70, 328, 681, 194, 27, 139, 1410, 1501, 605, 13, 591, 631, 1389, 576, 1030, 1038, 925, 1071, 518, 1414, 124, 1107, 634, 1383, 991, 344, 595, 661, 1310, 917, 1449, 707, 1051, 621, 401, 482, 629, 441, 140, 734, 1322, 1181, 26, 383, 125]. Teleosteii [665, 39]. teleosts [281, 1060]. Telmatherinidae [225]. telomeric [1361]. temensis [1130]. temperate [549, 684, 299, 218, 554, 686, 534, 898, 84]. Temperature [532, 524, 149, 885, 978, 1317, 1298, 536, 252, 1130, 1244, 80, 981, 1135, 593, 56, 948, 659, 1191, 794, 840, 1517, 323, 1175, 561, 231, 353, 846, 790, 722, 1280, 1150, 1072]. temperature-dependent [1191, 1072]. temperatures [101, 1118, 566, 908, 1306, 1451, 1149, 173]. Temporal [269, 178, 1378, 643, 1253, 1170]. tench [265]. tendency [1295]. Tennessee [1101]. Tenualosa [245, 847]. term [208, 214, 741, 912, 1213, 216, 1514, 379, 230, 305, 509, 1264, 78, 1169, 1119, 114]. Terminalia [1100]. Terrestrial [235, 341, 260]. territory [851]. tessellated [469, 470]. test [385, 1058, 1428, 172]. testable [665]. tested [1034]. Testing [183, 1174, 891]. testis [1172]. testosterone [797]. tetra [402]. Texas [1264]. Thailand [1471, 558, 955]. Thalassoma [297]. Their [763, 569, 450, 1323, 1117, 795, 812, 799, 967, 1025, 212, 1119, 958, 1236, 542, 410, 1375]. Theodore [465]. theoretical [289]. Theory [381]. Thermal [560, 1299, 540, 445, 38, 730, 788, 1426, 1167, 992, 979, 1191, 885, 925, 534, 190, 1360, 1249, 1220, 1150, 1357, 179]. thermal-resistant [560]. thermic [386]. thermo [1135]. thermo-velocity [1135]. thermophilic [210]. thespesios [124]. Thiamin [1441]. Thiersant [914]. third [336, 568, 225]. Thirty [903]. Thirty-five [903]. thornback [813, 1423]. thorny [1506, 1512]. thread [929]. threadfin [826, 219]. threat [301, 703, 119]. threatened [247, 214, 983, 1348, 71, 262, 1190, 1101, 401, 1397, 758]. threats [655, 880, 1006]. Three [1109, 833, 127, 383, 1436, 1233, 1045, 281, 28, 1151, 5, 138, 756, 91, 39, 814, 1177, 1071, 600, 1145, 1310, 287, 1069, 984, 790, 873, 784, 140]. Three-dimensional [1109, 140]. three-spined [1436, 1151, 790]. three-spot [28]. threonine-protein [1080]. thresher [1002, 207, 1379]. thresholds [231]. Thrissina [1501]. throughout [158, 841, 936]. throughput [1205]. Thunnus [714, 1318, 617, 1340, 1205, 768]. thurstoni [1518]. Thymichthys [1470]. thynnus [1318, 617]. Thysochromis [125]. Tibet [1013]. Tibetan [1405]. tiburo [729, 357, 178]. tidal [1250]. tide [989, 847, 275]. tides [414]. tiger [650, 1366, 1287, 1024, 903]. tigerfish [1325, 131]. Tigris [1389]. tilapia [1284, 800, 1082, 946, 913, 20, 514, 455, 1219, 691, 700, 609, 637, 79, 1100, 1475]. tilefish [1313]. Time [613, 341, 260, 227, 1012, 1517, 1077, 1121, 828]. times [1374, 657]. timing [1326, 1281, 892]. Tinca [265]. tissue [331, 809, 454, 901, 222]. tissues [980, 1233, 1263, 1155, 115, 646, 770, 19]. titan [335]. TNFR [20]. toadfish [241, 899]. tobianus [281]. Together [95]. Tol2 [669]. Tol2-mediated [669]. Tolerance [386, 894, 1130, 535, 190]. tolerant [43]. tolerate [445]. Toll [1180, 1266]. Toll-like [1180, 1266]. tomographic [194, 482]. tomography [324]. tompot [1119]. tongue [720, 706]. too [574]. tool [891, 930, 475, 227, 418, 411, 259, 427, 1255, 334, 457, 899]. tools [1439]. tooth [998, 1433, 495]. toothfish [862, 861]. top [811]. topknot [665]. topmouth [318, 1177, 1194]. topography [1490]. torrent [950]. torus [1014]. total [738, 1342]. Toxicity [1412, 1100]. trace [980]. tracers [1171]. Trachelyopterus [651]. Trachichthyidae [947]. Trachinocephalus [1393]. Trachinotus [50, 1468, 1472]. Trachinus [1516]. Trachurus [417, 1274, 83, 490, 724]. Tracking [892]. tract [699, 1380, 26]. Trade [628, 393, 1245, 775, 783]. Trade-off [628, 1245]. Trade-offs [393]. Traditional [319, 409]. traf [958]. Training [184]. trait [956, 402, 400, 262, 443, 1230, 1009]. traits [439, 1256, 951, 501, 153, 1296, 37, 816, 446, 1387, 697, 1043, 638, 226, 92, 174, 1056, 491, 456, 1182, 1308]. trajectories [14, 188]. trans [488]. trans-Atlantic [488]. transboundary [1129]. transcript [209]. transcriptional [81]. Transcriptome [1112, 331, 752, 713, 1217]. transcriptome-based [752]. Transcriptomic [1221]. transducers [65]. transfer [1214]. transformation [57]. Transgene [471]. Transgene-mediated [471]. transgenesis [669]. transgenic [478]. transgrediens [191]. transition [256, 586, 791, 985, 1236]. transitioning [1455]. transitions [1281]. translucent [1231]. transmitters [305, 290]. transmontanus [1390]. transparent [766]. transponder [741, 1037, 818]. transporter [725]. trap [973]. trap-sampled [973]. Travelling [1121]. trawl [696]. trawling [1507]. treatment [6, 306]. Trends [1297, 256, 1073, 643, 1320]. trevally [1388]. triacis [757]. Triakis [1150]. trial [1306]. tributary [1326, 1431, 195, 1352]. tributary-specific [1431, 195]. Trichogaster [1030]. trichomycterid [855]. Trichomycteridae [907, 106, 99, 327, 736, 710, 764, 1231, 345, 1496, 855, 874, 40, 774]. Trichomycterus [106, 99, 345, 855]. Trichopodus [1030]. trick [700]. Tridens [1231]. tridentatus [253, 1027]. Tridentiger [1189]. triggerfish [488, 335]. triggerfishes [1054]. TRIM39 [913]. trimaculata [446]. trimaculatus [28, 898]. Trimma [995]. triostegus [1125]. tripletail [1483]. Triploid [483, 181, 480, 925]. Triplophysa [1405, 621]. troglomorphic [106]. Trombetas [104]. Trophic [223, 845, 1377, 459, 1184, 1069, 1133, 239, 250, 1132, 1086, 74, 1439, 449, 957, 803, 1253, 695, 422, 11, 1078, 144, 66, 1404, 1168, 623, 1088, 676, 363]. Tropical [1270, 650, 484, 714, 1331, 904, 1055, 367, 630, 643, 610, 1512, 1377, 1420, 63, 671, 551, 393, 146, 15, 84, 654, 1282, 585, 503, 1168, 870, 1196, 893, 676, 627, 244, 283]. tropicus [585]. trout [589, 1382, 1341, 949, 292, 990, 560, 486, 1178, 1128, 1444, 755, 423, 856, 989, 593, 574, 155, 379, 505, 888, 525, 545, 840, 305, 817, 996, 885, 635, 925, 1110, 1451, 31, 721, 689, 632, 231, 1087, 1477, 626, 1154, 573, 1121, 36, 1220, 1246, 663, 1157, 1504, 152, 1074, 851, 296, 1199, 460, 172, 166]. true [227, 257]. trunk [887]. trutta [589, 1382, 949, 990, 1178, 1444, 755, 423, 989, 379, 888, 525, 545, 840, 817, 996, 635, 1110, 1087, 663, 1074, 851, 1199]. Trygonorrhinidae [1277]. Trypauchen [1032]. Tschudi [894]. tshawytscha [1141, 1127, 1314, 323, 78, 819]. tsinlingensis [1412]. Tube [242]. Tube-like [242]. tubercles [642, 965]. tumbil [591]. tumour [20]. tuna [1227, 714, 310, 299, 1318, 1469, 633, 1340, 1205]. tunicates [757]. Turan [1028]. Turbo [718]. turbot [285, 739]. turbulent [793]. Turcichondrostoma [734]. Turcinoemacheilus [1389]. Turkey [1279, 152, 547]. Türkiye [1322, 1181]. Turn [1166]. turnover [1420, 1207]. turns [262]. turquoise [325, 1206, 839]. turtle [372]. tuskfish [334]. Twilight [210]. Two [350, 1231, 823, 513, 629, 318, 113, 243, 440, 868, 1070, 1086, 387, 1179, 1318, 759, 618, 983, 566, 14, 593, 1410, 274, 1253, 47, 393, 1356, 817, 996, 959, 82, 1306, 553, 509, 405, 1469, 1089, 674, 176, 190, 356, 660, 905, 1152, 9, 707, 129, 1099, 92, 7, 623, 65, 337, 1344, 304, 275, 349, 390, 108, 711, 1373, 1114, 1321]. two-dimensional [1373]. twospot [1302]. Tympanopleura [4]. Type [255, 671, 953, 581]. Typha [1272]. Typha-dominated [1272]. typhoon [744]. typus [297, 688, 865, 461, 287, 1259]. tyrosinase [784].

U.K [93]. U.S. [841, 1313]. uarnak [1514]. ubidiai [1031]. UCrit [145]. ugly [517]. UK [526, 233, 1483, 635, 682]. ulmoides [960]. ultimate [985]. Ultra [648]. Ultra-acute [648]. ultraconserved [826]. ultrasonography [1247]. ultrastructural [469, 873, 1380]. ultrastructure [887]. Umbrina [557]. uncertainty [408]. uncovering [1121]. undecimalis [1278, 312]. underestimate [348]. underlying [1283, 1036, 1174]. Understanding [439, 343, 1327, 107, 614, 60, 1183, 533, 1195, 1514, 427, 936, 899]. understudied [1138]. Underwater [267, 1275, 28, 1247, 1012, 1038]. underyearling [1118]. undetected [896]. Unexpected [261, 714]. unexploited [299]. unfamiliar [850]. uninfected [1047]. Uniparental [1341, 1462]. Unique [887, 907]. United [903, 941, 1366]. units [1242]. University [465, 571]. unknown [350, 685]. unlabelled [1205]. unlabelled-probe [1205]. unlock [203]. unnoticed [395]. unoccupied [709]. unprecedentedly [1495]. unravel [1341]. Unravelling [551, 36, 1406, 383]. unregulated [8]. unreported [1022]. unstocked [1519].

Untangling [1430]. unusual [757, 301]. Unveiling [1320]. update [293, 198, 873]. updated [664, 440, 859, 433]. upland [1118]. upon [48]. Upper [76, 1068, 1203, 40, 555, 42, 504, 123, 1389, 885, 255, 97, 1494, 1324, 16, 950, 1027, 24, 1288, 753, 651, 124, 1271, 122, 1068, 1203]. upregulate [645]. upriver [822]. ups [534, 1301]. Upside [1448]. Upside-down [1448]. upstream [333, 187, 999]. uptake [1192, 464, 1508, 498, 462, 172, 662]. upward [498]. upwelling [484, 362]. urban [516, 1162]. Urobatis [1512, 1196]. urogenital [97, 1005]. Urogymnus [927, 1061]. urophthalmus [303]. Uropterygius [1273]. Urotrygon [1512]. Uruguay [668, 294]. USA [336, 1272, 64, 1027, 1363]. Use [1439, 1247, 842, 765, 1248, 709, 963, 1326, 102, 384, 1171, 368, 112, 761, 200, 665, 74, 163, 692, 1333, 1307, 979, 849, 155, 73, 1305, 2, 438, 82, 1062, 430, 910, 64, 353, 1394, 334, 75, 1069, 1446, 322, 1340, 35, 851, 457]. used [479, 1412]. useful [1023]. Using [567, 1046, 619, 598, 862, 359, 969, 214, 620, 498, 1398, 1318, 1431, 1440, 566, 692, 157, 826, 216, 47, 591, 396, 272, 88, 412, 1184, 289, 1005, 425, 1492, 1185, 338, 1409, 1101, 1205, 1313, 982, 896, 364, 1330, 1174]. uterine [563]. utilisation [425, 1150]. Utility [424, 1344, 454]. Utilization [254, 1011]. utilize [1160]. utilized [846]. Uwa [1274].

vagina [1032]. vaigiensis [1446]. Valenciennes [1500, 1374, 208, 360, 894, 1336, 345, 869, 1421, 77, 1242]. Valid [498, 1016, 1058]. validates [1268]. Validating [1234, 1313]. validation [891, 714]. Valley [621, 1184, 1078]. valuable [1083, 893]. value [1293, 674]. values [498, 505]. valve [207]. Vandelliinae [764]. vanicolensis [869]. Variability [815, 269, 130, 799, 1299, 959, 1272, 206, 670, 534, 1449, 1357, 1355, 330]. variable [402, 56, 1451]. variables [576]. variata [1147, 1143]. Variation [295, 442, 1145, 46, 378, 956, 1298, 1309, 343, 1382, 1458, 1434, 493, 904, 988, 1512, 329, 792, 641, 1281, 2, 471, 1245, 697, 403, 535, 1464, 632, 540, 683, 144, 886, 1170, 18, 1283, 663, 1036, 937, 1182]. variations [256, 1003, 1055, 1273, 1346, 1085]. varied [814]. variegata [914]. variegatus [556]. variipictus [555]. various [334]. vary [110]. Varying [972, 81]. vasculature [170]. vateritic [515, 972]. Vaupés [97, 1319]. vegetated [1372]. vehicle [709]. velezi [451]. velocity [1135, 1478, 353]. velox [910]. velvet [552]. vendace [1358]. Venezuela [919, 327]. ventral [510]. ventriosum [622]. Venturi [665]. Venturi-effect [665]. Venus [669]. Veracruz [569]. Verification [275, 1202]. vermiculate [587]. vermiculatus [587]. versatile [475]. versus [1496, 697]. vertebrae [1227, 1348, 811, 474]. Vertebral [480, 3, 692, 1379]. vertebrate [839]. vertebrates [468]. Vertical [979, 384, 554, 596, 1340, 1413]. vertically [1340]. vestigial [400]. vgll3 [871]. via [86, 170, 1219, 1269]. viability [751]. Vibration [402]. Vibrio [1099]. Victoria [30, 937]. victoriae [53]. video [267]. videos [628]. Vietnam [1351, 1466]. viewed [1492]. Vigour [491]. Vigour-related [491]. villages [495]. violaceus [1288]. viraemia [394]. viridensis [210]. viridescens [335]. Virus [45, 394]. visitor [1400]. visits [334]. visual [494, 959]. visualize [1155]. vitellogenin [797]. vitreus [1365, 911]. vitro [1503, 960]. vittatus [1325, 131]. viviparous [450, 664, 6]. vivo [1503, 1157]. vlaminghi [786]. vocalization [798]. volitional [1135]. voltai [291]. volumes [295]. vortices [708]. vrolikii [1172]. vs. [973]. vulgaris [1284]. vulnerability [210, 92, 1280]. vulnerable [423, 419]. vulpes [87, 719]. vulpinus [1379].

W [465]. Waite [1103, 590]. Walbaum [1001, 1186, 1220, 296]. Wales [963]. walking [341, 260]. wall [510]. walleye [1365, 911]. Wang [233]. wangchiachii [1413]. wapixana [710]. Warm [647, 1309, 1505]. warming [532, 237, 531, 530, 537, 529, 806, 981, 533, 236, 205, 539, 528, 535, 192, 1417, 1429, 1280, 38, 378, 1491]. warmingii [1491]. warmwater [759]. was [1076]. waste [497]. wastes [763]. Water [1298, 1478, 1517, 409, 1052, 1082, 1132, 163, 1135, 860, 1188, 81, 570, 1272, 645, 539, 1053, 353, 1011, 1005, 584, 370, 187, 846, 637, 165, 1149, 1504, 1042, 180, 55]. water-level [1272]. waters [529, 89, 963, 458, 249, 1400, 630, 449, 490, 1425, 890, 1228, 565, 1123, 205, 146, 381, 158, 616, 747, 1027, 314]. waterway [999]. waterways [580]. Waveform [116]. way [898, 58]. weakly [53, 119]. weaknesses [509]. weather [413]. weatherfish [141]. web [1017, 1088]. webs [410]. wedgefish [366]. weedy [842, 765]. weeks [553]. weever [1516]. weight [904, 1123, 146]. weir [82]. Welfare [1306, 1499, 614, 619, 945]. well [1166, 874]. well-known [874]. We’re [879]. West [644, 1460, 695, 684, 824, 1188, 1254, 596, 1053, 1472, 29, 724, 125, 763, 833, 548, 602]. West-Svalbard [548]. Western [423, 1248, 696, 1488, 727, 1331, 404, 1108, 1161, 444, 151, 979, 1044, 591, 428, 446, 568, 1081, 1469, 67, 422, 921, 1376, 947, 768, 432, 1243, 1019, 322, 640]. wetland [1272, 174]. Weyl [377]. Whale [297, 96, 688, 955, 744, 457, 1259]. Where [126, 835, 415, 229, 573]. Whether [1411, 1195]. whipray [927, 1514, 1061, 244]. White [288, 369, 988, 400, 1390, 1182, 799, 853]. white-blooded [400]. white-spotted [1182]. whitefish [939, 895, 179, 648]. whitei [712]. whitespotted [843, 384, 845, 804]. Who [785]. wide [720, 1151, 993, 967, 752, 958, 1375, 1074, 173, 19]. widely [1041]. Widespread [158, 971, 345, 869]. wild [902, 218, 912, 560, 479, 399, 948, 37, 712, 1287, 230, 1455, 1362, 1351, 442, 540, 425, 687, 1024, 815, 378, 883, 961, 1174, 844]. wild-caught [230, 815, 1174]. wild-sampled [442]. Wiley [233]. will [539]. willdenowi [383]. wilsoni [926]. wimple [240]. window [981]. winghead [1138]. Winter [1352, 939, 1264, 378]. winters [647]. Wisconsin [580]. within [1298, 1382, 17, 787, 694, 271, 218, 242, 904, 402, 847, 686, 132, 518, 156, 150, 1487, 289, 370, 1203, 786, 183, 663, 1459, 1239]. within-species [904]. Wolastoq [1131]. Wolastoq/Saint [1131]. wolffish [772, 1482, 749]. won’t [1444]. World [1519, 916, 1290, 1470, 1091, 446, 461, 882, 1276, 21, 166]. Wound [159, 1057]. wrasse [297, 1499, 816, 1306, 275]. wrasses [686, 670].

X [1107, 324]. X-ray [324]. xanthos [139]. Xenotoca [1147, 1143]. Xiangshan [1104]. Xingu [279, 876]. Xinjiang [728]. Xiphias [596, 1376, 905, 579, 547]. Xiphiidae [905]. Xiphophorus [1218].

Ya [1107]. Yangtze [71, 255, 1494]. year [785, 580, 811, 149, 895, 337, 1211]. year-class [895]. years [256, 1431, 712, 1026, 1144, 1449, 903]. yeast [258]. Yellow [351, 1187, 791, 967, 1201, 958, 503, 1196, 1201]. Yellow-eyed [351]. yellowfin [1214, 1469, 1340, 869]. yellowmouth [210]. yellowspotted [727]. Yellowtail [1511]. yields [172]. yolk [450]. yolk-sac [450]. York [1364]. young [386, 785, 580, 811]. young-of-the-year [785, 811]. young-of-year [580]. YOY [785].

Z [233]. Zapteryx [1277, 1075]. Závora [962]. Zealand [490, 1459]. Zebrafish [148, 475, 850, 189, 642, 751, 1299, 204, 477, 471, 213, 472, 474, 1252, 473, 346, 394]. Zeugopterus [665]. zonation [868]. zone [256, 103, 661, 1472, 29]. zones [1392]. Zongolica [569]. zonura [1123]. Zoogeography [465]. zooplankton [673]. zygaena [206].

References

Anonymous:2020:IIa

[1] Anonymous. Issue information. Journal of Fish Biology, 96(1):1–2, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kaiser:2020:RSV

[2] Michel J. Kaiser. Reconstructing seasonal variation in egg-laying strategies: a forensic use of stable isotopes. Journal of Fish Biology, 96(1): 3, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

James:2020:VGB

[3] Kelsey C. James. Vertebral growth and band-pair deposition in sexually mature little skates Leucoraja erinacea: is adult band-pair deposition annual? Journal of Fish Biology, 96(1):4–13, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hashimoto:2020:MAS

[4] Shizuka Hashimoto, Lúcia H. Rapp Py-Daniel, and Jacqueline S. Batista. A molecular assessment of species diversity in Tympanopleura and Ageneiosus catfishes (Auchenipteridae: Siluriformes). Journal of Fish Biology, 96(1):14–22, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dao:2020:PTS

[5] Wei Dao, Xu Li, Hong-Fu Yang, and Wei Zhou. Pareuchiloglanis (Teleostei: Sisoridae) from the Pearl River, China with description of three new species. Journal of Fish Biology, 96(1):23–36, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bhat:2020:DTA

[6] Shilpa K. Bhat and C. B. Ganesh. Domperidone treatment attenuates stress-induced suppression of reproduction in viviparous mosquitofish Gambusia affinis. Journal of Fish Biology, 96(1):37–48, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rocha:2020:CLO

[7] Maíra S. A. Rocha, Robson C. Silva, José C. E. Santos, Marianne Schorer, Maria P. Nascimento, and Marcelo M. Pedreira. Comparative larval ontogeny of two fish species (Characiformes and Siluriformes) endemic to the São Francisco River in Brazil. Journal of Fish Biology, 96(1):49– 58, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rabuffetti:2020:CFM

[8] Ana Pia Rabuffetti, Luis A. Espínola, Elie Abrial, Mario L. Amsler, Martin C. M. Blettler, Maria F. Eurich, and Eliana G. Eberle. Commercial fisheries in a mega unregulated floodplain river: Assessment of the most favourable hydrological conditions for its preservation. Journal of Fish Biology, 96(1):59–73, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Quimbayo:2020:CFA

[9] Juan P. Quimbayo, Vinicius J. Giglio, Carlos E. L. Ferreira, Alana Fraga, José A. C. C. Nunes, and Thiago C. Mendes. Contrasting feeding and agonistic behaviour of two blenny species on a small and remote island in the equatorial Atlantic Ocean. Journal of Fish Biology, 96(1):74–82, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Seo:2020:MFM

[10] Eunseok Seo, Gun-Young Yoon, Hae-Nyeok Kim, Jae H. Lim, Seonghan Kim, Bumju Kim, Ki H. Kim, and Sang J. Lee. Morphological features of mucous secretory organ and mucous secretion of loach Misgurnus anguillicaudatus skin for friction drag reduction. Journal of Fish Biology, 96(1):83–91, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mozzon:2020:TNS

[11] Christina M. Mozzon, Gretchen L. Lescord, Pascale-Laure Savage, and Thomas A. Johnston. The trophic niche of sculpins Cottus spp. in forage fish assemblages of boreal lakes. Journal of Fish Biology, 96(1):92–101, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hussein:2020:CSP

[12] Mona N. A. Hussein, Xiaojuan Cao, Abdelmotaleb A. Elokil, and Songqian Huang. Characterisation of stem and proliferating cells on the retina and lens of loach Misgurnus anguillicaudatus. Journal of Fish Biology, 96(1):102–110, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huang:2020:OGT

[13] Jin-Qing Huang, Jian Yang, Zhi-Qiang Wu, and Ya-Hui Zhao. Oreonectes guilinensis (Teleostei, Cypriniformes, Nemacheilidae), a new loach species from Guangxi, China. Journal of Fish Biology, 96(1):111– 119, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Geheber:2020:DMG

[14] Aaron D. Geheber and Destini C. Miller. Differences in morphological growth trajectories between greenside darter Etheostoma blennioides populations from two distinct drainages. Journal of Fish Biology, 96 (1):120–129, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maitra:2020:FSD

[15] Saswata Maitra, Mahadevan Harikrishnan, and Balachandran Nidhin. Feeding strategy, dietary overlap and resource partitioning among four mesopredatory catfishes of a tropical estuary. Journal of Fish Biology, 96(1):130–139, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Teixeira:2020:NSH

[16] Tulio F. Teixeira, Guilherme M. Dutra, Iago S. Penido, Sérgio A. dos Santos, and Tiago C. Pessali. A new species of Hyphessobrycon (Characiformes, Characidae) from the upper Rio Jequitinhonha basin, Minas Gerais, Brazil, with comments on morphological similarities with Stethaprioninae species from headwater environments. Journal of Fish Biology, 96(1):140–153, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barbosa:2020:EGD

[17] Crislaine Barbosa, Daiana K. Garcez, Matheus V. Volcan, and Lizandra J. Robe. Evidences for genetic differentiation within the highly endemic and endangered annual fish Austrolebias nigrofasciatus (Cyprinodontiformes: Rivulidae). Journal of Fish Biology, 96(1):154–167, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sawada:2020:IAR

[18] Hayato Sawada, Kanji Shigeta, Masaki Kawakami, Masahide Yuma, and Atsushi Maruyama. Isotope analysis reveals proportional change and site-selection variation of river- and lake-produced eggs of a landlocked migratory fish. Journal of Fish Biology, 96(1):168–174, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yang:2020:GWA

[19] Jing Yang, Yacheng Hu, Jilu Han, Kan Xiao, Xueqing Liu, Chun Tan, Qingkai Zeng, and Hejun Du. Genome-wide analysis of the Chinese sturgeon sox gene family: identification, characterisation and expression profiles of different tissues. Journal of Fish Biology, 96(1):175–184, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Han:2020:ICT

[20] Xue qing Han, Fengying Gao, Maixin Lu, Zhi gang Liu, Miao Wang, Xiao li Ke, and Yan xia Gao. Identification and characterisation of tumour necrosis factor receptor (TNFR) associated factor 3 from Nile tilapia, Oreochromis niloticus. Journal of Fish Biology, 96(1):185–193, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bloom:2020:MMI

[21] Devin D. Bloom, Matthew Kolmann, Kimberly Foster, and Helen Watrous. Mode of miniaturisation influences body shape evolution in New World anchovies (Engraulidae). Journal of Fish Biology, 96(1):194–201, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Colautti:2020:ELR

[22] Darío C. Colautti, Leandro Miranda, Mariano Gonzalez-Castro, Vanina Villanova, Carlos A. Strüssmann, Miguel Mancini, Tomas Maiztegui, Gustavo Berasain, Ricardo Hattori, Fabian Grosman, Pablo Sanzano, Ismael Lozano, Sabina L. Vegh, Victor Salinas, Omar Del Ponti, Pamela del Fresno, Priscila Minotti, Yoji Yamamoto, and Claudio R. M. Baigún. Evidence of a landlocked reproducing population of the marine pejerrey Odontesthes argentinensis (Actinopterygii; Atherinopsidae). Journal of Fish Biology, 96(1):202–216, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stevens:2020:DSD

[23] Darren W. Stevens, Matthew R. Dunn, Matthew H. Pinkerton, and Janet M. Bradford-Grieve. Diet of six deep-sea grenadiers (Macrouridae). Journal of Fish Biology, 96(1):217–229, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zawadzki:2020:TRH

[24] Cláudio H. Zawadzki, Luiz F. C. Tencatt, and Heraldo A. Britski. Taxonomic revision of Hypostomus albopunctatus (Siluriformes: Loricariidae) reveals a new piece of the Hypostomus jigsaw in the upper Rio Paraná basin. Journal of Fish Biology, 96(1):230–242, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rocha:2020:NSS

[25] Marcelo Rocha, Felipe Rossoni, Alberto Akama, and Jansen Zuanon. A new species of spiny driftwood catfish Spinipterus (Siluriformes: Auchenipteridae) from the Amazon basin. Journal of Fish Biology, 96 (1):243–250, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vidal:2020:MHC

[26] Mateus R. Vidal, Thalles F. R. Ruiz, Diego D. dos Santos, Mario V. B. Gardinal, Fernando L. de Jesus, Claudemir K. Faccioli, Irene B. F. Vicentini, and Carlos A. Vicentini. Morphological and histochemical characterisation of the mucosa of the digestive tract in matrinxã Brycon amazonicus (Teleostei: Characiformes). Journal of Fish Biology, 96(1): 251–260, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grabowska:2020:CIP

[27] Anna I. Grabowska, Alicja Boroń, Lech Kirtiklis, Aneta Spóz, Dorota Juchno, and Jan Kotusz. Chromosomal inheritance of parental rDNAs distribution pattern detected by FISH in diploid F_{1} hybrid progeny of Cobitis (Teleostei, Cobitidae) species has non-Mendelian character. Journal of Fish Biology, 96(1):261–273, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cohn:2020:RUL

[28] Emilie Cohn, Patrick Cole, Amanda Haymaker, Austin M. Garner, and Richard L. Londraville. Response to underwater laser pointer in the orange-finned anemonefish Amphiprion chrysopterus and three-spot damselfish Dascyllus trimaculatus. Journal of Fish Biology, 96(1):274– 277, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sawai:2020:BHS

[29] Etsuro Sawai and Morihiko Yamada. Bump-head sunfish Mola alexandrini photographed in the north-west Pacific Ocean mesopelagic zone. Journal of Fish Biology, 96(1):278–280, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2020:VAB

[30] Anonymous. Victoria Anne Braithwaite 1967–2019. Journal of Fish Biology, 96(1):281–282, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Levings:2020:CTE

[31] Colin D. Levings. Cutthroat trout: evolutionary biology and taxonomy. Journal of Fish Biology, 96(1):283–284, January 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2020:IIb

[32] Anonymous. Issue information. Journal of Fish Biology, 96(2):285–286, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kaiser:2020:ACF

[33] Michel J. Kaiser. Amphidromy can be a flexible life history strategy in some hawai’ian gobies. Journal of Fish Biology, 96(2):287, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nordeide:2020:ABM

[34] Jarle Tryti Nordeide. Accuracy of body mass estimates of formalinpreserved fish — a review. Journal of Fish Biology, 96(2):288–296, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Singer:2020:AUB

[35] Randal A. Singer, Shari Ellis, and Lawrence M. Page. Awareness and use of biodiversity collections by fish biologists. Journal of Fish Biology, 96 (2):297–306, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rashidabadi:2020:USS

[36] Fahimeh Rashidabadi, Asghar Abdoli, Fatemeh Tajbakhsh, Farshad Nejat, and Esteban Avigliano. Unravelling the stock structure of the Persian brown trout by otolith and scale shape. Journal of Fish Biology, 96 (2):307–315, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hanson:2020:HAW

[37] Nora Hanson, James Ounsley, Tim Burton, Sonya Auer, James H. Hunt, Brian Shaw, Jim Henderson, and Stuart J. Middlemas. Hierarchical analysis of wild Atlantic salmon (Salmo salar) fecundity in relation to body size and developmental traits. Journal of Fish Biology, 96(2):316– 326, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Strom:2020:THA

[38] John Fredrik Strøm, Eva Bonsak Thorstad, and Audun Håvard Rikardsen. Thermal habitat of adult Atlantic salmon Salmo salar in a warming ocean. Journal of Fish Biology, 96(2):327–336, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jaafar:2020:CGS

[39] Tun N. A. Mat Jaafar, Martin I. Taylor, Siti A. Mohd Nor, Mark de Bruyn, and Gary R. Carvalho. Comparative genetic stock structure in three species of commercially exploited Indo–Malay Carangidae (Teleosteii, Perciformes). Journal of Fish Biology, 96(2):337–349, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

dosReis:2020:NSC

[40] Renan B. dos Reis, Augusto Frota, Thomaz M. C. Fabrin, and Weferson J. da Graça. A new species of Cambeva (Siluriformes, Trichomycteridae) from the Rio Ivaí basin, Upper Rio Paraná basin, Paraná State, Brazil. Journal of Fish Biology, 96(2):350–363, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Saunders:2020:AGB

[41] Ryan A. Saunders, Silvia Lourenço, Rui P. Vieira, Martin A. Collins, Carlos A. Assis, and Jose C. Xavier. Age and growth of Brauer’s lanternfish Gymnoscopelus braueri and rhombic lanternfish Krefftichthys anderssoni (family Myctophidae) in the Scotia Sea, Southern Ocean. Journal of Fish Biology, 96(2):364–377, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bogutskaya:2020:DNS

[42] Nina G. Bogutskaya, Dušan Jelić, Matej Vucić, Mišel Jelić, Oleg A. Diripasko, Tihomir Stefanov, and Göran Klobučar. Description of a new species of Phoxinus from the upper Krka River (Adriatic Basin) in Croatia (Actinopterygii: Leuciscidae), first discovered as a molecular clade. Journal of Fish Biology, 96(2):378–393, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tunnicliffe:2020:HTF

[43] Verena Tunnicliffe, Ryan Gasbarro, Francis Juanes, Jessica Qualley, Nicole Soderberg, and Jackson W. F. Chu. An hypoxia-tolerant flatfish: consequences of sustained stress on the slender sole Lyopsetta exilis (Pleuronectidae) in the context of a changing ocean. Journal of Fish Biology, 96(2):394–407, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jacobson:2020:PSS

[44] Philip Jacobson, Anna Gårdmark, and Magnus Huss. Population and size-specific distribution of Atlantic salmon Salmo salar in the Baltic Sea over five decades. Journal of Fish Biology, 96(2):408–417, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wang:2020:VSN

[45] Yizhou Wang, Ting Xue, Qian Wang, Bilin Xia, Qihua Pan, and Tiansheng Chen. Virus susceptibility of a new cell line derived from the fin of black carp Mylopharyngodon piceus. Journal of Fish Biology, 96 (2):418–426, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tagliafico:2020:VCR

[46] Alejandro Tagliafico, Paul A. Butcher, Andrew P. Colefax, Graeme F. Clark, and Brendan P. Kelaher. Variation in cownose ray Rhinoptera neglecta abundance and group size on the central east coast of Australia. Journal of Fish Biology, 96(2):427–433, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jargowsky:2020:DDH

[47] Matthew B. Jargowsky, Pearce T. Cooper, Matthew J. Ajemian, Michael E. Colvin, and J. Marcus Drymon. Discerning the dietary habits of the smooth butterfly ray Gymnura lessae using two distinct methods, otolith identification and metagenetics. Journal of Fish Biology, 96(2): 434–443, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harris:2020:EGU

[48] Sarah J. Harris, Daniel P. Bray, Mikolaj Adamek, David R. Hulse, Dieter Steinhagen, and David Hoole. Effect of β-1/3,1/6-glucan upon immune responses and bacteria in the gut of healthy common carp (Cyprinus carpio). Journal of Fish Biology, 96(2):444–455, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heim-Ballew:2020:MFN

[49] Heidi Heim-Ballew, Kristine N. Moody, Michael J. Blum, Peter B. McIntyre, and James D. Hogan. Migratory flexibility in native hawai’ian amphidromous fishes. Journal of Fish Biology, 96(2):456–468, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brownscombe:2020:SOC

[50] Jacob W. Brownscombe, Lucas P. Griffin, Danielle Morley, Alejandro Acosta, John Hunt, Susan K. Lowerre-Barbieri, Glenn T. Crossin, Sara J. Iverson, Ross Boucek, Aaron J. Adams, Steven J. Cooke, and Andy J. Danylchuk. Seasonal occupancy and connectivity amongst nearshore flats and reef habitats by permit Trachinotus falcatus: considerations for fisheries management. Journal of Fish Biology, 96(2):469–479, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nakamura:2020:FEB

[51] Masahiro Nakamura, Reiji Masuda, Katsumi Tsukamoto, and Tsuguo Otake. Freshwater entry behaviour of a non-migratory stenohaline marine fish Takifugu snyderi. Journal of Fish Biology, 96(2):480–485, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kashiwagi:2020:IGI

[52] Tom Kashiwagi, Aaron J. DeLonay, Patrick J. Braaten, Kimberly A. Chojnacki, Rachel M. Gocker, and Edward J. Heist. Improved genetic identification of acipenseriform embryos with application to the endangered pallid sturgeon Scaphirhynchus albus. Journal of Fish Biology, 96 (2):486–495, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moulton:2020:EHA

[53] Tyler L. Moulton, Lauren J. Chapman, and Rüdiger Krahe. Effects of hypoxia on aerobic metabolism and active electrosensory acquisition in the African weakly electric fish Marcusenius victoriae. Journal of Fish Biology, 96(2):496–505, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ord:2020:PBS

[54] James Ord, Kelle E. Holmes, William V. Holt, Alireza Fazeli, and Penelope J. Watt. Premature birth stunts early growth and is a possible driver of stress-induced maternal effects in the guppy Poecilia reticulata. Journal of Fish Biology, 96(2):506–515, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wakiya:2020:EDM

[55] Ryoshiro Wakiya, Kenzo Kaifu, Kazuhisa Azechi, Katsumi Tsukamoto, and Noritaka Mochioka. Evaluation of downward movements of Japanese eel Anguilla japonica inhabiting brackish water areas. Journal of Fish Biology, 96(2):516–526, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goertler:2020:LES

[56] Pascale A. L. Goertler, Anjali W. Shakya, Alicia M. Seesholtz, Brian M. Schreier, S. Zoltan Matica, and K. Sheena Holley. Lamprey (Entosphenus sp. and Lampetra sp.) estuarine occupancy is regionally variable and constrained by temperature. Journal of Fish Biology, 96(2):527–532, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Battaglia:2020:LSL

[57] Pietro Battaglia, David Smith, Giovanni Ammendolia, Mauro Cavallaro, Teresa Maggio, Pierpaolo Consoli, Maria C. Guerrera, Franco Andaloro, and Teresa Romeo. The last stage in the life history of the European conger eel Conger conger: a transformation before death. Journal of Fish Biology, 96(2):533–538, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pimentel:2020:BRL

[58] Caio R. Pimentel, Ryan Andrades, Carlos E. L. Ferreira, Otto B. F. Gadig, Euan S. Harvey, Jean-Christophe Joyeux, and Tommaso Giarrizzo. BRUVS reveal locally extinct shark and the way for shark monitoring in Brazilian oceanic islands. Journal of Fish Biology, 96(2):539–542, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aldana-Moreno:2020:RDM

[59] Alejandro Aldana-Moreno, E. Mauricio Hoyos-Padilla, Rogelio González-Armas, Felipe Galván-Magaña, Alex Hearn, A. Peter Klimley, William Winram, Edgar Eduardo Becerril-García, and James T. Ketchum. Residency and diel movement patterns of the endangered scalloped hammerhead Sphyrna lewini in the Revillagigedo National Park. Journal of Fish Biology, 96(2):543–548, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Raj:2020:AIA

[60] Smrithy Raj, Appukuttannair Biju Kumar, Rajeev Raghavan, and Neelesh Dahanukar. Amazonian invaders in an Asian biodiversity hotspot: Understanding demographics for the management of the armoured sailfin catfish, Pterygoplichthys pardalis in Kerala, India. Journal of Fish Biology, 96(2):549–553, February 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2020:IIc

[61] Anonymous. Issue information. Journal of Fish Biology, 96(3):555–556, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hubbard:2020:SYL

[62] Peter Colin Hubbard. Spice up your life! Journal of Fish Biology, 96(3): 557, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hagihara:2020:RPC

[63] Seishi Hagihara, Jun Aoyama, Ryusuke Sudo, Daniel Limbong, Shigeho Ijiri, Shinji Adachi, and Katsumi Tsukamoto. Reproductive physiological characteristics of tropical Celebes eels Anguilla celebesensis in relation to downstream migration and ovarian development. Journal of Fish Biology, 96(3):558–569, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morgan:2020:RUD

[64] Clark Morgan, Oliver N. Shipley, and James Gelsleichter. Resource-use dynamics of co-occurring chondrichthyans from the First Coast, North Florida, USA. Journal of Fish Biology, 96(3):570–579, March 2020. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sadangi:2020:MCE

[65] Sushmita Sadangi, Arpita Mohanty, Mahismita Paichha, and Mrinal Samanta. Molecular characterization and expression analysis of two crucial MAPKs-jnk1 and erk1 as cellular signal transducers in Labeo rohita in response to PAMPs stimulation and pathogenic invasion. Journal of Fish Biology, 96(3):580–589, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Phillips:2020:DCS

[66] Sophy R. McCully Phillips, Alastair Grant, and Jim R. Ellis. Diet composition of starry smooth-hound Mustelus asterias and methodological considerations for assessing the trophic level of predatory fish. Journal of Fish Biology, 96(3):590–600, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mas:2020:NRR

[67] Federico Mas, Rodrigo Forselledo, Andrés Domingo, Oscar Pin, Pablo Troncoso, Eugenia Errico, Alejandro Marquez, Sho Tanaka, and Simon Weigmann. New records and range extension of the Portuguese dogfish Centroscymnus coelolepis in the south-western Atlantic Ocean, with comments on its morphology. Journal of Fish Biology, 96(3):601–616, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McLeish:2020:FGE

[68] Jenny McLeish, Rob A. Briers, Jennifer A. Dodd, and Sonja Rueckert. First genetic evidence that invasive bullhead (Cottus L. 1758) in Scotland is of English origin and the difficulty of resolving the European Cottus species taxonomy. Journal of Fish Biology, 96(3):617–630, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Uehara:2020:EDO

[69] Masato Uehara and Katsunori Tachihara. Early development and occurrence patterns of the Pagrinae seabream Argyrops bleekeri (Perciformes: Sparidae): Does early life history have implications for sparid relationships? Journal of Fish Biology, 96(3):631–641, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dey:2020:PNN

[70] Abhinit Dey, Hrishikesh Choudhury, Abhishek Mazumder, Sarbojit Thaosen, and Dandadhar Sarma. Psilorhynchus nahlongthai, a new fish species (Teleostei: Psilorhynchidae) from the Brahmaputra drainage, northeast India. Journal of Fish Biology, 96(3):642–650, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2020:NBC

[71] Chao Li, Junjie Wang, Jiaqi Chen, Kevin Schneider, Radhakrishnan K. Veettil, Kathryn R. Elmer, and Jun Zhao. Native bighead carp Hypophthalmichthys nobilis and silver carp Hypophthalmichthys molitrix populations in the Pearl River are threatened by Yangtze River introductions as revealed by mitochondrial DNA. Journal of Fish Biology, 96(3):651– 662, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

LaPlante:2020:MMC

[72] Lori H. LaPlante and Samantha Delaney. Male mate choice for a female ornament in a monogamous cichlid fish, Mikrogeophagus ramirezi. Journal of Fish Biology, 96(3):663–668, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henseler:2020:PRC

[73] Christina Henseler, Marie C. Nordström, Anna Törnroos, Martin Snickars, and Erik Bonsdorff. Predation risk and competition affect habitat use of adult perch, Perca fluviatilis. Journal of Fish Biology, 96(3):669–680, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davidsen:2020:MTN

[74] Jan G. Davidsen, Michael Power, Rune Knudsen, Aslak D. Sjursen, Gaute Kjærstad, Lars Rønning, and Jo V. Arnekleiv. Marine trophic niche use and life history diversity among Arctic charr Salvelinus alpinus in southwestern Greenland. Journal of Fish Biology, 96(3):681–692, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Purchase:2020:FOF

[75] Craig F. Purchase and Anna C. Rooke. Freezing ovarian fluid does not alter how it affects fish sperm swimming performance: creating a cryptic female choice ‘spice rack’ for use in split-ejaculate experimentation. Journal of Fish Biology, 96(3):693–699, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kisekelwa:2020:ACF

[76] Tchalondawa Kisekelwa, Jos Snoeks, and Emmanuel Vreven. An annotated checklist of the fish fauna of the river systems draining the Kahuzi-Biega National Park (Upper Congo: Eastern DR Congo). Journal of Fish Biology, 96(3):700–721, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Swaminathan:2020:ECC

[77] T. Raja Swaminathan, Arathi Dharmaratnam, S. Arun Raja, N. Sundar Raj, and Kuldeep K. Lal. Establishment and cryopreservation of a cell line derived from caudal fin of endangered catfish Clarias dussumieri Valenciennes, 1840. Journal of Fish Biology, 96(3):722–730, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lulijwa:2020:MIR

[78] Ronald Lulijwa, Andrea C. Alfaro, Fabrice Merien, Mark Burdass, Jill Meyer, Leonie Venter, and Tim Young. Metabolic and immune responses of Chinook salmon (Oncorhynchus tshawytscha) smolts to a short-term poly (I:C) challenge. Journal of Fish Biology, 96(3):731–746, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vijayalaxmi:2020:OPD

[79] Vijayalaxmi and C. B. Ganesh. The opioid peptide dynorphin suppresses pituitary-ovary axis in the tilapia Oreochromis mossambicus. Journal of Fish Biology, 96(3):747–754, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Collar:2020:FSE

[80] David C. Collar, Jessica S. Thompson, Tyler C. Ralston, and Trevor J. Hobbs. Fast-start escape performance across temperature and salinity gradients in mummichog Fundulus heteroclitus. Journal of Fish Biology, 96(3):755–767, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kolosov:2020:CSL

[81] Dennis Kolosov, Phuong Bui, Mike P. Wilkie, and Scott P. Kelly. Claudins of sea lamprey (Petromyzon marinus) — organ-specific expression and transcriptional responses to water of varying ion content. Journal of Fish Biology, 96(3):768–781, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Koster:2020:HUM

[82] Wayne M. Koster, David R. Dawson, Adrian Kitchingman, Paul D. Moloney, and Robin Hale. Habitat use, movement and activity of two large-bodied native riverine fishes in a regulated lowland weir pool. Journal of Fish Biology, 96(3):782–794, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Healey:2020:GAP

[83] Amy J. E. Healey, Matthew W. Farthing, Francis K. E. Nunoo, Warren M. Potts, Warwick H. H. Sauer, Ilze Skujina, Nathan King, Sophie de Becquevort, Paul W. Shaw, and Niall J. McKeown. Genetic analysis provides insights into species distribution and population structure in East Atlantic horse mackerel (Trachurus trachurus and T. capensis). Journal of Fish Biology, 96(3):795–805, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Paijmans:2020:PAF

[84] Kai C. Paijmans, David J. Booth, and Marian Y. L. Wong. Predation avoidance and foraging efficiency contribute to mixed-species shoaling by tropical and temperate fishes. Journal of Fish Biology, 96(3):806– 814, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Keeping:2020:CAP

[85] Jennifer A. Keeping, Katie E. Reeve-Arnold, Neil M. Burns, Nilza Catarina, David M. Bailey, and Deborah C. McNeill. Computer-aided photoidentification of a rare stingray, Megatrygon microps. Journal of Fish Biology, 96(3):815–819, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fost:2020:RCD

[86] Brooks Fost, Kevin Morey, Bryan Ferguson, Danielle Bourque, Amanda Naaum, Doug Bradley, and Robert Hanner. Rapid cooling via dry ice preserves the genetic and morphological integrity of fish embryos. Journal of Fish Biology, 96(3):820–824, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Halstead:2020:ISE

[87] Williams R. Halstead, Sahar Mejri, Anthony C. Cianciotto, Paul S. Wills, Travis E. Van Leeuwen, Aaron J. Adams, and Jonathan M. Shenker. Induced spawning and embryonic and early larval development of bonefish (Albula vulpes). Journal of Fish Biology, 96(3):825–830, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McInnes:2020:NMS

[88] Mark G. McInnes, Neil M. Burns, Charlotte R. Hopkins, Guy P. Henderson, Deborah C. McNeill, and David M. Bailey. A new model study species: high accuracy of discrimination between individual freckled hawkfish (Paracirrhites forsteri) using natural markings. Journal of Fish Biology, 96(3):831–834, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Armstrong:2020:GDR

[89] Amelia J. Armstrong, Asia O. Armstrong, Michael B. Bennett, Frazer McGregor, Kátya G. Abrantes, Adam Barnett, Anthony J. Richardson, Kathy A. Townsend, and Christine L. Dudgeon. The geographic distribution of reef and oceanic manta rays (Mobula alfredi and Mobula birostris) in Australian coastal waters. Journal of Fish Biology, 96(3): 835–840, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cabezas:2020:FGR

[90] M. Pilar Cabezas, Oscar M. Lasso-Alcalá, Raquel Xavier, and Michael J. Jowers. First genetic record of the non-native muzzled blenny Omobranchus punctatus (Teleostei: Blenniidae) in the Atlantic Coast of Central and South America. Journal of Fish Biology, 96(3):841–846, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hanzen:2020:SIR

[91] Céline Hanzen, Martyn C. Lucas, Gordon O’Brien, Peter Calverley, and Colleen T. Downs. Surgical implantation of radio tags in three eel species (Anguilla spp.) in South Africa. Journal of Fish Biology, 96(3):847– 852, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Riginella:2020:CLH

[92] Emilio Riginella, Vittoria Correale, Ilaria A. M. Marino, Maria B. Rasotto, Anamarija Vrbatovic, Lorenzo Zane, and Carlotta Mazzoldi. Contrasting life-history traits of two sympatric smooth-hound species: implication for vulnerability. Journal of Fish Biology, 96(3):853–857, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Peirson:2020:BRE

[93] Graeme Peirson. Book review: Eels Biology, Monitoring, Management, Culture and Exploitation: Proceedings of the First International Eel Science Symposium. By A. Don & P. Coulson. (Eds). 504 pp. Published by 5m Publishing, Sheffield, U.K., 2019. Price £150.00. ISBN: 978-1-78918-069-5. Journal of Fish Biology, 96(3):858–859, March 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2020:IId

[94] Anonymous. Issue information. Journal of Fish Biology, 96(4):861–862, April 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kaiser:2020:TB

[95] Michel J. Kaiser. Together is bolder. Journal of Fish Biology, 96(4): 863, April 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Araujo:2020:GMC

[96] Gonzalo Araujo, Abdul R. Ismail, Cat McCann, David McCann, Christine G. Legaspi, Sally Snow, Jessica Labaja, Mabel Manjaji-Matsumoto, and Alessandro Ponzo. Getting the most out of citizen science for endangered species such as whale shark. Journal of Fish Biology, 96(4): 864–867, April 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lima:2020:NHU

[97] Flávio C. T. Lima, Alexander Urbano-Bonilla, and Saúl Prada-Pedreros. A new Hemigrammus from the upper Río Vaupés, Colombia (Characiformes: Characidae), with a discussion on the presence of an enlarged urogenital papilla in the family. Journal of Fish Biology, 96(4):868–876, April 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dias:2020:HDC

[98] Murilo S. Dias, Ian F. de Faria, Paula C. P. Guarido, Fabrício B. Teresa, Pedro de Podestà Uchôa de Aquino, and Juan P. Quimbayo. Historical distribution and current drivers of guppy occurrence in Brazil. Journal of Fish Biology, 96(4):877–885, April 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Donin:2020:TST

[99] Laura M. Donin, Juliano Ferrer, and Tiago P. Carvalho. Taxonomical study of Trichomycterus (Siluriformes: Trichomycteridae) from the Ribeira de Iguape River basin reveals a new species recorded in the early 20th century. Journal of Fish Biology, 96(4):886–904, April 2020. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeCahsan:2020:PEB

[100] Binia De Cahsan, Rebecca Nagel, Ina-Maria Schedina, James J. King, Pier G. Bianco, Ralph Tiedemann, and Valerio Ketmaier. Phylogeography of the European brook lamprey (Lampetra planeri) and the European river lamprey (Lampetra fluviatilis) species pair based on mitochondrial data. Journal of Fish Biology, 96(4):905–912, April 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aidos:2020:HDR

[101] Lucia Aidos, Alessandra Cafiso, Daniela Bertotto, Chiara Bazzocchi, Giuseppe Radaelli, and Alessia Di Giancamillo. How different rearing temperatures affect growth and stress status of Siberian sturgeon Acipenser baerii larvae. Journal of Fish Biology, 96(4):913–924, April 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bouchard:2020:SMI

[102] Colin Bouchard, Frédéric Lange, François Guéraud, Jacques Rives, and Cédric Tentelier. Sexual maturity increases mobility and heterogeneity in individual space use in Atlantic salmon (Salmo salar) parr. Journal of Fish Biology, 96(4):925–938, April 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Izumiyama:2020:SZR

[103] Michael Izumiyama, Michael F. Westphal, and Karen D. Crow. In the surf zone: Reproductive strategy of the calico surfperch (Amphistichus koelzi) in a comparative context. Journal of Fish Biology, 96(4):939– 949, April 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garcia-Ayala:2020:NAS

[104] James R. Garcia-Ayala and Ricardo C. Benine. A new Amazonian species of Brachychalcinus (Characiformes: Characidae) from the Trombetas River basin, Brazil. Journal of Fish Biology, 96(4):950–955, April 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rojo:2020:PGD

[105] Javier Hernán Rojo, Daniel Alfredo Fernández, Daniel Enrique Figueroa, and Claudia Clementina Boy. Phenotypic and genetic differentiation between diadromous and landlocked puyen Galaxias maculatus. Journal of Fish Biology, 96(4):956–967, April 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DoNascimiento:2020:NTS

[106] Carlos DoNascimiento and Saúl Prada-Pedreros. A new troglomorphic species of Trichomycterus (Siluriformes: Trichomycteridae) from northeastern Colombia, with proposal of a new Trichomycterus subclade and remarks on some nominal species from Colombia. Journal of Fish Biology, 96(4):968–985, April 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goodrich:2020:UGP

[107] Harriet R. Goodrich, Mark Bayley, Lina Birgersson, William G. Davison, Ora E. Johannsson, Anne B. Kim, Phuong Le My, Tran H. Tinh, Phuong N. Thanh, Huong Do Thi Thanh, and Chris M. Wood. Understanding the gastrointestinal physiology and responses to feeding in air-breathing Anabantiform fishes. Journal of Fish Biology, 96(4):986– 1003, April 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamazaki:2020:DST

[108] Aya Yamazaki, Akino Ogino, and Hiroyuki Munehara. Dispersion and settlement of two sympatric sculpins of the genus Gymnocanthus. Journal of Fish Biology, 96(4):1004–1013, April 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Otsuka:2020:PCI

[109] Airi Otsuka, Kenta Shimomura, Haruka Niwa, and Nao Kagawa. The presence of a conspecific induces risk-taking behaviour and enlargement of somata size of dopaminergic neurons in the brain of male medaka fish. Journal of Fish Biology, 96(4):1014–1023, April 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schmucker:2020:AEP

[110] Andrew K. Schmucker, Nicholas Scott Johnson, Ugo Bussy, Ke Li, Heather S. Galbraith, Yu-Wen Chung-Davidson, and Weiming Li. American eels produce and release bile acid profiles that vary across life stage. Journal of Fish Biology, 96(4):1024–1033, April 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ebrahim:2020:CIS

[111] Ameer Ebrahim, Jude P. Bijoux, Peter J. Mumby, and Ian R. Tibbetts. The commercially important shoemaker spinefoot, Siganus sutor, connects coral reefs to neighbouring seagrass meadows. Journal of Fish Biology, 96(4):1034–1044, April 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Byrnes:2020:RBD

[112] Evan E. Byrnes, Karissa O. Lear, David L. Morgan, and Adrian C. Gleiss. Respirometer in a box: development and use of a portable field respirometer for estimating oxygen consumption of large-bodied fishes. Journal of Fish Biology, 96(4):1045–1050, April 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Banon:2020:ACR

[113] Rafael Bañón, Alejandro de Carlos, Alexandre Alonso-Fernández, Fernando Ramos, and Francisco Baldó. Apparently contradictory routes in the expansion of two fish species in the Eastern Atlantic. Journal of Fish Biology, 96(4):1051–1054, April 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skov:2020:FEL

[114] Christian Skov, Joan H. Hansen, Henrik Baktoft, Christer Brönmark, Jakob Brodersen, Ben B. Chapman, Lars-Anders Hansson, Kaj Hulthén, and P. Anders Nilsson. A field evaluation of long-term effects of PIT tagging. Journal of Fish Biology, 96(4):1055–1059, April 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Osgood:2020:DBE

[115] Geoffrey J. Osgood, Brian Timmer, Kieran Cox, Francis Juanes, and Julia K. Baum. Differences in δ ^{15}N and δ ^{13}C between embryonic and maternal tissues of the ovoviviparous bluntnose sixgill shark Hexanchus griseus. Journal of Fish Biology, 96(4):1060–1064, April 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Waddell:2020:WDP

[116] Joseph C. Waddell and Angel A. Caputi. Waveform discrimination in a pair of pulse-generating electric fishes. Journal of Fish Biology, 96 (4):1065–1071, April 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2020:IIe

[117] Anonymous. Issue information. Journal of Fish Biology, 96(5):1073– 1074, May 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weyl:2020:JFB

[118] Olaf Lf Weyl and Albert Chakona. Journal of Fish Biology special issue on The Biology and Ecology of African Freshwater Fishes. Journal of Fish Biology, 96(5):1075–1076, May 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tagliacollo:2020:SDT

[119] Victor A. Tagliacollo, Priscila Camelier, Angela M. Zanata, and Roberto E. Reis. A shocking discovery of threat risks on newly described species of weakly electric fishes. Journal of Fish Biology, 96(5):1077– 1086, May 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bronaugh:2020:BOH

[120] Whitcomb Meade Bronaugh, Ernst R. Swartz, and Brian L. Sidlauskas. Between an ocean and a high place: coastal drainage isolation generates endemic cryptic species in the Cape kurper Sandelia capensis (Anabantiformes: Anabantidae), Cape Region, South Africa. Journal of Fish Biology, 96(5):1087–1099, May 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baba:2020:HSD

[121] Taylor Mambo Baba, Tchalondawa Kisekelwa, Célestin Danadu Mizani, Eva Decru, and Emmanuel Vreven. Hidden species diversity in Marcusenius moorii (Teleostei: Mormyridae) from the Congo Basin. Journal of Fish Biology, 96(5):1100–1122, May 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mulelenu:2020:CMO

[122] Christian Mukweze Mulelenu, Bauchet Katemo Manda, Eva Decru, Auguste Chocha Manda, and Emmanuel Vreven. The Cyphomyrus Myers 1960 (Osteoglossiformes: Mormyridae) of the Lufira basin (Upper Lualaba: DR Congo): a generic reassignment and the description of a new species. Journal of Fish Biology, 96(5):1123–1141, May 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ilunga:2020:DNS

[123] Micheline Kasongo Ilunga, Emmanuel Abwe, Eva Decru, Auguste Chocha Manda, and Emmanuel Vreven. Description of a new small-sized Synodontis species (Siluriformes: Mochokidae) that is important for local subsistence fisheries in the middle Lufira (upper Congo River, DR Congo). Journal of Fish Biology, 96(5):1142–1159, May 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Manda:2020:ETT

[124] Bauchet Katemo Manda, Jos Snoeks, Eva Decru, Roger Bills, and Emmanuel Vreven. Enteromius thespesios (Teleostei: Cyprinidae): a new minnow species with a remarkable sexual dimorphism from the southeastern part of the Upper Congo River. Journal of Fish Biology, 96 (5):1160–1175, May 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Walsh:2020:RCG

[125] Gina Walsh, Anton Lamboj, and Melanie L. J. Stiassny. Review of the cichlid genus Thysochromis (Teleostei: Ovalentaria) with the description of a new species from the Kouilou Province in the Republic of Congo, west– Central Africa. Journal of Fish Biology, 96(5):1176–1185, May 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Decru:2020:WIP

[126] Eva Decru, Nathan Vranken, Pedro H. N. Bragança, Jos Snoeks, and Maarten Van Steenberge. Where ichthyofaunal provinces meet: the fish fauna of the Lake Edward system, East Africa. Journal of Fish Biology, 96(5):1186–1201, May 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vranken:2020:SED

[127] Nathan Vranken, Maarten Van Steenberge, and Jos Snoeks. Similar ecology, different morphology: Three new species of oral-mollusc shellers from Lake Edward. Journal of Fish Biology, 96(5):1202–1217, May 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mipounga:2020:EPS

[128] Hans Kevin Mipounga, Joseph Cutler, Jean Hervé Mve Beh, Benjamin Adam, and Brian L. Sidlauskas. Enteromius pinnimaculatus sp. nov. (Cypriniformes: Cyprinidae) from southern Gabon. Journal of Fish Biology, 96(5):1218–1233, May 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ramoejane:2020:MME

[129] Mpho Ramoejane, Gavin Gouws, Ernst R. Swartz, Brian L. Sidlauskas, and Olaf L. F. Weyl. Molecular and morphological evidence reveals hybridisation between two endemic cyprinid fishes. Journal of Fish Biology, 96(5):1234–1250, May 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Frankowski:2020:GVD

[130] Jens Frankowski, Kay Lübke, Mike Coke, and Olaf L. F. Weyl. Genetic variability and demographic history of Anguilla mossambica (Peters, 1852) from continental Africa and Madagascar. Journal of Fish Biology, 96(5):1251–1259, May 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jacobs:2020:IMB

[131] Francois J. Jacobs, Tor F. Næsje, Eva M. Ulvan, Olaf L. F. Weyl, Deon Tiyeho, Clinton J. Hay, Gordon C. O’Brien, and Colleen T. Downs. Implications of the movement behaviour of African tigerfish Hydrocynus vittatus for the design of freshwater protected areas. Journal of Fish Biology, 96(5):1260–1268, May 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kadye:2020:ENP

[132] Wilbert T. Kadye and Anthony J. Booth. Environmental niche patterns of native and non-native fishes within an invaded African river system. Journal of Fish Biology, 96(5):1269–1277, May 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Magellan:2020:BCS

[133] Kit Magellan. Behavioural crypsis in a South African galaxiid fish induced by predatory and non-predatory heterospecifics. Journal of Fish Biology, 96(5):1278–1283, May 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Magellan:2020:IRC

[134] Kit Magellan, Anthony J. Booth, and Olaf L. F. Weyl. Innate responses to conspecific and heterospecific alarm cues in the endangered Eastern Cape redfin Pseudobarbus afer. Journal of Fish Biology, 96(5):1284– 1290, May 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2020:IIf

[135] Anonymous. Issue information. Journal of Fish Biology, 96(6):1291– 1292, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kaiser:2020:SFR

[136] Michel J. Kaiser. Site fidelity is not revealed by genetic markers in stingrays. Journal of Fish Biology, 96(6):1293, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cunningham:2020:DSI

[137] Eoghan M. Cunningham, Konstadinos Kiriakoulakis, Jaimie T. A. Dick, Louise Kregting, Pia Schuchert, and Julia D. Sigwart. Driven by speculation, not by impact — the effects of plastic on fish species. Journal of Fish Biology, 96(6):1294–1297, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dwivedi:2020:DTI

[138] Arvind K. Dwivedi. Differentiating three Indian shads by applying shape analysis from digital images. Journal of Fish Biology, 96(6):1298–1308, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Guclu:2020:BXN

[139] Salim Serkan Güçlü, Gökhan Kalaycı, Fahrettin Küçük, and Davut Turan. Barbus xanthos, a new barbel from the Southern Aegean basin (Teleostei: Cyprinidae). Journal of Fish Biology, 96(6):1309–1319, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thieme:2020:OGG

[140] Philipp Thieme and Timo Moritz. The osteology of the golden grey mullet Liza aurata (Teleostei: Mugiliformes: Mugilidae) including interactive three-dimensional reconstructions. Journal of Fish Biology, 96 (6):1320–1340, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2020:MCE

[141] Jian H. Chen, Xue Li, Meng J. Wang, Chun M. Wang, Yan Q. Peng, Hai H. Wang, and Ming Zhu. Molecular cloning and expression analysis of myd88 from oriental weatherfish (Misgurnus anguillicaudatus) in response to bacterial challenge. Journal of Fish Biology, 96(6):1341–1348, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Birindelli:2020:NES

[142] José L. O. Birindelli, Heraldo A. Britski, and Jorge L. Ramirez. A new endangered species of Megaleporinus (Characiformes: Anostomidae) from the Rio de Contas basin, eastern Brazil. Journal of Fish Biology, 96 (6):1349–1359, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rodriguez-Silva:2020:NLF

[143] Rodet Rodriguez-Silva and Pablo F. Weaver. A new livebearing fish of the genus Limia (Cyprinodontiformes: Poeciliidae) from Lake Miragoane, Haiti. Journal of Fish Biology, 96(6):1360–1369, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perez-Miranda:2020:RHS

[144] Fabian Pérez-Miranda, Omar Mejía, Alfonso A. González-Díaz, Norberto Martínez-Méndez, Eduardo Soto-Galera, Gerardo Zúñiga, and Oldrich Říčan. The role of head shape and trophic variation in the diversification of the genus Herichthys in sympatry and allopatry. Journal of Fish Biology, 96(6):1370–1378, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kieffer:2020:RUE

[145] James D. Kieffer and Lindsay E. May. Repeat UCrit and endurance swimming in juvenile shortnose sturgeon (Acipenser brevirostrum). Journal of Fish Biology, 96(6):1379–1387, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lopez-Perez:2020:LWR

[146] Cristina López-Pérez, M. Pilar Olivar, Percy A. Hulley, and Víctor M. Tuset. Length–weight relationships of mesopelagic fishes from the equatorial and tropical Atlantic waters: influence of environment and body shape. Journal of Fish Biology, 96(6):1388–1398, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mohanty:2020:MCE

[147] Arpita Mohanty, Sushmita Sadangi, Mahismita Paichha, and Mrinal Samanta. Molecular characterization and expressional quantification of lgp2, a modulatory co-receptor of RLR-signalling pathway in the Indian major carp Labeo rohita following pathogenic challenges and PAMP stimulations. Journal of Fish Biology, 96(6):1399–1410, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alfonso:2020:ZDR

[148] Sébastien Alfonso, Manon Peyrafort, Xavier Cousin, and Marie-Laure Bégout. Zebrafish Danio rerio shows behavioural cross-context consistency at larval and juvenile stages but no consistency between stages. Journal of Fish Biology, 96(6):1411–1421, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Keller:2020:TEE

[149] Marieke Keller, Pieke Molenaar, Joep de Leeuw, Wolf Mooij, Adriaan Rijnsdorp, and Karen van de Wolfshaar. Temperature effects on egg and larval development rate in European smelt, Osmerus eperlanus, experiments and a 50 year hindcast. Journal of Fish Biology, 96(6):1422–1433, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McKeown:2020:DPS

[150] Niall J. McKeown, Michael P. Gwilliam, Amy J. E. Healey, Ilze Skujina, Warren M. Potts, Warwick H. H. Sauer, and Paul W. Shaw. Deep phylogeographic structure may indicate cryptic species within the Sparid genus Spondyliosoma. Journal of Fish Biology, 96(6):1434–1443, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garcia-Alzate:2020:NSH

[151] Carlos Arturo García-Alzate, Flavio Lima, Donald Charles Taphorn, Jose Ivan Mojica, Alexander Urbano-Bonilla, and Tulio Franco Teixeira. A new species of Hyphessobrycon Durbin (Characiformes: Characidae) from the western Amazon basin in Colombia and Peru. Journal of Fish Biology, 96(6):1444–1453, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turan:2020:NST

[152] Davut Turan, Gökhan Kalaycı, Yusuf Bektaş, Cüneyt Kaya, and Esra Bayçelebi. A new species of trout from the northern drainages of Euphrates River, Turkey (Salmoniformes: Salmonidae). Journal of Fish Biology, 96(6):1454–1462, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Colchen:2020:OPF

[153] Tatiana Colchen, Andréa Dias, Enric Gisbert, Fabrice Teletchea, Pascal Fontaine, and Alain Pasquet. The onset of piscivory in a freshwater fish species: analysis of behavioural and physiological traits. Journal of Fish Biology, 96(6):1463–1474, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schwanck:2020:LLM

[154] Tanja N. Schwanck, Maximilian Schweinsberg, Kathrin P. Lampert, Tristan L. Guttridge, Ralph Tollrian, and Owen O’Shea. Linking local movement and molecular analysis to explore philopatry and population connectivity of the southern stingray oHypanus americanus. Journal of Fish Biology, 96(6):1475–1488, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harris:2020:AMH

[155] Les N. Harris, Heidi K. Swanson, Matthew J. H. Gilbert, Brendan K. Malley, Aaron T. Fisk, and Jean-Sébastien Moore. Anadromy and marine habitat use of lake trout (Salvelinus namaycush) from the central Canadian Arctic. Journal of Fish Biology, 96(6):1489–1494, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maldonado-Marquez:2020:SDL

[156] Alan Maldonado-Márquez, Tamara Contador, Javier Rendoll-Cárcamo, Sabrina Moore, Carolina Pérez-Troncoso, Daniel Gomez-Uchida, and Chris Harrod. Southernmost distribution limit for endangered Peladillas (Aplochiton taeniatus) and non-native Coho salmon (Oncorhynchus kisutch) coexisting within the Cape Horn Biosphere Reserve, Chile. Journal of Fish Biology, 96(6):1495–1500, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fujinami:2020:NLA

[157] Yuki Fujinami and Yasuko Semba. Non-lethal assessment of reproductive stage for female blue sharks Prionace glauca using sex steroid hormones. Journal of Fish Biology, 96(6):1501–1504, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nielsen:2020:WOP

[158] Julius Nielsen, Aqqalu Rosing-Asvid, Lorenz Meire, and Rasmus Nygaard. Widespread occurrence of pink salmon (Oncorhynchus gorbuscha) throughout Greenland coastal waters. Journal of Fish Biology, 96(6): 1505–1507, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bouyoucos:2020:WHE

[159] Ian A. Bouyoucos, Oliver N. Shipley, Emily Jones, Edward J. Brooks, and John W. Mandelman. Wound healing in an elasmobranch fish is not impaired by high-CO_{2} exposure. Journal of Fish Biology, 96(6):1508– 1511, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Finucci:2020:FRA

[160] Brittany Finucci. First record of albinism in the lanternshark family, etmopteridae. Journal of Fish Biology, 96(6):1512–1515, June 2020. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Umemura:2020:NGS

[161] Keitaro Umemura, Yoshihisa Kurita, and Norio Onikura. Novel genotyping system for distinguishing among native, non-native and admixed individuals of rosy bitterling Rhodeus ocellatus subspecies. Journal of Fish Biology, 96(6):1516–1522, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Duchatelet:2020:IDG

[162] Laurent Duchatelet, Nicolas Oury, Jérôme Mallefet, and Hélène Magalon. In the intimacy of the darkness: Genetic polyandry in deep-sea luminescent lanternsharks Etmopterus spinax and Etmopterus molleri (Squaliformes, Etmopteridae). Journal of Fish Biology, 96(6):1523–1529, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Doan:2020:ABS

[163] Melanie D. Doan and Stephen M. Kajiura. Adult blacktip sharks (Carcharhinus limbatus) use shallow water as a refuge from great hammerheads (Sphyrna mokarran). Journal of Fish Biology, 96(6):1530–1533, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nilsson:2020:SSC

[164] P. Anders Nilsson, Ivi J. Pettersson, Carl Tamario, Erik Degerman, Jonas Elghagen, Johan Watz, and Olle Calles. Substrate-size choice in European eel (Anguilla anguilla) elvers is not altered by piscivore chemical cues. Journal of Fish Biology, 96(6):1534–1537, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Takahashi:2020:BPO

[165] Tetsumi Takahashi and Stephan Koblmüller. Brood parasitism of an open-water spawning cichlid by the cuckoo catfish. Journal of Fish Biology, 96(6):1538–1542, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ottewell:2020:BRT

[166] David Ottewell. Book review: Trout and Char of the World. Journal of Fish Biology, 96(6):1543–1544, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2020:Ca

[167] Anonymous. Corrigendum. Journal of Fish Biology, 96(6):1545, June 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2020:IIg

[168] Anonymous. Issue information. Journal of Fish Biology, 97(1):1–2, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kaiser:2020:LCS

[169] Michel J. Kaiser. Lionfish coming soon on the menu in the Mediterranean. Journal of Fish Biology, 97(1):3, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lawrence:2020:BPN

[170] Michael J. Lawrence, Graham D. Raby, Amy K. Teffer, Ken M. Jeffries, Andy J. Danylchuk, Erika J. Eliason, Caleb T. Hasler, Timothy D. Clark, and Steven J. Cooke. Best practices for non-lethal blood sampling of fish via the caudal vasculature. Journal of Fish Biology, 97(1):4–15, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2020:HCD

[171] Moya Smith, Esther Manzanares, Charlie Underwood, Chris Healy, Brett Clark, and Zerina Johanson. Holocephalan (Chondrichthyes) dental plates with hypermineralized dentine as a substitute for missing teeth through developmental plasticity. Journal of Fish Biology, 97(1):16–27, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhang:2020:MMO

[172] Yangfan Zhang, Matthew J. H. Gilbert, and Anthony P. Farrell. Measuring maximum oxygen uptake with an incremental swimming test and by chasing rainbow trout to exhaustion inside a respirometry chamber yields the same results. Journal of Fish Biology, 97(1):28–38, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

White:2020:EEC

[173] Dalon P. White, Michael A. Nannini, and David H. Wahl. Examining the effects of chronic, lake-wide elevated temperatures on behavioural expression in largemouth bass, Micropterus salmoides. Journal of Fish Biology, 97(1):39–50, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rind:2020:MPT

[174] Khalid Rind, Quentin Rodriguez-Barucg, Delphine Nicolas, Patricia Cucchi, and Jehan-Hervé Lignot. Morphological and physiological traits of Mediterranean sticklebacks living in the Camargue wetland (Rhône river delta). Journal of Fish Biology, 97(1):51–63, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Holubova:2020:SSS

[175] Michaela Holubová, Petr Blabolil, Martin Čech, Mojmír Vašek, and Jiří Peterka. Species-specific schooling behaviour of fish in the freshwater pelagic habitat: an observational study. Journal of Fish Biology, 97(1): 64–74, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nyalungu:2020:FMC

[176] Nonhlanhla P. Nyalungu and Vanessa Couldridge. Female mate choice and species recognition between two closely related cichlid fish of Lake Malaŵi, Metriaclima estherae and M. callainos. Journal of Fish Biology, 97(1):75–82, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Soo:2020:FCF

[177] Han-Jie Soo, Ka Kei Sam, Joey Chong, Nyok-Sean Lau, Seng Yeat Ting, Meng-Kiat Kuah, Sim Yee Kwang, Manogoran Ranjani, and Alexander Chong Shu-Chien. Functional characterisation of fatty acyl desaturase, Fads2, and elongase, Elovl5, in the Boddart’s goggle-eyed goby Boleophthalmus boddartia (Gobiidae) suggests an incapacity for longchain polyunsaturated fatty acid biosynthesis. Journal of Fish Biology, 97(1):83–99, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Palacios-Hernandez:2020:TLC

[178] Daniela Palacios-Hernández, José L. Castillo-Géniz, Iván Méndez-Loeza, and Juan C. Pérez-Jiménez. Temporal and latitudinal comparisons of reproductive parameters in a heavily exploited shark, the bonnethead, Sphyrna tiburo (L. 1758), in the southern Gulf of Mexico. Journal of Fish Biology, 97(1):100–112, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thome:2020:MEC

[179] Christopher Thome, Taylor Laframboise, Charles Mitz, Eric Clancy, Jacob Bates, Christopher M. Somers, Richard G. Manzon, Joanna Y. Wilson, John M. Gunn, and Douglas R. Boreham. Modifying effects of a cobble substrate on thermal environments and implications for embryonic development in lake whitefish (Coregonus clupeaformis). Journal of Fish Biology, 97(1):113–120, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Uehara:2020:CAS

[180] Masato Uehara, Akihiko Ebisawa, and Itaru Ohta. Comparative agespecific demography of four commercially important deep-water snappers: implication for fishery management of a long-lived lutjanid. Journal of Fish Biology, 97(1):121–136, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sambraus:2020:TAS

[181] Florian Sambraus, Tom Hansen, Britt S. Daae, Anders Thorsen, Roar Sandvik, Lars H. Stien, Thomas W. K. Fraser, and Per Gunnar Fjelldal. Triploid Atlantic salmon Salmo salar have a higher dietary phosphorus requirement for bone mineralization during early development. Journal of Fish Biology, 97(1):137–147, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Savva:2020:THS

[182] Ioannis Savva, Niki Chartosia, Charalampos Antoniou, Periklis Kleitou, Andreas Georgiou, Nir Stern, Louis Hadjioannou, Carlos Jimenez, Vasilis Andreou, Jason M. Hall-Spencer, and Demetris Kletou. They are here to stay: the biology and ecology of lionfish (Pterois miles) in the Mediterranean Sea. Journal of Fish Biology, 97(1):148–162, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

StJohn:2020:OSW

[183] Michelle E. St. John, Kristi E. Dixon, and Christopher H. Martin. Oral shelling within an adaptive radiation of pupfishes: Testing the adaptive function of a novel nasal protrusion and behavioural preference. Journal of Fish Biology, 97(1):163–171, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rae:2020:TFR

[184] David L. Rae, Benjamin Mos, Anna Scott, and Symon A. Dworjanyn. Training fish for restocking: refuge and predator training in the hatchery has limited benefits for a marine fish. Journal of Fish Biology, 97(1):172– 182, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pauers:2020:AAA

[185] Michael J. Pauers and Jacob A. Grudnowski. Assortative aggression among males in a sympatric pair of Labeotropheus from Lake Malaŵi, Africa. Journal of Fish Biology, 97(1):183–189, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kongsstovu:2020:IMH

[186] Sunnvør í Kongsstovu, Hans Atli Dahl, Hannes Gislason, Eydna Homrum, Jan Arge Jacobsen, Paul Flicek, and Svein-Ole Mikalsen. Identification of male heterogametic sex-determining regions on the Atlantic herring Clupea harengus genome. Journal of Fish Biology, 97(1):190– 201, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oto:2020:PSW

[187] Yumeki Oto. Preference for saline water of an amphidromous goby maintained during migration to upstream freshwater areas. Journal of Fish Biology, 97(1):202–211, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lopez-Romero:2020:GTP

[188] Faviel A. López-Romero, Claudia Klimpfinger, Sho Tanaka, and Jürgen Kriwet. Growth trajectories of prenatal embryos of the deep-sea shark Chlamydoselachus anguineus (Chondrichthyes). Journal of Fish Biology, 97(1):212–224, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cao:2020:ESI

[189] Xiaolong Cao and Weiming Li. Embryonic substances induce alarm response in adult zebrafish (Danio rerio). Journal of Fish Biology, 97(1): 225–230, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nyboer:2020:DAS

[190] Elizabeth A. Nyboer, Emmanuelle Chrétien, and Lauren J. Chapman. Divergence in aerobic scope and thermal tolerance is related to local thermal regime in two populations of introduced Nile perch (Lates niloticus). Journal of Fish Biology, 97(1):231–245, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yogurtcuoglu:2020:IEI

[191] Baran Yoğurtçuoğlu, Umut Uyan, and F. Güler Ekmekçi. The influence of environmental instability on the reproductive strategy of the critically endangered Acıgöl killifish (Aphanius transgrediens). Journal of Fish Biology, 97(1):246–256, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Musa:2020:OWH

[192] Syafiq M. Musa, Daniel M. Ripley, Timo Moritz, and Holly A. Shiels. Ocean warming and hypoxia affect embryonic growth, fitness and survival of small-spotted catsharks, Scyliorhinus canicula. Journal of Fish Biology, 97(1):257–264, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jacobi:2020:FAS

[193] Cristina M. Jacobi, Francisco Villamarín, João V. Campos-Silva, Timothy Jardine, and William E. Magnusson. Feeding of Arapaima sp.: integrating stomach contents and local ecological knowledge. Journal of Fish Biology, 97(1):265–272, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Germain:2020:TSH

[194] Damien Germain and François J. Meunier. A tomographic study of the histological structure of teeth in the gilthead sea bream, Sparus aurata (Teleostei, Perciformes, Sparidae). Journal of Fish Biology, 97(1):273– 278, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dyer:2020:SMT

[195] Joseph J. Dyer and Shannon K. Brewer. Seasonal movements and tributary-specific fidelity of blue sucker Cycleptus elongatus in a Southern Plains riverscape. Journal of Fish Biology, 97(1):279–292, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Campbell:2020:DRO

[196] Matthew A. Campbell, Pakorn Tongboonkua, Bruno Chanet, and Wei-Jen Chen. The distribution of the recessus orbitalis across flatfishes (order: Pleuronectiformes). Journal of Fish Biology, 97(1):293–297, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garcia:2020:EPD

[197] Daniel García and Martin Reichard. The effect of population density on growth and survival of a Neotropical annual killifish. Journal of Fish Biology, 97(1):298–301, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sales:2020:SRU

[198] Manuela Alves Nobre Sales, João Eduardo Pereira de Freitas, Carlos Costa Cavalcante, Jones Santander-Neto, Patricia Charvet, and Vicente Vieira Faria. The southernmost record and an update of the geographical range of the Atlantic chupare, Styracura schmardae (Chondrichthyes: Myliobatiformes). Journal of Fish Biology, 97(1):302–308, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martins:2020:RFM

[199] Mariana F. Martins, Fernanda A. Rolim, and Otto B. F. Gadig. Report on fetal mummification in the scalloped hammerhead shark Sphyrna lewini. Journal of Fish Biology, 97(1):309–313, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Correia:2020:FSS

[200] Miguel Correia, Diogo Paulo, Elina Samara, Panayota Koulouri, Vasilis Mentogiannis, and Costas Dounas. Field studies of seahorse population density, structure and habitat use in a semi-closed north-eastern Mediterranean marine area (Stratoni, North Aegean Sea). Journal of Fish Biology, 97(1):314–317, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2020:E

[201] Anonymous. Erratum. Journal of Fish Biology, 97(1):318–319, July 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2020:IIh

[202] Anonymous. Issue information. Journal of Fish Biology, 97(2):321–322, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kaiser:2020:OMH

[203] Michel J. Kaiser. Otolith microchemistry helps to unlock the chronology of age determination. Journal of Fish Biology, 97(2):323, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kelly:2020:IER

[204] Jeffrey R. Kelly and Scott A. Benson. Inconsistent ethical regulation of larval zebrafish in research. Journal of Fish Biology, 97(2):324–327, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Little:2020:WDW

[205] Alexander G. Little, Isabella Loughland, and Frank Seebacher. What do warming waters mean for fish physiology and fisheries? Journal of Fish Biology, 97(2):328–340, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lopez-Martinez:2020:IVD

[206] Juana López-Martínez, Carmen Cabanilla-Carpio, Walter Ruiz Choez, and Edgar Arnoldo Arzola-Sotelo. Interannual variability of distribution, abundance and population dynamics of the smooth hammerhead Sphyrna zygaena (Linnaeus, 1758) in the central-southeast Pacific Ocean. Journal of Fish Biology, 97(2):341–353, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Preti:2020:SVP

[207] Antonella Preti, Ken MacKenzie, Kate A. Spivey, Leslie R. Noble, Catherine S. Jones, Ralph G. Appy, and Graham J. Pierce. Spiral valve parasites of blue and common thresher sharks as indicators of shark feeding behaviour and ecology. Journal of Fish Biology, 97(2):354–361, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anderson:2020:PEI

[208] Antônio B. Anderson, Jodir Pereira da Silva, Raquel Sorvilo, Carlo Leopoldo B. Francini, Sergio R. Floeter, and João P. Barreiros. Population expansion of the invasive Pomacentridae chromis limbata (Valenciennes, 1833) in southern Brazilian coast: long-term monitoring, fundamental niche availability and new records. Journal of Fish Biology, 97 (2):362–373, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rahdari:2020:EKI

[209] Abdolali Rahdari, Majidreza Khoshkholgh, Mahtab Yarmohammadi, Maren Ortiz-Zarragoitia, Pieter Mark Lokman, Sobhan R. Akhavan, Oihane Diaz de Cerio, Ibon Cancio, and Bahram Falahatkar. The effects of 11-ketotestosterone implants on transcript levels of gonadotropin receptors, and foxl2 and dmrt1 genes in the previtellogenic ovary of cultured beluga (Huso huso). Journal of Fish Biology, 97(2):374–382, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Merciai:2020:TAP

[210] Roberto Merciai, Margarida Casadevall, Harold Villegas-Hernández, and Josep Lloret. Twilight activity patterns and angling vulnerability of yellowmouth barracuda (Sphyraena viridensis Cuvier, 1829), a rangeexpanding thermophilic fish. Journal of Fish Biology, 97(2):383–395, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rolim:2020:GDL

[211] Fernanda A. Rolim, Zachary A. Siders, Fabio P. Caltabellotta, Matheus M. Rotundo, and Teodoro Vaske-Júnior. Growth and derived life-history characteristics of the Brazilian electric ray Narcine brasiliensis. Journal of Fish Biology, 97(2):396–408, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Medina:2020:CDC

[212] Antonio Medina, Francisco J. Abascal, and José L. Varela. Common dolphinfish, Coryphaena hippurus, spawn at their easternmost Atlantic range adjacent to the Mediterranean Sea. Journal of Fish Biology, 97 (2):409–415, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lucon-Xiccato:2020:ICZ

[213] Tyrone Lucon-Xiccato and Cristiano Bertolucci. Inhibitory control in zebrafish, Danio rerio. Journal of Fish Biology, 97(2):416–423, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burgess:2020:SLT

[214] Katherine B. Burgess, Matt K. Broadhurst, Vincent Raoult, Betty J. L. Laglbauer, Melinda A. Coleman, and Michael B. Bennett. Short- and long-term diets of the threatened longhorned pygmy devil ray, Mobula eregoodoo determined using stable isotopes. Journal of Fish Biology, 97 (2):424–434, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rahi:2020:EPA

[215] Deepali Rahi, Borys Dzyuba, Miaomiao Xin, Yu Cheng, and Viktoriya Dzyuba. Energy pathways associated with sustained spermatozoon motility in the endangered Siberian sturgeon Acipenser baerii. Journal of Fish Biology, 97(2):435–443, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goutte:2020:MFF

[216] Aurélie Goutte, Noëlie Molbert, Sabrina Guérin, Robin Richoux, and Vincent Rocher. Monitoring freshwater fish communities in large rivers using environmental DNA metabarcoding and a long-term electrofishing survey. Journal of Fish Biology, 97(2):444–452, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Minder:2020:DSO

[217] Mario Minder, Emily R. Arsenault, Bolortsetseg Erdenee, and Mark Pyron. Dietary specificity and overlap in endorheic river fishes: How do native and nonnative species compare? Journal of Fish Biology, 97(2): 453–464, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cook:2020:IFL

[218] Denham G. Cook, Peter Jaksons, and Alistair Jerrett. Investigating food limitations in wild fisheries: the attendance and growth responses of fish at an anthropogenic feeding station within a temperate estuary. Journal of Fish Biology, 97(2):465–478, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhang:2020:PDT

[219] Kui Zhang, Yancong Cai, Baochao Liao, Yan’e Jiang, Mingshuai Sun, Li Su, and Zuozhi Chen. Population dynamics of threadfin porgy Evynnis cardinalis, an endangered species on the IUCN red list in the Beibu Gulf, South China Sea. Journal of Fish Biology, 97(2):479–489, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deOliveira:2020:NSN

[220] Renildo R. de Oliveira, Frank R. V. Ribeiro, André L. C. Canto, and Cláudio H. Zawadzki. A new species of the Neotropical loricariid Hypostomus cochliodon group (Hypostominae) from the lower Rio Tapajós basin, Brazilian Amazon. Journal of Fish Biology, 97(2):490–498, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sandoval:2020:SIA

[221] Luis Alejandro Sandoval, Jenny Leal-Flórez, Juan Felipe Blanco-Libreros, José Ernesto Mancera-Pineda, Antonio Delgado-Huertas, and Carlos Julio Polo-Silva. Stable-isotope analysis reveals sources of organic matter and ontogenic feeding shifts of a mangrove-dependent predator species, New Granada sea catfish, Ariopsis canteri. Journal of Fish Biology, 97(2):499–507, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tang:2020:CTD

[222] Ni Tang, Yanlin Liu, Zhengzhi Tian, Shaoqi Xu, Mei Wang, Hu Chen, Bin Wang, Ya Li, Yan Wang, Song Yang, Liulan Zhao, Defang Chen, and Zhiqiong Li. Characterization, tissue distribution of resistin gene and the effect of fasting and refeeding on resistin mRNA expression in Siberian sturgeon (Acipenser baerii). Journal of Fish Biology, 97(2):508– 514, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown-Vuillemin:2020:TEO

[223] Sarah Brown-Vuillemin, Thomas Barreau, Jean-Marie Caraguel, and Samuel P. Iglésias. Trophic ecology and ontogenetic diet shift of the blue skate (Dipturus cf. flossada). Journal of Fish Biology, 97(2):515– 526, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cohen:2020:DEE

[224] Karly E. Cohen, Amy E. George, Duane C. Chapman, John H. Chick, and L. Patricia Hernandez. Developmental ecomorphology of the epibranchial organ of the silver carp, Hypophthalmichthys molitrix. Journal of Fish Biology, 97(2):527–536, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wasiljew:2020:BRD

[225] Benjamin D. Wasiljew, Jobst Pfaender, Benjamin Wipfler, Ilham Vemandra Utama, and Fabian Herder. Book review: Do we need the third dimension? Quantifying the effect of the z-axis in 3D geometric morphometrics based on sailfin silversides (Telmatherinidae). Journal of Fish Biology, 97(2):537–545, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rangel:2020:EBT

[226] Bianca de Sousa Rangel, Douglas de Castro Ribeiro, Jumma M. A. Chagas, Lucas Spada, Renata G. Moreira, and Cristiéle da Silva Ribeiro. Effects of biological traits on capture-induced parturition in a freshwater stingray and perspectives for species management. Journal of Fish Biology, 97(2):546–551, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heimbrand:2020:STT

[227] Yvette Heimbrand, Karin E. Limburg, Karin Hüssy, Michele Casini, Rajlie Sjöberg, Anne-Marie Palmén Bratt, Svend-Erik Levinsky, Anastasia Karpushevskaia, Krzysztof Radtke, and Jill Öhlund. Seeking the true time: Exploring otolith chemistry as an age-determination tool. Journal of Fish Biology, 97(2):552–565, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Takegaki:2020:EAP

[228] Takeshi Takegaki, Yosuke Nakatake, and Noriko Amiya. Effect of the administration of prolactin-releasing peptide2 on feeding activity in the intertidal blenny Rhabdoblennius nitidus (Günther, 1861). Journal of Fish Biology, 97(2):566–571, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dressler:2020:LWH

[229] Terra L. Dressler, Kevin D. Lafferty, Christopher L. Jerde, and Tom Dudley. Looking where it’s hard to see: a case study documenting rare Eucyclogobius newberryi presence in a California lagoon. Journal of Fish Biology, 97(2):572–576, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Joyce:2020:STC

[230] Brendan J. Joyce and Grant E. Brown. Short-term captivity drives hypothalamic plasticity and asymmetry in wild-caught northern red bellied dace (Chrosomus eos). Journal of Fish Biology, 97(2):577–582, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morrison:2020:SLT

[231] Scott M. Morrison, Theresa E. Mackey, Travis Durhack, Jennifer D. Jeffrey, Lilian M. Wiens, Neil J. Mochnacz, Caleb T. Hasler, Eva C. Enders, Jason R. Treberg, and Ken M. Jeffries. Sub-lethal temperature thresholds indicate acclimation and physiological limits in brook trout Salvelinus fontinalis. Journal of Fish Biology, 97(2):583–587, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Papastamatiou:2020:EBP

[232] Yannis P. Papastamatiou, Deron Verbeck, Melanie Hutchinson, Heather D. Bracken-Grissom, and Demian Chapman. An encounter between a pelagic shark and giant cephalopod. Journal of Fish Biology, 97(2):588–589, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Orban:2020:BRS

[233] László Orbán. Book review: Sex Control in Aquaculture, Wang, H., Piferrer, F., Chen, S. and Shen, Z. G. (Eds.), 888 pp. Published by John Wiley & Sons, Ltd., Oxford, UK, 2019. Price £250.00. ISBN: 978-1-119-12726-0. Journal of Fish Biology, 97(2):590–591, August 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2020:IIi

[234] Anonymous. Issue information. Journal of Fish Biology, 97(3):593–594, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2020:TCI

[235] Anonymous. Terrestrial chemoreception in an invasive amphibious catfish. Journal of Fish Biology, 97(3):595, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Geffroy:2020:EGW

[236] Benjamin Geffroy and Claus Wedekind. Effects of global warming on sex ratios in fishes. Journal of Fish Biology, 97(3):596–606, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alix:2020:GSH

[237] Maud Alix, Olav S. Kjesbu, and Kelli C. Anderson. From gametogenesis to spawning: How climate-driven warming affects teleost reproductive biology. Journal of Fish Biology, 97(3):607–632, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Seraphim:2020:IBC

[238] Marie J. Seraphim, Katherine A. Sloman, Mhairi E. Alexander, Noel Janetski, Jamaluddin Jompa, Rohani Ambo-Rappe, Donna Snellgrove, Frank Mars, and Alastair R. Harborne. Interactions between coral restoration and fish assemblages: implications for reef management. Journal of Fish Biology, 97(3):633–655, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tschopp:2020:TNP

[239] Ayelen Tschopp, Franco Cristiani, Nestor Aníbal García, Enrique Alberto Crespo, and Mariano Alberto Coscarella. Trophic niche partitioning of five skate species of genus Bathyraja in northern and central Patagonia, Argentina. Journal of Fish Biology, 97(3):656–667, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mateussi:2020:MDT

[240] Nadayca T. Bonani Mateussi, Bruno F. Melo, and Claudio Oliveira. Molecular delimitation and taxonomic revision of the wimple piranha Catoprion (Characiformes: Serrasalmidae) with the description of a new species. Journal of Fish Biology, 97(3):668–685, September 2020. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deCarvalho:2020:HFD

[241] Barbara M. de Carvalho, Augusto L. Ferreira Junior, Luís F. Fávaro, Roberto F. Artoni, and Jean Vitule. Human-facilitated dispersal of the Gulf toadfish Opsanus beta (Goode & Bean, 1880) in the Guaratuba Bay, south-eastern Brazil. Journal of Fish Biology, 97(3):686–690, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crawford:2020:TLS

[242] Lydia Crawford, Henry Bart, Jr., Enrique L. Harold, Thomas M. Huckaba, and Ian R. Davenport. Tube-like structures within the ovarian follicle of Hydrolagus colliei indicate the early origins of follicle cell processes. Journal of Fish Biology, 97(3):691–695, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Biesack:2020:EPG

[243] Ellen E. Biesack, Binh T. Dang, Amanda S. Ackiss, Christopher E. Bird, Phen Chheng, Latsamy Phounvisouk, Oanh T. Truong, and Kent E. Carpenter. Evidence for population genetic structure in two exploited Mekong River fishes across a natural riverine barrier. Journal of Fish Biology, 97(3):696–707, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Temple:2020:LHE

[244] Andrew J. Temple, Selina M. Stead, Narriman Jiddawi, Nina Wambiji, Nicholas K. Dulvy, Ellen Barrowclift, and Per Berggren. Life-history, exploitation and extinction risk of the data-poor Baraka’s whipray (Maculabatis ambigua) in small-scale tropical fisheries. Journal of Fish Biology, 97(3):708–719, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

De:2020:BDC

[245] Debasis De, Panantharayil S. Shyne Anand, Suchita Mukherjee, Prem Kumar, Jagabathula Syama Dayal, Ramalingam Ananda Raja, Aritra Bera, Vettath Raghavan Suresh, and Koyadan Kizhakedath Vijayan. Broodstock development and captive maturation of hilsa (Tenualosa ilisha) in a brackishwater pond-based system. Journal of Fish Biology, 97 (3):720–733, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lopez-Unzu:2020:MHC

[246] Miguel A. López-Unzu, María Teresa Soto-Navarrete, Valentín Sans-Coma, Borja Fernández, and Ana Carmen Durán. Myosin heavy chain isoforms in the myocardium of the atrioventricular junction of Scyliorhinus canicula (Chondrichthyes, Carcharhiniformes). Journal of Fish Biology, 97(3):734–739, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arambam:2020:ILI

[247] Kalpana Arambam, Soibam Khogen Singh, Pradyut Biswas, Arun Bhai Patel, Alok Kumar Jena, and Pramod Kumar Pandey. Influence of light intensity and photoperiod on embryonic development, survival and growth of threatened catfish Ompok bimaculatus early larvae. Journal of Fish Biology, 97(3):740–752, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lallement:2020:LFM

[248] Mailén E. Lallement, Magalí Rechencq, María V. Fernández, Eduardo Zattara, Alejandro Sosnovsky, Pablo Vigliano, Gilda Garibotti, Marcelo F. Alonso, Gustavo Lippolt, and Patricio J. Macchi. Landscape factors modulating patterns of salmonid distribution during summer in north Patagonian rivers. Journal of Fish Biology, 97(3):753–762, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Catarino:2020:FRO

[249] Diana Catarino, Kirsten Jakobsen, Joachim Jakobsen, Eva Giacomello, Gui M. Menezes, Hugo Diogo, Ângela Canha, Filipe M. Porteiro, Octávio Melo, and Sergio Stefanni. First record of the opal chimaera, Chimaera opalescens (Holocephali: Chimaeridae) and revision of the occurrence of the rabbitfish Chimaera monstrosa in the Azores waters. Journal of Fish Biology, 97(3):763–775, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

doNascimento:2020:TMF

[250] Carolina P. do Nascimento, Natália C. L. dos Santos, Bianca M. Dal Vesco, and Luiz C. Gomes. Trophic morphology features allow Astyanax endemic species coexistence in a Neotropical river system. Journal of Fish Biology, 97(3):776–784, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Peng:2020:NSC

[251] Cheng Peng, Qing Wang, Herong Shi, Jiaxing Chen, Shuisheng Li, Huihong Zhao, Haoran Lin, Jianchun Yang, and Yong Zhang. Natural sex change in mature protogynous orange-spotted grouper (Epinephelus coioides): gonadal restructuring, sex hormone shifts and gene profiles. Journal of Fish Biology, 97(3):785–793, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Christensen:2020:CEB

[252] Emil A. F. Christensen, Morten B. S. Svendsen, and John F. Steffensen. The combined effect of body size and temperature on oxygen consumption rates and the size-dependency of preferred temperature in European perch Perca fluviatilis. Journal of Fish Biology, 97(3):794–803, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whitesel:2020:PLM

[253] Timothy A. Whitesel, Michelle McGree, and Gregory S. Silver. Predicting larval metamorphosis of Pacific lamprey Entosphenus tridentatus through measurements of length and mass. Journal of Fish Biology, 97 (3):804–816, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhang:2020:USC

[254] Qingtian Zhang and Guikun Hu. Utilization of species checklist data in revealing the spatial distribution of fish diversity. Journal of Fish Biology, 97(3):817–826, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2020:TLS

[255] Xu Li, Wei Dao, and Wei Zhou. Type locality and species identity of Pareuchiloglanis sinensis (Hora & Silas) with a description of a new species of the genus from the upper Yangtze River basin in southern China. Journal of Fish Biology, 97(3):827–844, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anderson:2020:SVD

[256] Antônio B. Anderson, Jean-Christophe Joyeux, and Sergio R. Floeter. Spatiotemporal variations in density and biomass of rocky reef fish in a biogeographic climatic transition zone: trends over 9 years, inside and outside the only nearshore no-take marine-protected area on the southern Brazilian coast. Journal of Fish Biology, 97(3):845–859, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Silva-Oliveira:2020:TIB

[257] Cárlison Silva-Oliveira, Cristiano R. Moreira, Flávio C. T. Lima, and Lúcia Rapp Py-Daniel. The true identity of Bryconops cyrtogaster (Norman), and description of a new species of Bryconops kner (Characiformes: Iguanodectidae) from the Rio Jari, lower Amazon basin. Journal of Fish Biology, 97(3):860–868, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Espinosa:2020:EDS

[258] Cristóbal Espinosa and Maria Ángeles Esteban. Effect of dietary supplementation with yeast Saccharomyces cerevisiae on skin, serum and liver of gilthead seabream (Sparus aurata l). Journal of Fish Biology, 97(3): 869–881, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mendoza-Portillo:2020:GDS

[259] Verónica Mendoza-Portillo, Carolina Galván-Tirado, David S. Portnoy, Fausto Valenzuela-Quiñonez, Omar Domínguez-Domínguez, Jean-Dominique Durand, Juan Carlos Pérez-Urbiola, and Francisco J. García-De León. Genetic diversity and structure of circumtropical almaco jack, Seriola rivoliana: tool for conservation and management. Journal of Fish Biology, 97(3):882–894, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bressman:2020:WDI

[260] Noah R. Bressman, Jeffrey E. Hill, and Miriam A. Ashley-Ross. Why did the invasive walking catfish cross the road? Terrestrial chemoreception described for the first time in a fish. Journal of Fish Biology, 97(3): 895–907, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Banon:2020:UDS

[261] Rafael Bañón, Alejandro de Carlos, Susana Ruiz-Pico, and Francisco Baldó. Unexpected deep-sea fish species on the Porcupine Bank (NE Atlantic): Biogeographical implications. Journal of Fish Biology, 97(3): 908–913, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Machado:2020:WAR

[262] Rodrigo Machado, Eduardo Gentil, and Jorge L. Rodrigues-Filho. When an advantageous reproductive trait turns bad: Eggs of the threatened fish Genidens barbus as a natural bait in recreational fisheries. Journal of Fish Biology, 97(3):914–918, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schooley:2020:FOI

[263] Jason D. Schooley, Adam Geik, and Dennis L. Scarnecchia. First observations of intersex development in paddlefish Polyodon spathula. Journal of Fish Biology, 97(3):919–925, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Buglass:2020:FRS

[264] Salome Buglass, Shannon Nagy, David Ebert, Paulina Sepa, Alan Turchik, Katherine L. C. Bell, Fernando Rivera, and Jonatha Giddens. First records of the seven-gilled Notorynchus cepedianus and six-gilled Hexanchus griseus sharks (Chondrichthyes: Hexanchiformes: Hexanchidae) found in the Galápagos Marine Reserve. Journal of Fish Biology, 97 (3):926–929, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Karaiskou:2020:GSD

[265] Nikoleta Karaiskou, Konstantinos Gkagkavouzis, Stella Minoudi, Dimitrios Botskaris, Kyriaki Markou, Stavros Kalafatakis, Efthimia Antonopoulou, and Alexandros Triantafyllidis. Genetic structure and divergence of tench Tinca tinca European populations. Journal of Fish Biology, 97(3):930–934, September 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2020:IIj

[266] Anonymous. Issue information. Journal of Fish Biology, 97(4):935–936, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:2020:UVI

[267] Culum Brown. Underwater video identifies individual sharks to aid population management. Journal of Fish Biology, 97(4):937, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Falahiehzadeh:2020:HEA

[268] Nadieh Falahiehzadeh and Negin Salamat. Histomorphological and endocrine assessment of female Arabian carpetshark, Chiloscyllium arabicum (Elasmobranchii: Hemiscylliidae) from the Persian Gulf during annual reproductive cycle. Journal of Fish Biology, 97(4):938–952, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Becker:2020:TVS

[269] James R. Becker, Matthew D. Cieri, David A. Libby, Adam St. Gelais, Graham Sherwood, and Yong Chen. Temporal variability in size and growth of Atlantic herring in the Gulf of Maine. Journal of Fish Biology, 97(4):953–963, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Griffiths:2020:FHD

[270] David Griffiths. Foraging habitat determines predator–prey size relationships in marine fishes. Journal of Fish Biology, 97(4):964–973, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Claassens:2020:LHP

[271] Louw Claassens and David Harasti. Life history and population dynamics of an endangered seahorse (Hippocampus capensis) within an artificial habitat. Journal of Fish Biology, 97(4):974–986, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lewis:2020:EPP

[272] Robert Lewis, Steve Dawson, and William Rayment. Estimating population parameters of broadnose sevengill sharks (Notorynchus cepedianus) using photo identification capture-recapture. Journal of Fish Biology, 97 (4):987–995, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kargenberg:2020:MPP

[273] Einar Kärgenberg, Finn Økland, Mart Thalfeldt, Eva B. Thorstad, Odd Terje Sandlund, and Meelis Tambets. Migration patterns of a potamodromous piscivore, asp (Leuciscus aspius), in a river–lake system. Journal of Fish Biology, 97(4):996–1008, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Helminen:2020:LMA

[274] Jani Helminen, Guillaume J. R. Dauphin, and Tommi Linnansaari. Length measurement accuracy of adaptive resolution imaging sonar and a predictive model to assess adult Atlantic salmon (Salmo salar) into two size categories with long-range data in a river. Journal of Fish Biology, 97(4):1009–1026, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Udagawa:2020:VDE

[275] Shingo Udagawa, Sung-Pyo Hur, Jun-Hwan Byun, Hiroki Takekata, Yuki Takeuchi, and Akihiro Takemura. Verification of differentially expressed genes in relation to hydrostatic pressure in the brain of two wrasse species with high-tide preference in spawning. Journal of Fish Biology, 97(4): 1027–1038, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garcia-Martinez:2020:CRD

[276] Rosa María García-Martínez, Irene de los Angeles Barriga-Sosa, Benjamín López, and Omar Mejía. “Colonization routes and demographic history of Chirostoma humboldtianum in the central Mexican plateau”. Journal of Fish Biology, 97(4):1039–1050, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carbia:2020:SDD

[277] Penelope S. Carbia, Culum Brown, Joo M. Park, Troy F. Gaston, Vincent Raoult, and Jane E. Williamson. Seasonal and developmental diet shifts in sympatric and allopatric intertidal gobies determined by stomach content and stable isotope analysis. Journal of Fish Biology, 97(4): 1051–1062, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cheal:2020:CCR

[278] Alistair J. Cheal and Michael J. Emslie. Counts of coral reef fishes by an experienced observer are not biased by the number of target species. Journal of Fish Biology, 97(4):1063–1071, October 2020. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tencatt:2020:JTA

[279] Luiz F. C. Tencatt, Janice Muriel-Cunha, Jansen Zuanon, Marlon F. C. Ferreira, and Marcelo R. Britto. A journey through the Amazon Middle Earth reveals Aspidoras azaghal (Siluriformes: Callichthyidae), a new species of armoured catfish from the Rio Xingu basin, Brazil. Journal of Fish Biology, 97(4):1072–1086, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cuevas-Gomez:2020:INA

[280] Gabriela Alejandra Cuevas-Gómez, Juan Carlos Pérez-Jiménez, Iván Méndez-Loeza, Maribel Carrera-Fernández, and José Leonardo Castillo-Géniz. Identification of a nursery area for the critically endangered hammerhead shark (Sphyrna lewini) amid intense fisheries in the southern Gulf of Mexico. Journal of Fish Biology, 97(4):1087–1096, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Canei:2020:CAI

[281] Jérôme Canei, Arnaud Trupia, and Denis Nonclercq. Cytological analysis of integumentary and muscular adaptations in three sand-dwelling marine teleosts, Ammodytes tobianus (Ammodytidae), Gorgasia preclara (Congridae) and Heteroconger hassi (Congridae) (Teleostei; Actinopterygii). Journal of Fish Biology, 97(4):1097–1112, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Preiszner:2020:SBS

[282] Bálint Preiszner, István Czeglédi, Gergely Boros, András Liker, Bernadett Kern, and Tibor Erős. Scavenging behaviour and sizedependent carcass consumption of the black bullhead (Ameiurus melas).

Journal of Fish Biology, 97(4):1113–1119, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Petean:2020:ITI

[283] Flávia F. Petean, Gavin J. P. Naylor, and Sergio M. Q. Lima. Integrative taxonomy identifies a new stingray species of the genus Hypanus Rafinesque, 1818 (Dasyatidae, Myliobatiformes), from the Tropical Southwestern Atlantic. Journal of Fish Biology, 97(4):1120–1142, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pinheiro:2020:FBS

[284] Hudson T. Pinheiro, Bruno C. L. Macena, Ronaldo B. Francini-Filho, Carlos E. L. Ferreira, Fernanda V. Albuquerque, Natalia P. A. Bezerra, Alfredo Carvalho-Filho, Romulo C. P. Ferreira, Osmar J. Luiz, Thayna J. Mello, Sibele A. Mendonça, Diogo M. Nunes, Caio R. Pimentel, Alessandra M. A. Pires, Abilio Soares-Gomes, Danielle L. Viana, Fabio H. V. Hazin, and Luiz A. Rocha. Fish biodiversity of Saint Peter and Saint Paul’s Archipelago, Mid-Atlantic Ridge, Brazil: new records and a species database. Journal of Fish Biology, 97(4):1143–1153, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Firidin:2020:PGS

[285] Sirin Firidin, Rafet Cagri Ozturk, Melike Alemdag, Oguzhan Eroglu, Yahya Terzi, Ilyas Kutlu, Zehra Duygu Duzgunes, Eyup Cakmak, and Ilhan Aydin. Population genetic structure of turbot (Scophthalmus maximus L., 1758) in the Black Sea. Journal of Fish Biology, 97(4):1154– 1164, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Payet:2020:CDC

[286] Samuel D. Payet, Jake R. Lowe, Bruce D. Mapstone, Morgan S. Pratchett, Tane H. Sinclair-Taylor, Brett M. Taylor, Peter A. Waldie, and Hugo B. Harrison. Comparative demography of commercially important species of coral grouper, Plectropomus leopardus and P. laevis, from Australia’s Great Barrier Reef and Coral Sea marine parks. Journal of Fish Biology, 97(4):1165–1176, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Petatan-Ramirez:2020:HSR

[287] David Petatán-Ramírez, Darren A. Whitehead, Triana Guerrero-Izquierdo, Miguel A. Ojeda-Ruiz, and Edgar E. Becerril-García. Habitat suitability of Rhincodon typus in three localities of the Gulf of California: Environmental drivers of seasonal aggregations. Journal of Fish Biology, 97(4):1177–1186, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Avigliano:2020:WMM

[288] Esteban Avigliano, Ana Ibañez, Nidia Fabré, Roberta Callicó Fortunato, Ana Méndez, Jorge Pisonero, and Alejandra V. Volpedo. White mullet Mugil curema population structure from Mexico and Brazil revealed by otolith chemistry. Journal of Fish Biology, 97(4):1187–1200, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neisse:2020:SBF

[289] Niklas Neiße, Matteo Santon, Pierre-Paul Bitton, and Nico K. Michiels. Small benthic fish strike at prey over distances that fall within theoretical predictions for active sensing using light. Journal of Fish Biology, 97(4): 1201–1208, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winter:2020:PFI

[290] Emily R. Winter, Andrew M. Hindes, Steve Lane, and J. Robert Britton. Predicting the factors influencing the inter- and intraspecific survival rates of riverine fishes implanted with acoustic transmitters. Journal of Fish Biology, 97(4):1209–1219, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oliveira:2020:DCE

[291] Marcos S. B. Oliveira, Raimundo N. G. Mendes-Júnior, and Marcos Tavares-Dias. Diet composition of the electric eel Electrophorus voltai (Pisces: Gymnotidae) in the Brazilian Amazon region. Journal of Fish Biology, 97(4):1220–1223, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Buchinger:2020:BRF

[292] Tyler J. Buchinger, Weiming Li, and Nicholas S. Johnson. Behavioural responses of female lake trout Salvelinus namaycush to male chemical stimuli and prostaglandin F_{2α}. Journal of Fish Biology, 97(4):1224–1227, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Becerril-Garcia:2020:SRC

[293] Edgar E. Becerril-García, Edgar M. Hoyos-Padilla, Brad Henning, and Pelayo Salinas-De León. Sharks, rays, and chimaeras of the Revillagigedo National Park: an update of new and confirmed records. Journal of Fish Biology, 97(4):1228–1232, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pachla:2020:FRM

[294] Lucas A. Pachla, Paula B. Hartmann, Marthoni V. Massaro, Rodrigo Bastian, Fernando M. Pelicice, and David A. Reynalte-Tataje. First record of the mating behaviour of the spotted surubim Pseudoplatystoma corruscans in the Uruguay River. Journal of Fish Biology, 97(4):1233– 1237, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ahn:2020:VJE

[295] Hyojin Ahn. Variation of Japanese eel eDNA in sequentially changing conditions and in different sample volumes. Journal of Fish Biology, 97 (4):1238–1241, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wood:2020:DRP

[296] Chris M. Wood, Eric F. Pane, and George J. F. Heigenhauser. Dichloroacetate reveals the presence of metabolic inertia at the start of exercise in rainbow trout (Oncorhynchus mykiss, Walbaum 1792). Journal of Fish Biology, 97(4):1242–1246, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Araujo:2020:WSR

[297] Gonzalo Araujo, Joni A. Miranda, Harriet L. Allen, Jessica Labaja, Sally Snow, Alessandro Ponzo, and Christine G. Legaspi. Whale sharks Rhincodon typus get cleaned by the blue-streak cleaner wrasse Labroides dimidiatus and the moon wrasse Thalassoma lunare in the Philippines. Journal of Fish Biology, 97(4):1247–1251, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mulas:2020:LNF

[298] Antonello Mulas, Andrea Bellodi, Cristina Porcu, Alessandro Cau, Elisabetta Coluccia, Riccardo Demurtas, Martina Francesca Marongiu, Paola Pesci, and Maria Cristina Follesa. Living naked: first case of lack of skinrelated structures in an elasmobranch, the blackmouth catshark (Galeus melastomus). Journal of Fish Biology, 97(4):1252–1256, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bradley:2020:AGS

[299] Kirsty A. Bradley and Alexander Arkhipkin. Age and growth of slender tuna (Allothunnus fallai) in an unexploited temperate population. Journal of Fish Biology, 97(4):1257–1261, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whitfield:2020:FSE

[300] Alan K. Whitfield. Fish species in estuaries — from partial association to complete dependency. Journal of Fish Biology, 97(4):1262–1264, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:2020:RLP

[301] Richard J. Kennedy, Warren Campbell, Kevin Gallagher, and Derek Evans. River lamprey present an unusual predation threat to Atlantic salmon smolts in Lough Neagh, Northern Ireland. Journal of Fish Biology, 97(4):1265–1267, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lamarca:2020:FRP

[302] Felipe Lamarca, Marcelo Vianna, and Anderson Vilasboa. The first reproductive parameters and evidence of multiple paternity in one new spiny dogfish species, Squalus albicaudus (Squaliformes, Squalidae). Journal of Fish Biology, 97(4):1268–1272, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lardizabal:2020:IRM

[303] Claudia C. Lardizábal, Dilenia E. Martínez, Lesbia E. Zamora, Caleb D. McMahan, and Wilfredo A. Matamoros. An insular record of the Mayan cichlid Mayaheros urophthalmus (Günther, 1862) (Cichliformes: Cichlidae) on an island 70 km offshore in the Honduran Caribbean. Journal of Fish Biology, 97(4):1273–1275, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Trnski:2020:SNC

[304] Lena Trnski, Armagan Sabetian, and Julian Lilkendey. Scaring Nemo: Contrasting effects of observer presence on two anemonefish species. Journal of Fish Biology, 97(4):1276–1280, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:2020:LTR

[305] Richard J. Kennedy, Derek Evans, and Michelle Allen. Long-term retention of dummy acoustic transmitters in adult brown trout. Journal of Fish Biology, 97(4):1281–1284, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

OToole:2020:ATA

[306] Christina O’Toole, Emily Weigum, Conor T. Graham, Philip White, Kurt Samways, Brian Hayden, and Deirdre Brophy. Acid treatment of Atlantic salmon (Salmo salar) scales prior to analysis has negligible effects on δ ^{13}C and δ ^{15}N isotope ratios. Journal of Fish Biology, 97 (4):1285–1290, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2020:Cb

[307] Anonymous. Corrigendum. Journal of Fish Biology, 97(4):1291, October 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2020:IIk

[308] Anonymous. Issue information. Journal of Fish Biology, 97(5):1293– 1294, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kaiser:2020:PNB

[309] Michel J. Kaiser. Pufferfish nest building reveals a skilled architect. Journal of Fish Biology, 97(5):1295, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blanco:2020:FEM

[310] Edurne Blanco, Patricia Reglero, Aurelio Ortega, Arild Folkvord, Fernando de la Gándara, Alma Hernández de Rojas, and Marta Moyano. First estimates of metabolic rate in Atlantic bluefin tuna larvae. Journal of Fish Biology, 97(5):1296–1305, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sanches:2020:AAL

[311] Paulo Vanderlei Sanches, Tatiane Mary Gogola, Rafael Olimpio Silva, Dhyego Amilton Topan, Pablo Henrique dos Santos Picapedra, and Pitágoras Augusto Piana. Arms as areas for larval development of migratory fish species in a Neotropical reservoir and the influence of rainfall over abundances. Journal of Fish Biology, 97(5):1306–1316, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Young:2020:MFC

[312] Joy M. Young, Beau G. Yeiser, James A. Whittington, and Jynessa Dutka-Gianelli. Maturation of female common snook Centropomus undecimalis: implications for managing protandrous fishes. Journal of Fish Biology, 97(5):1317–1331, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Steendam:2020:BBE

[313] Charlotte Steendam, Pieterjan Verhelst, Sam Van Wassenbergh, and Jens De Meyer. Burrowing behaviour of the European eel (Anguilla anguilla): Effects of life stage. Journal of Fish Biology, 97(5):1332–1342, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zanata:2020:NRS

[314] Angela M. Zanata, Willian M. Ohara, Osvaldo T. Oyakawa, and Fernando C. P. Dagosta. A new rheophilic South American darter (Crenuchidae: Characidium) from the Rio Juruena basin, Brazil, with comments on morphological adaptations to life in fast-flowing waters. Journal of Fish Biology, 97(5):1343–1353, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Murphy:2020:EEM

[315] Anna E. Murphy, Michael J. W. Stokesbury, and Russell H. Easy. Exploring epidermal mucus protease activity as an indicator of stress in Atlantic sturgeon (Acipenser oxyrinchus oxyrhinchus). Journal of Fish Biology, 97(5):1354–1362, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deOliveira:2020:EGM

[316] Rianne Caroline de Oliveira and Weferson Júnio da Graça. Encephalon gross morphology of the cichlid Geophagus sveni (Cichlidae: Geophagini): Comparative description and ecological perspectives. Journal of Fish Biology, 97(5):1363–1374, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Griffiths:2020:EDM

[317] Nathan P. Griffiths, Jonathan D. Bolland, Rosalind M. Wright, Leona A. Murphy, Robert K. Donnelly, Hayley V. Watson, and Bernd Hänfling. Environmental DNA metabarcoding provides enhanced detection of the European eel Anguilla anguilla and fish community structure in pumped river catchments. Journal of Fish Biology, 97(5):1375–1384, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Balzani:2020:ONB

[318] Paride Balzani, Rodolphe E. Gozlan, and Phillip J. Haubrock. Overlapping niches between two co-occurring invasive fish: the topmouth gudgeon Pseudorasbora parva and the common bleak Alburnus alburnus. Journal of Fish Biology, 97(5):1385–1392, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Soares:2020:TAE

[319] Bruno E. Soares, Thiago F. Barros, Danielly T. Hashiguti, Douglas C. Pereira, Karina C. F. Ferreira, and Érica P. Caramaschi. Traditional approaches to estimate length at first maturity (L_{50}) retrieve better results than alternative ones in a Neotropical heptapterid. Journal of Fish Biology, 97(5):1393–1400, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bond:2020:MPC

[320] Todd Bond, Robert J. Mueller, Matthew J. Birt, Jane Prince, Karen Miller, Julian C. Partridge, and Dianne L. McLean. Mystery pufferfish create elaborate circular nests at mesophotic depths in Australia. Journal of Fish Biology, 97(5):1401–1407, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dittman:2020:AAC

[321] Andrew H. Dittman and Thomas P. Quinn. Amino acid cues emanating from Pacific salmon eggs and ovarian fluid. Journal of Fish Biology, 97(5):1408–1414, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rohner:2020:MHU

[322] Christoph A. Rohner, Roy Bealey, Bernerd M. Fulanda, and Simon J. Pierce. Movement and habitat use of striped marlin Kajikia audax in the Western Indian Ocean. Journal of Fish Biology, 97(5):1415–1427, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lim:2020:EIT

[323] Dan Dohyung Lim, C. Louise Milligan, and Yolanda E. Morbey. Elevated incubation temperature improves later-life swimming endurance in juvenile Chinook salmon, Oncorhynchus tshawytscha. Journal of Fish Biology, 97(5):1428–1439, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ceballos-Francisco:2020:RCG

[324] Diana Ceballos-Francisco, Nuria G. Carrillo, Francisco J. Pardo-Fernández, Alberto Cuesta, and María Á. Esteban. Radiological characterization of gilthead seabream (Sparus aurata) by X-ray computed tomography. Journal of Fish Biology, 97(5):1440–1447, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thore:2020:CDE

[325] Eli S. J. Thoré, Luc Brendonck, and Tom Pinceel. Conspecific density and environmental complexity impact behaviour of turquoise killifish (Nothobranchius furzeri). Journal of Fish Biology, 97(5):1448–1461, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huang:2020:RGS

[326] Wen-Chien Huang, Anil Mohapatra, Pham The Thu, Hong-Ming Chen, and Te-Yu Liao. A review of the genus Strophidon (Anguilliformes: Muraenidae), with description of a new species. Journal of Fish Biology, 97(5):1462–1480, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henschel:2020:ANN

[327] Elisabeth Henschel, Nathan K. Lujan, and Jonathan N. Baskin. Ammoglanis natgeorum, a new miniature pencil catfish (Siluriformes: Trichomycteridae) from the lower Atabapo River, Amazonas, Venezuela. Journal of Fish Biology, 97(5):1481–1490, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dymek:2020:SIA

[328] Anna M. Dymek and Anna Pecio. Spermatogenesis in the inseminating African butterflyfish Pantodon buchholzi (Teleostei: Osteoglossiformes: Pantodontidae) with the revision of residual bodies formation. Journal of Fish Biology, 97(5):1491–1506, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gronkjaer:2020:IAV

[329] Peter Grønkjær, Rasmus Ottosen, Thor Joensen, Lee Reeve, Einar E. Nielsen, and Rasmus Hedeholm. Intra-annual variation in feeding of Atlantic cod Gadus morhua: the importance of ephemeral prey bursts. Journal of Fish Biology, 97(5):1507–1519, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wiecaszek:2020:FSE

[330] Beata Wiecaszek, Adam Nowosielski, Jarosław Dabrowski, Klaudia Górecka, Sławomir Keszka, and Agnieszka Strzelczak. Fish size effect on sagittal otolith outer shape variability in round goby Neogobius melanostomus (Pallas 1814). Journal of Fish Biology, 97(5):1520–1541, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bai:2020:STT

[331] Jie Bai, Xiucai Hu, Aijun Lü, Ruixia Wang, Rongrong Liu, Jingfeng Sun, and Yuchen Niu. Skin transcriptome, tissue distribution of mucin genes and discovery of simple sequence repeats in crucian carp (Carassius auratus). Journal of Fish Biology, 97(5):1542–1553, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vranken:2020:SHP

[332] Nathan Vranken, Maarten Van Steenberge, Akonkwa Balagizi, and Jos Snoeks. The synonymy of Haplochromis pharyngalis and Haplochromis petronius (Cichlidae). Journal of Fish Biology, 97(5):1554–1559, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ohkuma:2020:ASN

[333] Kazumasa Ohkuma, Kunio Abe, Manabu Kagaya, Kazuyuki Yamaya, Motokazu Okukawa, Masa aki Watanabe, and Takeo Yamamoto. Ascending speed and nocturnal activity of adult masu salmon (Oncorhynchus masou Brevoort) during upstream migration. Journal of Fish Biology, 97(5):1560–1563, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pryor:2020:FVV

[334] Kimberley J. Pryor. Frequent visits to various anvils during tool use by the graphic tuskfish Choerodon graphicus (Labridae). Journal of Fish Biology, 97(5):1564–1568, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tietbohl:2020:IPB

[335] Matthew D. Tietbohl, Royale S. Hardenstine, Lyndsey K. Tanabe, Ann Marie Hulver, and Michael L. Berumen. Intentional partial beaching in a coral reef fish: a newly recorded hunting behaviour of titan triggerfish, Balistoides viridescens. Journal of Fish Biology, 97(5):1569–1572, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Frable:2020:NMS

[336] Benjamin W. Frable, Merit H. Mccrea, Milton S. Love, and Matthew T. Craig. A new maximum size for the family Nomeidae and the third record of a longfin cigarfish, Cubiceps paradoxus (Stromateoidei: Nomeidae) from California, USA. Journal of Fish Biology, 97(5):1573–1575, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sattler:2020:ABS

[337] Jessica L. Sattler and Janette W. Boughman. Advancing breeding in stickleback (Gasterosteus aculeatus) to produce two reproductive cycles per year. Journal of Fish Biology, 97(5):1576–1581, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rohtla:2020:QOA

[338] Mehis Rohtla, Lagle Matetski, Imre Taal, Roland Svirgsden, Martin Kesler, Päärn Paiste, and Markus Vetemaa. Quantifying an overlooked aspect of partial migration using otolith microchemistry. Journal of Fish Biology, 97(5):1582–1585, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Flowers:2020:ABC

[339] Kathryn I. Flowers, Demian D. Chapman, Todd Kemp, Dave Wert, and Kevin A. Feldheim. Annual breeding in a captive smalltooth sawfish, Pristis pectinata. Journal of Fish Biology, 97(5):1586–1589, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lacoste:2020:FRS

[340] Francisco Firpo Lacoste, Julieta Andreoli Bize, and Luis Fernández. First record sheds light on the distribution of the cyprinodontiform genus Jenynsia (Günther, 1866) in the High Andean Plateau. Journal of Fish Biology, 97(5):1590–1594, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2020:WDI

[341] Anonymous. Why did the invasive walking catfish cross the road? Terrestrial chemoreception described for the first time in a fish. Journal of Fish Biology, 97(5):1595, November 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2020:IIl

[342] Anonymous. Issue information. Journal of Fish Biology, 97(6):1597– 1598, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2020:UVE

[343] Anonymous. Understanding variation in energy flows in host–parasite relationships. Journal of Fish Biology, 97(6):1599, December 2020. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mofu:2020:LHR

[344] Lubabalo Mofu, Darragh J. Woodford, Ryan J. Wasserman, and Olaf L. F. Weyl. Life history of the river goby Glossogobius callidus (Teleostei: Gobiidae). Journal of Fish Biology, 97(6):1600–1606, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reis:2020:ITR

[345] Vinícius J. C. Reis, Sergio A. dos Santos, Marcelo R. Britto, Thaís de Assis Volpi, and Mário C. C. de Pinna. Iterative taxonomy reveals a new species of Trichomycterus Valenciennes 1832 (Siluriformes, Trichomycteridae) widespread in Rio Doce basin: a pseudocryptic of T. immaculatus. Journal of Fish Biology, 97(6):1607–1623, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tang:2020:AAF

[346] Shulin Tang, Xu-Fang Liang, Shan He, Muhammad Shoaib Alam, Haocan Luo, and Yulan Kuang. Activin a affects feeding by promoting the inner diameter and muscle development of the pharynx and oesophagus in zebrafish (Danio rerio) larvae. Journal of Fish Biology, 97(6): 1624–1631, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Svozil:2020:MPS

[347] Daniel Phillip Svozil, Lee J. Baumgartner, Christopher J. Fulton, Richard Keller Kopf, and Robyn J. Watts. Morphological predictors of swimming speed performance in river and reservoir populations of Australian smelt Retropinna semoni. Journal of Fish Biology, 97(6): 1632–1643, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Andersson:2020:CAR

[348] Matilda L. Andersson, Fredrik Sundberg, and Peter Eklöv. Chasing away accurate results: exhaustive chase protocols underestimate maximum metabolic rate estimates in European perch Perca fluviatilis. Journal of Fish Biology, 97(6):1644–1650, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wu:2020:ARB

[349] Feng Wu, Richard Kindong, Xiaojie Dai, Ousmane Sarr, Jiangfeng Zhu, Siquan Tian, Yunkai Li, and Bruno T. N. Nsangue. Aspects of the reproductive biology of two pelagic sharks in the eastern Atlantic Ocean. Journal of Fish Biology, 97(6):1651–1661, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Agudelo-Zamora:2020:UDG

[350] Henry D. Agudelo-Zamora, José Tavera, Yiskar D. Murillo, and Armando Ortega-Lara. The unknown diversity of the genus Characidium (Characiformes: Crenuchidae) in the Chocó biogeographic region, Colombian Andes: Two new species supported by morphological and molecular data. Journal of Fish Biology, 97(6):1662–1675, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Raick:2020:YEP

[351] Xavier Raick, Alessia Huby, Gregório Kurchevski, Alexandre L. Godinho, and Éric Parmentier. Yellow-eyed piranhas produce louder sounds than red-eyed piranhas in an invasive population of Serrasalmus marginatus. Journal of Fish Biology, 97(6):1676–1680, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zanata:2020:NSC

[352] Angela M. Zanata and Willian M. Ohara. A new species of Characidium (Characiformes: Crenuchidae) from the Rio Madeira basin, Brazil. Journal of Fish Biology, 97(6):1681–1689, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mrnak:2020:EWV

[353] Joseph T. Mrnak, Laura B. Heironimus, Daniel A. James, and Steven R. Chipps. Effect of water velocity and temperature on energy use, behaviour and mortality of pallid sturgeon Scaphirhynchus albus larvae. Journal of Fish Biology, 97(6):1690–1700, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chuctaya:2020:NSO

[354] Junior Chuctaya, Willian M. Ohara, and Luiz R. Malabarba. A new species of Odontostilbe cope (Characiformes: Cheirodontinae) from Rio Madeira basin diagnosed based on morphological and molecular data. Journal of Fish Biology, 97(6):1701–1712, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Braganca:2020:ITR

[355] Pedro H. N. Braganca, Timothy G. Smith, Emmanuel J. W. M. N. Vreven, and Albert Chakona. Integrative taxonomy reveals hidden diversity in the southern African darters genus Nannocharax Günther 1867 (Characiformes: Distichodontidae). Journal of Fish Biology, 97(6):1713– 1723, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

ONeill:2020:DEC

[356] Helen L. O’Neill, Chris Avila, and William T. White. Description of the egg cases and juvenile colouration in two catsharks of the genus Atelomycterus (Carcharhiniformes: Scyliorhinidae). Journal of Fish Biology, 97(6):1724–1732, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

GonzalezDeAcevedo:2020:RCF

[357] Melissa Gonzalez De Acevedo, Bryan S. Frazier, Carolyn Belcher, and James Gelsleichter. Reproductive cycle and fecundity of the bonnethead Sphyrna tiburo L. from the northwest Atlantic Ocean. Journal of Fish Biology, 97(6):1733–1747, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Delapieve:2020:SDR

[358] Maria Laura S. Delapieve, Tiago P. Carvalho, and Roberto E. Reis. Species delimitation in a range-restricted group of cascudinhos (Loricariidae: Epactionotus) supports morphological and genetic differentiation across coastal rivers of southern Brazil. Journal of Fish Biology, 97(6): 1748–1769, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bani:2020:RMP

[359] Ali Bani, Razieh Abdollahi, Narjes Karimi, Jeremy M. Lyle, and Jay Thompson. Retracing migration pattern in reproductive and nonreproductive female kutum Rutilus frisii, in south Caspian Sea, using otolith microchemistry. Journal of Fish Biology, 97(6):1770–1779, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:2020:RER

[360] Amanda N. Brown, Bryan S. Frazier, and Jim Gelsleichter. Re-evaluation of reproductive cycle and fecundity of finetooth sharks Carcharhinus isodon (Valenciennes 1839) from the Northwest Atlantic Ocean, with new observations on ovarian cycle and reproductive endocrinology of biennially reproducing sharks. Journal of Fish Biology, 97(6):1780–1793, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Abe:2020:GEN

[361] Takashi Abe, Yosuke Koshino, Tomoharu Watanabe, Yasuyuki Miyakoshi, Yuji Yoshida, and Hideaki Kudo. Gene expression of neuronal soluble N-ethylmaleimide-sensitive factor attachment protein receptor complex components in the olfactory organ and brain during seaward and homeward migration by pink salmon (Oncorhynchus gorbuscha). Journal of Fish Biology, 97(6):1794–1807, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Galeano-Chavarria:2020:EDM

[362] Ana María Galeano-Chavarria, Mauricio F. Landaeta, Guido Plaza, Manuel I. Castillo, and Darly S. Alarcón. Environmental determinants in morphospace and diet of the larval blenny Calliclinus geniguttatus from an upwelling ecosystem. Journal of Fish Biology, 97(6):1808–1820, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Taccardi:2020:SIR

[363] Emma Y. Taccardi, Ian R. Bricknell, and Carrie J. Byron. Stable isotopes reveal contrasting trophic dynamics between host–parasite relationships: a case study of Atlantic salmon (Salmo salar) and parasitic lice (Lepeophtheirus salmonis and Argulus foliaceus). Journal of Fish Biology, 97(6): 1821–1832, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turcotte:2020:ASS

[364] Luc A. Turcotte and J. Mark Shrimpton. Assessment of spawning site fidelity in interior Fraser River Coho salmon Oncorhynchus kisutch using otolith microchemistry, in British Columbia, Canada. Journal of Fish Biology, 97(6):1833–1841, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hagihara:2020:MEP

[365] Seishi Hagihara, Jun Aoyama, Daniel Limbong, and Katsumi Tsukamoto. Morphological, ecological and physiological characteristics of downstreammigrating and non-migrating Pacific bicolor eels Anguilla bicolor pacifica. Journal of Fish Biology, 97(6):1842–1845, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McDavitt:2020:SMP

[366] Matthew T. McDavitt and Peter M. Kyne. Social media posts reveal the geographic range of the Critically Endangered clown wedgefish, Rhynchobatus cooki. Journal of Fish Biology, 97(6):1846–1851, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Diaz-Carballido:2020:PPF

[367] Pedro Luis Díaz-Carballido, Adair Adolfo Gutiérrez-Coria, Perla Elizabeth Carrasco-Bautista, Eduardo Juventino Ramírez-Chávez, and Ana María Torres-Huerta. Presence of pregnant females of the Gorgona guitarfish, Pseudobatos prahli, in the Mexican tropical Pacific. Journal of Fish Biology, 97(6):1852–1856, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burns:2020:EAH

[368] Echelle S. Burns, Alyssa J. Clevenstine, Ryan K. Logan, and Christopher G. Lowe. Evidence of artificial habitat use by a recovering marine predator in southern California. Journal of Fish Biology, 97(6): 1857–1860, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Becerril-Garcia:2020:ENW

[369] Edgar E. Becerril-García, Edgar M. Hoyos-Padilla, Omar Santana-Morales, Miguel A. Gutiérrez-Ortiz, Arturo Ayala-Bocos, and Felipe Galván-Magaña. An estimate of the number of white sharks Carcharodon carcharias interacting with ecotourism in Guadalupe Island. Journal of Fish Biology, 97(6):1861–1864, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oladipo:2020:MCM

[370] Segun O. Oladipo, Lotanna M. Nneji, Abass T. Anifowoshe, and Ifeanyi C. Nneji. Molecular characterization of Malapterurus minjiriya Sagua, 1987 and phylogenetic relationships within the genus Malapterurus (Silurifomes, Malapteruridae) from Nigerian inland water bodies. Journal of Fish Biology, 97(6):1865–1869, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Breve:2020:SDS

[371] Niels W. P. Brevé, Hendrik V. Winter, Paul A. D. M. Wijmans, Eleanor S. I. Greenway, and Leopold A. J. Nagelkerke. Sex differentiation in seasonal distribution of the starry smooth-hound Mustelus asterias. Journal of Fish Biology, 97(6):1870–1875, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bashir:2020:EPS

[372] Zalina Bashir, Maizah M. Abdullah, Mazlan Abd. Ghaffar, and Mohd Uzair Rusli. Exclusive predation of sea turtle hatchlings by juvenile blacktip reef sharks Carcharhinus melanopterus at a turtle nesting site in Malaysia. Journal of Fish Biology, 97(6):1876–1879, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fish:2020:CDO

[373] Teresa Fish, Nathan Wolf, Bradley P. Harris, and Josep V. Planas. A comprehensive description of oocyte developmental stages in Pacific halibut, Hippoglossus stenolepis. Journal of Fish Biology, 97(6):1880–1885, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2020:RL

[374] Anonymous. Referees list. Journal of Fish Biology, 97(6):1886–1891, December 2020. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2021:IIa

[375] Anonymous. Issue information. Journal of Fish Biology, 98(1):1–2, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kaiser:2021:SIB

[376] Michel J. Kaiser. Suppression of inappropriate behaviours is a demanding task for some guppies. Journal of Fish Biology, 98(1):3, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cowx:2021:POL

[377] Ian G. Cowx, Nicholas Mandrak, Paul Skelton, and Denis Tweddle. Professor Olaf Lawrence Fredrick Weyl (1972–2020). Journal of Fish Biology, 98(1):4–5, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Todd:2021:VPS

[378] Christopher D. Todd, Nora N. Hanson, Lars Boehme, Crawford W. Revie, and Ana R. Marques. Variation in the post-smolt growth pattern of wild one sea-winter salmon (Salmo salar L.), and its linkage to surface warming in the eastern North Atlantic Ocean. Journal of Fish Biology, 98(1):6–16, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jamieson:2021:STH

[379] Leanna Jamieson, Angel Waters, Kaitlyn E. Ho, Harvey Y. S. Chan, Jacky T. Hung, Sarah E. Webb, Ching Man Chan, Alan M. Shipley, John G. Williamson, Jon Beer, Chevonne Angus, and Andrew L. Miller. Short-term homeostatic regulation of blood/interstitial fluid Ca^{2+} concentration by the scales of anadromous sea trout Salmo trutta L. during smoltification and migration. Journal of Fish Biology, 98(1):17–32, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vivancos:2021:HCD

[380] Aurélien Vivancos, Konrad Górski, Aliro Manosalva, Barbara Toledo, Malcolm Reid, and Evelyn Habit. Hydrological connectivity drives longitudinal movement of endangered endemic Chilean darter Percilia irwini (Eigenmann, 1927). Journal of Fish Biology, 98(1):33–43, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Meyer:2021:GOL

[381] Kevin A. Meyer and Daniel J. Schill. The Gill–Oxygen Limitation Theory and size at maturity/maximum size relationships for salmonid populations occupying flowing waters. Journal of Fish Biology, 98(1):44–49, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coulson:2021:LHC

[382] Peter G. Coulson. The life history characteristics of Neosebastes pandus and the relationship between sexually dimorphic growth and reproductive strategy among scorpaeniformes. Journal of Fish Biology, 98(1):50–63, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wagner:2021:UTI

[383] Maximilian Wagner, Marcelo Kovačić, and Stephan Koblmüller. Unravelling the taxonomy of an interstitial fish radiation: Three new species of Gouania (Teleostei: Gobiesocidae) from the Mediterranean Sea and redescriptions of G. willdenowi and G. pigra. Journal of Fish Biology, 98(1):64–88, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brewster:2021:FIV

[384] Lauran R. Brewster, Brianna V. Cahill, Miranda N. Burton, Cassady Dougan, Jeffrey S. Herr, Laura Issac Norton, Samantha A. McGuire, Marisa Pico, Elizabeth Urban-Gedamke, Kim Bassos-Hull, John P. Tyminski, Robert E. Hueter, Bradley M. Wetherbee, Mahmood Shivji, Neil Burnie, and Matthew J. Ajemian. First insights into the vertical habitat use of the whitespotted eagle ray Aetobatus narinari revealed by pop-up satellite archival tags. Journal of Fish Biology, 98(1):89–101, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hvas:2021:HRA

[385] Malthe Hvas, Ole Folkedal, and Frode Oppedal. Heart rates of Atlantic salmon Salmo salar during a critical swim speed test and subsequent recovery. Journal of Fish Biology, 98(1):102–111, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baumann:2021:TYA

[386] Loïc Baumann, Joanna Vega, Joris Philip, Fabien Polese, Fabrice Vétillard, Maud Pierre, Romaric Le Barh, Philippe Jatteau, Agnès Bardonnet, and Marie-Laure Acolas. Tolerance of young allis shad Alosa alosa (Clupeidae) to oxy-thermic stress. Journal of Fish Biology, 98 (1):112–131, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cohen:2021:MBA

[387] Stefania Cohen, Alcira O. Díaz, and Marina V. Diaz. Morphological and biochemical approaches to assess the nutritional condition of the Argentine hake Merluccius hubbsi larvae from two different nursery areas. Journal of Fish Biology, 98(1):132–141, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Peixoto:2021:NSG

[388] Luiz A. W. Peixoto, Murilo N. L. Pastana, and Gustavo A. Ballen. New species of glass knifefish genus Eigenmannia (Gymnotiformes: Sternopygidae) with comments on the morphology and function of the enlarged cephalic lateral-line canals of Sternopygidae. Journal of Fish Biology, 98 (1):142–153, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Londono-Burbano:2021:LCS

[389] Alejandro Londoño-Burbano, Alexander Urbano-Bonilla, and Matthew R. Thomas. Loricaria cuffyi (Siluriformes: Loricariidae), a new species of loricariin catfish from the Guiana Shield. Journal of Fish Biology, 98 (1):154–167, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wueringer:2021:AML

[390] Barbara E. Wueringer, Marit Winther-Janson, Vincent Raoult, and Tristan L. Guttridge. Anatomy of the mechanosensory lateral line canal system and electrosensory ampullae of Lorenzini in two species of sawshark (fam. Pristiophoridae). Journal of Fish Biology, 98(1):168–177, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Teixeira:2021:NSC

[391] Tulio F. Teixeira and Marcelo R. S. Melo. A new species of Characidium reinhardt (Characiformes: Crenuchidae) from the Río Huallaga, central Peruvian Amazon, with a review of secondary sexual characters in the genus. Journal of Fish Biology, 98(1):178–188, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Castillo:2021:FPD

[392] Yuritzi Castillo and Elsah Arce. Female preference for dominant males in the Mexican mojarra cichlid fish, Cichlasoma istlanum. Journal of Fish Biology, 98(1):189–195, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Junior:2021:TOB

[393] Alexandre da Gama Fernandes Vieira Júnior, Diele Emele Pontes Carvalho de Lima, Natalice dos Santos Sales, Bianca de Freitas Terra, and André Pessanha. Trade-offs between ontogenetic changes and food consumption in Brazilian silverside Atherinella brasiliensis from two tropical estuaries. Journal of Fish Biology, 98(1):196–207, January 2021. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tian:2021:RES

[394] Jinbao Tian, Yong Zhang, Rong Zhu, Yeqing Wu, Xiaoxiao Liu, and Xingguo Wang. Red elemental selenium (Se^{0}) improves the immunoactivities of EPC cells, crucian carp and zebrafish against spring viraemia of carp virus. Journal of Fish Biology, 98(1):208–218, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Malabarba:2021:HUM

[395] Luiz R. Malabarba, Junior Chuctaya, Alice Hirschmann, Eduardo Bitencourt de Oliveira, and Andréa T. Thomaz. Hidden or unnoticed? Multiple lines of evidence support the recognition of a new species of Pseudocorynopoma (Characidae: Corynopomini). Journal of Fish Biology, 98(1):219–236, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Klinard:2021:AML

[396] Natalie V. Klinard, Jordan K. Matley, Silviya V. Ivanova, Sarah M. Larocque, Aaron T. Fisk, and Timothy B. Johnson. Application of machine learning to identify predators of stocked fish in Lake Ontario: using acoustic telemetry predation tags to inform management. Journal of Fish Biology, 98(1):237–250, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Treasurer:2021:RDM

[397] James W. Treasurer. Reproductive development and maturity stage benchmarking in a freshwater teleost ruffe Gymnocephalus cernuus L. in Loch Lomond. Journal of Fish Biology, 98(1):251–266, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Irigoyen:2021:PGA

[398] Alejo Irigoyen, Cecilia Riestra, Sofía Lopez, and Ana M. Parma. Patagonian grouper Acanthistius patachonicus (Jenyns, 1840) forms massive spawning aggregations in the Southwest Atlantic. Journal of Fish Biology, 98(1):267–276, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Geladakis:2021:DOS

[399] George Geladakis, Stylianos Somarakis, and George Koumoundouros. Differences in otolith shape and fluctuating-asymmetry between reared and wild gilthead seabream (Sparus aurata Linnaeus, 1758). Journal of Fish Biology, 98(1):277–286, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Joyce:2021:RSC

[400] William Joyce and Michael Axelsson. Regulation of splenic contraction persists as a vestigial trait in white-blooded Antarctic fishes. Journal of Fish Biology, 98(1):287–291, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sifundza:2021:DPH

[401] Delsy S. Sifundza, Albert Chakona, and Wilbert T. Kadye. Distribution patterns and habitat associations of Sandelia bainsii (Teleostei: Anabantidae), a highly threatened narrow-range endemic freshwater fish. Journal of Fish Biology, 98(1):292–303, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Espinasa:2021:VAR

[402] Luis Espinasa, Carly Heintz, Sylvie Rétaux, Masato Yoshisawa, François Agnès, Patricia Ornelas-Garcia, and Robert Balogh-Robinson. Vibration attraction response is a plastic trait in blind Mexican tetra (Astyanax mexicanus), variable within subpopulations inhabiting the same cave. Journal of Fish Biology, 98(1):304–316, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Macario:2021:IVI

[403] Alessandro Macario, Safi K. Darden, Frederick Verbruggen, and Darren P. Croft. Intraspecific variation in inhibitory motor control in guppies, Poecilia reticulata. Journal of Fish Biology, 98(1):317–328, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Devine:2021:LGR

[404] Brynn Devine and Lou Van Guelpen. Loss of gill rakers and teeth in adult specimens of barracudina Arctozenus risso (Aulopiformes: Paralepididae) from the western North Atlantic. Journal of Fish Biology, 98 (1):329–332, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marquez-Velasquez:2021:LDT

[405] Viviana Márquez-Velásquez, Rosseval Galdino Leite, Andres Hernandez-Serna, and Fredy Alvarado. Larval diet of two Amazonian goliath catfish species. Journal of Fish Biology, 98(1):333–336, January 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2021:IIb

[406] Anonymous. Issue information. Journal of Fish Biology, 98(2):337–338, February 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hanfling:2021:EDB

[407] Bernd Hänfling and Lori Lawson-Handley. Environmental DNA-based approaches for the monitoring of fish populations have come of age. Journal of Fish Biology, 98(2):339–340, February 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jerde:2021:CWM

[408] Christopher L. Jerde. Can we manage fisheries with the inherent uncertainty from eDNA? Journal of Fish Biology, 98(2):341–353, February 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pont:2021:FFB

[409] Didier Pont, Alice Valentini, Mathieu Rocle, Anthony Maire, Olivier Delaigue, Pauline Jean, and Tony Dejean. The future of fish-based ecological assessment of European rivers: from traditional EU Water Framework Directive compliant methods to eDNA metabarcoding-based approaches. Journal of Fish Biology, 98(2):354–366, February 2021. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Traugott:2021:FPP

[410] Michael Traugott, Bettina Thalinger, Corinna Wallinger, and Daniela Sint. Fish as predators and prey: DNA-based assessment of their role in food webs. Journal of Fish Biology, 98(2):367–382, February 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lacoursiere-Roussel:2021:EDT

[411] Anaïs Lacoursière-Roussel and Kristy Deiner. Environmental DNA is not the tool by itself. Journal of Fish Biology, 98(2):383–386, February 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Milhau:2021:SDR

[412] Tristan Milhau, Alice Valentini, Nicolas Poulet, Nicolas Roset, Pauline Jean, Coline Gaboriaud, and Tony Dejean. Seasonal dynamics of riverine fish communities using eDNA. Journal of Fish Biology, 98(2):387–398, February 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brys:2021:RED

[413] Rein Brys, David Halfmaerten, Sabrina Neyrinck, Quentin Mauvisseau, Johan Auwerx, Michael Sweet, and Joachim Mergeay. Reliable eDNA detection and quantification of the European weather loach (Misgurnus fossilis). Journal of Fish Biology, 98(2):399–414, February 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lafferty:2021:PAF

[414] Kevin D. Lafferty, Ana E. Garcia-Vedrenne, John P. McLaughlin, Jasmine N. Childress, Marisa F. Morse, and Christopher L. Jerde. At Palmyra Atoll, the fish-community environmental DNA signal changes across habitats but not with tides. Journal of Fish Biology, 98(2):415– 425, February 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2021:LWH

[415] Anonymous. Looking where it’s hard to see: a case study documenting rare Eucyclogobius newberryi presence in a California lagoon. Journal of Fish Biology, 98(2):426, February 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dressler:2021:EDM

[416] Terra L. Dressler, Kevin D. Lafferty, Christopher L. Jerde, and Tom Dudley. Environmental DNA metabarcoding provides enhanced detection of the European eel Anguilla anguilla and fish community structure in pumped river catchments. Journal of Fish Biology, 98(2):427, February 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Griffiths:2021:BDE

[417] Nathan P. Griffiths, Jonathan D. Bolland, Rosalind M. Wright, Leona A. Murphy, Robert K. Donnelly, Hayley V. Watson, and Bernd Hänfling. Biomass-dependent emission of environmental DNA in jack mackerel Trachurus japonicus juveniles. Journal of Fish Biology, 98(2):428, February 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Horiuchi:2021:EDM

[418] Tomoya Horiuchi, Reiji Masuda, Hiroaki Murakami, Satoshi Yamamoto, and Toshifumi Minamoto. Environmental DNA metabarcoding as an effective and rapid tool for fish monitoring in canals. Journal of Fish Biology, 98(2):429, February 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mandal:2021:GMA

[419] Sangeeta Mandal, Achal Singh, Priyanka Sah, Rajeev K. Singh, Raj Kumar, Kuldeep K. Lal, and Vindhya Mohindra. Genetic and morphological assessment of a vulnerable large catfish, Silonia silondia (Hamilton, 1822), in natural populations from India. Journal of Fish Biology, 98 (2):430–444, February 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Paillard:2021:FRE

[420] Adele Paillard, Kenshu Shimada, and Catalina Pimiento. The fossil record of extant elasmobranchs. Journal of Fish Biology, 98(2):445–455, February 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kumar:2021:CGS

[421] Rahul Girish Kumar, Ravi Charan, Nadumury Pradeep Krishnaprasoon, and Valaparambil Saidumohammad Basheer. Catfishes of the genus Sperata (Pisces: Bagridae) in India. Journal of Fish Biology, 98(2):456–469, February 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Matich:2021:NIT

[422] Philip Matich, Jeffrey D. Plumlee, Ornella C. Weideli, and Mark Fisher. New insights into the trophic ecology of blacktip sharks (Carcharhinus limbatus) from a subtropical estuary in the western Gulf of Mexico. Journal of Fish Biology, 98(2):470–484, February 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gargan:2021:SLV

[423] Patrick G. Gargan, Finn Økland, Samuel Shephard, Colm Fitzgerald, and Martin O’Grady. The spawning location of vulnerable ferox trout (Salmo trutta L.) in the Lough Corrib and Lough Mask catchments, Western Ireland. Journal of Fish Biology, 98(2):485–497, February 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mojekwu:2021:UDB

[424] Tonna O. Mojekwu, Michael J. Cunningham, Roger I. Bills, Petrus C. Pretorius, and Thierry B. Hoareau. Utility of DNA barcoding in native Oreochromis species. Journal of Fish Biology, 98(2):498–506, February 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Noda:2021:MRH

[425] Takuji Noda, Toshihiro Wada, Hiromichi Mitamura, Manabu Kume, Takuhei Komaki, Tsuneo Fujita, Tatsuma Sato, Kaoru Narita, Manabu Yamada, Akira Matsumoto, Tomoya Hori, Junichi Takagi, Alisa Kutzer, Nobuaki Arai, and Yoh Yamashita. Migration, residency and habitat utilisation by wild and cultured Japanese eels (Anguilla japonica) in a shallow brackish lagoon and inflowing rivers using acoustic telemetry. Journal of Fish Biology, 98(2):507–525, February 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eldoy:2021:BFE

[426] Sindre H. Eldøy, Jan G. Davidsen, Matthias Vignon, and Michael Power. The biology and feeding ecology of Arctic charr in the Kerguelen Islands. Journal of Fish Biology, 98(2):526–536, February 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moya:2021:GET

[427] Ana Carolina Moya, Francisco Acuña, María Constanza Díaz Andrade, Claudio Gustavo Barbeito, and Elena Juana Galíndez. Glycan expression as a tool for a deeper understanding of a reproductive gland in a skate of economic importance. Journal of Fish Biology, 98(2):537–547, February 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kikuchi:2021:DRP

[428] Eidi Kikuchi, Sebastían García, Paulo Alberto Silva da Costa, Luis Gustavo Cardoso, and Manuel Haimovici. Discrimination of red porgy Pagrus pagrus (Sparidae) potential stocks in the south-western Atlantic by otolith shape analysis. Journal of Fish Biology, 98(2):548–556, February 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dhamelincourt:2021:IGC

[429] Marius Dhamelincourt, Mathieu Buoro, Jacques Rives, Stellia Sebihi, and Cédric Tentelier. Individual and group characteristics affecting nest building in sea lamprey (Petromyzon marinus L. 1758). Journal of Fish Biology, 98(2):557–565, February 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Meyer:2021:UML

[430] Lauren Meyer, Sherrie Chambers, Connor Gervais, Heidi Pethybridge, Crystal Beckmann, Barry Bruce, and Charlie Huveneers. The use of muscle lipids and fatty acids to assess shark diet and condition. Journal of Fish Biology, 98(2):566–571, February 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ghorbankhah:2021:IIP

[431] Maedeh Ghorbankhah and Ali Bani. Inherent immunological parameters in kutum, Rutilus frisii, larvae obtained from the adult females with orange and green eggs. Journal of Fish Biology, 98(2):572–576, February 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weber:2021:GES

[432] Helen K. Weber, Christian M. Jones, Matthew J. Ajemian, Michael P. McCallister, Brent L. Winner, Gregg R. Poulakis, Dana M. Bethea, Lisa D. Hollensead, Daniel Zapf, John D. Swenson, Jill M. Hendon, Toby S. Daly-Engel, and Nicole M. Phillips. Genetic evidence supports a range extension for the Brazilian cownose ray Rhinoptera brasiliensis in the western North Atlantic. Journal of Fish Biology, 98(2):577–582, February 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rutledge:2021:FRG

[433] Kelsi M. Rutledge. First record of Gorgona guitarfish (Pseudobatos prahli) off the Baja California peninsula with updated key to the guitarfishes of the North Eastern Pacific. Journal of Fish Biology, 98(2): 583–586, February 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2021:IIc

[434] Anonymous. Issue information. Journal of Fish Biology, 98(3):589–590, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2021:EDI

[435] Anonymous. Elasmobranch decline and the importance of good data: the case of Angel sharks. Journal of Fish Biology, 98(3):591, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ellis:2021:ASS

[436] Jim R. Ellis, Joanna Barker, Sophy R. McCully Phillips, Eva K. M. Meyers, and Michelle Heupel. Angel sharks (Squatinidae): a review of biological knowledge and exploitation. Journal of Fish Biology, 98(3): 592–621, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Allan:2021:REC

[437] Hugh Allan, Richard P. Duncan, Peter Unmack, Duanne White, and Mark Lintermans. Reproductive ecology of a critically endangered alpine galaxiid. Journal of Fish Biology, 98(3):622–633, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Keep:2021:LLI

[438] John K. Keep, Jabin R. Watson, Rebecca L. Cramp, Matthew J. Jones, Matthew A. Gordos, Patrick J. Ward, and Craig E. Franklin. Low light intensities increase avoidance behaviour of diurnal fish species: implications for use of road culverts by fish. Journal of Fish Biology, 98(3): 634–642, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Albanesi:2021:UEO

[439] Camila Albanesi, Mariano González-Castro, and Alejandra López-Mañanes. Understanding the early ontogenetic stages of Mugil liza (Mugilidae): Morphological traits and digestive/metabolic profile of prejuveniles after recruitment. Journal of Fish Biology, 98(3):643–654, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Braganca:2021:DTN

[440] Pedro H. N. Bragança, Jouke R. van der Zee, Rainer Sonnenberg, and Emmanuel J. W. M. N. Vreven. Description of two new miniature species of Hylopanchax Poll & Lambert, 1965 (Cyprinodontiformes: Procatopodidae), from northeastern Gabon, with an updated diagnosis of the genus based on morphology, colouration and osteology. Journal of Fish Biology, 98(3):655–667, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Souza:2021:MMD

[441] Camila S. Souza, Gabriel S. C. Silva, Luz E. Ochoa, Fabio F. Roxo, Guilherme J. Costa-Silva, Fausto Foresti, Bruno F. Melo, and Claudio Oliveira. Molecular and morphological diversity in species of Kronichthys (Teleostei, Loricariidae) from Atlantic coastal rivers of Brazil. Journal of Fish Biology, 98(3):668–679, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moorhead:2021:VBC

[442] Shannon G. Moorhead, Austin J. Gallagher, Liza Merly, and Neil Hammerschlag. Variation of body condition and plasma energy substrates with life stage, sex, and season in wild-sampled nurse sharks Ginglymostoma cirratum. Journal of Fish Biology, 98(3):680–693, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rauber:2021:SSR

[443] Rafaela G. Rauber, Larissa Strictar, Luiz C. Gomes, Harumi I. Suzuki, and Angelo A. Agostinho. Spatial segregation in the reproductive activity of Neotropical fish species as an indicator of the migratory trait. Journal of Fish Biology, 98(3):694–706, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Funk:2021:FFG

[444] Steffen Funk, Romain Frelat, Christian Möllmann, Axel Temming, and Uwe Krumme. The forgotten feeding ground: patterns in seasonal and depth-specific food intake of adult cod Gadus morhua in the western Baltic Sea. Journal of Fish Biology, 98(3):707–722, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nay:2021:RTT

[445] Tiffany J. Nay, Rohan J. Longbottom, Connor R. Gervais, Jacob L. Johansen, John F. Steffensen, Jodie L. Rummer, and Andrew S. Hoey. Regulate or tolerate: Thermal strategy of a coral reef flat resident, the epaulette shark, Hemiscyllium ocellatum. Journal of Fish Biology, 98 (3):723–732, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kunishima:2021:RTD

[446] Taiga Kunishima, Hirotoshi Saimaru, and Katsunori Tachihara. Reproductive traits of the dwarf gobies Pandaka trimaculata and Pandaka lidwilli in the western Pacific Ocean: histological evidence from one of the smallest fishes in the world. Journal of Fish Biology, 98(3):733–742, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Guedes:2021:ERS

[447] Gustavo Henrique Soares Guedes, Iracema David Gomes, Aparecida Alves do Nascimento, Fernanda Silva de Aguiar, and Francisco Gerson Araújo. Equilibrium reproductive strategy of the peacock bass Cichla kelberi facilitates invasion into a Neotropical reservoir. Journal of Fish Biology, 98(3):743–755, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Holubova:2021:FPF

[448] Michaela Holubová, Josef Hejzlar, Martin Čech, Mojmír Vašek, Petr Blabolil, and Jiří Peterka. Fluctuations in pelagic fish density linked to ambient conditions. Journal of Fish Biology, 98(3):756–767, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gonzalez-Pestana:2021:DTI

[449] Adriana Gonzalez-Pestana, Jeffrey C. Mangel, Eliana Alfaro-Córdova, Nicolas Acuña-Perales, Francisco Córdova-Zavaleta, Eduardo Segura-Cobeña, Diego Benites, Maximiliano Espinoza, Javier Coasaca-Céspedes, Astrid Jiménez, Sergio Pingo, Victor Moscoso, Joanna Alfaro-Shigueto, and Pepe Espinoza. Diet, trophic interactions and possible ecological role of commercial sharks and batoids in northern Peruvian waters. Journal of Fish Biology, 98(3):768–783, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baro-Camarasa:2021:IRP

[450] Isis Baró-Camarasa, Ana J. Marmolejo-Rodríguez, Todd M. O’Hara, Fernando R. Elorriaga-Verplancken, Abel Trejo-Ramírez, Raúl O. Martínez-Rincón, and Felipe Galván-Magaña. Isotopic (δ ^{15}N) relationship of pregnant females and their embryos: Comparing placental and yolk-sac viviparous elasmobranchs. Journal of Fish Biology, 98(3):784–790, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Soto-Lopez:2021:RBR

[451] Katherin Soto-López, Rosa Isabel Ochoa-Báez, Felipe Galván-Magaña, and Maria Cristina Oddone. Reproductive biology of the Rasptail skate Rostroraja velezi (Rajiformes: Rajidae). Journal of Fish Biology, 98 (3):791–802, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Peixoto:2021:NSS

[452] Luiz Antônio Wanderley Peixoto, Aléssio Datovo, Naércio Aquino Menezes, and Carlos David de Santana. A new species of sexually dimorphic and rheophilic ghost knifefish (Apteronotidae: Gymnotiformes) from the Amazon basin. Journal of Fish Biology, 98(3):803–816, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Puncher:2021:LSD

[453] Gregory Neils Puncher, Sherrylynn Rowe, George A. Rose, Geneviève J. Parent, Yanjun Wang, and Scott A. Pavey. Life-stage-dependent supergene haplotype frequencies and metapopulation neutral genetic patterns of Atlantic cod, Gadus morhua, from Canada’s Northern cod stock region and adjacent areas. Journal of Fish Biology, 98(3):817–828, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pahl:2021:DAP

[454] K. Blue Pahl, David J. Yurkowski, Sabine P. Wintner, Geremy Cliff, Matthew L. Dicken, and Nigel E. Hussey. Determining the appropriate pretreatment procedures and the utility of liver tissue for bulk stable isotope (δ ^{13}C and δ ^{15}N) studies in sharks. Journal of Fish Biology, 98(3):829–841, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Long:2021:ECC

[455] Yong Long, Ran Liu, Guili Song, Qing Li, and Zongbin Cui. Establishment and characterization of a cold-sensitive neural cell line from the brain of tilapia (Oreochromis niloticus). Journal of Fish Biology, 98(3): 842–854, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Williams:2021:GDL

[456] Collin T. Williams, Ashlie J. McIvor, Elizabeth M. Wallace, Yu-Jia Lin, and Michael L. Berumen. Genetic diversity and life-history traits of bonefish Albula spp. from the Red Sea. Journal of Fish Biology, 98(3): 855–864, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whitehead:2021:UML

[457] Darren A. Whitehead, Felipe G. Magaña, James T. Ketchum, Edgar M. Hoyos, Rogelio G. Armas, Francesca Pancaldi, and Damien Olivier. The use of machine learning to detect foraging behaviour in whale sharks: a new tool in conservation. Journal of Fish Biology, 98(3):865–869, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bustamante:2021:OCA

[458] Carlos Bustamante, Ana M. García-Cegarra, and Carolina Vargas-Caro. Observations of coastal aggregations of the broadnose sevengill shark (Notorynchus cepedianus) in Chilean waters. Journal of Fish Biology, 98(3):870–873, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gouveia:2021:TRB

[459] Élida J. Gouveia, Lidiany D. Cavalcanti, Fernanda C. Leal, Sarah G. F. Mendes, and Márcia R. Russo. Trophic relationship between the Patinga hybrid (Piaractus mesopotamicus × Piaractus brachypomus) and Echinorhynchus gomesi Machado Filho, 1948 in fish farms. Journal of Fish Biology, 98(3):874–877, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zarco:2021:PIR

[460] Agustín Zarco, Valeria Corbalán, and Guillermo Debandi. Predation by invasive rainbow trout on the critically endangered Pehuenche spinychest frog. Journal of Fish Biology, 98(3):878–880, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Miranda:2021:DIR

[461] Joni A. Miranda, Natasha Yates, Ariana Agustines, Nonie P. Enolva, Jessica Labaja, Christine Legaspi, Emer McCoy, Alessandro Ponzo, Sally Snow, and Gonzalo Araujo. Donsol: an important reproductive habitat for the world’s largest fish Rhincodon typus? Journal of Fish Biology, 98(3):881–885, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hachim:2021:OUH

[462] Mouniboudine Hachim, Tristan Rouyer, Gilbert Dutto, Vincent Kerzerho, Serge Bernard, Jérôme Bourjea, and David J. McKenzie. Oxygen uptake, heart rate and activities of locomotor muscles during a critical swimming speed protocol in the gilthead sea bream Sparus aurata. Journal of Fish Biology, 98(3):886–890, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gardiner:2021:EBB

[463] Jayne M. Gardiner and Tonya R. Wiley. Escaped bait: Bull shark Carcharhinus leucas with an intentionally attached harness rig. Journal of Fish Biology, 98(3):891–894, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baer:2021:QAR

[464] Jan Baer, Cornelius Becke, Helga M. Bentele, Friederike Habedank, and Alexander Brinker. Quantification of alizarin red S uptake in coregonid eggs after mass-marking. Journal of Fish Biology, 98(3):895–897, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harasti:2021:BRF

[465] David Harasti. Book review: Frogfishes: Biodiversity, Zoogeography and Behavioural Ecology, Theodore W. Pietsch and Rachel J. Arnold, John Hopkins University Press, Baltimore, 586 pp. ISBN: 978-1-4214-3252-6. Journal of Fish Biology, 98(3):898, March 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2021:IId

[466] Anonymous. Issue information. Journal of Fish Biology, 98(4):901–902, April 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Witten:2021:BFS

[467] P. Eckhard Witten, Ann Huysseune, John G. Maisey, Christoph Winkler, and Zhiyuan Gong. A boost for fish skeletal research. Journal of Fish Biology, 98(4):903–905, April 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Berio:2021:EDG

[468] Fidji Berio and Mélanie Debiais-Thibaud. Evolutionary developmental genetics of teeth and odontodes in jawed vertebrates: a perspective from the study of elasmobranchs. Journal of Fish Biology, 98(4):906– 918, April 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maisey:2021:AUF

[469] John G. Maisey, John S. S. Denton, Carole Burrow, and Alan Pradel. Architectural and ultrastructural features of tessellated calcified cartilage in modern and extinct chondrichthyan fishes. Journal of Fish Biology, 98(4):919–941, April 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Seidel:2021:MAT

[470] Ronald Seidel, Aravind K. Jayasankar, and Mason N. Dean. The multiscale architecture of tessellated cartilage and its relation to function. Journal of Fish Biology, 98(4):942–955, April 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kimmel:2021:TMS

[471] Charles B. Kimmel, Alexander L. Wind, Whitney Oliva, Samuel D. Ahlquist, Charline Walker, John Dowd, Bernardo Blanco-Sánchez, Tom A. Titus, Peter Batzel, Jared C. Talbot, John H. Postlethwait, and James T. Nichols. Transgene-mediated skeletal phenotypic variation in zebrafish. Journal of Fish Biology, 98(4):956–970, April 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martini:2021:PSS

[472] Arianna Martini, Ann Huysseune, P. Eckhard Witten, and Clara Boglione. Plasticity of the skeleton and skeletal deformities in zebrafish (Danio rerio) linked to rearing density. Journal of Fish Biology, 98(4): 971–986, April 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Printzi:2021:EIL

[473] Alice Printzi, Stefanos Fragkoulis, Anastasia Dimitriadi, Kleoniki Keklikoglou, Christos Arvanitidis, P. Eckhard Witten, and George Koumoundouros. Exercise-induced lordosis in zebrafish Danio rerio (Hamilton, 1822). Journal of Fish Biology, 98(4):987–994, April 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ofer:2021:CSC

[474] Lior Ofer, Paul Zaslansky, and Ron Shahar. A comparison of the structure, composition and mechanical properties of anosteocytic vertebrae of medaka (o. latipes) and osteocytic vertebrae of zebrafish (d. rerio). Journal of Fish Biology, 98(4):995–1006, April 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bek:2021:PFP

[475] Jan Willem Bek, Adelbert De Clercq, Hanna De Saffel, Mieke Soenens, Ann Huysseune, P. Eckhard Witten, Paul J. Coucke, and Andy Willaert. Photoconvertible fluorescent proteins: a versatile tool in zebrafish skeletal imaging. Journal of Fish Biology, 98(4):1007–1017, April 2021. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Imangali:2021:DAD

[476] Nurgul Imangali, Quang Tien Phan, Gail Mahady, and Christoph Winkler. The dietary anthocyanin delphinidin prevents bone resorption by inhibiting rankl-induced differentiation of osteoclasts in a medaka (Oryzias latipes) model of osteoporosis. Journal of Fish Biology, 98(4):1018–1030, April 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Khajuria:2021:NMR

[477] Deepak Kumar Khajuria and David Karasik. Novel model of restricted mobility induced osteopenia in zebrafish. Journal of Fish Biology, 98 (4):1031–1038, April 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pham:2021:IRB

[478] Cuong V. Pham, Thanh T. Pham, Thuy T. Lai, Dat C. Trinh, Huong V. M. Nguyen, Tam T. M. Ha, Thuong T. Phuong, Long D. Tran, Christoph Winkler, and Thuy T. To. Icariin reduces bone loss in a rankl-induced transgenic medaka (Oryzias latipes) model for osteoporosis. Journal of Fish Biology, 98(4):1039–1048, April 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fjelldal:2021:SDW

[479] Per Gunnar Fjelldal, Angelico Madaro, Malthe Hvas, Lars Helge Stien, Frode Oppedal, and Thomas Wk Fraser. Skeletal deformities in wild and farmed cleaner fish species used in Atlantic salmon Salmo salar aquaculture. Journal of Fish Biology, 98(4):1049–1058, April 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fraser:2021:VDI

[480] Thomas W. K. Fraser, Tom J. Hansen, Florian Sambraus, and Per Gunnar Fjelldal. Vertebral deformities in interspecific diploid and triploid salmonid hybrids. Journal of Fish Biology, 98(4):1059–1070, April 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2021:HCD

[481] Anonymous. Holocephalan (Chondrichthyes) dental plates with hypermineralized dentine as a substitute for missing teeth through developmental plasticity. Journal of Fish Biology, 98(4):1071, April 2021. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2021:TSH

[482] Moya Smith, Esther Manzanares, Charlie Underwood, Chris Healy, Brett Clark, and Zerina Johanson. A tomographic study of the histological structure of teeth in the gilthead sea bream, Sparus aurata (Teleostei, Perciformes, Sparidae). Journal of Fish Biology, 98(4):1072, April 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Germain:2021:TAS

[483] Damien Germain and François J. Meunier. Triploid Atlantic salmon Salmo salar have a higher dietary phosphorus requirement for bone mineralization during early development. Journal of Fish Biology, 98(4): 1073, April 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eisele:2021:DRF

[484] Marius H. Eisele, Sergio Madrigal-Mora, and Mario Espinoza. Drivers of reef fish assemblages in an upwelling region from the Eastern Tropical Pacific Ocean. Journal of Fish Biology, 98(4):1074–1090, April 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Urbano-Bonilla:2021:NSC

[485] Alexander Urbano-Bonilla and Gustavo A. Ballen. A new species of Chaetostoma (Siluriformes: Loricariidae) from the Orinoco basin with comments on Amazonian species of the genus in Colombia. Journal of Fish Biology, 98(4):1091–1104, April 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Currier:2021:GSB

[486] Megan Currier, Jon Rouse, and David J. Coughlin. Group swimming behaviour and energetics in bluegill Lepomis macrochirus and rainbow trout Oncorhynchus mykiss. Journal of Fish Biology, 98(4):1105–1111, April 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stenevik:2021:DSD

[487] Erling K. Stenevik, Richard D. M. Nash, and Arild Folkvord. Development stage distribution as a proxy for feeding success and growth for first feeding Norwegian spring spawning herring larvae. Journal of Fish Biology, 98(4):1112–1119, April 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shervette:2021:AGG

[488] Virginia R. Shervette, Jesús M. Rivera Hernández, and Francis Kofi Ewusie Nunoo. Age and growth of grey triggerfish Balistes capriscus from trans-Atlantic populations. Journal of Fish Biology, 98(4):1120– 1136, April 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Valipour:2021:EDE

[489] AbdolMajid Valipour, Behrooz Heidari, S. Mohsen Asghari, Saeed Balalaie, Hanieh Rabouti, and Navid Omidian. The effect of different exogenous kisspeptins on sex hormones and reproductive indices of the goldfish (Carassius auratus) broodstock. Journal of Fish Biology, 98 (4):1137–1143, April 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Horn:2021:GAS

[490] Peter L. Horn and Caoimhghin Ó Maolagáin. The growth and age structure of Chilean jack mackerel (Trachurus murphyi) following its influx to New Zealand waters. Journal of Fish Biology, 98(4):1144–1154, April 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Simkova:2021:VRT

[491] Andrea Šimková, Michal Janáč, Pavel Hyršl, Vadym Krasnovyd, and Lukáš Vetešník. Vigour-related traits and immunity in hybrids of evolutionary divergent cyprinoid species: advantages of hybrid heterosis? Journal of Fish Biology, 98(4):1155–1171, April 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ostbye:2021:MHM

[492] Tone-Kari K. Østbye, Nardos T. Woldemariam, Camilla E. Lundberg, Gerd M. Berge, Bente Ruyter, and Rune Andreassen. Modulation of hepatic miRNA expression in Atlantic salmon (Salmo salar) by family background and dietary fatty acid composition. Journal of Fish Biology, 98(4):1172–1185, April 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chuctaya:2021:NER

[493] Junior Chuctaya, Andrea C. Encalada, Karla S. Barragán, Maria L. Torres, Karla E. Rojas, Valeria Ochoa-Herrera, and Tiago P. Carvalho. New Ecuadorian records of the eyeless banjo catfish Micromyzon akamai (Siluriformes: Aspredinidae) expand the species range and reveal intraspecific morphological variation. Journal of Fish Biology, 98(4):1186–1191, April 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Backstrom:2021:PCV

[494] Tobias Backström, Eva Brännäs, Jan Nilsson, Hanna Carlberg, Kajsa Johansson, and Carin Magnhagen. Proof of concept: visual categorization of carotenoid pigmentation in Arctic charr (Salvelinus alpinus L.) can predict stress response. Journal of Fish Biology, 98(4):1192–1195, April 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stuart:2021:NIP

[495] Courtney E. Stuart, Stephanie J. Green, Oksana Vernygora, Aaron R. H. LeBlanc, Mary H. Bertschi, and Maria Brown. New insights into patterns and rates of tooth replacement in serrasalmid and characid fishes, with implications for the subsistence fishery of Peru’s remote ribereños villages. Journal of Fish Biology, 98(4):1196–1201, April 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2021:IIe

[496] Anonymous. Issue information. Journal of Fish Biology, 98(5):1203– 1204, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2021:FPI

[497] William Bernard Perry. Foul play: the impact of dog waste on freshwater fish. Journal of Fish Biology, 98(5):1205, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chabot:2021:VOU

[498] Denis Chabot, Yangfan Zhang, and Anthony P. Farrell. Valid oxygen uptake measurements: using high r^{2} values with good intentions can bias upward the determination of standard metabolic rate. Journal of Fish Biology, 98(5):1206–1216, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

daSilva:2021:PEC

[499] Melina Coelho da Silva, Adelino Vicente Mendonça Canário, Peter Colin Hubbard, and David Manuel Flores Gonçalves. Physiology, endocrinology and chemical communication in aggressive behaviour of fishes. Journal of Fish Biology, 98(5):1217–1233, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ericsson:2021:PDI

[500] Philip Ericsson, Anders Persson, Jane W. Behrens, Tomas Brodin, Philipp Emanuel Hirsch, Anna Sundelin, Mikael van Deurs, Lisa W. von Friesen, and P. Anders Nilsson. Personality-dependent inter- and intraspecific foraging competition in the invasive round goby, Neogobius melanostomus. Journal of Fish Biology, 98(5):1234–1241, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chretien:2021:DSI

[501] Emmanuelle Chrétien, Steven J. Cooke, and Daniel Boisclair. Does shelter influence the metabolic traits of a teleost fish? Journal of Fish Biology, 98(5):1242–1252, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beltran-Lopez:2021:GDA

[502] Rosa Gabriela Beltrán-López, Omar Domínguez-Domínguez, Kyle R. Piller, Humberto Mejía-Mojica, Adán Fernando Mar-Silva, and Ignacio Doadrio. Genetic differentiation among populations of the blackfin goodea Goodea atripinnis (Cyprinodontiformes: Goodeidaea): implications for its evolutionary history. Journal of Fish Biology, 98(5):1253– 1266, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reguera-Rouzaud:2021:DGS

[503] Nicole Reguera-Rouzaud, Noé Díaz-Viloria, Ricardo Pérez-Enríquez, Elaine Espino-Barr, Mailin Isabel Rivera-Lucero, and Adrián Munguía-Vega. Drivers for genetic structure at different geographic scales for Pacific red snapper (Lutjanus peru) and yellow snapper (Lutjanus argentiventris) in the tropical eastern Pacific. Journal of Fish Biology, 98 (5):1267–1280, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garcia-Ayala:2021:PFN

[504] James R. Garcia-Ayala and Ricardo C. Benine. Poptella fuscata, a new Stethaprionini from the upper Amazon basin, Peru (Characiformes: Characidae). Journal of Fish Biology, 98(5):1281–1288, May 2021. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jan:2021:HSB

[505] Kousar Jan, Imtiaz Ahmed, and Nazir A. Dar. Haematological and serum biochemical reference values of snow trout, Schizothorax labiatus habiting in river Sindh of Indian Himalayan region. Journal of Fish Biology, 98 (5):1289–1302, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Raick:2021:AHP

[506] Xavier Raick, Rodney Rountree, Gregório Kurchevski, Francis Juanes, Alessia Huby, Alexandre L. Godinho, and Éric Parmentier. Acoustic homogeneity in the piranha Serrasalmus maculatus. Journal of Fish Biology, 98(5):1303–1307, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2021:PIF

[507] Jiaxing Chen, Cheng Peng, Jingjun Huang, Herong Shi, Ling Xiao, Lin Tang, Haoran Lin, Shuisheng Li, and Yong Zhang. Physical interactions facilitate sex change in the protogynous orange-spotted grouper, Epinephelus coioides. Journal of Fish Biology, 98(5):1308–1320, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kinziger:2021:RSB

[508] Andrew P. Kinziger, Jason L. White, Rodney J. Nakamoto, and Bret C. Harvey. Recent, small beginnings: genetic analysis suggests Catostomus rimiculus (Klamath smallscale sucker) in the Smith River, California, are introduced. Journal of Fish Biology, 98(5):1321–1328, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lin:2021:SWL

[509] Yu-Jia Lin, Lotfi Rabaoui, Rommel H. Maneja, Abdu Rahiman Kambrath Pulikkoden, Panickan Premlal, Zahid Nazeer, Mohammad A. Qurban, Khaled Abdulkader, Perdana K. Prihartato, Ali M. Qasem, Nabil Fita, and Ruben H. Roa-Ureta. Strengths and weaknesses in the longterm sustainability of two sympatric seabreams (Argyrops spinifer and Rhabdosargus haffara, Sparidae). Journal of Fish Biology, 98(5):1329– 1341, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mones:2021:ICV

[510] Alissa B. Mones, Craig A. Harms, and Michael K. Stoskopf. Implications of comparative ventral body wall histology on selection of abdominal surgical approach and closure in 12 species of fish. Journal of Fish Biology, 98(5):1342–1348, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pena:2021:MBE

[511] Renato Peña, Ivette Moguel-Hernández, and Gretchen M. Haro-Ballesteros. Morphological and biochemical effects of food deprivation during the early development of Pacific red snapper Lutjanus peru. Journal of Fish Biology, 98(5):1349–1362, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mohapatra:2021:NSG

[512] Anil Mohapatra, Swarup Ranjan Mohanty, Dipanjan Ray, Subhrendu Sekhar Mishra, and Jaya Kishor Seth. A new species of the genus Cirrhimuraena (Anguilliformes: Ophichthidae) from the Bay of Bengal, India. Journal of Fish Biology, 98(5):1363–1370, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Penido:2021:WDC

[513] Iago S. Penido, Tiago C. Pessali, and Cláudio H. Zawadzki. When destruction comes first: Two new species of Hypostomus Lacépède, 1803 (Siluriformes: Loricariidae) from a deeply-impacted river in the Rio São Francisco basin in Brazil. Journal of Fish Biology, 98(5):1371–1384, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Levina:2021:TPF

[514] Aleksandra D. Levina, Elena S. Mikhailova, and Alexander O. Kasumyan. Taste preferences and feeding behaviour in the facultative herbivorous fish, Nile tilapia Oreochromis niloticus. Journal of Fish Biology, 98(5):1385–1400, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Austad:2021:FVO

[515] Benedikte Austad, L. Asbjørn Vøllestad, and Anders Foldvik. Frequency of vateritic otoliths and potential consequences for marine survival in hatchery-reared Atlantic salmon. Journal of Fish Biology, 98(5):1401– 1409, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bunt:2021:ECU

[516] Christopher M. Bunt and Bailey Jacobson. Exposure to a common urban pollutant affects the survival and swimming behaviour of creek chub (Semotilus atromaculatus). Journal of Fish Biology, 98(5):1410–1420, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Appleyard:2021:ADF

[517] Sharon A. Appleyard, Safia Maher, John J. Pogonoski, Stephen J. Bent, Xin-Yi Chua, and Annette McGrath. Assessing DNA for fish identifications from reference collections: the good, bad and ugly shed light on formalin fixation and sequencing approaches. Journal of Fish Biology, 98(5):1421–1432, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lorenzoni:2021:CDW

[518] Massimo Lorenzoni, Antonella Carosi, Silvia Quadroni, Vanessa De Santis, Isabella Vanetti, Giovanni B. Delmastro, and Serena Zaccara. Cryptic diversity within endemic Italian barbels: revalidation and description of new Barbus species (Teleostei: Cyprinidae). Journal of Fish Biology, 98(5):1433–1449, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Charmpila:2021:FDO

[519] Eleni A. Charmpila, Maximilian Wagner, and Bettina Reichenbacher. First documentation of the otoliths of the species of Gouania (Teleostei: Gobiesocidae) in the Mediterranean Sea. Journal of Fish Biology, 98 (5):1450–1455, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garzon-Pena:2021:IER

[520] Lina Vanesa Garzón-Peña, Angelica Barrera-García, Antonio Delgado-Huertas, and Carlos Julio Polo-Silva. Isotopic enrichment in Rhizoprionodon porosus (Poey, 1861) and Carcharhinus porosus (Ranzani, 1840) embryos. Journal of Fish Biology, 98(5):1456–1458, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Matsuda:2021:CAA

[521] Keishi Matsuda. A comparison of avoidance to acoustic stimuli in fish with different auditory capabilities: juvenile chum salmon (Oncorhynchus keta) and common carp (Cyprinus carpio). Journal of Fish Biology, 98 (5):1459–1464, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Santos:2021:CSR

[522] Paulo R. S. Santos, Marcelo N. Alves, and Magnus M. Severo. Coastal sharks and rays in a marine protected area in southern Brazil. Journal of Fish Biology, 98(5):1465–1470, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carnovali:2021:MHE

[523] Marta Carnovali, Massimo Mariotti, and Giuseppe Banfi. Molecular hydrogen enhances osteogenesis in Danio rerio embryos. Journal of Fish Biology, 98(5):1471–1474, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dettleff:2021:TEO

[524] Phillip Dettleff, Rodrigo Zuloaga, Marcia Fuentes, Juan M. Estrada, Alfredo Molina, and Juan A. Valdés. Temperature effect on oxidative stress and egg quality-related genes on post-ovulatory eggs and ovary of red cusk-eel (Genypterus chilensis). Journal of Fish Biology, 98(5):1475– 1480, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jonsson:2021:COS

[525] Bror Jonsson and Nina Jonsson. Continuous outmigration and sequential encountering of environmental cues are important for successful homing of hatchery-reared, anadromous brown trout Salmo trutta. Journal of Fish Biology, 98(5):1481–1484, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moritz:2021:BRB

[526] Timo Moritz. Book review: The Bichir Handbook, Joshua Pickett, Published by Independent Publishing Network, London, UK, 138 pp., price £24.95, ISBN: 978-1-78972-770-8. Journal of Fish Biology, 98(5):1485– 1486, May 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2021:IIf

[527] Anonymous. Issue information. Journal of Fish Biology, 98(6):1487– 1488, June 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McKenzie:2021:EGW

[528] David J. McKenzie, Benjamin Geffroy, and Anthony P. Farrell. Effects of global warming on fishes and fisheries. Journal of Fish Biology, 98 (6):1489–1492, June 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2021:WDW

[529] Anonymous. What do warming waters mean for fish physiology and fisheries? Journal of Fish Biology, 98(6):1493, June 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2021:GSH

[530] Anonymous. From gametogenesis to spawning: How climate-driven warming affects teleost reproductive biology. Journal of Fish Biology, 98(6):1494, June 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2021:EGW

[531] Anonymous. Effects of global warming on sex ratios in fishes. Journal of Fish Biology, 98(6):1495, June 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alfonso:2021:TIE

[532] Sébastien Alfonso, Manuel Gesto, and Bastien Sadoul. Temperature increase and its effects on fish stress physiology in the context of global warming. Journal of Fish Biology, 98(6):1496–1508, June 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Feidantsis:2021:AUI

[533] Konstantinos Feidantsis, Hans O. Pörtner, Ioannis A. Giantsis, and Basile Michaelidis. Advances in understanding the impacts of global warming on marine fishes farmed offshore: Sparus aurata as a case study. Journal of Fish Biology, 98(6):1509–1523, June 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morash:2021:PUD

[534] Andrea J. Morash, Ben Speers-Roesch, Sean Andrew, and Suzanne Currie. The physiological ups and downs of thermal variability in temperate freshwater ecosystems. Journal of Fish Biology, 98(6):1524–1535, June 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McKenzie:2021:IVT

[535] David J. McKenzie, Yangfan Zhang, Erika J. Eliason, Patricia M. Schulte, Guy Claireaux, Felipe R. Blasco, Julie J. H. Nati, and Anthony P. Farrell. Intraspecific variation in tolerance of warming in fishes. Journal of Fish Biology, 98(6):1536–1555, June 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2021:CEB

[536] Anonymous. The combined effect of body size and temperature on oxygen consumption rates and the size-dependency of preferred temperature in European perch Perca fluviatilis. Journal of Fish Biology, 98(6):1556, June 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2021:OWH

[537] Anonymous. Ocean warming and hypoxia affect embryonic growth, fitness and survival of small-spotted catsharks, Scyliorhinus canicula. Journal of Fish Biology, 98(6):1557, June 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vargas-Chacoff:2021:FEO

[538] Luis Vargas-Chacoff, Francisco Dann, Kurt Paschke, Ricardo Oyarzún-Salazar, Daniela Nualart, Danixa Martínez, Jonathan M. Wilson, Pedro M. Guerreiro, and Jorge M. Navarro. Freshening effect on the osmotic response of the Antarctic spiny plunderfish Harpagifer antarcticus. Journal of Fish Biology, 98(6):1558–1571, June 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lloret:2021:PHD

[539] Josep Lloret, Alba Serrat, Gudmundur Thordarson, Kristin Helle, Angelique Jadaud, Isabel Bruno, Francesc Ordines, Paolo Sartor, Pierluigi Carbonara, and Hans-Joachim Rätz. The poor health of deep-water species in the context of fishing activity and a warming climate: will populations of Molva species rebuild or collapse? Journal of Fish Biology, 98(6):1572–1584, June 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morissette:2021:TVN

[540] Jenna Morissette, Sula Swart, Tyson J. MacCormack, Suzanne Currie, and Andrea J. Morash. Thermal variation near the thermal optimum does not affect the growth, metabolism or swimming performance in wild Atlantic salmon Salmo salar. Journal of Fish Biology, 98(6):1585– 1589, June 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2021:IIg

[541] Anonymous. Issue information. Journal of Fish Biology, 99(1):1–2, July 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2021:SRG

[542] William Bernard Perry. Some rays get extra help from their mothers. Journal of Fish Biology, 99(1):3, July 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Laursen:2021:PRM

[543] Danielle Caroline Laursen and Franck L. B. Meijboom. Pausing for reflection to make progress: an assessment framework for ethical discussions on innovations in fisheries. Journal of Fish Biology, 99(1):4–8, July 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sun:2021:MCF

[544] Minmin Sun, Shipeng Jiang, Weihao Song, Jiangbo Qu, and Jie Qi. Molecular characterization and functional analysis of cyp11a and cyp11b in black rockfish (Sebastes schlegelii). Journal of Fish Biology, 99(1): 9–17, July 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jonsson:2021:DGB

[545] Bror Jonsson and Nina Jonsson. Differences in growth between offspring of anadromous and freshwater brown trout Salmo trutta. Journal of Fish Biology, 99(1):18–24, July 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Xu:2021:IEE

[546] Yuxin Xu, Hongxia Li, Dihui Xu, Jianlin Li, Fan Yu, Meiyao Wang, Qin Wang, Yunsheng Wu, Qiyuan Zhang, Yongkai Tang, and Juhua Yu. Identification, expression and enzyme activity of the group III sPLA_{2} s in Cyprinus carpio L. Journal of Fish Biology, 99(1):25–36, July 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yuncu:2021:DSX

[547] Eren Yüncü, Sinan Can Açan, Vedat Onar, F. Saadet Karakulak, Mehmet Gökoğlu, Turgut Zahit Alıçlı, Fidel Chiriboga, İnci Togan, and Füsun Özer. Demography of swordfish (Xiphias gladius Linneus) populations from the coasts of Turkey, based on mitochondrial DNA and microsatellites. Journal of Fish Biology, 99(1):37–48, July 2021. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maes:2021:HGF

[548] Sarah M. Maes, Henrik Christiansen, Felix C. Mark, Magnus Lucassen, Anton Van de Putte, Filip A. M. Volckaert, and Hauke Flores. High gene flow in polar cod (Boreogadus saida) from West-Svalbard and the Eurasian Basin. Journal of Fish Biology, 99(1):49–60, July 2021. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Amtstaetter:2021:ERD

[549] Frank Amtstaetter, Jian D. L. Yen, Robin Hale, Wayne Koster, Justin O’Connor, Ivor Stuart, and Zeb Tonkin. Elevated river discharge enhances the immigration of juvenile catadromous and amphidromous fishes into temperate coastal rivers. Journal of Fish Biology, 99(1):61– 72, July 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yoon:2021:CMQ

[550] Gwangseok R. Yoon, Forrest Bjornson, David Deslauriers, and W. Gary Anderson. Comparison of methods to quantify metabolic rate and its relationship with activity in larval lake sturgeon Acipenser fulvescens. Journal of Fish Biology, 99(1):73–86, July 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hammer:2021:UTI

[551] Michael P. Hammer, Laura Taillebois, Alison J. King, David A. Crook, Dion Wedd, Mark Adams, Peter J. Unmack, Doug F. Hoese, and Terry Bertozzi. Unravelling the taxonomy and identification of a problematic group of benthic fishes from tropical rivers (Gobiidae: Glossogobius). Journal of Fish Biology, 99(1):87–100, July 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kousteni:2021:SLD

[552] Vasiliki Kousteni. Shedding light on the deep: The case of the velvet belly lanternshark in the North Aegean Sea. Journal of Fish Biology, 99(1):101–117, July 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kupprat:2021:IIO

[553] Franziska Kupprat, Franz Hölker, Klaus Knopf, Torsten Preuer, and Werner Kloas. Innate immunity, oxidative stress and body indices of Eurasian perch Perca fluviatilis after two weeks of exposure to artificial light at night. Journal of Fish Biology, 99(1):118–130, July 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gomi:2021:DVM

[554] Kazuma Gomi, Yohei Nakamura, Masaru Kanda, Kentaro Honda, Masahiro Nakaoka, Chiho Honma, and Masao Adachi. Diel vertical movements and feeding behaviour of blue humphead parrotfish Scarus ovifrons in a temperate reef of Japan. Journal of Fish Biology, 99(1): 131–142, July 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Azevedo:2021:ITR

[555] Filipe Manoel Azevedo, Cláudio Henrique Zawadzki, Thatiana Vanessa Soria, Thomaz Mansini Carrenho Fabrin, Alessandra Valéria De Oliveira, Sônia Maria Alves Pinto Prioli, and Alberto Jose Prioli. Integrative taxonomy reveals the historically poorly defined armoured catfish Hypostomus variipictus (Ihering 1911), from the upper Rio Paraná basin, Brazil (Siluriformes, Loricariidae). Journal of Fish Biology, 99(1):143–152, July 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pascucci:2021:HDC

[556] Daniel Pascucci, Christopher J. Gobler, and Konstantine J. Rountos. Harmful dinoflagellate Cochlodinium polykrikoides impairs the feeding behavior of larval sheepshead minnows (Cyprinodon variegatus). Journal of Fish Biology, 99(1):153–163, July 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Basusta:2021:SDO

[557] Nuri Başusta and Umar Khan. Sexual dimorphism in the otolith shape of shi drum, Umbrina cirrosa (L.), in the eastern Mediterranean Sea: Fish size–otolith size relationships. Journal of Fish Biology, 99(1):164– 174, July 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Khudamrongsawat:2021:PGS

[558] Jenjit Khudamrongsawat, Jes Kettratad, Pituk Intasorn, Nipon Pinyo, Sornpasok Tapcheewin, and Ponchai Wanusrut. Pattern of genetic structure of the common stream fish, Neolissochilus soroides (Pisces: Cyprinidae), addresses the importance of protected areas in eastern Thailand. Journal of Fish Biology, 99(1):175–185, July 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scarponi:2021:PRC

[559] Valentina Scarponi, Enrico Gennari, and William Hughes. Physiological response to capture stress in endemic Southern African catsharks (family Scyliorhinidae). Journal of Fish Biology, 99(1):186–196, July 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crichigno:2021:TER

[560] Sonia A. Crichigno, Mabel Orellana, Rodrigo Larraza, Guillermo Mirenna, and Víctor E. Cussac. Thermal effects in rainbow trout (Oncorhynchus mykiss ) F1 embryos (farmed female × wild thermal-resistant male). Journal of Fish Biology, 99(1):197–205, July 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mohamad:2021:HET

[561] Suhaini Mohamad, Hon Jung Liew, Rabiatul Adawiyyah Zainuddin, Sharifah Rahmah, Khor Waiho, Mazlan Abd Ghaffar, Hua Thai Nhan, Jiun-Yan Loh, Leong-Seng Lim, Yumei Chang, Liqun Liang, and Gudrun De Boeck. High environmental temperature and low pH stress alter the gill phenotypic plasticity of Hoven’s carp Leptobarbus hoevenii. Journal of Fish Biology, 99(1):206–218, July 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mazungula:2021:ITR

[562] Daniel Nkosinathi Mazungula and Albert Chakona. An integrative taxonomic review of the Natal mountain catfish, Amphilius natalensis Boulenger 1917 (Siluriformes, Amphiliidae), with description of four new species. Journal of Fish Biology, 99(1):219–239, July 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kina:2021:CUM

[563] Taito Kina, Tetsuya Masuda, Kiyoshi Asahina, and Miwa Suzuki. Composition of uterine milk and its changes with gestational period in red stingrays (Hemitrygon akajei). Journal of Fish Biology, 99(1):240–252, July 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Catarino:2021:FRP

[564] Diana Catarino, Sergio Stefanni, Filipe M. Porteiro, Alexandra Rosa, and Eva Giacomello. First record of the pencil cardinal Epigonus denticulatus (Perciformes: Epigonidae) in the Azores archipelago. Journal of Fish Biology, 99(1):253–257, July 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kousteni:2021:MDB

[565] Vasiliki Kousteni. Morphometric description and biological notes on the rare kitefin shark Dalatias licha (Chondrichthyes: Dalatidae) from the Hellenic waters. Journal of Fish Biology, 99(1):258–263, July 2021. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ern:2021:AHA

[566] Rasmus Ern and Andrew J. Esbaugh. Assessment of hypoxia avoidance behaviours in a eurythermal fish at two temperatures using a modified shuttlebox system. Journal of Fish Biology, 99(1):264–270, July 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burke:2021:UCB

[567] Patrick J. Burke and Jane E. Williamson. Using cone beam CT scans to reveal headfirst ingestion and possible prey manipulation tactics in sawsharks. Journal of Fish Biology, 99(1):271–274, July 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Laurrabaquio-Alvarado:2021:MDG

[568] Nadia Sandoval Laurrabaquio-Alvarado, Píndaro Díaz-Jaimes, Silvia Hinojosa-Álvarez, Maria del Pilar Blanco-Parra, Douglas H. Adams, Juan Carlos Pérez-Jiménez, and J. Leonardo Castillo-Géniz. Mitochondrial DNA genome evidence for the existence of a third divergent lineage in the western Atlantic Ocean for the bull shark (Carcharhinus leucas).

Journal of Fish Biology, 99(1):275–282, July 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arroyave:2021:CDP

[569] Jairo Arroyave and Dalia A. De La Cruz Fernández. Cave-dwelling populations of Rhamdia (Siluriformes: Heptapteridae) from the Sierra de Zongolica, Veracruz, Mexico: an evidence-based checklist with comments on their evolutionary history and taxonomy. Journal of Fish Biology, 99 (1):283–287, July 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kumai:2021:SFA

[570] Yusuke Kumai and Mari Kuroki. Salinity, freshwater and agricultural water preferences of glass eels of the Japanese eel Anguilla japonica collected in southern Japan. Journal of Fish Biology, 99(1):288–292, July 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sternes:2021:BRF

[571] Phillip C. Sternes. Book review: Field Guide to Sharks, Rays, & Chimaeras of Europe and the Mediterranean, D. A. Ebert & M. Dando, Published by Princeton University Press, Princeton, NJ, 2021, 383 pp., Price EUR 28/$33. ISBN: 978-0-691-20598-4. Journal of Fish Biology, 99(1):293–294, July 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2021:IIh

[572] Anonymous. Issue information. Journal of Fish Biology, 99(2):295–296, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2021:WIC

[573] William Bernard Perry. When it comes to brown trout and Arctic charr migrants, home is where the heart is. Journal of Fish Biology, 99(2): 297, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Guinand:2021:BTP

[574] Bruno Guinand, Münevver Oral, and Christelle Tougard. Brown trout phylogenetics: a persistent mirage towards (too) many species. Journal of Fish Biology, 99(2):298–307, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pensinger:2021:RBH

[575] Lucas G. Pensinger, Nancy J. Brown-Peterson, Christopher C. Green, and Stephen R. Midway. Reproductive biology of hardhead catfish Ariopsis felis: evidence for overwintering oocytes. Journal of Fish Biology, 99(2):308–320, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kisekelwa:2021:ABL

[576] Tchalondawa Kisekelwa, Jos Snoeks, Armel Ibala Zamba, Gaston Sefu Amzati, Mwapu Isumbisho, Pascal Mulungula Masilya, Pieter Lemmens, and Emmanuel Vreven. Association between Labeobarbus spp. (Teleostei: Cyprinidae) and environmental variables in the Luhoho basin (Eastern Congo River basin; DRC). Journal of Fish Biology, 99(2):321–334, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Andrews:2021:EBS

[577] Kimberly R. Andrews, Iria Fernandez-Silva, John E. Randall, and Hsuan-Ching Ho. Etelis boweni sp. nov., a new cryptic deepwater eteline snapper from the Indo–Pacific (Perciformes: Lutjanidae). Journal of Fish Biology, 99(2):335–344, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Qiu:2021:CTC

[578] Reng Qiu, Yuan Y. Sun, Cui C. Guan, Yun C. Kan, and Lun G. Yao. Characterization of TCR^{+} and CD8^{+} head kidney leucocytes in Japanese flounder (Paralichthys olivaceus) with antisera against TCRα and CD8α. Journal of Fish Biology, 99(2):345–353, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wells:2021:NOP

[579] R. J. David Wells, Veronica A. Quesnell, Robert L. Humphreys, Jr., Heidi Dewar, Jay R. Rooker, Jaime Alvarado Bremer, and Owyn E. Snodgrass. Nursery origin and population connectivity of swordfish Xiphias gladius in the North Pacific Ocean. Journal of Fish Biology, 99(2):354– 363, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cottrell:2021:POY

[580] Amy M. Cottrell, Solomon R. David, and Patrick S. Forsythe. Production and outmigration of young-of-year northern pike Esox lucius from natural and modified waterways connected to Lower Green Bay, Wisconsin. Journal of Fish Biology, 99(2):364–372, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sun:2021:TRR

[581] Zhi-Xian Sun, Seigo Kawase, Rui Zhang, and Ya-Hui Zhao. Taxonomic revision and redescription of Microphysogobio hsinglungshanensis, the type species of Microphysogobio Mori, 1934 (Cypriniformes: Cyprinidae). Journal of Fish Biology, 99(2):373–383, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lulijwa:2021:HMP

[582] Ronald Lulijwa, Andrea C. Alfaro, Leonie Venter, Tim Young, Paul Decker, Fabrice Merien, and Jill Meyer. Haematological and metabolic profiles associated with age and sex in giant kokopu (Galaxias argenteus) (Gmelin 1789) broodstock. Journal of Fish Biology, 99(2):384–395, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Calixto-Rojas:2021:PRE

[583] Miguel Calixto-Rojas, Andrés Lira-Noriega, Miguel Rubio-Godoy, Gerardo Pérez-Ponce de León, and Carlos D. Pinacho-Pinacho. Phylogenetic relationships and ecological niche conservatism in killifish (Profundulidae) in Mesoamerica. Journal of Fish Biology, 99(2):396–410, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Okyere:2021:SHE

[584] Isaac Okyere, Juliet Afrah Obeng, Samuel Ayitey, Jouke Rients Van der Zee, and Herman Meeus. Spawning habits and embryonic development of the banded lampeye killifish Aplocheilichthys spilauchen (Duméril 1861) in ex situ fresh and brackish water environments. Journal of Fish Biology, 99(2):411–417, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pinion:2021:LAO

[585] Amanda K. Pinion, Dustin S. Siegel, Ralf Britz, Rafael Martínez-García, Carlos A. Álvarez-González, and Kevin W. Conway. The larval attachment organ of the tropical gar Atractosteus tropicus Gill, 1863 (Lepisosteiformes: Lepisosteidae). Journal of Fish Biology, 99(2):418–424, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Franz:2021:OGC

[586] George P. Franz, Lars Lewerentz, and Bianka Grunow. Observations of growth changes during the embryonic-larval-transition of pikeperch (Sander lucioperca) under near-natural conditions. Journal of Fish Biology, 99(2):425–436, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Torres-Huerta:2021:SPD

[587] Ana María Torres-Huerta, Alicia Cruz-Martínez, and Eduardo Juventino Ramírez-Chávez. Spatial pattern of distribution and reproductive strategy of vermiculate electric-rays Narcine vermiculatus. Journal of Fish Biology, 99(2):437–449, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Choudhury:2021:DNS

[588] Hrishikesh Choudhury, Rajdeep Das, Ratul Ch. Bharali, Kangkan Sarma, Lalit K. Tyagi, Kuldeep K. Lal, and Dandadhar Sarma. Description of a new species of Mustura kottelat (Teleostei: Nemacheilidae) from the Brahmaputra drainage, India. Journal of Fish Biology, 99(2):450– 461, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Atencio:2021:KCR

[589] Benjamin J. Atencio, Eva B. Thorstad, Audun H. Rikardsen, and Jenny L. A. Jensen. Keeping close to the river, shore and surface: the first marine migration of brown trout (Salmo trutta) and Arctic charr (Salvelinus alpinus) post-smolts. Journal of Fish Biology, 99(2):462–471, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vaz:2021:SPW

[590] Diego F. B. Vaz. Scymnodon plunketi (Waite, 1910): a junior synonym of Scymnodon macracanthus (Regan, 1906) (Somniosidae: Elasmobranchii). Journal of Fish Biology, 99(2):472–494, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jawad:2021:SDG

[591] Laith A. Jawad, Jassim M. Abed, and Ana L. Ibáňez. Stock differentiation of the greater lizardfish Saurida tumbil (Teleostei: Synodontidae) collected along the western coast of the Arabian Gulf and Sea of Oman using meristic characters. Journal of Fish Biology, 99(2):495–501, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Makaras:2021:AEF

[592] Tomas Makaras, Milda Stankevičiūtė, Eglė Šidagytė-Copilas, Tomas Virbickas, and Julija Razumienė. Acclimation effect on fish behavioural characteristics: determination of appropriate acclimation period for different species. Journal of Fish Biology, 99(2):502–512, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Giroux:2021:ETS

[593] Marissa Giroux and Daniel Schlenk. The effects of temperature and salinity on the endocrinology in two life stages of juvenile rainbow/steelhead trout (Oncorhynchus mykiss ). Journal of Fish Biology, 99(2):513–523, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kim:2021:MMC

[594] Jin-Seok Kim, Hwan-Sung Ji, and Jin-Koo Kim. Molecular and morphological comparison of larvae and juveniles of five species of callionymids, with ontogenetic evidence on the monophyly of Repomucenus (Callionymidae). Journal of Fish Biology, 99(2):524–542, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mutizwa:2021:DGM

[595] Tadiwa I. Mutizwa, Wilbert T. Kadye, and Albert Chakona. Deep genetic and morphological divergence in the Hippopotamyrus ansorgii species complex (Teleostei: Mormyridae) in southern Africa. Journal of Fish Biology, 99(2):543–556, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Logan:2021:BSX

[596] John M. Logan, Walt Golet, Sean C. Smith, John Neilson, and Lou Van Guelpen. Broadbill swordfish (Xiphias gladius) foraging and vertical movements in the north-west Atlantic. Journal of Fish Biology, 99(2): 557–568, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Geffen:2021:LSM

[597] Audrey J. Geffen, Jon Albretsen, Bastian Huwer, and Richard D. M. Nash. Lemon sole Microstomus kitt in the northern North Sea: a multidisciplinary approach to the early life-history dynamics. Journal of Fish Biology, 99(2):569–580, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Saboret:2021:UPT

[598] Grégoire Saboret, Philip Dermond, and Jakob Brodersen. Using PIT-tags and portable antennas for quantification of fish movement and survival in streams under different environmental conditions. Journal of Fish Biology, 99(2):581–595, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fleury:2021:RBE

[599] Aharon G. Fleury, Eva M. MacLennan, Rylan J. Command, and Francis Juanes. Reproductive biology and ecology of Pacific hagfish (Eptatretus stoutii) and black hagfish (Eptatretus deani). Journal of Fish Biology, 99(2):596–606, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lindstrom:2021:SAR

[600] Kai Lindström, Jonathan Havenhand, Erica Leder, Sofie Schöld, Ola Svensson, and Charlotta Kvarnemo. Sperm adaptation in relation to salinity in three goby species. Journal of Fish Biology, 99(2):607–613, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tencatt:2021:CFS

[601] Luiz Fernando Caserta Tencatt, Sérgio Alexandre dos Santos, Hans-Georg Evers, and Marcelo R. Britto. Corydoras fulleri (Siluriformes: Callichthyidae), a new catfish species from the Rio Madeira basin, Peru. Journal of Fish Biology, 99(2):614–628, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Simier:2021:BES

[602] Monique Simier, Olaloudé Judicaël Franck Osse, Oumar Sadio, and Jean-Marc Ecoutin. Biology and ecology of sea catfish (Ariidae) of estuarine, lagoon and coastal ecosystems in West Africa. Journal of Fish Biology, 99(2):629–643, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lira:2021:FGP

[603] Mateus G. S. Lira, Waldir M. Berbel-Filho, Helder M. V. Espírito-Santo, Andrei Tatarenkov, John C. Avise, Carlos Garcia de Leaniz, Sofia Consuegra, and Sergio M. Q. Lima. Filling the gaps: phylogeography of the self-fertilizing Kryptolebias species (Cyprinodontiformes: Rivulidae) along South American mangroves. Journal of Fish Biology, 99(2):644– 655, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sobri:2021:LSM

[604] Noor Zahidah Ahmad Sobri, Sébastien Lavoué, Fazimah Aziz, Siti Azizah Mohd Nor, Noor Adelyna Mohammed Akib, and Md. Zain Khaironizam. To lump, to split or to maintain? Molecular taxonomy of the spotted barb Barbodes binotatus (Cyprinidae) and closely related species in Peninsular Malaysia. Journal of Fish Biology, 99(2):656–668, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

HoffmannCostaeSilva:2021:RBF

[605] Gisleine Hoffmann Costa e Silva, Matheus Oliveira Freitas, and Vinícius Abilhoa. Reproductive biology of the fat snook Centropomus parallelus Poey, 1860 (Teleostei, Centropomidae) and implications for its management in the southern Atlantic Ocean. Journal of Fish Biology, 99(2): 669–672, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mabruk:2021:RSM

[606] Sara A. A. Al Mabruk, Abdulghani Abdulghani, Ola Mohamed Nour, Mohammed Adel, Fabio Crocetta, Nikolaos Doumpas, Periklis Kleitou, and Francesco Tiralongo. The role of social media in compensating for the lack of field studies: Five new fish species for Mediterranean Egypt. Journal of Fish Biology, 99(2):673–678, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arvizu:2021:IPE

[607] Brittany Arvizu, Bridie J. M. Allan, and Justin R. Rizzari. Indirect predator effects influence behaviour but not morphology of juvenile coral reef Ambon damselfish Pomacentrus amboinensis. Journal of Fish Biology, 99(2):679–683, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lulijwa:2021:EAF

[608] Ronald Lulijwa, Andrea C. Alfaro, Tim Young, Leonie Venter, Paul Decker, Fabrice Merien, and Jill Meyer. Effect of anticoagulants on farmed giant kokopu, Galaxias argenteus (Gmelin 1789) haematological parameters and erythrocyte fragility. Journal of Fish Biology, 99 (2):684–689, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shen:2021:PME

[609] Yawei Shen, Huiyang Li, Min Zhou, Shoujie Tang, Yan Zhao, Jinliang Zhao, and Xiaowu Chen. Promoter methylation and expression of the rassf4 gene in Nile tilapia (Oreochromis niloticus). Journal of Fish Biology, 99(2):690–694, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Freret-Meurer:2021:EFS

[610] Natalie Villar Freret-Meurer, Amanda do Carmo Vaccani, and Gabriela dos Santos Cabiró. Evidence of feminization in seahorses from a tropical estuary. Journal of Fish Biology, 99(2):695–699, August 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2021:IIi

[611] Anonymous. Issue information. Journal of Fish Biology, 99(3):701–702, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2021:LCF

[612] William Bernard Perry. Life in captivity: future research in physical enrichment for fish. Journal of Fish Biology, 99(3):703, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jones:2021:PER

[613] Nick A. R. Jones, Mike M. Webster, and Anne Gro Vea Salvanes. Physical enrichment research for captive fish: Time to focus on the DE-TAILS. Journal of Fish Biology, 99(3):704–725, September 2021. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

LombardiBrandao:2021:UBI

[614] Manuela Lombardi Brandão, Felipe Dorigão-Guimarães, Marcela Cesar Bolognesi, Ana Carolina Dos Santos Gauy, André Vitor Salinas Pereira, Lethicia Vian, Thais Billalba Carvalho, and Eliane Gonçalves de Freitas. Understanding behaviour to improve the welfare of an ornamental fish. Journal of Fish Biology, 99(3):726–739, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Koenig:2021:SCS

[615] Lauren A. Koenig and Jason R. Gallant. Sperm competition, sexual selection and the diverse reproductive biology of osteoglossiformes. Journal of Fish Biology, 99(3):740–754, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Paramo:2021:MSP

[616] Jorge Paramo, Rodrigo Wiff, and Rina González. A matter of size: the population structure of the smallest known living shark, Etmopterus perryi (Springer & Burgess, 1985), from deep waters off the Colombian Caribbean coast. Journal of Fish Biology, 99(3):755–764, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Medina:2021:ABS

[617] Antonio Medina, Ana Magro, José Antonio Paullada-Salmerón, and José Luis Varela. An autofluorescence-based survey of late follicular atresia in ovaries of a teleost fish (Thunnus thynnus). Journal of Fish Biology, 99(3):765–772, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dominguez-Castanedo:2021:AIA

[618] Omar Domínguez-Castanedo. Agonistic interactions with asymmetric body size in two adult-age groups of the annual killifish Millerichthys robustus (Miller & Hubbs, 1974). Journal of Fish Biology, 99(3):773–781, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rey:2021:UMS

[619] Sonia Rey, Jim Treasurer, Connie Pattillo, and Bruce J. McAdam. Using model selection to choose a size-based condition index that is consistent with operational welfare indicators. Journal of Fish Biology, 99(3):782– 795, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carlson:2021:DMA

[620] Lindsay G. Carlson, Timothy F. Sheehan, Michael D. Tillotson, and Katherine E. Mills. Date of marine annulus formation in Atlantic salmon (Salmo salar) and implications for retrospective growth analyses using scales. Journal of Fish Biology, 99(3):796–806, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sheraliev:2021:TFN

[621] Bakhtiyor Sheraliev and Zuogang Peng. Triplophysa ferganaensis, a new loach species from Fergana Valley in Central Asia (Teleostei: Nemacheilidae). Journal of Fish Biology, 99(3):807–817, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bernal-Perez:2021:RBS

[622] Stephanie Bernal-Pérez, Rosa Isabel Ochoa-Báez, Felipe Galván-Magaña, and Katherin Soto-López. Reproductive biology of the swell shark Cephaloscyllium ventriosum (Carcharhiniformes: Scyliorhinidae). Journal of Fish Biology, 99(3):818–830, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Romero:2021:DTP

[623] Joana Romero, Carolina Vieira, Susana Garrido, Margarida Hermida, Paulo Catry, Graça Faria, and José Pedro Granadeiro. Diet and trophic position of two mackerel species in the archipelago of Madeira, Portugal. Journal of Fish Biology, 99(3):831–843, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tang:2021:DNP

[624] Chi-Ngai Tang and Hsuan-Ching Ho. Description of a new perchlet, Chelidoperca formosa, from southwestern Taiwan (Peciformes: Serranidae). Journal of Fish Biology, 99(3):844–855, September 2021. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Babin:2021:OMB

[625] Amanda B. Babin, Mouhamed Ndong, Katy Haralampides, Stephan Peake, Ross Jones, R. Allen Curry, and Tommi Linnansaari. Overwintering and migration behaviour of post-spawned Atlantic salmon Salmo salar in a large hydropower-regulated river and reservoir. Journal of Fish Biology, 99(3):856–874, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Paterson:2021:FIR

[626] Rachel A. Paterson, Henrik H. Berntsen, Tor F. Næsje, Marius Berg, and Bengt Finstad. Factors influencing return rate and marine residence duration in sea trout populations in Central Norway. Journal of Fish Biology, 99(3):875–887, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tavera:2021:NSP

[627] Jose Tavera. A new species of pikeblenny (Blenniiformes: Chaenopsidae: Chaenopsis) from the tropical eastern Pacific and a key to all species. Journal of Fish Biology, 99(3):888–895, September 2021. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garcia:2021:TBN

[628] Gabriel S. Garcia, Murilo S. Dias, and Guilherme O. Longo. Trade-off between number and length of remote videos for rapid assessments of reef fish assemblages. Journal of Fish Biology, 99(3):896–904, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Soares:2021:TNS

[629] Yan F. F. Soares, Pedro De Podestà Uchôa de Aquino, Justin C. Bagley, Francisco Langeani, and Guarino R. Colli. Two new species of Hypostomus suckermouth-armoured catfishes (Teleostei: Loricariidae) from central Brazil. Journal of Fish Biology, 99(3):905–920, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fahmi:2021:AGT

[630] Fahmi, Wanwan Kurniawan, Ian R. Tibbetts, Selvia Oktaviyani, Christine L. Dudgeon, and Michael B. Bennett. Age and growth of the tropical oviparous shark, Chiloscyllium punctatum in Indonesian waters. Journal of Fish Biology, 99(3):921–930, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kambikambi:2021:ADE

[631] Manda J. Kambikambi, Wilbert T. Kadye, and Albert Chakona. Allopatric differentiation in the Enteromius anoplus complex in South Africa, with the revalidation of Enteromius cernuus and Enteromius oraniensis, and description of a new species, Enteromius mandelai (Teleostei: Cyprinidae). Journal of Fish Biology, 99(3):931–954, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Monnet:2021:BVB

[632] Gauthier Monnet, Jordan S. Rosenfeld, and Jeffrey G. Richards. Behavioural variation between piscivore and insectivore rainbow trout Oncorhynchus mykiss. Journal of Fish Biology, 99(3):955–963, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rodriguez:2021:EAB

[633] Jose M. Rodriguez, Carolina Johnstone, and Diego Lozano-Peral. Evidence of Atlantic bluefin tuna spawning in the Bay of Biscay, northeastern Atlantic. Journal of Fish Biology, 99(3):964–969, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McMahan:2021:MSA

[634] Caleb D. McMahan, Diego J. Elías, Yue Li, Omár Domínguez-Domínguez, Sheila Rodriguez-Machado, Alejandra Morales-Cabrera, Diana Velásquez-Ramírez, Kyle R. Piller, Prosanta Chakrabarty, and Wilfredo A. Matamoros. Molecular systematics of the Awaous banana complex (river gobies; Teleostei: Oxudercidae). Journal of Fish Biology, 99 (3):970–979, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

King:2021:SCI

[635] R. Andrew King, Andrew L. Miller, and Jamie R. Stevens. Has stocking contributed to an increase in the rod catch of anadromous trout (Salmo trutta L.) in the Shetland Islands, UK? Journal of Fish Biology, 99(3): 980–989, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wong:2021:GSA

[636] Serena Wong, Jennifer S. Bigman, Kara E. Yopak, and Nicholas K. Dulvy. Gill surface area provides a clue for the respiratory basis of brain size in the blacktip shark (Carcharhinus limbatus). Journal of Fish Biology, 99(3):990–998, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shnaf:2021:CHP

[637] Anwaar S. M. Abu Shnaf, Shaban H. Abd El-Aziz, and Abdeltawab M. Ata. Cyto-histopathological and protein polymorphism alterations in five populations of Nile tilapia (Oreochromis niloticus) as biomonitor for water heavy metal pollution. Journal of Fish Biology, 99(3):999–1009, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Podder:2021:PMT

[638] Anupam Podder, Soumyadip Panja, Atreyee Chaudhuri, Anwesha Roy, Missidona Biswas, and Sumit Homechaudhuri. Patterns of morphological traits shaping the feeding guilds in the intertidal mudflat fishes of the Indian sundarbans. Journal of Fish Biology, 99(3):1010–1031, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bergstad:2021:NAD

[639] Odd Aksel Bergstad, Rebecca H. Hunter, Nicola J. Cousins, David M. Bailey, and Terje Jørgensen. Notes on age determination, size and age structure, longevity and growth of co-occurring macrourid fishes. Journal of Fish Biology, 99(3):1032–1043, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rohner:2021:MEB

[640] Christoph A. Rohner, Roy Bealey, Bernerd M. Fulanda, Jason D. Everett, Anthony J. Richardson, and Simon J. Pierce. Movement ecology of black marlin Istiompax indica in the Western Indian Ocean. Journal of Fish Biology, 99(3):1044–1059, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Irving:2021:DGH

[641] Kate Irving, Maren Wellenreuther, and Peter A. Ritchie. Description of the growth hormone gene of the Australasian snapper, Chrysophrys auratus, and associated intra- and interspecific genetic variation. Journal of Fish Biology, 99(3):1060–1070, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dai:2021:EBT

[642] Xiangyan Dai, Deyong Pu, Liping Wang, Xinkai Cheng, Xiaoqin Liu, Zhan Yin, and Zhijian Wang. Emergence of breeding tubercles and puberty onset in male zebrafish: evidence for a dependence on body growth. Journal of Fish Biology, 99(3):1071–1078, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fontrodona-Eslava:2021:NAB

[643] Ada Fontrodona-Eslava, Amy E. Deacon, Indar W. Ramnarine, and Anne E. Magurran. Numerical abundance and biomass reveal different temporal trends of functional diversity change in tropical fish assemblages. Journal of Fish Biology, 99(3):1079–1086, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hoarau:2021:AGB

[644] Florian Hoarau, Audrey Darnaude, Thomas Poirout, Lou-Anne Jannel, Maylis Labonne, and Sébastien Jaquemet. Age and growth of the bull shark (Carcharhinus leucas) around Reunion Island, South West Indian Ocean. Journal of Fish Biology, 99(3):1087–1099, September 2021. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lim:2021:IPW

[645] YouRee Lim, Vivienne Lee, Ayelen Blanco, Scott P. Kelly, and Suraj Unniappan. Ion-poor water and dietary salt deprivation upregulate the ghrelinergic system in the goldfish (Carassius auratus). Journal of Fish Biology, 99(3):1100–1109, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Parzanini:2021:FHS

[646] Camilla Parzanini, Michael T. Arts, Mehis Rohtla, Janet Koprivnikar, Michael Power, Anne Berit Skiftesvik, Howard I. Browman, Dino Milotic, and Caroline M. F. Durif. Feeding habitat and silvering stage affect lipid content and fatty acid composition of European eel Anguilla anguilla tissues. Journal of Fish Biology, 99(3):1110–1124, September 2021. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marsh:2021:WWC

[647] Jessica E. Marsh, Rasmus B. Lauridsen, William D. Riley, Olivia M. Simmons, Céline Artero, Luke J. Scott, William R. C. Beaumont, William A. Beaumont, John Davy-Bowker, Thomas Lecointre, Dylan E. Roberts, and Stephen D. Gregory. Warm winters and cool springs negatively influence recruitment of Atlantic salmon (Salmo salar L.) in a southern England chalk stream. Journal of Fish Biology, 99(3):1125–1129, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yaripour:2021:UAE

[648] Sareh Yaripour, Jukka Kekäläinen, Hannu Huuskonen, Matti Janhunen, and Raine Kortet. Ultra-acute exposure to cadmium does not impair whitefish sperm motility. Journal of Fish Biology, 99(3):1130–1134, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deMello:2021:ODS

[649] Paulo Henrique de Mello, Lícia Maria Lundstedt, Gilberto Moraes, Bruno Cavalheiro Araújo, Rossana Luiza Leite Venturieri, and Renata Guimarães Moreira. Ontogeny of the digestive system and the profile of proteases in larvae of cachara (Pseudoplatystoma reticulatum Siluriformes: Pimelodidae) and its hybrid (Pseudoplatystoma corruscans × Pseudoplatystoma reticulatum). Journal of Fish Biology, 99(3):1135– 1139, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cambra:2021:FRP

[650] Marta Cambra, Sergio Madrigal-Mora, Isaac Chinchilla, Geiner Golfín-Duarte, Christopher G. Lowe, and Mario Espinoza. First record of a potential neonate tiger shark (Galeocerdo cuvier) at a remote oceanic island in the Eastern Tropical Pacific. Journal of Fish Biology, 99(3):

1140–1144, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ferraz:2021:MDI

[651] João D. Ferraz, Lucas R. Jarduli, Diego A. Z. Garcia, Armando C. R. Casimiro, Fernanda S. de Almeida, and Mário L. Orsi. Morphological deformities in the invasive driftwood catfish Trachelyopterus galeatus in the Upper Paraná River basin, Brazil: a sign of human impact? Journal of Fish Biology, 99(3):1145–1149, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Benjamins:2021:FCC

[652] Steven Benjamins, Georgina Cole, Adam Naylor, James A. Thorburn, and Jane Dodd. First confirmed complete incubation of a flapper skate (Dipturus intermedius) egg in captivity. Journal of Fish Biology, 99(3): 1150–1154, September 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2021:IIj

[653] Anonymous. Issue information. Journal of Fish Biology, 99(4):1155– 1156, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2021:CKF

[654] William Bernard Perry. Complexity is key: fish diversity and tropical inshore seascapes. Journal of Fish Biology, 99(4):1157, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aguirre:2021:CTF

[655] Windsor E. Aguirre, Gabriela Alvarez-Mieles, Fernando Anaguano-Yancha, Ricardo Burgos Morán, Roberto V. Cucalón, Daniel Escobar-Camacho, Iván Jácome-Negrete, Pedro Jiménez Prado, Enrique Laaz, Katherin Miranda-Troya, Ronald Navarrete-Amaya, Fredy Nugra Salazar, Willan Revelo, Juan F. Rivadeneira, Jonathan Valdiviezo Rivera, and Edwin Zárate Hugo. Conservation threats and future prospects for the freshwater fishes of Ecuador: a hotspot of Neotropical fish diversity. Journal of Fish Biology, 99(4):1158–1189, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Graham:2021:AGE

[656] Patrick M. Graham, James S. Franks, Evan J. Anderson, Robert T. Leaf, and Jason D. Tilley. Age and growth of early-life-stage Atlantic tarpon (Megalops atlanticus) from the northcentral Gulf of Mexico. Journal of Fish Biology, 99(4):1190–1200, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aprahamian:2021:CTE

[657] Miran W. Aprahamian, Derek W. Evans, Cedric Briand, Alan M. Walker, Yvonne McElarney, and Michelle Allen. The changing times of Europe’s largest remaining commercially harvested population of eel Anguilla anguilla L. Journal of Fish Biology, 99(4):1201–1221, October 2021. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alves:2021:MDS

[658] Angélica P. C. Alves, Raquel T. Pereira, and Priscila V. Rosa. Morphology of the digestive system in carnivorous freshwater dourado Salminus brasiliensis. Journal of Fish Biology, 99(4):1222–1235, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Guo:2021:DSM

[659] Lian W. Guo, Stephen D. McCormick, Eric T. Schultz, and Adrian Jordaan. Direct and size-mediated effects of temperature and rationdependent growth rates on energy reserves in juvenile anadromous alewives (Alosa pseudoharengus). Journal of Fish Biology, 99(4):1236– 1246, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pardee:2021:AGM

[660] Cassandra Pardee, John Wiley, and Sarah Springer. Age, growth and maturity for two highly targeted jack species: Caranx ignobilis and Caranx melampygus. Journal of Fish Biology, 99(4):1247–1255, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ng:2021:CSG

[661] Casey Keat-Chuan Ng and Ji Tan. Cryptic species and grey zone speciation of the Barbodes binotatus complex (Teleostei, Cyprinidae) in Sundaland. Journal of Fish Biology, 99(4):1256–1273, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deCassiodaSilva:2021:ECU

[662] Rodrigo de Cássio da Silva, Rafael Rosas Oliveira, Sandra Maria Feliciano de Oliveira Azevedo, Raquel Moares Soares, and Valéria Freitas De Magalhães. Evidence of cylindrospermopsin uptake and clearance in fish (Oreochromis niloticus) under laboratory conditions. Journal of Fish Biology, 99(4):1274–1279, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic). See correction [1192].

Strom:2021:STS

[663] John Fredrik Strøm, Jenny L. A. Jensen, Anna Nikolopoulos, Eivind Nordli, Pål Arne Bjørn, and Thomas Bøhn. Sea trout Salmo trutta in the subarctic: home-bound but large variation in migratory behaviour between and within populations. Journal of Fish Biology, 99(4):1280– 1291, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bauer:2021:NSV

[664] Arthur B. Bauer, Werner W. Schwarzhans, Rodrigo L. Moura, José Anchieta C. C. Nunes, and Michael M. Mincarone. A new species of viviparous brotula genus Pseudogilbia (Ophidiiformes: Dinematichthyidae) from Brazilian reefs, with an updated diagnosis of the genus. Journal of Fish Biology, 99(4):1292–1298, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davenport:2021:DCT

[665] John Davenport, Vincent Maran, and Bruno Chanet. Does the common topknot Zeugopterus punctatus (Teleosteii: Pleuronectiformes: Scophthalmidae) use a novel Venturi-effect attachment mechanism? A testable hypothesis. Journal of Fish Biology, 99(4):1299–1306, October 2021. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schons:2021:EHS

[666] Rieke F. Schons, Simon Vitt, and Timo Thünken. Environmental habituation and sexual composition affect juveniles’ shoaling activity in a cichlid fish (Pelvicachromis taeniatus). Journal of Fish Biology, 99(4): 1307–1317, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Plaul:2021:GMM

[667] Silvia E. Plaul, Alcira O. Díaz, and Claudio G. Barbeito. Gill morphology and morphometry of the facultative air-breathing armoured catfish, Corydoras paleatus, in relation on aquatic respiration. Journal of Fish Biology, 99(4):1318–1327, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barbini:2021:AGM

[668] Santiago A. Barbini, David E. Sabadin, Jorge M. Román, Pablo A. Scarabotti, and Luis O. Lucifora. Age, growth, maturity and extinction risk of an exploited and endangered skate, Atlantoraja castelnaui, from off Uruguay and northern Argentina. Journal of Fish Biology, 99 (4):1328–1340, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Higaki:2021:GCS

[669] Shogo Higaki, Tomomi Nishie, Takaaki Todo, Reiko Teshima, Kenichiro Kusumi, Risa Mitsumori, Ikuo Tooyama, Yasuhiro Fujioka, Toshihiro Kawasaki, Noriyoshi Sakai, and Tatsuyuki Takada. Germ cell-specific expression of Venus by Tol2-mediated transgenesis in endangered endemic cyprinid Honmorokoo (Gnathopogon caerulescens). Journal of Fish Biology, 99(4):1341–1347, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lowe:2021:GVG

[670] Jake R. Lowe, Garry R. Russ, Abner A. Bucol, Rene A. Abesamis, and John H. Choat. Geographic variability in the gonadal development and sexual ontogeny of Hemigymnus, Cheilinus and Oxycheilinus wrasses among Indo–Pacific coral reefs. Journal of Fish Biology, 99(4):1348– 1363, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hall:2021:HTC

[671] April E. Hall and Michael J. Kingsford. Habitat type and complexity drive fish assemblages in a tropical seascape. Journal of Fish Biology, 99 (4):1364–1379, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reia:2021:MAC

[672] Lais Reia, Claudio Oliveira, and Ricardo C. Benine. Moenkhausia andrica (Characiformes: Characidae): a new species from the Rio Tapajós basin, Brazil, with minute fin hooklets in females. Journal of Fish Biology, 99 (4):1380–1392, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Melnikov:2021:DSS

[673] Victor Melnikov, Falk Pollehne, Natalia Minkina, and Lidia Melnik. Distribution of Sprattus sprattus phalericus (Risso, 1827) and zooplankton near the Black Sea redoxcline. Journal of Fish Biology, 99(4):1393–1402, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neves:2021:EKL

[674] Ana Neves, Inês Sousa, Vera Sequeira, Ana Rita Vieira, Elisabete Silva, Frederica Silva, Ana Marta Duarte, Susana Mendes, Rui Ganhão, Carlos Assis, Rui Rebelo, Maria Filomena Magalhães, Maria Manuel Gil, and Leonel Serrano Gordo. Enhancing knowledge on low-value fishing species: the distinct reproductive strategy of two gurnard species. Journal of Fish Biology, 99(4):1403–1414, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stewart:2021:REF

[675] Sarah I. D. Stewart, Aaron D. Spares, José L. Varela, Nic R. McLellan, and Michael J. W. Stokesbury. Running on empty? Freshwater feeding by spawning anadromous alewife Alosa pseudoharengus. Journal of Fish Biology, 99(4):1415–1429, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Soe:2021:FHS

[676] Kay Khine Soe, Siriporn Pradit, and Sukree Hajisamae. Feeding habits and seasonal trophic guilds structuring fish community in the bay mouth region of a tropical estuarine habitat. Journal of Fish Biology, 99(4): 1430–1445, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Collins:2021:MFL

[677] Rupert A. Collins, Giulia Trauzzi, Katherine M. Maltby, Thomas I. Gibson, Frances C. Ratcliffe, Jane Hallam, Sophie Rainbird, James Maclaine, Peter A. Henderson, David W. Sims, Stefano Mariani, and Martin J. Genner. Meta-fish-lib: a generalised, dynamic DNA reference library pipeline for metabarcoding of fishes. Journal of Fish Biology, 99(4): 1446–1454, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bass:2021:RMP

[678] Nathan Charles Bass, Joanna Day, Tristan L. Guttridge, Johann Mourier, Nathan A. Knott, Catarina Vila Pouca, and Culum Brown. Residency and movement patterns of adult Port Jackson sharks (Heterodontus portusjacksoni) at a breeding aggregation site. Journal of Fish Biology, 99(4):1455–1466, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ramos:2021:PJN

[679] Telton P. A. Ramos, Silvia Y. Lustosa-Costa, Luciano de F. Barros-Neto, and José E. L. Barbosa. Parotocinclus jacksoni, a new hypoptopomatine catfish (Siluriformes: Loricariidae) from the Rio Mamanguape basin, north-eastern Brazil. Journal of Fish Biology, 99(4):1467–1475, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Godoy:2021:RED

[680] Robson S. Godoy, Vinicius Weber, Luis Esteban Krause Lanés, Martin Reichard, Tanise Gemelli, Raquel Von Hohendorff, and Leonardo Maltchik. Recognizing the enemy: do predator cues influence hatching in Neotropical annual killifish? Journal of Fish Biology, 99(4):1476–1484, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

El-Regal:2021:SPN

[681] Mohamed A. Abu El-Regal, Mohsen M. El-Sherbiny, Mohamed H. Gabr, and Ronald Fricke. Schindleria parva, a new species of Schindler’s fish (Teleostei: Schindleriidae: Schindleria) from Jeddah, Saudi Arabia, Red Sea. Journal of Fish Biology, 99(4):1485–1491, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Phillips:2021:EEL

[682] Natasha D. Phillips, Amy Garbett, Daniel Wise, Sophie L. Loca, Olivia Daly, Lawrence E. Eagling, Jonathan D. R. Houghton, Peter Verhoog, James Thorburn, and Patrick C. Collins. Evidence of egg-laying grounds for critically endangered flapper skate (Dipturus intermedius) off Orkney, UK. Journal of Fish Biology, 99(4):1492–1496, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Noda:2021:ARS

[683] Shohei Noda, Rui Ueda, Tatsuya Tanaka, Kotaro Shirai, Daisuke Kishi, and Takuya Sato. Anadromous red-spotted masu salmon (Oncorhynchus masou ishikawae), a southernmost sea-migration form of salmonid, displays low variation in both age at seaward migration and sea age. Journal of Fish Biology, 99(4):1497–1502, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bartes:2021:PES

[684] Saia Bartes and Matías Braccini. Potential expansion in the spatial distribution of subtropical and temperate west Australian sharks. Journal of Fish Biology, 99(4):1503–1506, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lacoste:2021:ULH

[685] Francisco Firpo Lacoste, Luis Fernández, and Cristina Scioscia. On an unknown lamprey holotype (Petromyzon macrostomus Burmeister 1868) and the genus Exomegas Gill 1883 from South America. Journal of Fish Biology, 99(4):1507–1512, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Halvorsen:2021:MPT

[686] Kim Tallaksen Halvorsen, Torkel Larsen, Howard I. Browman, Caroline Durif, Nicolai Aasen, Leif Asbjørn Vøllestad, Alessandro Cresci, Tonje Knutsen Sørdalen, Reidun M. Bjelland, and Anne Berit Skiftesvik. Movement patterns of temperate wrasses (Labridae) within a small marine protected area. Journal of Fish Biology, 99(4):1513–1518, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Novacovsky:2021:EWM

[687] Gabriela N. Novacovsky, María G. Palacios, and María C. Sueiro. Epitheliocystis in wild marine fishes and its relation with anthropogenic pollution. Journal of Fish Biology, 99(4):1519–1523, October 2021. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Becerril-Garcia:2021:GDD

[688] Edgar E. Becerril-García, Francesca Pancaldi, Ariel A. Cruz-Villacorta, Alma R. Rivera-Camacho, Carlos A. Aguilar-Cruz, Darren A. Whitehead, Rogelio González-Armas, Marcial Arellano-Martínez, and Felipe Galván-Magaña. General descriptions of the dermis structure of a juvenile whale shark Rhincodon typus from the Gulf of California. Journal of Fish Biology, 99(4):1524–1528, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marina:2021:SAR

[689] Giacomin Marina, Chris N. Glover, Greg G. Goss, and Alex M. Zimmer. The skin of adult rainbow trout is not a significant site of ammonia clearance from the blood. Journal of Fish Biology, 99(4):1529–1534, October 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2021:IIk

[690] Anonymous. Issue information. Journal of Fish Biology, 99(5):1535– 1536, November 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2021:SSS

[691] William Bernard Perry. Sex, smells and sickness: deceitful mate choice in Nile tilapia. Journal of Fish Biology, 99(5):1537, November 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Feitosa:2021:HUN

[692] Leonardo M. Feitosa, Aristóteles P. N. Queiroz, Maylis Labonne, Valderi L. Dressler, and Rosangela P. Lessa. Habitat use and nursery evaluation for the longnose stingray Hypanus guttatus (Bloch & Schneider, 1801) using vertebral microchemistry. Journal of Fish Biology, 99(5): 1538–1549, November 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tavera:2021:NSG

[693] José Tavera, Stephania Rojas-Vélez, and Edgardo Londoño-Cruz. A new species of the genus Acyrtus on the eastern Pacific: a cornerstone for the evolution and biogeography of the genus. Journal of Fish Biology, 99(5): 1550–1560, November 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bucol:2021:DRP

[694] Abner A. Bucol, Rene A. Abesamis, Brian L. Stockwell, Jake R. Lowe, and Garry R. Russ. Development of reproductive potential in protogynous coral reef fishes within Philippine no-take marine reserves. Journal of Fish Biology, 99(5):1561–1575, November 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martins:2021:ATI

[695] Karla Martins, Latifa Pelage, Anne K. S. Justino, Flavia Lucena Frédou, Teodoro Vaske Júnior, François Le Loc’h, and Paulo Travassos. Assessing trophic interactions between pelagic predatory fish by gut content and stable isotopes analysis around Fernando de Noronha Archipelago (Brazil), Equatorial West Atlantic. Journal of Fish Biology, 99(5):1576– 1590, November 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Belleggia:2021:ECD

[696] Mauro Belleggia, Jorge Colonello, Federico Cortés, and Daniel Enrique Figueroa. Eating catch of the day: the diet of porbeagle shark Lamna nasus (Bonnaterre 1788) based on stomach content analysis, and the interaction with trawl fisheries in the south-western Atlantic (52^{o}S-56^{o} S). Journal of Fish Biology, 99(5):1591–1601, November 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lowe:2021:RVL

[697] Jake R. Lowe, Samuel D. Payet, Hugo B. Harrison, Jean-Paul A. Hobbs, Andrew S. Hoey, Brett M. Taylor, Tane H. Sinclair-Taylor, and Morgan S. Pratchett. Regional versus latitudinal variation in the life-history traits and demographic rates of a reef fish, Centropyge bispinosa, in the Coral Sea and Great Barrier Reef Marine Parks, Australia. Journal of Fish Biology, 99(5):1602–1612, November 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lin:2021:DFS

[698] Tingting Lin, Xin Liu, and Dong Zhang. Does the female seahorse still prefer her mating partner after a period of separation? Journal of Fish Biology, 99(5):1613–1621, November 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Matheus:2021:MHF

[699] Valquiria Aparecida Matheus, Claudemir Kuhn Faccioli, Renata Alari Chedid, José Augusto Senhorini, Irene Bastos Franceschini-Vicentini, and Carlos Alberto Vicentini. Morphological and histochemical features of the digestive tract of Leiarius marmoratus (Gill, 1870). Journal of Fish Biology, 99(5):1622–1631, November 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rossi:2021:ATM

[700] Vanessa Stramantinoli Rossi, Isabela Maria de Mello, Assaf Barki, and Percilia Cardoso Giaquinto. The attraction trick: males in early stages of disease become more chemically attractive to Nile tilapia (Oreochromis niloticus) females. Journal of Fish Biology, 99(5):1632–1639, November 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zheng:2021:EFC

[701] Yu-Hui Zheng and Shi-Jian Fu. Effects of fasting on collective movement and fission–fusion dynamics in both homogeneous and heterogeneous shoals of a group-living cyprinid fish species. Journal of Fish Biology, 99(5):1640–1649, November 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rincon-Diaz:2021:FFD

[702] Martha Patricia Rincón-Díaz, Nelson D. Bovcon, Pablo D. Cochia, María Eva Góngora, and David E. Galván. Fish functional diversity as an indicator of resilience to industrial fishing in Patagonia Argentina. Journal of Fish Biology, 99(5):1650–1667, November 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lee:2021:DMD

[703] Chien-I Lee, Feng-Yu Wang, Min-Yun Liu, Tak-Kei Chou, and Te-Yu Liao. DNA metabarcoding for dietary analysis of Holland’s carp (Spinibarbus hollandi) to evaluate the threat to native fishes in Taiwan. Journal of Fish Biology, 99(5):1668–1676, November 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Miyake:2021:GDK

[704] Takuya Miyake, Jun Nakajima, Keitaro Umemura, Norio Onikura, Takayoshi Ueda, Carl Smith, and Kouichi Kawamura. Genetic diversification of the Kanehira bitterling Acheilognathus rhombeus inferred from mitochondrial DNA, with comments on the phylogenetic relationship with its sister species Acheilognathus barbatulus. Journal of Fish Biology, 99(5):1677–1695, November 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhang:2021:PPM

[705] Guoqiang Zhang, Zhenbo Mou, Wenhua Xue, and Haiping Liu. Phosphorylated protein modification analysis on normal liver and Exo-celiac liver of Glyptosternum maculatum. Journal of Fish Biology, 99(5):1696–1707, November 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jia:2021:OCA

[706] Qian Jia, Yuan Ni, Shengnan Min, Liang Ming, Yiwen Qian, Xueyan Cen, Jialian Wang, and Xuehong Tong. The ontogenesis of catabolic abilities and energy metabolism during endogenous nutritional periods of tongue sole, Cynoglossus semilaevis. Journal of Fish Biology, 99(5): 1708–1718, November 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Quirino:2021:GMD

[707] Patricia Postingel Quirino, Diógenes Henrique de Siqueira-Silva, Maira da Silva Rodrigues, Amanda Pereira dos Santos-Silva, Maria Luiza Ribeiro Delgado, José Augusto Senhorini, Alexandre Ninhaus-Silveira, and Rosicleire Veríssimo-Silveira. Gonadal morphology and difference in reproductive development of two isolated populations of Astyanax rivularis (Teleostei, Characidae). Journal of Fish Biology, 99(5):1719–1728, November 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Seamone:2021:ORS

[708] Scott G. Seamone and Douglas A. Syme. The ocellate river stingray (Potamotrygon motoro) exploits vortices of sediment to bury into the substrate. Journal of Fish Biology, 99(5):1729–1734, November 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ayres:2021:UUA

[709] Kathryn A. Ayres, James T. Ketchum, Rogelio González-Armas, Felipe Galván-Magaña, Alex Hearn, Fernando R. Elorriaga-Verplancken, Edgar M. Hoyos-Padilla, and Stephen M. Kajiura. The use of an unoccupied aerial vehicle to survey shark species over sand and rocky-reef habitats in a marine protected area. Journal of Fish Biology, 99(5): 1735–1740, November 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henschel:2021:REM

[710] Elisabeth Henschel and Nathan K. Lujan. Range extension of the miniature pencil-catfish Potamoglanis wapixana (Siluriformes: Trichomycteridae) into the Essequibo River basin, Guyana. Journal of Fish Biology, 99(5):1741–1745, November 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zheng:2021:IFE

[711] Yan jing Zheng, Xiu qin Li, Zhen xiong Yang, Wei xu Cai, Quan sheng Lou, and Wei Tao. The identification of fish eggs of two species, the ovate sole Solea ovata and black porgy Acanthopagrus schlegelii. Journal of Fish Biology, 99(5):1746–1751, November 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harasti:2021:GOE

[712] David Harasti. Getting old: an endangered seahorse (Hippocampus whitei) lives for up to 7 years in the wild. Journal of Fish Biology, 99(5):1752–1754, November 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhou:2021:GEP

[713] Min Zhou, Ziwei Zhao, Jinliang Zhao, Minglin Wu, and Xiaowu Chen. Gene expression profiling of DNA methyltransferase genes in Siniperca chuatsi based on transcriptome sequencing. Journal of Fish Biology, 99(5):1755–1760, November 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bielsa:2021:GVU

[714] Nuria Pérez Bielsa, Judith Ollé, David Macías, Sámar Saber, and Jordi Viñas. Genetic validation of the unexpected presence of a tropical tuna, bigeye tuna (Thunnus obesus), in the Mediterranean. Journal of Fish Biology, 99(5):1761–1764, November 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Takahashi:2021:NLG

[715] Tetsumi Takahashi. A new linkage group involved in sex determination in haplotilapiine cichlids. Journal of Fish Biology, 99(5):1765–1768, November 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bessudo:2021:EMF

[716] Sandra Bessudo, Felipe Ladino, Edgar Eduardo Becerril-García, Charles M. Shepard, Pelayo Salinas-De-León, and Edgar Mauricio Hoyos-Padilla. The elasmobranchs of Malpelo Flora and Fauna Sanctuary, Colombia. Journal of Fish Biology, 99(5):1769–1774, November 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2021:IIl

[717] Anonymous. Issue information. Journal of Fish Biology, 99(6):1775– 1776, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2021:TCS

[718] William Bernard Perry. Turbo charged salmon and stream ecosystem function: the plight of aquaculture escapees. Journal of Fish Biology, 99(6):1777, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Uribe:2021:CBC

[719] Victoria Uribe, Paul S. Wills, Jon M. Shenker, Aaron J. Adams, and Sahar Mejri. A comprehensive biochemical characterization of settlement stage leptocephalus larvae of bonefish (Albula vulpes). Journal of Fish Biology, 99(6):1778–1785, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2021:GWS

[720] Zhangfan Chen, Xiwen Xu, Jie Wang, Qian Zhou, and Songlin Chen. A genome-wide survey of NOD-like receptors in Chinese tongue sole (Cynoglossus semilaevis): Identification, characterization and expression analysis in response to bacterial infection. Journal of Fish Biology, 99(6): 1786–1797, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liu:2021:MMP

[721] Xiaoxia Liu, Yongjie Wang, Zhe Liu, Yujun Kang, Fang Ma, Zhicheng Luo, Jianfu Wang, and Jinqiang Huang. miR-434 and miR-242 have a potential role in heat stress response in rainbow trout (Oncorhynchus mykiss). Journal of Fish Biology, 99(6):1798–1803, December 2021. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rioual:2021:EDE

[722] Fanny Rioual, Claudia Ofelio, Maryandrea Rosado-Salazar, Jhon Dionicio-Acedo, Myron A. Peck, and Arturo Aguirre-Velarde. Embryonic development and effect of temperature on larval growth of the Peruvian anchovy Engraulis ringens. Journal of Fish Biology, 99(6):1804–1821, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tazunoki:2021:LHE

[723] Yuhei Tazunoki, Kasumi Akashi, Sumire Haramoto, Akihito Kita, Yukari Mochioka, Hiroki Matsuda, Kazuki Ohta, and Makoto Tokuda. Life history of the endangered Japanese striped loach, Cobitis kaibarai (Cypriniformes: Cobitidae), with special reference to its reproductive ecology and the influence of creek reshapings on its population density. Journal of Fish Biology, 99(6):1822–1831, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Temperoni:2021:ODS

[724] Brenda Temperoni, Santiago Aldo Barbini, Paula Orlando, and Claudio César Buratti. Ontogenetic diet shifts of rough scad Trachurus lathami in the North Patagonian shelf (south-west Atlantic Ocean). Journal of Fish Biology, 99(6):1832–1842, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yang:2021:MIG

[725] Liping Yang, Shaoyang Zhi, Guokun Yang, Chaobin Qin, Xiao Yan, Mingming Niu, Wenlei Zhang, Mingyu Liu, Mengjuan Zhao, and Guoxing Nie. Molecular identification of glucose transporter 4: the responsiveness to starvation, glucose, insulin and glucagon on glucose transporter 4 in common carp (Cyprinus carpio L.). Journal of Fish Biology, 99(6): 1843–1856, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Radchenko:2021:SMC

[726] Olga A. Radchenko, Irina N. Moreva, and Anna V. Petrovskaya. The subfamily Myoxocephalinae of cottid fishes (Cottidae): Genetic divergence and phylogenetic relationships. Journal of Fish Biology, 99(6): 1857–1868, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Boddington:2021:AGR

[727] Dion K. Boddington, Corey B. Wakefield, Emily A. Fisher, David V. Fairclough, Euan S. Harvey, and Stephen J. Newman. Age, growth and reproductive life-history characteristics infer a high population productivity for the sustainably fished protogynous hermaphroditic yellowspotted rockcod (Epinephelus areolatus) in north-western Australia. Journal of Fish Biology, 99(6):1869–1886, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhang:2021:MIP

[728] Yan-Zhen Zhang, Qing-Dian Han, Long-Wei Fu, Yong-Xing Wang, Zhi-Hai Sui, and Yun-Guo Liu. Molecular identification and phylogenetic analysis of fungal pathogens isolated from diseased fish in Xinjiang, China. Journal of Fish Biology, 99(6):1887–1898, December 2021. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gonzalez:2021:PPG

[729] Cindy Gonzalez, Bautisse Postaire, Rodrigo R. Domingues, Kevin A. Feldheim, Susana Caballero, and Demian Chapman. Phylogeography and population genetics of the cryptic bonnethead shark Sphyrna aff. tiburo in Brazil and the Caribbean inferred from mtDNA markers. Journal of Fish Biology, 99(6):1899–1911, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bugg:2021:EPT

[730] William S. Bugg, Gwangseok R. Yoon, Catherine Brandt, Madison L. Earhart, W. Gary Anderson, and Ken M. Jeffries. The effects of population and thermal acclimation on the growth, condition and cold responsive mRNA expression of age-0 lake sturgeon (Acipenser fulvescens). Journal of Fish Biology, 99(6):1912–1927, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Franklin:2021:CPE

[731] Paul A. Franklin, Cindy F. Baker, and Kathryn A. Reeve. A comparison of passage efficiency for native and exotic fish species over an artificial baffled ramp. Journal of Fish Biology, 99(6):1928–1939, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Friday:2021:EEC

[732] Mike Friday and Tim Haxton. Evaluating the effects of controlled flows on historical spawning site access, reproduction and recruitment of lake sturgeon Acipenser fulvescens. Journal of Fish Biology, 99(6):1940–1957, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gonzalez-Silvera:2021:IDM

[733] Daniel González-Silvera, Alberto Cuesta, and Maria Ángeles Esteban. Immune defence mechanisms presented in liver homogenates and bile of gilthead seabream (Sparus aurata). Journal of Fish Biology, 99(6): 1958–1967, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turan:2021:TNG

[734] Davut Turan, Fahrettin Küçük, Salim Serkan Güçlü, and I.smail Aksu. Turcichondrostoma, a new genus for the Leuciscidae (Teleostei: Cypriniformes) from Southwestern Anatolia. Journal of Fish Biology, 99(6): 1968–1977, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cucherousset:2021:GES

[735] Julien Cucherousset, Line E. Sundt-Hansen, Mathieu Buoro, Libor Závorka, Rémy Lassus, Knut A. E. Bækkelie, Ian A. Fleming, Björn Thrandur Björnsson, Jörgen I. Johnsson, and Kjetil Hindar. Growth-enhanced salmon modify stream ecosystem functioning. Journal of Fish Biology, 99(6):1978–1989, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henschel:2021:NCG

[736] Elisabeth Henschel, Axel M. Katz, and Wilson J. E. M. Costa. A new candiru of the genus Paracanthopoma (Siluriformes: Trichomycteridae) from the Araguaia River basin, central Brazil. Journal of Fish Biology, 99(6):1990–1997, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huang:2021:SIP

[737] Shaojun Huang, Ruonan Jia, Ruiqin Hu, Wanying Zhai, Shouwen Jiang, Wenhao Li, Faxiang Wang, and Qianghua Xu. Specific immunity proteomic profile of the skin mucus of Antarctic fish Chionodraco hamatus and Notothenia coriiceps. Journal of Fish Biology, 99(6):1998–2007, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Deng:2021:ETD

[738] Yuxuan Deng, Cuixia Feng, Qingjie Yin, Xiaoqing Liu, Haoran Shi, Yao Yang, and Yufeng Wu. Effect of total dissolved gas supersaturation on the survival of common carp (Cyprinus carpio) and silver carp (Hypophthalmichthys molitrix). Journal of Fish Biology, 99(6):2008–2017, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Landeira-Dabarca:2021:CMT

[739] Andrea Landeira-Dabarca, Cristina S. R. Abreu, Maruxa Álvarez, and Pilar Molist. Changes in marine turbot (Scophthalmus maximus) epidermis and skin mucus composition during development from bilateral larvae to juvenile flat fish. Journal of Fish Biology, 99(6):2018–2029, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anderson:2021:SHP

[740] Antônio B. Anderson, Thiago M. J. Fiuza, Gabriel S. Araujo, Angela M. Canterle, Luiza M. C. Canto, Renato H. A. Freitas, Otto B. F. Gadig, and Sergio R. Floeter. A safe haven for potential reproductive aggregations of the critically endangered Brazilian guitarfish (Pseudobatos horkelii). Journal of Fish Biology, 99(6):2030–2034, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chrysafi:2021:EPI

[741] Anna Chrysafi, Niels Jepsen, Diego del Villar-Guerra, Martin H. Larsen, and Christian Skov. Effects of passive integrated transponder tags on short-term feeding patterns in European perch (Perca fluviatilis). Journal of Fish Biology, 99(6):2035–2039, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fischer:2021:BRF

[742] Abraham J. Fischer and William A. Dew. Behavioural responses of fathead minnows to carbohydrates found in aquatic environments. Journal of Fish Biology, 99(6):2040–2043, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sharifuzzaman:2021:ACP

[743] S. M. Sharifuzzaman, Istiuk Ahammed Rubby, Kazi Ahsan Habib, Seishi Kimura, Md. Harun Rasid, Md. Jayedul Islam, Amit Kumer Neogi, and M. Shahadat Hossain. Annotated checklist of ponyfishes (Perciformes: Leiognathidae) from Bangladesh, the northern Bay of Bengal. Journal of Fish Biology, 99(6):2044–2051, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nakamura:2021:BRW

[744] Itsumi Nakamura, Hiroyuki Tsuchida, Kazuki Tone, Kazuyoshi Komeyama, Akira Sasaki, and Ryo Kawabe. Behavioural response of a whale shark during the passage of a typhoon. Journal of Fish Biology, 99(6):2052–2055, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oliveira:2021:LDF

[745] Lucas S. Oliveira, Ruineris A. Cajado, Fabíola K. S. Silva, Andréa Bialetzki, and Diego M. Zacardi. Larval development of the freshwater croaker Pachypops fourcroi (La Cépède 1802) (Perciformes: Sciaenidae). Journal of Fish Biology, 99(6):2056–2059, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morgan:2021:NDS

[746] David L. Morgan, Karissa O. Lear, Brendan C. Ebner, and Stephen J. Beatty. Net design for selective control of the “plague minnow” Gambusia holbrooki that minimises impact on native Australian fishes. Journal of Fish Biology, 99(6):2060–2065, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

VanLeeuwen:2021:ISL

[747] Travis E. Van Leeuwen, David Cote, Christina Pretty, Joseph Townley, Rebecca Poole, Brian Dempson, Tomas J. Bird, Charlene Kippenhuck, and Corey Morris. Incursions of sea lamprey, Petromyzon marinus, and striped bass, Morone saxatilis, in Labrador waters: Episodic events or evidence of a northward range expansion? Journal of Fish Biology, 99(6): 2066–2070, December 2021. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2022:IIa

[748] Anonymous. Issue information. Journal of Fish Biology, 100(1):1–2, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2022:DDB

[749] William Bernard Perry. The devastating duo: bycatch and climate change hit the Atlantic wolffish Anarhichas lupus. Journal of Fish Biology, 100(1):3, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burdett:2022:TTG

[750] Heidi L. Burdett, Ingrid Kelling, and Marylyn Carrigan. #TimesUp: tackling gender inequities in marine and fisheries science. Journal of Fish Biology, 100(1):4–9, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Delomas:2022:AVR

[751] Thomas A. Delomas and Konrad Dabrowski. Asymmetric viability in reciprocal crosses of zebrafish Danio rerio and pearl danio Danio albolineatus. Journal of Fish Biology, 100(1):10–14, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liu:2022:GWI

[752] Fujiang Liu, Li Zhou, Jing Zhang, Yilei Wang, Zhiyong Wang, Xiande Liu, and Mingyi Cai. Genome-wide identification and transcriptomebased expression profiling of the Sox gene family in the spinyhead croaker (Collichthys lucidus). Journal of Fish Biology, 100(1):15–24, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Faustino-Fuster:2022:NSC

[753] Dario R. Faustino-Fuster and Lesley S. de Souza. A new species of Cetopsorhamdia (Siluriformes: Heptapteridae) from the Upper Amazon River basin. Journal of Fish Biology, 100(1):25–39, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hayashi:2022:LHA

[754] Kina Hayashi, Tetsuo Kuwamura, Katsunori Tachihara, and James Davis Reimer. Large host anemones can be shelters of a diverse assemblage of fish species, not just anemonefish. Journal of Fish Biology, 100(1):40– 50, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fetherman:2022:HAR

[755] Eric R. Fetherman and Brian W. Avila. Habitat associations of rainbow trout Oncorhynchus mykiss and brown trout Salmo trutta fry. Journal of Fish Biology, 100(1):51–61, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fujiwara:2022:RLL

[756] Kyoji Fujiwara and Hiroyuki Motomura. Review of the Lepadichthys lineatus complex (Gobiesocidae: Diademichthyinae) with descriptions of three new species. Journal of Fish Biology, 100(1):62–81, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Awata:2022:FRE

[757] Satoshi Awata, Takeshi Ito, Karen D. Crow, Yasunori Koya, and Hiroyuki Munehara. The first record of egg masses in tunicates deposited by the snubnose sculpin, Orthonopias triacis, from the Northeastern Pacific: evidence for convergent evolution of an unusual reproductive strategy. Journal of Fish Biology, 100(1):82–91, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Walter:2022:CIR

[758] Ryan P. Walter, Clare J. Venney, Nicholas E. Mandrak, and Daniel D. Heath. Conservation implications of revised genetic structure resulting from new population discovery: the threatened eastern sand darter (Ammocrypta pellucida) in Canada. Journal of Fish Biology, 100(1):92– 98, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Doherty:2022:ERB

[759] Claire L. J. Doherty, Aaron T. Fisk, Steven J. Cooke, Trevor E. Pitcher, and Graham D. Raby. Exploring relationships between oxygen consumption and biologger-derived estimates of heart rate in two warmwater piscivores. Journal of Fish Biology, 100(1):99–106, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bowen:2022:DAB

[760] Stephen H. Bowen. Digestion and assimilation of benthic biofilm by the Sábalo, Prochilodus lineatus. Journal of Fish Biology, 100(1):107–116, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Corgos:2022:NHU

[761] Antonio Corgos and Alejandro Rosende-Pereiro. Nursery habitat use patterns of the scalloped hammerhead shark, Sphyrna lewini, in coastal areas of the central Mexican Pacific. Journal of Fish Biology, 100(1):117– 133, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Klein:2022:SFS

[762] Juliana D. Klein, Tamaryn A. Asbury, Charlene da Silva, Kelvin L. Hull, Matthew L. Dicken, Enrico Gennari, Simo N. Maduna, and Aletta E. Bester van der Merwe. Site fidelity and shallow genetic structure in the common smooth-hound shark Mustelus mustelus confirmed by tagrecapture and genetic data. Journal of Fish Biology, 100(1):134–149, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Friedman:2022:DPM

[763] Ivana S. Friedman, Leonel A. Behrens, Nair de los Angeles Pereira, Edgardo M. Contreras, and Analia V. Fernández-Gimenez. Digestive proteinases from the marine fish processing wastes of the South-West Atlantic Ocean: Their partial characterization and comparison. Journal of Fish Biology, 100(1):150–160, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henschel:2022:OFR

[764] Elisabeth Henschel, Maxwell J. Bernt, Jonathan N. Baskin, Robert E. Schmidt, and Nathan K. Lujan. Osteology-focused redescription and description of the blood-feeding candirus Paracanthopoma parva and Paravandellia alleynei sp.n. (Trichomycteridae: Vandelliinae). Journal of Fish Biology, 100(1):161–174, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Allan:2022:SUE

[765] Sam J. Allan, Max J. O’Connell, David Harasti, O. Selma Klanten, and David J. Booth. Space use by the endemic common (weedy) seadragon (Phyllopteryx taeniolatus): influence of habitat and prey. Journal of Fish Biology, 100(1):175–183, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Griffioen:2022:DTT

[766] Arie Benjamin Griffioen, Woody Janssen, Timon Menke, Tony Wilkes, and Hendrik Volken Winter. Does tagging transparent fish increase predation risk? A laboratory study with glass eel (Anguilla anguilla) and sea bass (Dicentrarchus labrax). Journal of Fish Biology, 100(1):184– 191, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhao:2022:IFR

[767] Minghui Zhao, Yazhen Liu, Yajun Gao, Xinyan Wang, Hong Zhou, and Anying Zhang. Insights into the functional role of grass carp IL-8 in head kidney leukocytes: pro-inflammatory effects and signalling mechanisms. Journal of Fish Biology, 100(1):192–202, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Valls:2022:DFB

[768] Maria Valls, Sámar Saber, María José Gómez, Patricia Reglero, and Salvatore Mele. Diet and feeding behaviour of albacore Thunnus alalunga in the western Mediterranean Sea during the spawning period. Journal of Fish Biology, 100(1):203–217, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Paulsen:2022:GIP

[769] Tora Paulsen, Odd Terje Sandlund, Gunnel Østborg, Eva B. Thorstad, Peder Fiske, Rune Muladal, and Stig Tronstad. Growth of invasive pink salmon (Oncorhynchus gorbuscha) at sea assessed by scale analysis. Journal of Fish Biology, 100(1):218–228, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roberts:2022:SSI

[770] Karling N. Roberts, Taylor Lund, Brian Hayden, and Mark Poesch. Season and species influence stable isotope ratios between lethally and nonlethally sampled tissues in freshwater fish. Journal of Fish Biology, 100 (1):229–241, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rawat:2022:HSM

[771] Prashant Rawat, Peng Liu, Chaohui Zhang, Shuaicheng Guo, Laith A. Jawad, Zahra Sadighzadeh, and Deju Zhu. Hierarchical structure and mechanical properties of fish scales from Lutjanidae with different habitat depths. Journal of Fish Biology, 100(1):242–252, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bluemel:2022:DAW

[772] Joanna K. Bluemel, Simon H. Fischer, David W. Kulka, Christopher P. Lynam, and Jim R. Ellis. Decline in Atlantic wolffish Anarhichas lupus in the North Sea: Impacts of fishing pressure and climate change. Journal of Fish Biology, 100(1):253–267, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Singh:2022:ITB

[773] Mahender Singh, Thazhathe K. Teena Jayakumar, Thipramalai Thangappan A. Kumar, Sanjeev Murali, Akhilesh Mishra, Achal Singh, and Kuldeep K. Lal. Integrative taxonomy–based discovery of Dussumieria modakandai sp. nov. from India. Journal of Fish Biology, 100 (1):268–278, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

dosReis:2022:NSC

[774] Renan B. dos Reis, Juliano Ferrer, and Weferson J. da Graça. A new species of Cambeva (Siluriformes, Trichomycteridae) from the Rio Iguaçu basin, Paraná State, Brazil and redescription of Cambeva stawiarski (Miranda Ribeiro 1968). Journal of Fish Biology, 100(1):279–299, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Msukwa:2022:OFE

[775] Amulike V. Msukwa, Ian G. Cowx, and Jonathan P. Harvey. Ornamental fish export trade in Malawi. Journal of Fish Biology, 100(1):300–314, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moreira:2022:GML

[776] Inês Moreira, Ivone Figueiredo, Inês Farias, Neide Lagarto, Catarina Maia, Joana Robalo, and Teresa Moura. Growth and maturity of the lesser-spotted dogfish Scyliorhinus canicula (Linnaeus, 1758) in the southern Portuguese continental coast. Journal of Fish Biology, 100(1): 315–319, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reismann:2022:IRF

[777] Theresa Reismann and Jens Frankowski. Impaired reproductive fitness despite high fecundity in European eel (Anguilla anguilla L.) from a Baltic Sea drainage area. Journal of Fish Biology, 100(1):320–324, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cuevas:2022:FER

[778] Juan M. Cuevas, Mirta L. García, Sebastián Gómez, and Gustavo E. Chiaramonte. First evidence of regional migration of the copper shark Carcharhinus brachyurus (Günther, 1870) in the Southwest Atlantic. Journal of Fish Biology, 100(1):325–328, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gajic:2022:CRR

[779] Andrej A. Gajić, Suvad Lelo, Aleksandar Joksimović, Ana Pešić, Jovana Tomanić, Hajrudin Beširović, and Branko Dragičević. Contemporary records of the rare and critically endangered angular rough shark, Oxynotus centrina (Linnaeus, 1758), from the eastern Adriatic Sea. Journal of Fish Biology, 100(1):329–334, January 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2022:IIb

[780] Anonymous. Issue information. Journal of Fish Biology, 100(2):335–336, February 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2022:LSS

[781] William Bernard Perry. Into the lion’s den: spawning success of lionfish (Pterois miles) in the Mediterranean Sea. Journal of Fish Biology, 100 (2):337–338, February 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zbinden:2022:NHA

[782] Zachery D. Zbinden. A needle in the haystack? Applying species cooccurrence frameworks with fish assemblage data to identify species associations and sharpen ecological hypotheses. Journal of Fish Biology, 100 (2):339–351, February 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wood:2022:BOF

[783] Louisa E. Wood, James Guilder, Marnie L. Brennan, Nicola J. Birland, Virginia Taleti, Nicholas Stinton, Nick G. H. Taylor, and Mark A. Thrush. Biosecurity and the ornamental fish trade: a stakeholder perspective in England. Journal of Fish Biology, 100(2):352–365, February 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sheng:2022:APC

[784] Junqing Sheng, Le Guan, Bin Sheng, Shiyu Geng, Di Wu, Beijuan Hu, Zhixiong Li, Shunlong Le, and Yijiang Hong. Analysis of pigment cell composition, pigment content, tyrosinase content and activity of three kinds of loaches Misgurnus anguillicaudatus from Poyang Lake. Journal of Fish Biology, 100(2):366–377, February 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Belleggia:2022:WLY

[785] Mauro Belleggia, Gustavo L. Alvarez-Colombo, Betina A. Santos, Martin Castelletta, and Belen Mattera. Who let the young-of-the-year (YOY) up? The pelagic habitat as a nursery and feeding ground area for the YOY Argentine hake Merluccius hubbsi Marini, 1933. Journal of Fish Biology, 100(2):378–389, February 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2022:ELL

[786] Kimberley A. Smith, Peter G. Coulson, and S. Alex Hesp. Exceptional longevity in a lightly exploited, semi-anadromous clupeid (Perth herring Nematalosa vlaminghi) within a degraded estuarine environment. Journal of Fish Biology, 100(2):390–405, February 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Berendzen:2022:MDW

[787] Peter B. Berendzen, Sam R. Holmes, Jeremy R. Abels, and Corinthia R. Black. Morphological diversity within the Ozark minnow (Notropis nubilus: Leuciscidae). Journal of Fish Biology, 100(2):406–415, February 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burbank:2022:ITC

[788] Jacob Burbank, D. Andrew R. Drake, and Michael Power. The influence of thermal cues on the reproductive phenology of Silver Shiner, Notropis photogenis. Journal of Fish Biology, 100(2):416–424, February 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Guimaraes:2022:NSH

[789] Karen L. A. Guimarães, Juan J. Rosso, Mariano González-Castro, Mendelsohn F. B. Souza, Juan M. Díaz de Astarloa, and Luís R. R. Rodrigues. A new species of Hoplias malabaricus species complex (Characiformes: Erythrinidae) from the Crepori River, Amazon basin, Brazil. Journal of Fish Biology, 100(2):425–443, February 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ressel:2022:ASM

[790] Kirsten N. Ressel, Louise Cominassi, Jon Sarrimanolis, and Kristin M. O’Brien. Aerobic scope is not maintained at low temperature and is associated with cardiac aerobic capacity in the three-spined stickleback Gasterosteus aculeatus. Journal of Fish Biology, 100(2):444–453, February 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hatakeyama:2022:DFJ

[791] Rui Hatakeyama, Ryusuke Sudo, Takashi Yatabe, Keisuke Yamano, and Kazuharu Nomura. Developmental features of Japanese eels, Anguilla japonica, from the late leptocephalus to the yellow eel stages: an early metamorphosis to the eel-like form and a prolonged transition to the juvenile. Journal of Fish Biology, 100(2):454–473, February 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hegarty:2022:GVA

[792] Anne-Marie Hegarty, John Stewart, and William Gladstone. Geographical variation in age and growth of the endemic Australian sciaenid Atractoscion atelodus. Journal of Fish Biology, 100(2):474–485, February 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ke:2022:SBS

[793] Senfan Ke, Zhiying Tu, Elsa Goerig, Junjun Tan, Bixin Cheng, Zhimin Li, and Xiaotao Shi. Swimming behaviour of silver carp (Hypophthalmichthys molitrix) in response to turbulent flow induced by a Dcylinder. Journal of Fish Biology, 100(2):486–497, February 2022. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jin:2022:SSI

[794] Yan Jin, Zun lei Liu, Xing wei Yuan, and Ya zhou Jiang. Stage-specific influence of temperature on the growth rate of Japanese Spanish mackerel (Scomberomorus niphonius) in early life. Journal of Fish Biology, 100 (2):498–506, February 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Delrieu-Trottin:2022:DRM

[795] Erwan Delrieu-Trottin, Hans Hartmann-Salvo, Pablo Saenz-Agudelo, Mauricio F. Landaeta, and Alejandro Pérez-Matus. DNA reconciles morphology and colouration in the drunk blenny genus Scartichthys (Teleostei: Blenniidae) and provides insights into their evolutionary history. Journal of Fish Biology, 100(2):507–518, February 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rodrigues-Ribeiro:2022:EDP

[796] Marcelo Rodrigues-Ribeiro, Kazuko Suzuki, and Rodrigo Silvestre Martins. Early development of pearlside Maurolicus stehmanni off southeastern Brazil. Journal of Fish Biology, 100(2):519–531, February 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roosta:2022:CSA

[797] Zahra Roosta, Bahram Falahatkar, Mirmasoud Sajjadi, Hamed Paknejad, Syaghalirwa N. M. Mandiki, and Patrick Kestemont. Comparative study on accuracy of mucosal estradiol-17β, testosterone and 11ketotestosterone, for maturity, and cutaneous vitellogenin gene expression in goldfish (Carassius auratus). Journal of Fish Biology, 100(2): 532–542, February 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Seigel:2022:APR

[798] Alex R. Seigel, Isabel G. DeVriendt, Madisen C. Strand, Ananda Shastri, and Brian D. Wisenden. Association of predation risk with a heterospecific vocalization by an anabantoid fish. Journal of Fish Biology, 100 (2):543–548, February 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gilbey:2022:CGV

[799] John Gilbey, Valeriia A. Soshnina, Alexander A. Volkov, and Daria A. Zelenina. Comparative genetic variability of pink salmon from different parts of their range: native Pacific, artificially introduced White Sea and naturally invasive Atlantic Scottish rivers. Journal of Fish Biology, 100 (2):549–560, February 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Akian:2022:FRA

[800] Dieudonné Djétouan Akian, Kouakou Yao, Eric Parmentier, Frédéric Clota, Jean-François Baroiller, and Marie-Laure Bégout. Familiarity reduces aggression but does not modify acoustic communication in pairs of Nile tilapia (Oreochromis niloticus) and black-chinned tilapia (Sarotherodon melanotheron). Journal of Fish Biology, 100(2):561–573, February 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mouchlianitis:2022:RDI

[801] Foivos A. Mouchlianitis, Georgia Kalaitzi, Periklis Kleitou, Ioannis Savva, Demetris Kletou, and Kostas Ganias. Reproductive dynamics of the invasive lionfish (Pterois miles) in the Eastern Mediterranean Sea. Journal of Fish Biology, 100(2):574–581, February 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bass:2022:POM

[802] Nathan Charles Bass, Joanna Day, Tristan L. Guttridge, Nathan A. Knott, and Culum Brown. Preliminary observations on the movement ecology of a crested horn shark (Heterodontus galeatus). Journal of Fish Biology, 100(2):582–585, February 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic). See erratum [942].

Hoving:2022:POM

[803] Henk-Jan T. Hoving and Rui Freitas. Pelagic observations of the midwater scorpionfish Ectreposebastes imus (Setarchidae) suggests a role in trophic coupling between deep-sea habitats. Journal of Fish Biology, 100 (2):586–589, February 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rangel:2022:EHF

[804] Bianca S. Rangel, Roberta Viegas, Vanessa B. Bettcher, and Ricardo C. Garla. Eye healing in a free-ranging whitespotted eagle ray (Aetobatus narinari) following shark-inflicted bite injuries. Journal of Fish Biology, 100(2):590–593, February 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bonhomme:2022:SEC

[805] François Bonhomme, Laura Meyer, Christine Arbiol, Daniela Bănaru, Lilia Bahri-Sfar, Karima Fadhlaoui-Zid, Petr Strelkov, Marco Arculeo, Laurent Soulier, Jean-Pierre Quignard, and Pierre-Alexandre Gagnaire. Systematics of European coastal anchovies (genus Engraulis Cuvier). Journal of Fish Biology, 100(2):594–600, February 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Astruch:2022:DSM

[806] Patrick Astruch, Bruno Belloni, Elodie Rouanet, Thomas Schohn, Mireille Harmelin-Vivien, Jean-Georges Harmelin, and Charles-François Boudouresque. Dominance of Scorpaena maderensis among scorpaenids of littoral rocky reefs in Corsica (NW Mediterranean): Further evidence of Mediterranean Sea warming. Journal of Fish Biology, 100(2):601–604, February 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2022:IIc

[807] Anonymous. Issue information. Journal of Fish Biology, 100(3):605– 606, March 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2022:HIY

[808] William Bernard Perry. Holding it in: yet another consequence of our fragmented rivers. Journal of Fish Biology, 100(3):607–608, March 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Drechsel:2022:AEC

[809] Victoria Drechsel, Gabriel Schneebauer, Birgit Fiechtner, Christopher P. Cutler, and Bernd Pelster. Aquaporin expression and cholesterol content in eel swimbladder tissue. Journal of Fish Biology, 100(3):609–618, March 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Asuquo:2022:MDE

[810] Philomena E. Asuquo and Honor T. Ifon. Morphological discreteness of the estuarine croaker Pseudotolithus elongatus (Teleostei: Sciaenidae). Journal of Fish Biology, 100(3):619–624, March 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jaureguizar:2022:GEY

[811] Andrés Javier Jaureguizar, Federico Cortés, J. Matias Braccini, Rodrigo Wiff, and Andrés C. Milessi. Growth estimates of young-of-the-year broadnose sevengill shark, Notorynchus cepedianus, a top predator with poorly calcified vertebrae. Journal of Fish Biology, 100(3):625–631, March 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Filous:2022:LHL

[812] Alexander Filous, Charles Daxboeck, Teva Beguet, and Chuck Cook. The life history of longnose emperors (Lethrinus olivaceus) and a data-limited assessment of their stock to support fisheries management at Rangiroa Atoll, French Polynesia. Journal of Fish Biology, 100(3):632–644, March 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Albert:2022:IBR

[813] Luana Albert, Frédéric Olivier, Aurélie Jolivet, Laurent Chauvaud, and Sylvain Chauvaud. Insights into the behavioural responses of juvenile thornback ray Raja clavata to alternating and direct current magnetic fields. Journal of Fish Biology, 100(3):645–659, March 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jac:2022:TSO

[814] Romaric Jac, Hannes Höffle, Jon Albretsen, Klara Jakobsdóttir, Arved Staby, Guldborg Søvik, and Claudia Junge. Of three sharks and one chimaera: varied habitat preferences across a latitudinal range revealed by coastal and offshore surveys. Journal of Fish Biology, 100(3):660– 674, March 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ritz:2022:VED

[815] Thornton A. Ritz and Jill B. K. Leonard. Variability in early development and growth of Great Lakes burbot: developmental patterns in cultured and wild-caught larval burbot. Journal of Fish Biology, 100 (3):675–686, March 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hvas:2022:IPP

[816] Malthe Hvas. Influence of photoperiod and protocol length on metabolic rate traits in ballan wrasse Labrus bergylta. Journal of Fish Biology, 100 (3):687–696, March 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:2022:IPJ

[817] Richard J. Kennedy, Robert Rosell, and Michelle Allen. Investigating the phenology of juvenile potamodromous brown trout (Salmo trutta L.) in two large lake catchments. Journal of Fish Biology, 100(3):697– 704, March 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Swarr:2022:TRE

[818] Tyler R. Swarr, Christopher A. Myrick, and Ryan M. Fitzpatrick. Tag retention in and effects of passive integrated transponder tagging on survival and swimming performance of a small-bodied darter. Journal of Fish Biology, 100(3):705–714, March 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Twardek:2022:HER

[819] William M. Twardek, Nicolas W. R. Lapointe, and Steven J. Cooke. High egg retention in Chinook salmon Oncorhynchus tshawytscha carcasses sampled downstream of a migratory barrier. Journal of Fish Biology, 100(3):715–726, March 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic). See correction [1141].

Kenzaki:2022:OFB

[820] Akira Kenzaki, Suguru Okunishi, Tsutomu Tomoda, Yuuya Shioura, Motoharu Uchida, Naoaki Tezuka, and Hiroto Maeda. Observation of the feeding behaviour of reared Japanese eel Anguilla japonica leptocephali fed picocyanobacteria Synechococcus spp. Journal of Fish Biology, 100 (3):727–737, March 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tveit:2022:AMS

[821] Guro M. Tveit, Neil Anders, Morten S. Bondø, John R. Mathiassen, and Mike Breen. Atlantic mackerel (Scomber scombrus) change skin colour in response to crowding stress. Journal of Fish Biology, 100(3):738– 747, March 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nobata:2022:SUM

[822] Shigenori Nobata, Katsufumi Sato, Shouji Houki, Motohiro Ito, Yoshinori Aoki, Takashi Kitagawa, and Susumu Hyodo. Straightforward upriver migration to spawning sites by chum salmon Oncorhynchus keta homing to coastal short rivers in the Sanriku region. Journal of Fish Biology, 100(3):748–757, March 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Matsunuma:2022:TNS

[823] Mizuki Matsunuma and Hiroyuki Motomura. Two new species of the butterflyfish genus Roa (Perciformes: Chaetodontidae). Journal of Fish Biology, 100(3):758–774, March 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kodeeswaran:2022:NSC

[824] Paramasivam Kodeeswaran, Anil Mohapatra, Arulraj Dhinakaran, Thipramalai Thangappan Pillai Ajith Kumar, and Kuldeep Kumar Lal. A new species of the congrid eel genus Ariosoma (Anguilliformes: Congridae) from the south-west coast of India. Journal of Fish Biology, 100 (3):775–782, March 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lawson:2022:FFR

[825] Christopher L. Lawson, Christine L. Dudgeon, Anthony J. Richardson, Matt K. Broadhurst, and Michael B. Bennett. Flexibility for fuelling reproduction in a pelagic ray (Mobula eregoodoo) suggested by bioenergetic modelling. Journal of Fish Biology, 100(3):783–792, March 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Girard:2022:MPT

[826] Matthew G. Girard, Matthew P. Davis, Carole C. Baldwin, Agnès Dettaï, Rene P. Martin, and W. Leo Smith. Molecular phylogeny of the threadfin fishes (Polynemidae) using ultraconserved elements. Journal of Fish Biology, 100(3):793–810, March 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mayer:2022:SBP

[827] Guilherme Burg Mayer, Elisa Cristina Santos de Souza, Simon-Pierre Gilson, and Renato Hajenius Aché de Freitas. South Brazil pre-colonial sharks: Insights into biodiversity and species distributions. Journal of Fish Biology, 100(3):811–819, March 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Samourdani:2022:FDB

[828] Antigoni Samourdani and Evangelos Tzanatos. Fish distribution and behaviour with regard to the time of day and the anthropogenic structural modification of the shallow littoral. Journal of Fish Biology, 100(3):820– 830, March 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Riva-Rossi:2022:IRL

[829] Carla Riva-Rossi, Claude B. Renaud, Francisco J. Neira, Claudio Baigún, Cindy F. Baker, Pamela Quiroga, and Ian Potter. On the invalid resurrection of the lamprey genus Exomegas Gill, 1883. Journal of Fish Biology, 100(3):831–834, March 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morales:2022:NYS

[830] Naití A. Morales, Amy Rose Coghlan, Erin E. Easton, Alan M. Friedlander, James Herlan, and Carlos F. Gaymer. Now you see me: “first” records of the greater amberjack Seriola dumerili at Rapa Nui range extension or increased scientific effort? Journal of Fish Biology, 100(3): 835–842, March 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ludt:2022:FDB

[831] William B. Ludt and Todd R. Clardy. First detection of biofluorescence in a deep-sea anglerfish. Journal of Fish Biology, 100(3):843–846, March 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Boardman:2022:NLS

[832] Rose M. Boardman, Adrian C. Pinder, Adam T. Piper, Catherine Gutmann Roberts, Rosalind M. Wright, and J. Robert Britton. Non-lethal sampling for the stable isotope analysis of the critically endangered European eel Anguilla anguilla: how fin and mucus compare to dorsal muscle. Journal of Fish Biology, 100(3):847–851, March 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kubicek:2022:TSO

[833] Kole M. Kubicek, Ralf Britz, Amanda K. Pinion, Luke M. Bower, and Kevin W. Conway. Three scleral ossicles in the West African Denticle herring Denticeps clupeoides (Clupeiformes: Denticipitidae). Journal of Fish Biology, 100(3):852–855, March 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2022:IId

[834] Anonymous. Issue information. Journal of Fish Biology, 100(4):857– 858, April 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2022:WMS

[835] William Bernard Perry. Where the mysid shrimp go, dragons follow. Journal of Fish Biology, 100(4):859, April 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Adams:2022:OPE

[836] Colin E. Adams, Louise Chavarie, Jessica R. Rodger, Hannele M. Honkanen, Davide Thambithurai, and Matthew P. Newton. An opinion piece: the evolutionary and ecological consequences of changing selection pressures on marine migration in Atlantic salmon. Journal of Fish Biology, 100(4):860–867, April 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Trivedi:2022:MCI

[837] Sunil P. Trivedi, Shefalee Singh, Abha Trivedi, and Manoj Kumar. Mercuric chloride-induced oxidative stress, genotoxicity, haematological changes and histopathological alterations in fish Channa punctatus (Bloch, 1793). Journal of Fish Biology, 100(4):868–883, April 2022. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yang:2022:RCA

[838] Fan Yang, Yuhong Zeng, Wenxin Huai, Wei Zha, and Yunjiao Wan. Responses of cyprinid (Ancherythroculter nigrocauda) to flow with a semicircular cylinder patch. Journal of Fish Biology, 100(4):884–893, April 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zak:2022:SFC

[839] Jakub Žák, Koushik Roy, Iva Dyková, Jan Mráz, and Martin Reichard. Starter feed for carnivorous species as a practical replacement of bloodworms for a vertebrate model organism in ageing, the turquoise killifish Nothobranchius furzeri. Journal of Fish Biology, 100(4):894–908, April 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jonsson:2022:EIT

[840] Bror Jonsson and Larry Greenberg. Egg incubation temperature influences the population-specific outmigration rate of juvenile brown trout Salmo trutta. Journal of Fish Biology, 100(4):909–917, April 2022. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Slesinger:2022:RDE

[841] Emily Slesinger, Kiernan Bates, Mark Wuenschel, and Grace K. Saba. Regional differences in energy allocation of black sea bass (Centropristis striata) along the U.S. Northeast Shelf (36^{o}N to 42^{o}N) and throughout the spawning season. Journal of Fish Biology, 100(4):918–934, April 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Allan:2022:SSP

[842] Sam J. Allan, Max J. O’Connell, David Harasti, O. Selma Klanten, and David J. Booth. Searching for seadragons: predicting micro-habitat use for the common (weedy) seadragon (Phyllopteryx taeniolatus) based on habitat and prey. Journal of Fish Biology, 100(4):935–943, April 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Araujo:2022:RBW

[843] Priscila Rocha Vasconcelos Araújo, Maria C. Oddone, Joaquim Evêncio-Neto, and Rosangela Lessa. Reproductive biology of the whitespotted eagle ray Aetobatus narinari (Myliobatiformes) captured in the coast of Paraíba and Pernambuco, Brazil. Journal of Fish Biology, 100(4):944– 957, April 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ying:2022:APA

[844] Congping Ying, Xin Fang, Honglan Wang, Yanping Yang, Pao Xu, Kai Liu, and Guojun Yin. Anisakidae parasitism activated immune response and induced liver fibrosis in wild anadromous Coilia nasus. Journal of Fish Biology, 100(4):958–969, April 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chan:2022:TNA

[845] Alysha J. Chan, Vincent Raoult, Fabrice R. A. Jaine, Victor M. Peddemors, Matt K. Broadhurst, and Jane E. Williamson. Trophic niche of Australian cownose rays (Rhinoptera neglecta) and whitespotted eagle rays (Aetobatus ocellatus) along the east coast of Australia. Journal of Fish Biology, 100(4):970–978, April 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Parisi:2022:CSR

[846] Monique A. Parisi, Craig E. Franklin, and Rebecca L. Cramp. Can slowing the rate of water temperature decline be utilized to reduce the impacts of cold water pollution from dam releases on fish physiology and performance? Journal of Fish Biology, 100(4):979–987, April 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Giri:2022:RTL

[847] Sandip Giri, Abhra Chanda, Sourav Maity, Kunal Chakraborty, and Sugata Hazra. Role of tide and lunar phases on the migration pattern of juvenile Hilsa shad (Tenualosa ilisha) within a meso-macrotidal estuary. Journal of Fish Biology, 100(4):988–996, April 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moore:2022:ABL

[848] Bradley R. Moore. Age-based life history of Pacific longnose parrotfish Hipposcarus longiceps from New Caledonia. Journal of Fish Biology, 100 (4):997–1008, April 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harborne:2022:DFS

[849] Alastair R. Harborne, David P. Kochan, Melanie M. Esch, Robert Y. Fidler, Matthew D. Mitchell, Drew W. Butkowski, and Manuel González-Rivero. Drivers of fine-scale diurnal space use by a coral-reef mesopredatory fish. Journal of Fish Biology, 100(4):1009–1024, April 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blonder:2022:MZD

[850] Aviva F. Blonder and Keith A. Tarvin. Male zebrafish (Danio rerio) do not preferentially associate with familiar over unfamiliar conspecifics. Journal of Fish Biology, 100(4):1025–1032, April 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wallerius:2022:ACS

[851] Magnus Lovén Wallerius, Vilhem Moran, Libor Závorka, and Johan Höjesjö. Asymmetric competition over space use and territory between native brown trout (Salmo trutta) and invasive brook trout (Salvelinus fontinalis). Journal of Fish Biology, 100(4):1033–1043, April 2022. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chakona:2022:DDE

[852] Albert Chakona, Martine S. Jordaan, Domitilla C. Raimondo, Roger I. Bills, Paul H. Skelton, and Dewidine van der Colff. Diversity, distribution and extinction risk of native freshwater fishes of South Africa. Journal of Fish Biology, 100(4):1044–1061, April 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jin:2022:DSN

[853] Jinjin Jin, Chao Li, and Jun Zhao. Descriptions of six new species in White Cloud Mountain minnow Tanichthys albonubes complex (Cypriniformes: Tanichthyidae) in southern China. Journal of Fish Biology, 100 (4):1062–1087, April 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamada:2022:DMP

[854] Taihei Yamada, Hirokazu Urabe, and Futoshi Nakamura. Diel migration pattern of pink salmon fry in small streams. Journal of Fish Biology, 100 (4):1088–1092, April 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vilardo:2022:PPT

[855] Paulo J. Vilardo, Axel M. Katz, and Wilson J. E. M. Costa. Phylogenetic position of Trichomycterus astromycterus (Siluriformes: Trichomycteridae), an enigmatic trichomycterid from eastern Brazil, inferred from molecular data. Journal of Fish Biology, 100(4):1093–1096, April 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Giacomin:2022:SAR

[856] Marina Giacomin, Chris N. Glover, Greg G. Goss, and Alex M. Zimmer. The skin of adult rainbow trout is not a significant site of ammonia clearance from the blood. Journal of Fish Biology, 100(4):1097, April 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2022:IIe

[857] Anonymous. Issue information. Journal of Fish Biology, 100(5):1099– 1100, May 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2022:BHC

[858] William Bernard Perry. The bitterling’s hunt for a cozy abode and the impact of freshwater habitat degradation. Journal of Fish Biology, 100 (5):1101, May 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reid:2022:URC

[859] Connor H. Reid, Paul H. Patrick, Trina Rytwinski, Jessica J. Taylor, William G. Willmore, Bob Reesor, and Steven J. Cooke. An updated review of cold shock and cold stress in fish. Journal of Fish Biology, 100 (5):1102–1137, May 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jiang:2022:WRB

[860] Shengjiang Jiang, Peiqi Hong, and Satoshi Katayama. What is the relationship between hypoxia, water chemistry and otolith manganese content? Journal of Fish Biology, 100(5):1138–1149, May 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grilly:2022:LDM

[861] Emily Grilly, Keith Reid, and Stéphane Thanassekos. Long-distance movements of Antarctic toothfish (Dissostichus mawsoni) as inferred from tag-recapture data. Journal of Fish Biology, 100(5):1150–1157, May 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arkhipkin:2022:TRA

[862] Alexander I. Arkhipkin, Paul Brickle, Brendon Lee, Paul W. Shaw, and Niall J. McKeown. Taxonomic re-appraisal for toothfish (Dissostichus: Notothenioidea) across the Antarctic Polar Front using genomic and morphological studies. Journal of Fish Biology, 100(5):1158–1170, May 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schedel:2022:NPI

[863] Frederic D. B. Schedel, Albert Chakona, Brian L. Sidlauskas, Michael O. Popoola, Nadine Usimesa Wingi, Dirk Neumann, Emmanuel J. W. M. N. Vreven, and Ulrich K. Schliewen. New phylogenetic insights into the African catfish families Mochokidae and Austroglanididae. Journal of Fish Biology, 100(5):1171–1186, May 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Judd:2022:CIS

[864] Timothy M. Judd, Sara J. Tripp, and David P. Herzog. Cause of increased size of acipenseriform eggs infected with Polypodium hydriforme. Journal of Fish Biology, 100(5):1187–1194, May 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fujimoto:2022:REA

[865] Yasufumi Fujimoto, Hiroaki Chiba, Kentaro Shindo, Junya Kitazima, and Munehico Iwata. Reproductive ecology and adaptive host choice correlated with body size in an autumn-spawning bitterling Acheilognathus typus. Journal of Fish Biology, 100(5):1195–1204, May 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hage:2022:LRN

[866] Jan Häge, Matthew J. Hansen, Korbinian Pacher, Félicie Dhellemmes, Paolo Domenici, John F. Steffensen, Michael Breuker, Stefan Krause, Thomas B. Hildebrandt, Guido Fritsch, Pascal Bach, Philippe S. Sabarros, Paul Zaslansky, Kristin Mahlow, Maria Schauer, Johannes Müller, and Jens Krause. Lacunae rostralis: a new structure on the rostrum of sailfish Istiophorus platypterus. Journal of Fish Biology, 100(5):1205– 1213, May 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Balanov:2022:FDL

[867] Andrey A. Balanov, Irina V. Epur, Vladimir A. Shelekhov, and Sergey V. Turanov. The first description of larvae and comments on the taxonomy of Stichaeus ochriamkini Taranetz, 1935 (Perciformes: Stichaeidae). Journal of Fish Biology, 100(5):1214–1222, May 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Charitonidou:2022:BBF

[868] Katerina Charitonidou, Emmanuel Panteris, and Kostas Ganias. Balbiani body formation and cytoplasmic zonation during early oocyte development in two Clupeiform fishes. Journal of Fish Biology, 100(5):

1223–1232, May 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Samejima:2022:AGR

[869] Shota Samejima and Katsunori Tachihara. Age, growth and reproductive biology of a widespread coral reef fish, yellowfin goatfish Mulloidichthys vanicolensis (Valenciennes, 1831). Journal of Fish Biology, 100(5):1233– 1244, May 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Resende:2022:EHS

[870] Anna Carolina Resende, Diego Mauro Carneiro Pereira, Ieda Cristina Schleger, Maria Rosa Dmengeon Pedreiro de Souza, Ananda Karla Alvez Neundorf, Silvia Romão, Tatiana Herrerias, and Lucélia Donatti. Effects of heat shock on energy metabolism and antioxidant defence in a tropical fish species Psalidodon bifasciatus. Journal of Fish Biology, 100(5):1245– 1263, May 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bangura:2022:LVG

[871] Paul Bai Bangura, Katriina Tiira, Petri T. Niemelä, Jaakko Erkinaro, Petra Liljeström, Anna Toikkanen, and Craig R. Primmer. Linking vgll3 genotype and aggressive behaviour in juvenile Atlantic salmon (Salmo salar). Journal of Fish Biology, 100(5):1264–1271, May 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baer:2022:SMS

[872] Jan Baer, Christina Spiessl, and Alexander Brinker. Size matters? Species- and size-specific fish predation on recently established invasive quagga mussels Dreissena rostriformis bugensis Andrusov 1897 in a large, deep oligotrophic lake. Journal of Fish Biology, 100(5):1272–1282, May 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Senarat:2022:UEO

[873] Sinlapachai Senarat, Jes Kettratad, Supanut Pairohakul, Sumate Ampawong, Brian P. Huggins, Melissa M. Coleman, and Gen Kaneko. An update on the evolutionary origin of aglomerular kidney with structural and ultrastructural descriptions of the kidney in three fish species. Journal of Fish Biology, 100(5):1283–1298, May 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Villa-Verde:2022:SFD

[874] Leandro Villa-Verde, Mario C. C. de Pinna, Vinícius J. C. Reis, and Osvaldo T. Oyakawa. Secretive fish diversity: a new species of Listrura (Siluriformes, Trichomycteridae) from a supposedly well-known river in south-eastern Brazil. Journal of Fish Biology, 100(5):1299–1310, May 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kajiura:2022:SKC

[875] Stephen M. Kajiura, John C. Loyer, Cassandra Ruddy, and Marianne E. Porter. Swimming kinematics of the Caribbean reef shark, Carcharhinus perezi. Journal of Fish Biology, 100(5):1311–1314, May 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Torres:2022:EMP

[876] Yan Torres, Patricia Charvet, Vicente V. Faria, and Andrey L. F. de Castro. Evidence of multiple paternity for the endemic Xingu River stingray. Journal of Fish Biology, 100(5):1315–1318, May 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huang:2022:EPC

[877] Yuwei Huang, Yunbang Zhang, Jian Gao, and Xiaojuan Cao. Expression profiles of cdca9 related to ovarian development in loach (Misgurnus anguillicaudatus). Journal of Fish Biology, 100(5):1319–1322, May 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2022:IIf

[878] Anonymous. Issue information. Journal of Fish Biology, 100(6):1323– 1324, June 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2022:BFM

[879] William Bernard Perry. Better fisheries management? We’re all (inner) ears! Journal of Fish Biology, 100(6):1325–1326, June 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alfaro-Shigueto:2022:RTP

[880] Joanna Alfaro-Shigueto, Eliana Alfaro-Cordova, and Jeffrey C. Mangel. Review of threats to the Pacific seahorse Hippocampus ingens (Girard 1858) in Peru. Journal of Fish Biology, 100(6):1327–1334, June 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mohapatra:2022:NCE

[881] Anil Mohapatra, Hsuan-Ching Ho, Smrutirekha Acharya, Dipanjan Ray, and Subhrendu Sekhar Mishra. A new Congrid eel, Rhynchoconger smithi sp. nov. (Anguilliformes: Congridae), from the Bay of Bengal, India. Journal of Fish Biology, 100(6):1335–1344, June 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sawai:2022:RGE

[882] Etsuro Sawai and Marianne Nyegaard. A review of giants: Examining the species identities of the world’s heaviest extant bony fishes (ocean sunfishes, family Molidae). Journal of Fish Biology, 100(6):1345–1364, June 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wakiya:2022:ABW

[883] Ryoshiro Wakiya, Hikaru Itakura, Yuji Imayoshi, and Kenzo Kaifu. Agonistic behaviour of wild eels and depressed survival and growth of farmed eels in mixed rearing experiments. Journal of Fish Biology, 100(6):1365– 1374, June 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Datta:2022:PSR

[884] Netri Datta, Pradip K. Kar, and Samar K. Saha. Primary stress response and biochemical profile of Labeo rohita (Hamilton, 1822) experimentally parasitized with Argulus bengalensis (Ramakrishna, 1951). Journal of Fish Biology, 100(6):1375–1387, June 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Khan:2022:TMS

[885] Umar Khan. Temperature, meal size and body size effects on the gastric evacuation of rainbow trout: modelling optimum and upper thermal limits. Journal of Fish Biology, 100(6):1388–1398, June 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rendon-Herrera:2022:RVM

[886] J. Juan Rendón-Herrera, Juan Carlos Pérez-Jiménez, and Nancy C. Saavedra-Sotelo. Regional variation in multiple paternity in the brown smooth-hound shark Mustelus henlei from the northeastern Pacific. Journal of Fish Biology, 100(6):1399–1406, June 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Flerova:2022:UAF

[887] Ekaterina A. Flerova, Alexey A. Morozov, Victoria V. Yurchenko, Anastasia S. Fomina, and Victoria V. Konovalova. Unique and ancestral features in trunk kidney microanatomy and ultrastructure of omul Coregonus migratorius. Journal of Fish Biology, 100(6):1407–1418, June 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jensen:2022:ESB

[888] Arne J. Jensen, Ola H. Diserud, Bengt Finstad, Peder Fiske, and Eva B. Thorstad. Early-season brown trout (Salmo trutta) migrants grow and survive better at sea. Journal of Fish Biology, 100(6):1419–1431, June 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2022:OEM

[889] Rosie Smith, Eric Hitkolok, Tracey Loewen, Amanda Dumond, Kent Kristensen, and Heidi Swanson. Overwintering ecology and movement of anadromous Arctic char (Salvelinus alpinus) in a large, ice-covered river in the Canadian Arctic. Journal of Fish Biology, 100(6):1432–1446, June 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kodeeswaran:2022:AIS

[890] Paramasivam Kodeeswaran, Ayyathurai Kathirvelpandian, Smrutirekha Acharya, Swarup Ranjan Mohanty, Anil Mohapatra, Thipramalai Thangappan Pillai Ajith Kumar, and Kuldeep Kumar Lal. Ariosoma indicum sp. nov., a new species of congrid eel (Anguilliformes: Congridae: Bathymyrinae) from the Indian waters. Journal of Fish Biology, 100(6):1447–1454, June 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baltazar-Soares:2022:NET

[891] Miguel Baltazar-Soares, Adrian C. Pinder, Andrew J. Harrison, Will Oliver, Jessica Picken, J. Robert Britton, and Demetra Andreou. A noninvasive eDNA tool for detecting sea lamprey larvae in river sediments: Analytical validation and field testing in a low-abundance ecosystem. Journal of Fish Biology, 100(6):1455–1463, June 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tronbol:2022:TOD

[892] Frida Tronbøl, Edda Johannesen, Maud Alix, Thassya C. dos Santos Schmidt, Katerina Charitonidou, Arild Folkvord, and Olav Sigurd Kjesbu. Tracking oocyte development and the timing of skipped spawning for north-east Arctic haddock (Melanogrammus aeglefinus). Journal of Fish Biology, 100(6):1464–1474, June 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sih:2022:HRO

[893] Tiffany Lorraine Sih, Ashley John Williams, Yi Hu, and Michael John Kingsford. High-resolution otolith elemental signatures in eteline snappers from valuable deepwater tropical fisheries. Journal of Fish Biology, 100(6):1475–1496, June 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Montero-Taboada:2022:TJP

[894] Rebeca Montero-Taboada, Giovanna Sotil, Jhon Dionicio-Acedo, Maryandrea Rosado-Salazar, and Arturo Aguirre-Velarde. Tolerance of juvenile Peruvian rock seabass (Paralabrax humeralis Valenciennes, 1828) and Peruvian grunt (Anisotremus scapularis Tschudi, 1846) to low-oxygen conditions. Journal of Fish Biology, 100(6):1497–1509, June 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2022:FAY

[895] Robert C. Perry and John M. Casselman. Factors affecting year-class strength and growth of lake whitefish (Coregonus clupeaformis) in impounded lakes as revealed by otolith chronologies. Journal of Fish Biology, 100(6):1510–1527, June 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Teal:2022:SSM

[896] Chad N. Teal, D. Katharine Coykendall, Matthew R. Campbell, Daniel L. Eardley, Thomas A. Delomas, James T. Shira, Daniel J. Schill, Scott A. Bonar, and Melanie Culver. Sex-specific markers undetected in green sunfish Lepomis cyanellus using restriction-site associated DNA sequencing. Journal of Fish Biology, 100(6):1528–1540, June 2022. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pardee:2022:LHH

[897] Cassandra Pardee and John Wiley. Life history of the humpnose big-eye bream Monotaxis grandoculis. Journal of Fish Biology, 100(6):1541– 1547, June 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

OConnell:2022:RLW

[898] Max J. O’Connell, Ashley M. Fowler, Sam J. Allan, Giglia A. Beretta, and David J. Booth. Recruiting a long way from home: Domino damselfish Dascyllus trimaculatus can detect new temperate coral habitat and conspecifics. Journal of Fish Biology, 100(6):1548–1552, June 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deCarvalho:2022:CST

[899] Barbara Maichak de Carvalho, Matheus O. Freitas, Acácio R. G. Tomás, Rodrigo Caires, Patricia Charvet, and Jean Vitule. Citizen science as a tool for understanding the silent dispersion of toadfish Opsanus beta (Goode and Bean, 1880). Journal of Fish Biology, 100(6):1553–1558, June 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2022:IIg

[900] Anonymous. Issue information. Journal of Fish Biology, 101(1):1–2, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2022:HMN

[901] William Bernard Perry. Is having more neural tissue really a no-brainer? Journal of Fish Biology, 101(1):3, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brownscombe:2022:EPF

[902] Jacob W. Brownscombe, Graham D. Raby, Karen J. Murchie, Andy J. Danylchuk, and Steven J. Cooke. An energetics–performance framework for wild fishes. Journal of Fish Biology, 101(1):4–12, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smukall:2022:TFY

[903] Matthew J. Smukall, John Carlson, Steven T. Kessel, Tristan L. Guttridge, Félicie Dhellemmes, Andrew C. Seitz, and Samuel Gruber. Thirty-five years of tiger shark Galeocerdo cuvier relative abundance near Bimini, The Bahamas, and the Southeastern United States with a comparison across jurisdictional bounds. Journal of Fish Biology, 101(1): 13–25, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Czudaj:2022:LWR

[904] Stephanie Czudaj, Christian Möllmann, and Heino O. Fock. Length– weight relationships of 55 mesopelagic fishes from the eastern tropical North Atlantic: Across- and within-species variation (body shape, growth stanza, condition factor). Journal of Fish Biology, 101(1):26–41, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pazzaglia:2022:CRM

[905] Ugo E. Pazzaglia, Marcella Reguzzoni, Marco Saroglia, Renata Manconi, Guido Zarattini, and Mario Raspanti. The complex rostral morphology and the endoskeleton ossification process of two adult samples of Xiphias gladius (Xiphiidae). Journal of Fish Biology, 101(1):42–54, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Honji:2022:DOM

[906] Renato Massaaki Honji, Bruno Cavalheiro Araújo, Danilo Caneppele, Fabiana Laura Lo Nostro, and Renata Guimarães Moreira. Dynamics of ovarian maturation during the annual reproductive cycle of the endangered fish Steindachneridion parahybae (Siluriformes: Pimelodidae) in captivity. Journal of Fish Biology, 101(1):55–68, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Costa:2022:NCS

[907] Wilson J. E. M. Costa, Vinicius Abilhoa, Roger Henrique Dalcin, and Axel M. Katz. A new catfish species of the genus Cambeva (Siluriformes: Trichomycteridae) from the Rio Iguaçu drainage, southern Brazil, with a remarkable unique colour pattern. Journal of Fish Biology, 101(1): 69–76, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hubalek:2022:HDS

[908] Martin Hubálek, Vojtěch Kašpar, Tomáš Tichopád, Marek Rodina, and Martin Flajšhans. How do suboptimal temperatures affect polyploid sterlet Acipenser ruthenus during early development? Journal of Fish Biology, 101(1):77–91, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stark:2022:LEB

[909] Gavin Stark. Large and expensive brain comes with a short lifespan: the relationship between brain size and longevity among fish taxa. Journal of Fish Biology, 101(1):92–99, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Miller:2022:MHU

[910] Andrew D. Miller and Shannon K. Brewer. Movement and habitat use by smallmouth bass Micropterus dolomieu velox in a dynamic Ozark Highlands riverscape. Journal of Fish Biology, 101(1):100–114, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reid:2022:CAW

[911] Connor H. Reid, Graham D. Raby, Matthew D. Faust, Steven J. Cooke, and Christopher S. Vandergoot. Cardiac activity in walleye (Sander vitreus) during exposure to and recovery from chemical anaesthesia, electroanaesthesia and electrostunning. Journal of Fish Biology, 101(1):115– 127, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cotter:2022:LTC

[912] Deirdre Cotter, Louise Vaughan, Nigel Bond, Mary Dillane, Roxanne Duncan, Russell Poole, Gerard Rogan, and Niall Ó. Maoiléidigh. Longterm changes and effects of significant fishery closures on marine survival and biological characteristics of wild and hatchery-reared Atlantic salmon Salmo salar. Journal of Fish Biology, 101(1):128–143, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gao:2022:NTT

[913] Feng ying Gao, Xin Zhou, Mai xin Lu, Miao Wang, Zhi gang Liu, Jiang meng Cao, Xiao li Ke, and Meng meng Yi. Nile tilapia TRIM39 recruits I3K413 and I3KL45 as adaptors and is involved in the NF-κ B pathway. Journal of Fish Biology, 101(1):144–153, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liu:2022:RDL

[914] Yi Liu, Liang Cao, and E. Zhang. Re-description of the loach species Homatula variegata (Dabry de Thiersant, 1874) (Pisces: Nemacheilidae) from the middle Huang-He basin in Shaanxi Province of Central China. Journal of Fish Biology, 101(1):154–167, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wu:2022:ESR

[915] Ping Wu, Lin Chen, Jia Cheng, Yaxiong Pan, Xin Zhu, Wuying Chu, and Jianshe Zhang. Effect of starvation and refeeding on reactive oxygen species, autophagy and oxidative stress in Chinese perch (Siniperca chuatsi) muscle growth. Journal of Fish Biology, 101(1):168–178, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anderson:2022:NAH

[916] Antônio B. Anderson, Manuela B. Bernardes, Hudson T. Pinheiro, Helder C. Guabiroba, Caio R. Pimentel, Ciro C. Vilar, Luiz E. O. Gomes, Angelo F. Bernardino, Stephanie D. T. Delfino, Tommaso Giarrizzo, Carlos E. L. Ferreira, and Jean-Christophe Joyeux. Niche availability and habitat affinities of the red porgy Pagrus pagrus (Linnaeus, 1758): an important ecological player on the world’s largest rhodolith beds. Journal of Fish Biology, 101(1):179–189, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ponte:2022:DBG

[917] Isabela de Abreu Rodrigues Ponte, Jonas Eloi de Vasconcelos Filho, Caroline Vieira Feitosa, and Beatrice Padovani Ferreira. Demography of the black grouper, Mycteroperca bonaci (Poey, 1860) (Teleostei: Epinephelidae) from the North Brazil Shelf. Journal of Fish Biology, 101(1):190– 203, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2022:EDN

[918] Xiao chun Chen, Xiao quan Huang, Yi wen Tang, Lei Zhang, and Feng Lin. Effects of dietary nucleotides on growth performance, immune response, intestinal morphology and disease resistance of juvenile largemouth bass, Micropterus salmoides. Journal of Fish Biology, 101(1): 204–212, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ehemann:2022:MDR

[919] Nicolás Ehemann, Edilia Acosta-Rodríguez, Alejandro Tagliafico, Nicole Pelletier, and Guy Stevens. Manta and devil ray species occurrence and distribution in Venezuela, assessed through fishery landings and citizen science data. Journal of Fish Biology, 101(1):213–225, July 2022. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Poscai:2022:MSO

[920] Aline N. Poscai, João Paulo C. B. da Silva, André Luis S. Casas, Phillip Lenktaitis, and Otto B. F. Gadig. Morphological study of the oral denticles of the porbeagle shark Lamna nasus. Journal of Fish Biology, 101(1):226–235, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Matich:2022:DBS

[921] Philip Matich, Camryn L. Bigelow, Barrett Chambers, Jillian J. Dodds, Jessica A. Hebert, Alexis Lemieux, Christy M. Pittman, Julianna Trapp, Brooke Bianco, Carolina P. Cadena, Emily I. Castillo, Gabriela I. Castillo, Alexandra Dawdy, Alina I. Dominguez, Nicholas Dominique, Donavon R. French, Callie F. Glenn, Elena C. H. Jackson, Breidon Johnson, Gunnar Kohl, Cameron Manka, Jared K. Martin, Matthew Pappas, Audrey J. Reedholm, Kailey M. Snead, Matthew K. Tyree, and Mark Fisher. Delineation of blacktip shark (Carcharhinus limbatus) nursery habitats in the north-western Gulf of Mexico. Journal of Fish Biology, 101(1):236–248, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

John:2022:EPS

[922] Venisza Cathy John, Ajit Kumar Verma, Kishore Kumar Krishnani, Mallikarjuan Hittinahalli Chandrakant, Tincy Varghese, and Madhuri Shrikant Pathak. Effect of potassium supplementation on osmoregulatory and stress response of Pangasianodon hypophthalmus (Sauvage, 1878) with Spinacia oleracea L. in aquaponics. Journal of Fish Biology, 101(1):249–261, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liu:2022:AII

[923] Hua Liu, Ying Yang, Ying-Ying Yu, Jing-Jing Feng, Xiao-Xue Bao, Jun Zhao, and Hui Yu. Astragaloside IV improved antioxidative stress capacity and related gene expression of the Keap1–Nrf2 pathway in grass carp (Ctenopharyngodon idella) hepatocytes under heat stress. Journal of Fish Biology, 101(1):262–268, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Konishi:2022:OBN

[924] Junnosuke Konishi, Takashi Abe, Atsushi Ogihara, Daisuke Adachi, Takashi Denboh, and Hideaki Kudo. Olfactory behavioural and neural responses of planktivorous lacustrine sockeye salmon (Oncorhynchus nerka) to prey odours. Journal of Fish Biology, 101(1):269–275, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lahnsteiner:2022:SDT

[925] Franz Lahnsteiner. Seasonal differences in thermal stress susceptibility of diploid and triploid brook trout, Salvelinus fontinalis (Teleostei, Pisces). Journal of Fish Biology, 101(1):276–288, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yang:2022:BPP

[926] Qingyuan Yang, Keith Reid, and Guoping Zhu. Biological–physical processes regulate autumn prey availability of spiny icefish Chaenodraco wilsoni in the Bransfield Strait, Antarctic. Journal of Fish Biology, 101 (1):289–301, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grant:2022:MSS

[927] Michael I. Grant, Anthony W. J. Bicknell, Thaung Htut, Antt Maung, Thu Maung, Khin Myo Myo, Thu Rein, Min Khan San, William T. White, Kyaw Zay Ya, and Meira Mizrahi. Market surveys and social media provide confirmation of the endangered giant freshwater whipray Urogymnus polylepis in Myanmar. Journal of Fish Biology, 101(1):302– 307, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kuparinen:2022:GAE

[928] Anna Kuparinen, Darby Gielewski, and Jeffrey A. Hutchings. Gill area explains deviations from body size–metabolic rate relationship in teleost fishes. Journal of Fish Biology, 101(1):308–311, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jarrin:2022:ELH

[929] Jose R. Marin Jarrin, Solange Andrade-Vera, Carla Brito, Lana Ray, and Michelle A. Schuiteman. Early life history of the middling thread herring (Opisthonema medirastre) in the central Galápagos Islands. Journal of Fish Biology, 101(1):312–316, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bartlett:2022:RCT

[930] Alanah D. Bartlett, J. Michelle Lavery, Jennifer Nafziger, and Richard A. Cunjak. A rapid classification tool for deformities in Atlantic salmon (Salmo salar) embryos. Journal of Fish Biology, 101(1):317–322, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anderson:2022:OSR

[931] James M. Anderson, Patrick T. Rex, Kelsey Maloney, Mark Johnston, Deron Verbeck, Neil Allen, and Kim Holland. Observations of a speciesrecord deep dive by a central Pacific female scalloped hammerhead shark (Sphyrna lewini). Journal of Fish Biology, 101(1):323–327, July 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2022:IIh

[932] Anonymous. Issue information. Journal of Fish Biology, 101(2):329– 330, August 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coscia:2022:IHA

[933] Ilaria Coscia and Michel J. Kaiser. The impact of the human activities on aquatic ecosystems. Journal of Fish Biology, 101(2):331–332, August 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pauly:2022:WDF

[934] Daniel Pauly. Why do fish reach first maturity when they do? Journal of Fish Biology, 101(2):333–341, August 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dickey-Collas:2022:EEB

[935] Mark Dickey-Collas, Jason S. Link, Paul Snelgrove, J. Murray Roberts, M. Robin Anderson, Ellen Kenchington, Alida Bundy, Margaret M. (Peg) Brady, Rebecca L. Shuford, Howard Townsend, Anna Rindorf, Murray A. Rudd, David Johnson, and Ellen Johannesen. Exploring ecosystem-based management in the North Atlantic. Journal of Fish Biology, 101(2):342– 350, August 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thurstan:2022:PHE

[936] Ruth H. Thurstan. The potential of historical ecology to aid understanding of human–ocean interactions throughout the Anthropocene. Journal of Fish Biology, 101(2):351–364, August 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wilwert:2022:COC

[937] Elodie Wilwert, Rampal S. Etienne, Louis van de Zande, and Martine E. Maan. Contribution of opsins and chromophores to cone pigment variation across populations of Lake Victoria cichlids. Journal of Fish Biology, 101(2):365–377, August 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Simmons:2022:PHE

[938] Olivia Meredith Simmons, J. Robert Britton, Phillipa K. Gillingham, Marie Nevoux, William D. Riley, Etienne Rivot, and Stephen D. Gregory. Predicting how environmental conditions and smolt body length when entering the marine environment impact individual Atlantic salmon Salmo salar adult return rates. Journal of Fish Biology, 101(2):378–388, August 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hayden:2022:WES

[939] Brian Hayden, Chris Harrod, Stephen Thomas, and Kimmo K. Kahilainen. Winter ecology of specialist and generalist morphs of European whitefish, Coregonus lavaretus, in subarctic northern Europe. Journal of Fish Biology, 101(2):389–399, August 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2022:RLF

[940] Zhaomin Chen, Jennifer Bigman, Weiwei Xian, Cui Liang, Elaine Chu, and Daniel Pauly. The ratio of length at first maturity to maximum length across marine and freshwater fishes. Journal of Fish Biology, 101 (2):400–407, August 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Antognazza:2022:HGP

[941] Caterina Maria Antognazza, Stephen J. Sabatino, Robert J. Britton, Rob J. Hillman, Miran Aprahamian, Emilie A. Hardouin, and Demetra Andreou. Hybridization and genetic population structure of Alosa population in the United Kingdom. Journal of Fish Biology, 101(2):408– 413, August 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bass:2022:EPO

[942] Nathan Charles Bass, Joanna Day, Tristan L. Guttridge, Nathan A. Knott, and Culum Brown. Erratum: Preliminary observations on the movement ecology of a crested horn shark (Heterodontus galeatus). Journal of Fish Biology, 101(2):414, August 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic). See [802].

Anonymous:2022:IIi

[943] Anonymous. Issue information. Journal of Fish Biology, 101(3):415–416, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2022:EFD

[944] William Bernard Perry. Emptying the file drawer: the importance of publishing non-significance results in research. Journal of Fish Biology, 101(3):417–418, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Runde:2022:ESM

[945] Brendan J. Runde, Jeffrey A. Buckel, Nathan M. Bacheler, Ryan M. Tharp, Paul J. Rudershausen, Craig A. Harms, and Tal Ben-Horin. Evaluation of six methods for external attachment of electronic tags to fish: assessment of tag retention, growth and fish welfare. Journal of Fish Biology, 101(3):419–430, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gao:2022:MCC

[946] Along Gao, Liting Wu, Lan Li, Biao Han, and Jianmin Ye. Molecular cloning, characterization and expression analysis of CXCR3a and CXCR3b from Nile tilapia (Oreochromis niloticus). Journal of Fish Biology, 101(3):431–440, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Su:2022:HRN

[947] Yo Su, Hsiu-Chin Lin, and Hsuan-Ching Ho. Hoplostethus roseus, a new roughy fish from the western Pacific based on morphology and DNA barcoding (family Trachichthyidae). Journal of Fish Biology, 101(3): 441–452, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Granroth-Wilding:2022:NSA

[948] Hanna M. V. Granroth-Wilding and Ulrika Candolin. No strong associations between temperature and the host–parasite interaction in wild stickleback. Journal of Fish Biology, 101(3):453–463, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blyth:2022:ASH

[949] Samuel A. Blyth and Shannon D. Bower. After the spawn and on the hook: Sea trout Salmo trutta biophysical responses to different components of catch and release in a coastal fishery. Journal of Fish Biology, 101(3):464–477, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Xie:2022:LBN

[950] Rui-Xia Xie, Liang Cao, and E. Zhang. Liobagrus brevispina, a new species of torrent catfish (Siluriformes: Amblycipitidae) from the upper Chang-Jiang basin, South China. Journal of Fish Biology, 101(3):478– 490, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barnes:2022:SMT

[951] Lachlan M. Barnes, Charles A. Gray, and Jane E. Williamson. Spawning and maturity traits of coexisting Platycephalidae (Platycephalus caeruleopunctatus, Platycephalus grandispinis, Platycephalus richardsoni) from southeast Australia. Journal of Fish Biology, 101(3):491–504, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bektas:2022:DBS

[952] Yusuf Bektas, İsmail Aksu, Cüneyt Kaya, Esra Bayçelebi, and Davut Turan. DNA barcoding and species delimitation of the genus Oxynoemacheilus (Teleostei: Nemacheilidae) in Anatolia. Journal of Fish Biology, 101(3):505–514, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lennox:2022:ETT

[953] Robert J. Lennox, Elisabeth Stöger, Lotte S. Dahlmo, Turid Helle, Tore Wiers, Erlend M. Hanssen, and Knut Wiik Vollset. Effects of tag type and surgery on migration of Atlantic salmon (Salmo salar) smolts. Journal of Fish Biology, 101(3):515–521, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Agostinho:2022:RAC

[954] Lorena S. Agostinho, Marina V. Loeb, and Fabio Di Dario. Redescription of Anchoviella cayennensis (Puyo, 1945) (Clupeiformes: Engraulidae), with the synonymization of Anchoviella sanfranciscana Barbosa, Gomes da Silva, da Rocha Araújo & Carvalho, 2017 and remarks on Anchoviella perfasciata (Poey, 1860). Journal of Fish Biology, 101(3): 522–539, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Magson:2022:CSR

[955] Kirsty Magson, Emily Monacella, Chad Scott, Noémie Buffat, Sirachai Arunrugstichai, Metavee Chuangcharoendee, Simon J. Pierce, Jason Holmberg, and Gonzalo Araujo. Citizen science reveals the population structure and seasonal presence of whale sharks in the Gulf of Thailand. Journal of Fish Biology, 101(3):540–549, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Abecia:2022:SMB

[956] Janine E. Abecia, Osmar J. Luiz, David A. Crook, Sam C. Banks, Dion Wedd, and Alison J. King. Sex and male breeding state predict intraspecific trait variation in mouth-brooding fishes. Journal of Fish Biology, 101(3):550–559, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gul:2022:OSD

[957] Güzin Gül and Nazli Demirel. Ontogenetic shift in diet and trophic role of Raja clavata inferred by stable isotopes and stomach content analysis in the Sea of Marmara. Journal of Fish Biology, 101(3):560–572, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nie:2022:GWI

[958] Zhiwei Nie, Bo Wang, Zhixuan Zhang, Zeming Jia, Runjie Xu, Heyu Wang, Wei Zhou, and Yifu Gong. Genome-wide identification of the traf gene family in yellow catfish (Pelteobagrus fulvidraco) and analysis of their expression in response to bacterial challenge. Journal of Fish Biology, 101(3):573–583, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kondrashev:2022:PVP

[959] Sergei L. Kondrashev. Photoreceptors, visual pigments and intraretinal variability in spectral sensitivity in two species of smelts (Pisces, Osmeridae). Journal of Fish Biology, 101(3):584–596, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yang:2022:VEE

[960] Hang Yang, Zhen Xu, Sumei Tan, Chunyan Zhang, Xiaoqin Li, and Xiangjun Leng. In vitro effects of Eucommia ulmoides and its active components on the growth, lipid metabolism and collagen metabolism of grass carp (Ctenopharyngodon idellus) hepatocyte and intramuscular fibroblast. Journal of Fish Biology, 101(3):597–612, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wakiya:2022:SGF

[961] Ryoshiro Wakiya, Hikaru Itakura, Tatsumu Hirae, Tadamitsu Igari, Miyuki Manabe, Noriaki Matsuya, Katsushi Miyata, Masayuki K. Sakata, Toshifumi Minamoto, Takashi Yada, and Kenzo Kaifu. Slower growth of farmed eels stocked into rivers with higher wild eel density. Journal of Fish Biology, 101(3):613–627, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carpenter:2022:EZB

[962] Michelle Carpenter, Nakia Cullain, Stephanie Kathleen Venables, Yara Tibiriçá, Charles Griffiths, and Andrea Denise Marshall. Evidence of Závora Bay as a critical site for reef manta rays, Mobula alfredi, in southern Mozambique. Journal of Fish Biology, 101(3):628–639, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barker:2022:DEP

[963] Joanna Barker, Jake Davies, Monika Goralczyk, Surshti Patel, John O’Connor, Jim Evans, Rowland Sharp, Matthew Gollock, Fenella R. Wood, James Rosindell, Charlie Bartlett, Brett J. Garner, Dafydd Jones, Declan Quigley, and Ben Wray. The distribution, ecology and predicted habitat use of the Critically Endangered angelshark (Squatina squatina) in coastal waters of Wales and the central Irish Sea. Journal of Fish Biology, 101(3):640–658, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

White:2022:FDK

[964] William T. White, Helen L. O’Neill, Jaimie Cleeland, and Tim D. Lamb. Further description of the Kerguelen sandpaper skate Bathyraja irrasa (Rajiformes: Arhynchobatidae) based on additional specimens, including egg cases and embryos. Journal of Fish Biology, 101(3):659–675, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kim:2022:CHM

[965] Hyeong Su Kim, Hyun Tae Kim, Jun Seong Park, and Jae Hyun Im. Changes in the height of minute tubercles on the skin of Korean bitterling embryos (Acheilognathus signifer) and embryo movement in the host mussels. Journal of Fish Biology, 101(3):676–685, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Peltonen:2022:RCF

[966] Heikki Peltonen and Benjamin Weigel. Responses of coastal fishery resources to rapid environmental changes. Journal of Fish Biology, 101(3): 686–698, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2022:GWI

[967] Ronghui Li, Mingyu Shu, Xuanxuan Liu, Zhiwei Nei, Ben Ye, Heyu Wang, and Yifu Gong. Genome-wide identification of mitogen-activated protein kinase (MAPK) gene family in yellow catfish (Pelteobagrus fulviadraco) and their expression profiling under the challenge of Aeromonas hydrophila. Journal of Fish Biology, 101(3):699–710, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Guo:2022:SIA

[968] Haoyu Guo, Joacim Näslund, Søren T. Thomassen, and Martin H. Larsen. Social isolation affects intra-specific interaction behaviour and reduces the size of the cerebellar brain region in juvenile Atlantic salmon Salmo salar. Journal of Fish Biology, 101(3):711–721, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barco:2022:DFS

[969] Andrea Barco, Björn Kullmann, Thomas Knebelsberger, Victoria Sarrazin, Vanessa Kuhs, Axel Kreutle, Christian Pusch, and Ralf Thiel. Detection of fish species from marine protected areas of the North Sea using environmental DNA. Journal of Fish Biology, 101(3):722–727, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gul:2022:OFC

[970] Güzin Gül, Mehmet Baki Yokeş, and Nazli Demirel. The occurrence and feeding of a critically endangered shark species, Oxynotus centrina in the Sea of Marmara. Journal of Fish Biology, 101(3):728–735, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mos:2022:REW

[971] Benjamin Mos and Daniel Mos. Range expansion of a widespread Indo– Pacific haemulid, the barred javelin Pomadasys kaakan (Cuvier, 1830), in a climate change hotspot. Journal of Fish Biology, 101(3):736–740, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Makinen:2022:VFV

[972] Katja Mäkinen, Marjut Rajasilta, Ermei Mäkilä, Sami Jokinen, and Jari Hänninen. Varying frequency of vateritic otoliths in the Baltic herring Clupea harengus membras. Journal of Fish Biology, 101(3):741–744, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:2022:CBS

[973] Richard Kennedy, Robert Rosell, Warren Campbell, Michelle Allen, and Diego Del Villar-Guerra. A comparison of the behaviour and survival of angling vs. trap-sampled Salmo salar smolts. Journal of Fish Biology, 101(3):745–748, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hagihara:2022:MGH

[974] Seishi Hagihara, Ryoshiro Wakiya, Tatsuhiko Maeda, and Shingo Kimura. Morphological and gonadal histological characteristics of the silver-phase male Indo–Pacific eel Anguilla marmorata. Journal of Fish Biology, 101(3):749–752, September 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2022:IIj

[975] Anonymous. Issue information. Journal of Fish Biology, 101(4):753–754, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2022:BHB

[976] William Bernard Perry. Bringing down the hammer on bycatch. Journal of Fish Biology, 101(4):755, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cooke:2022:MEF

[977] Steven J. Cooke, Jordanna N. Bergman, William M. Twardek, Morgan L. Piczak, Grace A. Casselberry, Keegan Lutek, Lotte S. Dahlmo, Kim Birnie-Gauvin, Lucas P. Griffin, Jacob W. Brownscombe, Graham D. Raby, Emily M. Standen, Andrij Z. Horodysky, Sönke Johnsen, Andy J. Danylchuk, Nathan B. Furey, Austin J. Gallagher, Elodie J. I. Lédée, Jon D. Midwood, Lee F. G. Gutowsky, David M. P. Jacoby, Jordan K. Matley, and Robert J. Lennox. The movement ecology of fishes. Journal of Fish Biology, 101(4):756–779, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scharsack:2022:TET

[978] Jörn Peter Scharsack and Frederik Franke. Temperature effects on teleost immunity in the light of climate change. Journal of Fish Biology, 101 (4):780–796, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Guttridge:2022:VSU

[979] Tristan L. Guttridge, Lukas Müller, Bryan A. Keller, Mark E. Bond, R. Dean Grubbs, William Winram, Lucy A. Howey, Bryan S. Frazier, and Samuel H. Gruber. Vertical space use and thermal range of the great hammerhead (Sphyrna mokarran), (Rüppell, 1837) in the western North Atlantic. Journal of Fish Biology, 101(4):797–810, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aydin:2022:DTE

[980] Sevinç Aydın, H. Turan Akkoyun, Mahire Bayramoglu Akkoyun, Aydın Ş. Bengu, Suat Ekin, and Ayşe G. Harlioğlu. Determination of trace element and fatty acid levels in tissues of mirror and scaly carp (Cyprinus carpio L., 1758) living in Keban Dam Reservoir (Elazig). Journal of Fish Biology, 101(4):811–821, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dahlke:2022:BEW

[981] Flemming Dahlke, Velmurugu Puvanendran, Atle Mortensen, Hans-Otto Pörtner, and Daniela Storch. Broodstock exposure to warming and elevated pCO_{2} impairs gamete quality and narrows the temperature window of fertilisation in Atlantic cod. Journal of Fish Biology, 101(4):822–833, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Swain:2022:OSO

[982] Himanshu S. Swain, Ambily Mn, Amir K. Samal, Mitesh H. Ramteke, and Basanta K. Das. Ontogeny of skeletal ossification in Labeo calbasu (Hamilton, 1822) using differential fluorescence staining technique. Journal of Fish Biology, 101(4):834–845, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ducos:2022:OSP

[983] Salomé Ducos, Sally Pugliese, Mikaël Demolliens, Louisa Beraud, Alizée Boussard, Alban Delmas, Sylvia Agostini, Jessica Garcia, Antoine Aiello, and Eric D. H. Durieux. Ontogeny of swimming performance of hatcheryreared post-larvae and juvenile fish: a case of two threatened Mediterranean species. Journal of Fish Biology, 101(4):846–856, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Renan:2022:ABL

[984] Ximena Renán, Thierry Brulé, Gabriela Galindo-Cortes, and Teresa Colás-Marrufo. Age-based life history of three groupers in the southern Gulf of Mexico. Journal of Fish Biology, 101(4):857–873, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morbey:2022:JAT

[985] Yolanda E. Morbey and Daniel Pauly. Juvenile-to-adult transition invariances in fishes: Perspectives on proximate and ultimate causation. Journal of Fish Biology, 101(4):874–884, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pombo-Ayora:2022:AON

[986] Lucía Pombo-Ayora, Viktor N. Peinemann, Collin T. Williams, Song He, Yu Jia Lin, Yukio Iwatsuki, Donal D. C. Bradley, and Michael L. Berumen. Acanthopagrus oconnorae, a new species of seabream (Sparidae) from the Red Sea. Journal of Fish Biology, 101(4):885–897, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liyandja:2022:ECB

[987] Tobit L. D. Liyandja, Jonathan W. Armbruster, Michael O. Poopola, and Melanie L. J. Stiassny. Evolutionary convergence in body shape obscures taxonomic diversity in species of the African Labeo forskalii group: Case study of L. parvus Boulenger 1902 and L. ogunensis Boulenger 1910. Journal of Fish Biology, 101(4):898–913, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Esin:2022:LHV

[988] Evgeny V. Esin, Nikolay O. Melnik, and Grigorii N. Markevich. Lifehistory variation as a source of diversity for endemic white charr (Salmonidae) of the lower Kamchatka River. Journal of Fish Biology, 101 (4):914–924, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Giese:2022:STH

[989] Adriana Carolina Giese, Martín Ignacio García-Asorey, Miguel Ángel Casalinuovo, María Marcela Amaya-Santi, Brian Patrick Kennedy, and Miguel Alberto Pascual. Surfing the tide: Homeward migration of sea trout (Salmo trutta) in a Patagonian river. Journal of Fish Biology, 101 (4):925–936, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chat:2022:STS

[990] Joelle Chat, Amaia Lamarins, and Olivier Lepais. Sea trout (Salmo trutta) straying rate decreases as distance from river mouth increases. Journal of Fish Biology, 101(4):937–944, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mincarone:2022:NRR

[991] Michael Maia Mincarone, Leandro Nolé Eduardo, Fabio Di Dario, Thierry Frédou, Arnaud Bertrand, and Flávia Lucena-Frédou. New records of rare deep-sea fishes (Teleostei) collected from off north-eastern Brazil, including seamounts and islands of the Fernando de Noronha Ridge. Journal of Fish Biology, 101(4):945–959, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Freret-Meurer:2022:IAO

[992] Natalie V. Freret-Meurer, Tatiane C. Fernández, and Amanda C. Vaccani. Influence of the Atlantic Ocean thermal anomaly on the Longsnout seahorse Hippocampus reidi in a Brazilian estuary. Journal of Fish Biology, 101(4):960–971, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Duan:2022:GWI

[993] Yongqiang Duan, Wenping Zhang, Xiaohui Chen, Minghua Wang, Liqiang Zhong, Ju Liu, Wenji Bian, and Shiyong Zhang. Genome-wide identification and expression analysis of mitogen-activated protein kinase (MAPK) genes in response to salinity stress in channel catfish (Ictalurus punctatus). Journal of Fish Biology, 101(4):972–984, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stounberg:2022:EEC

[994] Jonathan Stounberg, Tonny Bernt Thomsen, Benjamin Dominguez Heredia, and Karin Hüssy. Eyes and ears: a comparative approach linking the chemical composition of cod otoliths and eye lenses. Journal of Fish Biology, 101(4):985–995, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goldsworthy:2022:LHC

[995] Nisha C. Goldsworthy, Maya Srinivasan, Patrick Smallhorn-West, Lit-Chien Cheah, Philip L. Munday, and Geoffrey P. Jones. Life-history constraints, short adult life span and reproductive strategies in coral reef gobies of the genus Trimma. Journal of Fish Biology, 101(4):996–1007, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:2022:RBF

[996] Richard James Kennedy, James Barry, William Roche, Robert Rosell, and Michelle Allen. In-river behaviour and freshwater return rates of sea trout, Salmo trutta L., from two coastal river populations. Journal of Fish Biology, 101(4):1008–1020, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Isla:2022:ELD

[997] Adolfo Isla, Patricio Sánchez, Pamela Ruiz, Romina Albornoz, Juan P. Pontigo, María Cecilia Rauch, Christopher Hawes, Luis Vargas-Chacoff, and Alejandro J. Yáñez. Effect of low-dose Piscirickettsia salmonis infection on haematological-biochemical blood parameters in Atlantic salmon (Salmo salar). Journal of Fish Biology, 101(4):1021–1032, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goodman:2022:OCT

[998] Kyle Goodman, Yuri Niella, Taylor Bliss-Henaghan, Robert Harcourt, Amy F. Smoothey, and Victor M. Peddemors. Ontogenetic changes in the tooth morphology of bull sharks (Carcharhinus leucas). Journal of Fish Biology, 101(4):1033–1046, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

VanWichelen:2022:GEA

[999] Jeroen Van Wichelen, Pieterjan Verhelst, Michiel Perneel, Charlotte Van Driessche, David Buysse, Claude Belpaire, Johan Coeck, and Marleen De Troch. Glass eel (Anguilla anguilla L. 1758) feeding behaviour during upstream migration in an artificial waterway. Journal of Fish Biology, 101(4):1047–1057, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Juhasz-Dora:2022:FRB

[1000] Thomas Juhasz-Dora, Jonathan Teague, Thomas K. Doyle, and Julie Maguire. First record of biofluorescence in lumpfish (Cyclopterus lumpus), a commercially farmed cleaner fish. Journal of Fish Biology, 101 (4):1058–1062, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Erkinaro:2022:DIP

[1001] Jaakko Erkinaro, Panu Orell, Jan-Peter Pohjola, Mikko Kytökorpi, Henni Pulkkinen, and Jorma Kuusela. Development of invasive pink salmon (Oncorhynchus gorbuscha Walbaum) eggs in a large Barents Sea river. Journal of Fish Biology, 101(4):1063–1066, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kousteni:2022:RCR

[1002] Vasiliki Kousteni, Vasilis Resaikos, Magdalene Papatheodoulou, and Marios Papageorgiou. Review of catch records of the bigeye thresher shark Alopias superciliosus Lowe, 1841 (Lamniformes: Alopiidae) in the Mediterranean Sea: historical and recent data with biological notes. Journal of Fish Biology, 101(4):1067–1077, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Daupagne:2022:RVS

[1003] Léa Daupagne, Marius Dhamelincourt, Anne Michaud, Jacques Rives, Stellia Sebihi, and Cédric Tentelier. Realistic variations in substrate composition affect spawning preference and egg retention in river lamprey (Lampetra fluviatilis). Journal of Fish Biology, 101(4):1078–1083, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gao:2022:FOM

[1004] Guang Gao, Zhizhong Xiao, Guanglei Ji, Yongshuang Xiao, Daoyuan Ma, and Jun Li. First observation of the mating, egg-laying and hatching behaviour of a captive female Kong skate, Okamejei kenojei (Müller & Henle, 1841). Journal of Fish Biology, 101(4):1084–1091, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nichols:2022:NBA

[1005] Ryan S. Nichols, Edward E. DeMartini, and Erik C. Franklin. No butts about it: using urogenital disparity in a deep-water snapper, Etelis carbunculus (Lutjanidae), for field-based sexual identification. Journal of Fish Biology, 101(4):1092–1097, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bertaco:2022:RTI

[1006] Vinicius Araújo Bertaco, Fernando G. Becker, Marco A. Azevedo, Juliano Ferrer, Everton R. Behr, Tiago R. de Moraes, Nelson J. R. Fagundes, and Luiz R. Malabarba. The record and threats of the invasion of palometa Serrasalmus maculatus (Characiformes: Serrasalmidae) in the Patos lagoon drainage, southern Brazil. Journal of Fish Biology, 101(4): 1098–1103, October 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2022:IIk

[1007] Anonymous. Issue information. Journal of Fish Biology, 101(5):1105– 1106, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2022:SSD

[1008] William Bernard Perry. Sex or snack? Drivers behind the behavioural mystery of basking shark circles revealed. Journal of Fish Biology, 101 (5):1107, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zvavahera:2022:ESI

[1009] Munetsi Zvavahera, Sanet Hugo, Niall G. Vine, and Olaf L. F. Weyl. Exploratory stock identification through morphometric trait analysis of the estuarine round herring, Gilchristella aestuaria (Pisces: Clupeidae). Journal of Fish Biology, 101(5):1108–1118, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Salas-Singh:2022:RBB

[1010] Carolina Salas-Singh, Enrique Morales-Bojórquez, and Hugo Aguirre-Villaseñor. Reproductive biology of the bullseye puffer Sphoeroides annulatus: Gonadosomatic index and its suitability for estimating length at maturity. Journal of Fish Biology, 101(5):1119–1133, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mundodu:2022:EUD

[1011] Shubha Mundodu, Krishnamurthy S. Venakatarama Bhatta, and Reddy R. Sarjapur. Energy utilization during ovarian development of the cichlid fish Oreochromis mossambicus (Peters) inhabiting a perennial lentic water body in Bangalore, India. Journal of Fish Biology, 101(5): 1134–1141, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kawai:2022:STB

[1012] Kentaro Kawai, Hiroki Fujita, Diego Deville, and Tetsuya Umino. Spawning time of black sea bream Acanthopagrus schlegelii, related to underwater photoperiodism in oyster farms. Journal of Fish Biology, 101(5): 1142–1149, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wei:2022:ECL

[1013] Fulei Wei, Jian Liang, Miao Yue, Wengen Tian, Luxian Yu, Zhaohui Feng, and Qiang Hua. Establishment of a cell line derived from the gills of Gymnocypris przewalskii, an endemic Schizothoracine fish from Qinghai Lake of Tibet Plateau. Journal of Fish Biology, 101(5):1150– 1159, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sims:2022:CSA

[1014] David W. Sims, Simon D. Berrow, Ken M. O’Sullivan, Nicholas J. Pfeiffer, Richard Collins, Kev L. Smith, Brianna M. Pfeiffer, Paul Connery, Shane Wasik, Lois Flounders, Nuno Queiroz, Nicolas E. Humphries, Freya C. Womersley, and Emily J. Southall. Circles in the sea: annual courtship “torus” behaviour of basking sharks Cetorhinus maximus identified in the eastern North Atlantic Ocean. Journal of Fish Biology, 101(5):1160–1181, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Saavedra:2022:PBD

[1015] Margarida Saavedra, Teresa G. Pereira, Marisa Barata, Cláudia Aragão, Bárbara Requeijo, Luís E. C. Conceição, and Pedro Pousão-Ferreira. Plant-based diets fed to juvenile meagre Argyrosomus regius with low methionine and taurine supplementation led to an overall reduction in fish performance and to an increase in muscle fibre recruitment. Journal of Fish Biology, 101(5):1182–1188, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Praveenraj:2022:SGH

[1016] Jayasimhan Praveenraj, Rajendran Kiruba-Sankar, Kandasamy Saravanan, Tejas Thackeray, Pratima Singh, John Daniel Marcus Knight, and Philippe Keith. Sicyopterus garra Hora, 1925, a valid species of sicydiine goby from the Andaman Islands, India. Journal of Fish Biology, 101(5):1189–1198, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rowe:2022:FWE

[1017] Stacy Rowe and Brian E. Smith. Food web ecology of Gulf Stream flounder (Citharichthys arctifrons): a continental shelf perspective. Journal of Fish Biology, 101(5):1199–1209, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Buckland-Nicks:2022:ISA

[1018] John Buckland-Nicks and Thomas E. Reimchen. Innervation and structure of the adipose fin of a lanternfish. Journal of Fish Biology, 101(5): 1210–1216, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Braccini:2022:QTD

[1019] Matias Braccini, Saia Bartes, and Peter G. Coulson. Quantifying teleost discards in the shark fisheries of Western Australia. Journal of Fish Biology, 101(5):1217–1224, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vucic:2022:NSM

[1020] Matej Vucić, Mišel Jelić, Göran Klobučar, Dušan Jelić, Han Ming Gan, Chris Austin, Daniel Guyonnet, Isabelle Giraud, Thomas Becking, and Frederic Grandjean. A new set of microsatellite markers for Phoxinus lumaireul senso lato, Phoxinus marsilii and Phoxinus krkae for population and molecular taxonomic studies. Journal of Fish Biology, 101(5): 1225–1234, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fuke:2022:CGD

[1021] Yusuke Fuke, Yuichi Kano, Sein Tun, Lkc Yun, Seint Seint Win, and Katsutoshi Watanabe. Cryptic genetic divergence of the red dwarf rasbora, Microrasbora rubescens, in and around Inle Lake: implications for the origin of endemicity in the ancient lake in Myanmar. Journal of Fish Biology, 101(5):1235–1247, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aguilera:2022:IAM

[1022] Gastón Aguilera, Guillermo E. Terán, Juan Marcos Mirande, Felipe Alonso, Guido Miranda Chumacero, Yamila Cardoso, Sergio Bogan, and Dario R. Faustino-Fuster. An integrative approach method reveals the presence of a previously unreported species of Imparfinis eigenmann and Norris 1900 (Siluriformes: Heptapteridae) in Argentina. Journal of Fish Biology, 101(5):1248–1261, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arroyo-Zuniga:2022:LOS

[1023] Karla Isela Arroyo-Zúñiga, Renata Pacheco-Ovando, Esperanza Granados-Amores, Jasmín Granados-Amores, Javier González Ramírez, and Mariana Díaz-Santana-Iturrios. Lapillus otolith shape, a useful taxonomic feature for the identification of sea catfishes (Ariidae: Siluriformes) from the north-eastern Pacific. Journal of Fish Biology, 101(5):1262–1269, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ogino:2022:RLF

[1024] Yoshimi Ogino and Atsuko Yamaguchi. Reduced lifetime fitness (growth, body condition and survivability) of hatchery-reared tiger pufferfish Takifugu rubripes compared to wild counterparts. Journal of Fish Biology, 101(5):1270–1284, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lilly:2022:IBA

[1025] Jessie Lilly, Hannele M. Honkanen, David M. Bailey, Colin W. Bean, Ruaidhri Forrester, Jessica R. Rodger, and Colin E. Adams. Investigating the behaviour of Atlantic salmon (Salmo salar L.) post-smolts during their early marine migration through the Clyde Marine Region. Journal of Fish Biology, 101(5):1285–1300, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lackmann:2022:OSL

[1026] Alec R. Lackmann, Ewelina S. Bielak-Lackmann, Malcolm G. Butler, and Mark E. Clark. Otoliths suggest lifespans more than 30 years for free-living bowfin Amia calva: Implications for fisheries management in the bowfishing era. Journal of Fish Biology, 101(5):1301–1311, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Young:2022:BSE

[1027] Michael K. Young, Daniel J. Isaak, David Nagel, Dona L. Horan, Kellie J. Carim, Thomas W. Franklin, Victoria A. Zeller, Brett Roper, and Michael K. Schwartz. Broad-scale eDNA sampling for describing aquatic species distributions in running waters: Pacific lamprey Entosphenus tridentatus in the upper Snake River, USA. Journal of Fish Biology, 101(5): 1312–1325, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cicek:2022:CET

[1028] Erdogan Cicek, Soheil Eagderi, Sevil Sungur, and Burak Secer. Capoeta ekmekciae Turan, Kottelat, Kirankaya & Engin, 2006, a junior synonym of Capoeta capoeta (Güldenstädt, 1773) (Teleostei: Cyprinidae). Journal of Fish Biology, 101(5):1326–1332, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schedel:2022:TPP

[1029] Frederic D. B. Schedel, Zuzana Musilova, Adrian Indermaur, Arnold Roger Bitja-Nyom, Walter Salzburger, and Ulrich K. Schliewen. Towards the phylogenetic placement of the enigmatic African genus Prolabeops Schultz, 1941. Journal of Fish Biology, 101(5):1333–1342, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Knight:2022:IBG

[1030] J. D. Marcus Knight, Moulitharan Nallathambi, Balaji Vijayakrishnan, and Praveenraj Jayasimhan. On the identity of the banded gourami Trichogaster fasciata with notes on the taxonomic status of Trichopodus bejeus (Teleostei: Perciformes: Osphronemidae). Journal of Fish Biology, 101(5):1343–1347, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mena-Valenzuela:2022:FOC

[1031] Patricio Mena-Valenzuela, Jonathan Valdiviezo-Rivera, Javier Mena-Olmedo, and Windsor E. Aguirre. The first observation of copulation in Andean catfish Astroblepus ubidiai (Siluriformes, Astroblepidae), in Lago San Pablo, Imbabura, Ecuador. Journal of Fish Biology, 101(5): 1348–1352, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Trevisan:2022:FRB

[1032] João Pedro Trevisan, Rodrigo Antunes Caires, Vinícius Estrella Silva Carvalho, Matheus Marcos Rotundo, and João Alberto Paschoa dos Santos. The first record of burrowing goby Trypauchen vagina (Bloch & Schneider, 1801) (Gobiiformes: Gobiidae: Amblyopinae) from the Atlantic Ocean. Journal of Fish Biology, 101(5):1353–1357, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ugrin:2022:ROR

[1033] Nika Ugrin and Svjetlana Krstulović Šifner. A rare occurrence of reversal in the four-spotted megrim Lepidorhombus boscii (Risso 1810) (Pleuronectiformes: Scophthalmidae) in the central Adriatic Sea. Journal of Fish Biology, 101(5):1358–1360, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jarvis:2022:ECD

[1034] Dominic M. Jarvis, Edward C. Pope, Mathieu Duteil, Ines Fürtbauer, M. Rowan Brown, Richard J. Davis, and Andrew J. King. Elevated CO_{2} does not alter behavioural lateralization in free-swimming juvenile European sea bass (Dicentrarchus labrax) tested in groups. Journal of Fish Biology, 101(5):1361–1365, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Spector:2022:NAR

[1035] Michael Spector, Alyssa J. Clevenstine, Mari Cajandig, Chris Caldow, Elizabeth A. Duncan, Lindsey E. Peavey Reeves, and Ryan M. Freedman. Network analysis reveals aggregation behaviour for an endangered predator at an offshore island. Journal of Fish Biology, 101(5):1366–1370, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Taggart:2022:DPU

[1036] John B. Taggart, Michael J. Leaver, and Michaël Bekaert. DNA polymorphism underlying allozyme variation at a malic enzyme locus (mMEP-2*) in Atlantic salmon (Salmo salar L.). Journal of Fish Biology, 101(5): 1371–1374, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jepsen:2022:SGT

[1037] Niels Jepsen, Luise Richter, Michael Ingemann Pedersen, and Zhiqun (Daniel) Deng. Survival, growth and tag retention of juvenile European eel (Anguilla anguilla L.) with implanted 12 mm passive integrated transponder tags and acoustic tags. Journal of Fish Biology, 101(5):1375–1380, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kovacic:2022:IEM

[1038] Marcelo Kovačić, Roberto Pillon, and Julien P. Renoult. Identification of enigmatic Mediterranean fish Gobius ater Bellotti, 1888 (Teleostei: Gobiidae) based on morphology from underwater photographs. Journal of Fish Biology, 101(5):1381–1384, November 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2022:IIl

[1039] Anonymous. Issue information. Journal of Fish Biology, 101(6):1385– 1386, December 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2022:SFL

[1040] William Bernard Perry. Sleeping with the fishes: Looking under the bonnet of melatonin synthesis. Journal of Fish Biology, 101(6):1387, December 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chuctaya:2022:NSG

[1041] Junior Chuctaya, Pedro Nitschke, Marcelo C. Andrade, Juliana Wingert, and Luiz R. Malabarba. A new species of Geophagus (Teleostei: Cichlidae): Naming a cichlid species widely known in the aquarium hobby as ‘Geophagus sp. Tapajos red head’. Journal of Fish Biology, 101(6): 1388–1404, December 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turner:2022:NSD

[1042] George F. Turner. A new species of deep-water Lethrinops (Cichlidae) from Lake Malawi. Journal of Fish Biology, 101(6):1405–1410, December 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maeda:2022:LSR

[1043] Ken Maeda, Nozomi Hanahara, Masato Uehara, and Katsunori Tachihara. Larval study revealed diversity and life-history traits of cryptobenthic eel gobies. Journal of Fish Biology, 101(6):1411–1427, December 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huang:2022:DSM

[1044] Amy T. Huang, Katharina Alter, Patrick Polte, and Myron A. Peck. Disentangling seasonal from maternal effects on egg characteristics in western Baltic spring-spawning herring Clupea harengus. Journal of Fish Biology, 101(6):1428–1440, December 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burke:2022:ETM

[1045] Teah Grace Burke, Harri Pettitt-Wade, Jack P. W. Hollins, Colin Gallagher, Ellen Lea, Lisa Loseto, and Nigel E. Hussey. Evidence for three morphotypes among anadromous Arctic char (Salvelinus alpinus) sampled in the marine environment. Journal of Fish Biology, 101(6):1441– 1451, December 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Khan:2022:UOB

[1046] Umar Khan, Habib Bal, Zeynep Selva Battal, and Kadir Seyhan. Using otolith and body shape to discriminate between stocks of European anchovy (Engraulidae: Engraulis encrasicolus) from the Aegean, Marmara and Black Seas. Journal of Fish Biology, 101(6):1452–1465, December 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Molbert:2022:NIN

[1047] Noëlie Molbert and Aurélie Goutte. Narrower isotopic niche size in fish infected by the intestinal parasite Pomphorhynchus sp. compared to uninfected ones. Journal of Fish Biology, 101(6):1466–1473, December 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lee:2022:NDF

[1048] Hye-Lyang Lee, Hyo-Jae Yu, and Jin-Koo Kim. New descriptions of four larval lanternfish species and cryptic diversity of Benthosema pterotum (Pisces: Myctophidae) from the Northwest Pacific. Journal of Fish Biology, 101(6):1474–1500, December 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2022:SHP

[1049] Carl Smith, Grzegorz Zieba, Rowena Spence, and Mirosław Przybylski. Spatial heterogeneity in pH, body size and habitat size generates ecological opportunity in an evolutionary radiation. Journal of Fish Biology, 101(6):1501–1508, December 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Parsons:2022:PHA

[1050] Darren Parsons, Rikki Taylor, Richard Hughes, Crispin Middleton, Yann Gublin, and Demi Levell. Predators and habitat association of postsettlement snapper (Chrysophrys auratus). Journal of Fish Biology, 101(6):1509–1521, December 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rajakrishnan:2022:NSB

[1051] Rajeev Rajakrishnan, Meleppura Rajeeshkumar, Sherine Sonia Cubelio, and Artem M. Prokofiev. A new species of Bathysphyraenops from the Indian Ocean with comments on the genus (Teleostei: Acropomatiformes: Howellidae). Journal of Fish Biology, 101(6):1522–1529, December 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aguiar-Santos:2022:INA

[1052] Jamerson Aguiar-Santos, Pieter A. P. deHart, Bruce R. Forsberg, and Carlos E. C. Freitas. Isotopic niche alteration of a predator fish in a dammed Amazonian black water river. Journal of Fish Biology, 101(6): 1530–1539, December 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moore:2022:HGP

[1053] Glenn I. Moore, Corey B. Wakefield, Joseph D. DiBattista, and Stephen J. Newman. Hyporthodus griseofasciatus (Perciformes: Epinephelidae), a new species of deep-water grouper from the west coast of Australia. Journal of Fish Biology, 101(6):1540–1556, December 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shervette:2022:ION

[1054] Virginia R. Shervette and Jesús M. Rivera Hernández. Illuminating otoliths: New insights for life history of Balistes triggerfishes. Journal of Fish Biology, 101(6):1557–1568, December 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Devi:2022:IPV

[1055] Haobijam Sanjita Devi, Chongtham Rajiv, Gopinath Mondal, Zeeshan Ahmad Khan, Sijagurumayum Dharmajyoti Devi, Rupjyoti Bharali, and Asamanja Chattoraj. Influence of photoperiod variations on the mRNA expression pattern of melatonin bio-synthesizing enzyme genes in the pineal organ and retina: a study in relation to the serum melatonin profile in the tropical carp Catla catla. Journal of Fish Biology, 101(6): 1569–1581, December 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Silva-Junior:2022:NSP

[1056] Dario E. Silva-Junior and Angela M. Zanata. A new species of Parotocinclus (Loricariidae: Hypoptopomatinae) from the rio Pardo basin, Bahia State, Brazil, with comments on the sexually dimorphic traits of the nares and olfactory lamellae. Journal of Fish Biology, 101(6):1582–1590, December 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rangel:2022:FWR

[1057] Bianca S. Rangel, André S. Afonso, and Ricardo Garla. Female wound records suggest mating periods for the Caribbean reef shark at an insular marine protected area from the Equatorial Atlantic Ocean. Journal of Fish Biology, 101(6):1591–1594, December 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skeeles:2022:CEH

[1058] Michael R. Skeeles, Hanna Scheuffele, and Timothy D. Clark. Chronic experimental hyperoxia elevates aerobic scope: a valid method to test for physiological oxygen limitations in fish. Journal of Fish Biology, 101(6): 1595–1600, December 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Takeuchi:2022:PIP

[1059] Aya Takeuchi, Eitaro Sawayama, Mari Kuroki, Michael J. Miller, Shun Watanabe, and Katsumi Tsukamoto. Preliminary insight into parental contributions to Japanese eel (Anguilla japonica) preleptocephali spawned on different nights. Journal of Fish Biology, 101(6): 1601–1605, December 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhang:2022:SDF

[1060] Feiran Zhang, Jiahao Liu, Qingzhu Bi, Qiang Ma, Yuliang Wei, Mengqing Liang, and Houguo Xu. Sex difference in fatty acid composition of six marine teleosts. Journal of Fish Biology, 101(6):1606–1610, December 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sen:2022:NOE

[1061] Swatipriyanka Sen, Sujitha Thomas, Shoba Joe Kizhakudan, Gyanaranjan Dash, Rajesh Kumar Pradhan, Shubhadeep Ghosh, Madhumita Das, Biswajit Dash, and Pariyappanal Ulahannan Zacharia. New observations of the endangered giant freshwater whipray, Urogymnus polylepis, provide further evidence for its distribution and breeding in the north-east coast of India. Journal of Fish Biology, 101(6):1611–1616, December 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Matsushige:2022:CRD

[1062] Kazuki Matsushige, Yoshiya Yasutake, and Noritaka Mochioka. Contrasting riverine distribution and habitat use of the Japanese eel, Anguilla japonica, and the giant mottled eel, Anguilla marmorata, in a sympatric river. Journal of Fish Biology, 101(6):1617–1622, December 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Favela:2022:PNA

[1063] Juan P. Rodriguez-Arana Favela, Sebastián Hernández, Rogelio González-Armas, Felipe Galván-Magaña, Arturo Tripp-Valdez, Mauricio Hoyos-Padilla, and James T. Ketchum. A priority nursery area for the conservation of the scalloped hammerhead shark Sphyrna lewini in Mexico. Journal of Fish Biology, 101(6):1623–1627, December 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hankison:2022:ERS

[1064] Shala Hankison, Eric J. Gangloff, Breanna Fry, Alena Arnold, A. J. Lashway, Jenell M. Betts, Sandra D. Otap, Katherine Walter, Makenna Y. Juergens, and Alax Crawford. Effects of reliance on stored sperm on reproduction in the sailfin molly Poecilia latipinna. Journal of Fish Biology, 101(6):1628–1633, December 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Abdelsalam:2022:EEM

[1065] Ehdaa Eltayeb Eltigani Abdelsalam, Pavlina Hucková, and Veronika Piačková. Evaluation of establishment and maintenance of primary cell cultures from several strains of common carp (Cyprinus carpio L.). Journal of Fish Biology, 101(6):1634–1643, December 2022. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2023:IIa

[1066] Anonymous. Issue information. Journal of Fish Biology, 102(1):1–2, January 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2023:ARL

[1067] William Bernard Perry. All roads lead to… fish egg and larvae literature? Journal of Fish Biology, 102(1):3, January 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mutambala:2023:NPP

[1068] Pacifique Kiwele Mutambala, Emmanuel Abwe, Frederic D. B. Schedel, Auguste Chocha Manda, Ulrich K. Schliewen, and Emmanuel J. W. M. N. Vreven. A new Parakneria poll 1965 (Gonorhynchiformes: Kneriidae), ‘Mikinkidi’ from the Upper Lufira Basin (Upper Congo: DRC): Evidence from a morphologic and DNA barcoding integrative approach. Journal of Fish Biology, 102(1):4–26, January 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Queiroz:2023:TET

[1069] Aristóteles Philippe Nunes Queiroz, Maria Lúcia Góes Araújo, Nigel E. Hussey, and Rosângela P. T. Lessa. Trophic ecology of three stingrays (Myliobatoidei: Dasyatidae) off the Brazilian north-eastern coast: Habitat use and resource partitioning. Journal of Fish Biology, 102(1):27–43, January 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Charitonidou:2023:OPE

[1070] Katerina Charitonidou, Emmanuel Panteris, and Kostas Ganias. Oogonial proliferation and early oocyte dynamics during the reproductive cycle of two Clupeiform fish species. Journal of Fish Biology, 102(1):44– 52, January 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2023:RGG

[1071] Xu Li, Wei Dao, and Wei Zhou. Redefinition of the Glyptonsternine genus Pareuchiloglanis (Teleostei: Sisoridae), with descriptions of three new genera. Journal of Fish Biology, 102(1):53–74, January 2023. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

delFresno:2023:SRP

[1072] Pamela S. del Fresno, Javier R. Garcia de Souza, Darío C. Colautti, Yoji Yamamoto, Masashi Yokota, Carlos A. Strüssmann, and Leandro A. Miranda. Sex reversal of pejerrey (Odontesthes bonariensis), a species with temperature-dependent sex determination, in a seminatural environment. Journal of Fish Biology, 102(1):75–82, January 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carnicer:2023:GTB

[1073] Cleide Carnicer, Luciano B. Lima, Fernando Mayer Pelicice, and Dilermando Pereira Lima-Junior. Global trends, biases and gaps in the scientific literature about freshwater fish eggs and larvae. Journal of Fish Biology, 102(1):83–95, January 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Valette:2023:RAG

[1074] Théo Valette, Maëva Leitwein, Jean-Marc Lascaux, Erick Desmarais, Patrick Berrebi, and Bruno Guinand. Redundancy analysis, genome-wide association studies and the pigmentation of brown trout (Salmo trutta L.). Journal of Fish Biology, 102(1):96–118, January 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maganhe:2023:BGS

[1075] Bruna Larissa Maganhe, Laura de Oliveira Camilo, Rodrigo E. Kurokawa, Hugo Gallo Neto, and Eduardo Gomes Sanches. Birth and growth of the shortnose guitarfish Zapteryx brevirostris (Müller & Henle, 1841) (Chondrichthyes, Rhinobatidae) in captivity. Journal of Fish Biology, 102(1):119–127, January 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wu:2023:FRM

[1076] Jing Wu, Yifan Nie, Jingya Wang, Guoqing Feng, Le Hao, Yanping Ma, Yugu Li, and Zhenxing Liu. Fcγ receptor-mediated phagocytosis pathway was involved in phagocytosis of mIgM^{+} B lymphocytes from largemouth bass (Micropterus salmoides). Journal of Fish Biology, 102(1): 128–140, January 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2023:EHG

[1077] Ying-Ying Li, Xi Chen, Jin-Xian Yang, Qiang Chen, Tie-Ying Song, and Jun-Qing Ge. Evaluation of housekeeping genes as references for quantitative real-time PCR analysis of European eel, Anguilla anguilla. Journal of Fish Biology, 102(1):141–154, January 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nakamoto:2023:MTP

[1078] Bobby J. Nakamoto, Carson A. Jeffres, Nicholas J. Corline, Mollie Ogaz, Christina J. Bradley, Joshua H. Viers, and Marilyn L. Fogel. Multiple trophic pathways support fish on floodplains of California’s Central Valley. Journal of Fish Biology, 102(1):155–171, January 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bettcher:2023:FRM

[1079] Vanessa B. Bettcher, Nayara Bucair, Marta Granville, Amanda Góes, Liana de Figueiredo Mendes, Fabio Di Dario, Rodrigo Leão de Moura, Ricardo Garla, and Bianca S. Rangel. First record of mating behaviour and induced parturition of the Brazilian endemic Lutz’s stingray Hypanus berthalutzae. Journal of Fish Biology, 102(1):172–177, January 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lu:2023:IEP

[1080] Lu Lu, Shaotong Dai, Hao Zhu, Xin Zhang, Zhiqiong Li, Xiaoli Huang, Ping Ouyang, Yi Geng, and Defang Chen. Identification and expression profiling of receptor-interacting serine/threonine-protein kinase 2 in Siberian sturgeon (Acipenser baerii). Journal of Fish Biology, 102 (1):178–187, January 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Londono-Burbano:2023:NSS

[1081] Alejandro Londoño-Burbano and Marcelo R. Britto. A new species of Sturisoma swainson, 1838 (Loricariidae: Loricariinae), from the Madeira River basin, with a discussion of historical biogeography of western Amazonas and Paraguay River basins. Journal of Fish Biology, 102(1):188– 203, January 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anwar:2023:MCA

[1082] Gehan M. Anwar and Anwaar S. M. Abu Shnaf. Mitotic chromosomal abnormalities and DNA polymorphism in the Nile tilapia (Oreochromis niloticus, Linnaeus, 1758) as a biomarker for water pollution by heavy metals. Journal of Fish Biology, 102(1):204–213, January 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Santos:2023:EOC

[1083] Zaqueu Santos, Lucas S. Oliveira, Ruineris A. Cajado, Fabíola K. S. Silva, Lenise V. F. Silva, and Diego M. Zacardi. Early ontogeny of the commercially valuable fish red-bellied pacu Piaractus brachypomus (Characiformes, Serrasalmidae) from the Amazon, Brazil. Journal of Fish Biology, 102(1):214–224, January 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Randall:2023:RSL

[1084] Zachary S. Randall, Gabriel A. Somarriba, Sampan Tongnunui, and Lawrence M. Page. Review of the spotted lizard loaches, Pseudohomaloptera (Cypriniformes: Balitoridae) with a re-description of Pseudohomaloptera sexmaculata and description of a new species from Sumatra. Journal of Fish Biology, 102(1):225–240, January 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yazici:2023:SLV

[1085] Ramazan Yazici. Sex-linked variations in the sagittal otolith biometry of Nemipterus randalli (Russell, 1986) from the eastern Mediterranean Sea. Journal of Fish Biology, 102(1):241–247, January 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chelotti:2023:CTE

[1086] Luiza de David Chelotti and Otto Bismarck Fazzano Gadig. Comparative trophic ecology of two sympatric guitarfishes Pseudobatos (Chondrichthyes, Rhinobatidae) from Southeast Brazil, southwestern Atlantic. Journal of Fish Biology, 102(1):248–257, January 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Osmond:2023:LDS

[1087] Daniel R. Osmond, R. Andrew King, Bruce Stockley, Sophie Launey, and Jamie R. Stevens. A low-density single nucleotide polymorphism panel for brown trout (Salmo trutta L.) suitable for exploring genetic diversity at a range of spatial scales. Journal of Fish Biology, 102(1):258–270, January 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Santoro:2023:PPN

[1088] Mario Santoro, Bruno Bellisario, Fernando Ángel Fernández-Álvarez, Fabio Crocetta, and Marialetizia Palomba. Parasites and prey of the nursehound shark Scyliorhinus stellaris (Linnaeus, 1758): Insights into hidden trophic web interactions in the Mediterranean Sea. Journal of Fish Biology, 102(1):271–280, January 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Muhala:2023:FRM

[1089] Valdemiro Muhala, Aurycéia Guimarães-Costa, Isadola Eusébio Macate, Sara Tembe, Yolanda Mula, Érica Tóvela, Adam Rick Bessa-Silva, Marcelo Vallinoto, and Iracilda Sampaio. First record of mudskipper Boleophthalmus dussumieri (Gobiidae: Oxudercinae) on the coast of Mozambique and evidence of two putative lineages along its known distribution range. Journal of Fish Biology, 102(1):281–286, January 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shipley:2023:OHN

[1090] Oliver N. Shipley, Jamie Fitzgerald, Bryan Horne, Steven Crowe, and Austin J. Gallagher. Observations of hypomelanosis in the nurse shark Ginglymostoma cirratum. Journal of Fish Biology, 102(1):287–289, January 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gomes-Pereira:2023:HBF

[1091] José Nuno Gomes-Pereira, Christopher K. Pham, Joana Miodonski, Marco A. R. Santos, Gisela Dionísio, Diana Catarino, Marianne Nyegaard, Etsuro Sawai, Gilberto P. Carreira, and Pedro Afonso. The heaviest bony fish in the world: a 2744-kg giant sunfish Mola alexandrini (Ranzani, 1839) from the North Atlantic. Journal of Fish Biology, 102 (1):290–293, January 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Peinemann:2023:FRB

[1092] Viktor Nunes Peinemann, Lucía Pombo-Ayora, Jesse E. M. Cochran, Fabio Marchese, Giovanni Chimienti, Mattie Rodrigue, Ameer A. Eweida, Paul A. Marshall, Francesca Benzoni, and Michael L. Berumen. First record of Boulenger’s anthias Sacura boulengeri (Heemstra 1973) in the Red Sea. Journal of Fish Biology, 102(1):294–298, January 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2023:IIb

[1093] Anonymous. Issue information. Journal of Fish Biology, 102(2):299–300, February 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2023:PBS

[1094] William Bernard Perry. The power of bringing science to the people. Journal of Fish Biology, 102(2):301, February 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chuphal:2023:CBR

[1095] Bhawna Chuphal, Priyanka Sathoria, Umesh Rai, and Brototi Roy. Crosstalk between reproductive and immune systems: the teleostean perspective. Journal of Fish Biology, 102(2):302–316, February 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Broom:2023:HAI

[1096] Casey J. Broom, Olaf L. F. Weyl, and Josie South. Habitat associations of imperilled fishes after conservation intervention in the Cape Fold Ecoregion, South Africa. Journal of Fish Biology, 102(2):317–327, February 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cao:2023:GCG

[1097] Liu Cao, Pan Chen, Xingrong Hou, Jun Ma, Ning Yang, Yanping Xu, Yanjin Zhang, Anqi Zhao, Jiyuan Zhang, Xinyu Li, and Hai Huang. Genetic characteristics and growth patterns of the hybrid grouper derived from the hybridization of Epinephelus fuscoguttatus (female) × Epinephelus polyphekadion (male). Journal of Fish Biology, 102(2):328– 339, February 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bahiense:2023:FLR

[1098] Raphael Nogueira Bahiense, Leandro Santos Costa, Helder de Oliveira Guilherme, Camila Gomes de Oliveira, Verônica Guimarães Landa Prado, Rodrigo Fortes-Silva, and Paula Adriane Perez Ribeiro. Feeding and locomotor rhythm of the ornamental cichlid fish Amatitlania sp. in different social groups. Journal of Fish Biology, 102(2):340–348, February 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ren:2023:CET

[1099] Tianjiao Ren, Jinxia Liu, Ke Liu, Zining Zhang, Zihang Ma, Solomon F. Dan, Zhenyu Lan, Min Lu, Huaiyi Fang, Yan Zhang, Peng Zhu, and Yongyan Liao. Cloning and expression of two anti-lipopolysaccharide factors in Eriocheir hepuensis under Vibrio alginolyticus-induced stress. Journal of Fish Biology, 102(2):349–357, February 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yakubu:2023:PCT

[1100] Yahaya Yakubu, Muhammad T. Ahmad, Chou M. Chong, Intan S. Ismail, and Khozirah Shaari. Phenolic content of Terminalia catappa L. leaf and toxicity evaluation on red hybrid tilapia (Oreochromis sp.). Journal of Fish Biology, 102(2):358–372, February 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shollenberger:2023:DTS

[1101] Kurtis R. Shollenberger, Alexis M. Janosik, and Carol Johnston. Detection of the threatened snail darter Percina tanasi in the Tennessee River system using environmental DNA. Journal of Fish Biology, 102 (2):373–379, February 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhang:2023:RDE

[1102] Qi Zhang, Yumeng Wu, Weiyuan Li, Jia Wang, Huiting Zhou, Lei Zhang, Qi Liu, Liu Ying, and Hongwei Yan. Retinal development and the expression profiles of opsin genes during larval development in Takifugu rubripes. Journal of Fish Biology, 102(2):380–394, February 2023. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Corso:2023:LSA

[1103] Andrew D. Corso, Jan R. McDowell, Ellen E. Biesack, Sarah C. Muffelman, and Eric J. Hilton. Larval stages of the Antarctic dragonfish Akarotaxis nudiceps (Waite, 1916), with comments on the larvae of the morphologically similar species Prionodraco evansii Regan 1914 (Notothenioidei: Bathydraconidae). Journal of Fish Biology, 102(2):395– 402, February 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhang:2023:BAO

[1104] Shuo Zhang, Xiao Zhang, Shike Gao, Ruilin Shu, Guanghui Fu, and Jikun Lu. Bilateral asymmetry of otoliths from Collichthys lucidus of different sizes in Haizhou Bay and Xiangshan Bay. Journal of Fish Biology, 102 (2):403–412, February 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pena:2023:AGL

[1105] Renato Peña, Carlos U. McGregor-Bravo, and Mauricio Contreras-Olguín. Allometric growth and larval development in Pacific red snapper Lutjanus peru under culture conditions. Journal of Fish Biology, 102 (2):413–425, February 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ondrackova:2023:PGS

[1106] Markéta Ondračková, Maria Yu Tkachenko, Veronika Bartáková, Anna Bryjová, Michal Janáč, Grzegorz Zieba, Kacper Pyrzanowski, and Yuriy Kvach. Population genetic structure, parasite infection and somatic condition of pumpkinseed Lepomis gibbosus (Actinopterygii: Centrarchidae) in the Oder river basin. Journal of Fish Biology, 102(2):426–442, February 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marajo:2023:CRF

[1107] Leandro Marajó, Patrik Ferreira Viana, Alex Matheus Viana Ferreira, Lúcia Helena Rapp Py-Daniel, Marcelo de Bello Cioffi, Alexandr Sember, and Eliana Feldberg. Chromosomal rearrangements and the first indication of an [female-sign]X_{1} X_{1} X_{2} X_{2} /[male-sign]X_{1} X_{2} Ya sex chromosome system in Rineloricaria fishes (Teleostei: Siluriformes). Journal of Fish Biology, 102(2):443–454, February 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fainburg:2023:PSM

[1108] Leandro Fainburg, David Sabadin, Juan Martín Díaz de Astarloa, and Pedro Fernández Iriarte. Population structure of mud flounder Paralichthys orbignyanus from the south-western Atlantic Ocean. Journal of Fish Biology, 102(2):455–464, February 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kjaeraas:2023:TDM

[1109] Halvor Kjærås, Henrik Baktoft, Ana T. Silva, Karl Øystein Gjelland, Finn Økland, Torbjørn Forseth, Marcell Szabó-Mészáros, and Olle Calles. Three-dimensional migratory behaviour of European silver eels (Anguilla anguilla) approaching a hydropower plant. Journal of Fish Biology, 102(2):465–478, February 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lahteenmaki:2023:SMB

[1110] Linus Lähteenmäki, Panu Orell, Atso Romakkaniemi, and Martin Snickars. Spawning migration behaviour of sea trout (Salmo trutta L.) in a boreal river system: effects of flow conditions and obstacles on migratory activity. Journal of Fish Biology, 102(2):479–491, February 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rambahiniarison:2023:DRM

[1111] Joshua Rambahiniarison, Ariana Agustines, Konstantinos Alexopoulos, Gonzalo Araujo, Asia O. Armstrong, Shannon Arnold, Aldrin Barruga, Titus Cañete, Segundo Conales, Jr., Kymry Delijero, Nonie P. Enolva, Anna L. Flam, Eliya Keane, Jessica Labaja, Christine G. Legaspi, Calum Murie, Ryan Murray, Simon P. Oliver, Simon J. Pierce, Alessandro Ponzo, Christoph A. Rohner, Rosanna Schifferer, Sally Snow, Martina Spakowski, Guy M. W. Stevens, Timur Tilgel, Jessica N. C. Wong, Arnel Andrew Yaptinchay, and Yotam Barr. Distribution of the reef manta ray Mobula alfredi and the oceanic manta ray Mobula birostris in the Philippines: a collaborative effort for conservation. Journal of Fish Biology, 102(2):492–503, February 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Xie:2023:TAM

[1112] Xi Xie, Xianping Liao, Zhengsheng Xu, Wenlang Liang, Yilin Su, Li Lin, Jungang Xie, and Weiqiang Lin. Transcriptome analysis of the muscle of fast- and slow-growing phoenix barb (Spinibarbus denticulatus denticulatus). Journal of Fish Biology, 102(2):504–515, February 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Machida:2023:ESN

[1113] Yoshiyasu Machida, Hiroshi Takahashi, Tetsuya Tsuruta, and Akira Goto. Effect of salinity on nest building behaviour in the nine-spined stickleback Pungitius sinensis. Journal of Fish Biology, 102(2):516–519, February 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deSousa:2023:CNC

[1114] Rodrigo Petry Corrêa de Sousa, Jaymeson Lielson Anjos dos Santos, Gláucia Caroline Silva-Oliveira, Ivanete de Oliveira Furo, Edivaldo Herculano Corrêa de Oliveira, and Marcelo Vallinoto. Characterization of a new cytotype and ocurrence of a B microchromosome in two spot astyanax, Astyanax bimaculatus Linnaeus, 1758 (Characiformes: Characidae). Journal of Fish Biology, 102(2):520–524, February 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Feldheim:2023:PEP

[1115] Kevin A. Feldheim, Jean Dubach, and Lise Watson. Parthenogenesis in an elasmobranch in the presence of conspecific males. Journal of Fish Biology, 102(2):525–527, February 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Armada-Tapia:2023:FEM

[1116] Sofía Armada-Tapia, José Leonardo Castillo-Geniz, Nelva Victoria-Cota, Luis Rodrigo Arce-Valdés, and Luis Manuel Enríquez-Paredes. First evidence of multiple paternity in the blue shark (Prionace glauca). Journal of Fish Biology, 102(2):528–531, February 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cowan:2023:TCL

[1117] Zara-Louise Cowan, Rohan M. Brooker, Julien Gasc, Camille Gache, David Lecchini, and William E. Feeney. Targeted census of lionfishes (Scorpaenidae) reveals high densities in their native range. Journal of Fish Biology, 102(2):532–536, February 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bernthal:2023:HST

[1118] Fionn R. Bernthal, Ben W. Seaman, Edward Rush, John D. Armstrong, Darryl McLennan, Keith H. Nislow, and Neil B. Metcalfe. High summer temperatures are associated with poorer performance of underyearling Atlantic salmon (Salmo salar) in upland streams. Journal of Fish Biology, 102(2):537–541, February 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Naylor:2023:SBM

[1119] Paul Naylor, Teresa Naylor, Louise Hammond, and David M. P. Jacoby. From sneaker to bourgeois male: Long-term observations on recognized tompot blenny Parablennius gattorugine individuals reveal new information on their biology. Journal of Fish Biology, 102(2):542–545, February 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2023:IIc

[1120] Anonymous. Issue information. Journal of Fish Biology, 102(3):547– 548, March 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2023:TOT

[1121] William Bernard Perry. Travelling in the omics time machine: uncovering mysteries of the extinct Moroccan trout (Salmo pallaryi). Journal of Fish Biology, 102(3):549, March 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pushchin:2023:SDR

[1122] Igor Pushchin, Sergei Kondrashev, and Tatiana Borshcheva. The structure and diversity of retinal ganglion cells in the masked greenling Hexagrammos octogrammus Pallas, 1814 (Pisces: Scorpaeniformes: Hexagrammidae). Journal of Fish Biology, 102(3):550–563, March 2023. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leung:2023:RBL

[1123] Amanda Jhu-Xhin Leung, Amy Yee-Hui Then, and Kar-Hoe Loh. Reproductive biology, length–weight relationship and diet of co-occurring butterfly rays, Gymnura poecilura and Gymnura zonura, in Malaysian waters. Journal of Fish Biology, 102(3):564–574, March 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nyqvist:2023:PTI

[1124] Daniel Nyqvist, Alfredo Schiavon, Alessandro Candiotto, Gloria Mozzi, Florian Eggers, and Claudio Comoglio. PIT-tagging Italian spined loach (Cobitis bilineata): Methodology, survival and behavioural effects. Journal of Fish Biology, 102(3):575–580, March 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coleman:2023:GAL

[1125] Richard R. Coleman, Derek W. Kraft, Mykle L. Hoban, Robert J. Toonen, and Brian W. Bowen. Genomic assessment of larval odyssey: selfrecruitment and biased settlement in the Hawaiian surgeonfish Acanthurus triostegus sandvicensis. Journal of Fish Biology, 102(3):581–595, March 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huang:2023:AGP

[1126] Xiang Huang, Cheng Zhang, Jiabao Hu, Shunshun Tao, Jie Tang, Lin Huang, Chunjing Zheng, Shanliang Xu, and Yajun Wang. Analyses of growth performance and realized heritability of the second generation of Indo–Pacific Pampus argenteus. Journal of Fish Biology, 102(3):596– 604, March 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elvy:2023:RBS

[1127] Jordan E. Elvy, Jane E. Symonds, Zoë Hilton, Seumas P. Walker, Louis A. Tremblay, and Neill A. Herbert. The relationships between specific dynamic action, nutrient retention and feed conversion ratio in farmed freshwater Chinook salmon (Oncorhynchus tshawytscha). Journal of Fish Biology, 102(3):605–618, March 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Delling:2023:MRP

[1128] Bo Delling, Filip Thörn, Michael Norén, and Martin Irestedt. Museomics reveals the phylogenetic position of the extinct Moroccan trout Salmo pallaryi. Journal of Fish Biology, 102(3):619–627, March 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marshall:2023:SRE

[1129] Andrea D. Marshall, Anna L. Flam, Nakia Cullain, Michelle Carpenter, Janneman Conradie, and Stephanie K. Venables. Southward range extension and transboundary movements of reef manta rays Mobula alfredi along the east African coastline. Journal of Fish Biology, 102(3):628– 634, March 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coffill-Rivera:2023:ETM

[1130] Manuel E. Coffill-Rivera, J. Wesley Neal, and Peter J. Allen. Effects of temperature on metabolic rate and lower dissolved oxygen tolerance of juvenile speckled peacock bass Cichla temensis. Journal of Fish Biology, 102(3):635–642, March 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zelman:2023:REM

[1131] Kaleb Zelman, Phillip Harrison, Antóin M. O’Sullivan, Samuel Andrews, Stephen Peake, Tommi Linnansaari, Scott A. Pavey, and R. Allen Curry. Reproductive ecology of muskellunge (Esox masquinongy), an introduced predator, in the lower Wolastoq/Saint John River, New Brunswick, Canada. Journal of Fish Biology, 102(3):643–654, March 2023. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aranha:2023:GTE

[1132] Sofia Graça Aranha, Alexandra Teodósio, Vânia Baptista, Karim Erzini, and Ester Dias. A glimpse into the trophic ecology of deep-water sharks in an important crustacean fishing ground. Journal of Fish Biology, 102 (3):655–668, March 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Quiroga:2023:TEJ

[1133] Edna De la Llata Quiroga, Randall Arauz, Arturo Tripp Valdez, Laura Porras Murillo, Manuel Spinola Parallada, Ricardo Sánchez-Murillo, and Elpis J. Chávez. Trophic ecology of juvenile bull sharks (Carcharhinus leucas) in the Coyote estuary, Costa Rica. Journal of Fish Biology, 102 (3):669–679, March 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shalu:2023:DBP

[1134] Kannan Shalu, Liju Thomas, Ghosh Ramvilas, Kadapurathillam S. Shabeena, Siby Philip, Sivanpillai Sureshkumar, Rajeev Raghavan, and Kutty Ranjeet. DNA barcodes for the pipefish genus Corythoichthys (Actinopterygii: Syngnathiformes) from the Indian Ocean provide insights into cryptic diversity. Journal of Fish Biology, 102(3):680–688, March 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garcia-Vega:2023:ETV

[1135] Ana García-Vega, Jorge Ruiz-Legazpi, Juan Francisco Fuentes-Pérez, Francisco Javier Bravo-Córdoba, and Francisco Javier Sanz-Ronda. Effect of thermo-velocity barriers on fish: influence of water temperature, flow velocity and body size on the volitional swimming capacity of northern straight-mouth nase (Pseudochondrostoma duriense). Journal of Fish Biology, 102(3):689–706, March 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:2023:PAT

[1136] Richard Kennedy, Robert Rosell, Ewan Hunter, and Diego del Villar-Guerra. Programmed acoustic tags reveal novel information on late-phase marine life in Atlantic salmon, Salmo salar. Journal of Fish Biology, 102 (3):707–711, March 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ghinter:2023:FGL

[1137] Léopold Ghinter, Christophe Anderson, Dominique Robert, Gesche Winkler, Louis Bernatchez, and Céline Audet. A first glimpse of larval ecology of halibut species in the Gulf of St. Lawrence, Canada. Journal of Fish Biology, 102(3):712–717, March 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lubitz:2023:PRE

[1138] Nicolas Lubitz and Adam Barnett. Possible range extension in the endangered and understudied winghead shark (Eusphyra blochii). Journal of Fish Biology, 102(3):718–720, March 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skora:2023:EPE

[1139] Michał E. Skóra, John Iwan Jones, Alan F. Youngson, Sean Robertson, Alan Wells, Rasmus B. Lauridsen, and Gordon H. Copp. Evidence of potential establishment of pink salmon Oncorhynchus gorbuscha in Scotland. Journal of Fish Biology, 102(3):721–726, March 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Khalil:2023:DBF

[1140] Andrew M. Khalil, Ashton Gainsford, and Lynne van Herwerden. DNA barcoding of fresh seafood in Australian markets reveals misleading labelling and sale of endangered species. Journal of Fish Biology, 102(3): 727–733, March 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2023:CHE

[1141] Anonymous. Correction to ‘High egg retention in Chinook salmon Oncorhynchus tshawytscha carcasses sampled downstream of a migratory barrier’. Journal of Fish Biology, 102(3):734–735, March 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic). See [819].

Anonymous:2023:IId

[1142] Anonymous. Issue information. Journal of Fish Biology, 102(4):737– 738, April 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2023:KSA

[1143] William Bernard Perry. A knight in shining armour: iridescence in male jewelled splitfin (Xenotoca variata). Journal of Fish Biology, 102(4): 739, April 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Locatelli:2023:SAL

[1144] Anna C. P. Locatelli, Rodrigo F. Bastos, Matheus A. Oliveira, and Beatrice P. Ferreira. Scientometric analysis and literature synthesis of 60 years of science on the Atlantic goliath grouper (Epinephelus itajara). Journal of Fish Biology, 102(4):740–756, April 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Motson:2023:VPC

[1145] Katie Motson, Kate S. Hutson, and Andrew S. Hoey. Variation in the parasite communities of three co-occurring herbivorous coral reef fishes.

Journal of Fish Biology, 102(4):757–772, April 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harris:2023:FDA

[1146] Jalaldeen Mohamed Harris, Periyathamby Vinobaba, Ranil Kavindra Asela Kularatne, and Syed Ajmal Khan. Fish diversity and assemblage in the Batticaloa Lagoon, Sri Lanka. Journal of Fish Biology, 102(4):773–793, April 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moyaho:2023:ISS

[1147] Alejandro Moyaho, José Luis Salazar-Bautista, Evelyn Beristain-Castillo, and Flores-Urbina Amira. Iridescent scales signal male fighting ability to access females in the jewelled splitfin Xenotoca variata. Journal of Fish Biology, 102(4):794–802, April 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Guilherme:2023:FBL

[1148] Helder de Oliveira Guilherme, Paula Adriane Perez Ribeiro, Verônica Guimarães Landa Prado, Raphael Nogueira Bahiense, Pedro Gomes Gamarano, Camila Gomes de Oliveira, Débora de Almeida Freitas, and Leandro Santos Costa. Feeding behaviour, locomotion rhythms and blood biochemistry of the neotropical red-tail catfish (Phractocephalus hemioliopterus). Journal of Fish Biology, 102(4):803–815, April 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tasbozan:2023:FAB

[1149] Oğuz Taşbozan, Celal Erbaş, Mehtap Bayır, Erdal Özdemir, Gökhan Arslan, and Abdulkadir Bayır. Fatty acid-binding protein genes in gilthead seabream: molecular cloning and nutritional regulation under low water temperatures. Journal of Fish Biology, 102(4):816–828, April 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skelton:2023:LBM

[1150] Zachary R. Skelton, Tanya S. Prinzing, Philip A. Hastings, and Nicholas C. Wegner. Laboratory-based measures of temperature preference and metabolic thermal sensitivity provide insight into the habitat utilisation of juvenile California horn shark (Heterodontus francisci) and leopard shark (Triakis semifasciata). Journal of Fish Biology, 102(4): 829–843, April 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Currey:2023:GWA

[1151] Mark C. Currey, Charline Walker, Susan Bassham, Hope M. Healey, Emily A. Beck, and William A. Cresko. Genome-wide analysis facilitates estimation of the amount of male contribution in meiotic gynogenetic three-spined stickleback (Gasterosteus aculeatus). Journal of Fish Biology, 102(4):844–855, April 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Quera:2023:MLL

[1152] Vicenç Quera, Francesc S. Beltran, Elisabet Gimeno, and Ruth Dolado. Motion leadership and local interaction in two species of freshwater fish (Danio rerio and Hyphessobrycon herbertaxelrodi). Journal of Fish Biology, 102(4):856–869, April 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Quindazzi:2023:OMA

[1153] Micah J. Quindazzi, Leigh P. Gaffney, Emma Polard, Nick Bohlender, Will Duguid, and Francis Juanes. Otolith mineralogy affects otolith shape asymmetry: a comparison of hatchery and natural origin Coho salmon (Oncorhynchus kisutch). Journal of Fish Biology, 102(4):870– 882, April 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pengal:2023:CDR

[1154] Polona Pengal, Blaž Cokan, Finn Økland, Johan Höjesjö, Meelis Tambets, and Eva Bonsak Thorstad. Comparing dominance relationships and movement of native marble trout (Salmo marmoratus) and introduced rainbow trout (Oncorhynchus mykiss). Journal of Fish Biology, 102(4):883–892, April 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kolmann:2023:DFR

[1155] Matthew A. Kolmann, Ramon S. Nagesan, James V. Andrews, Samuel R. Borstein, Rodrigo Tinoco Figueroa, Randal A. Singer, Matt Friedman, and Hernán López-Fernández. DiceCT for fishes: recommendations for pairing iodine contrast agents with μ CT to visualize soft tissues in fishes. Journal of Fish Biology, 102(4):893–903, April 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bakshi:2023:RPD

[1156] Amrita Bakshi and Umesh Rai. Reproductive phase-dependent and sexually dimorphic expression of leptin and its receptor in different parts of brain of spotted snakehead Channa punctata. Journal of Fish Biology, 102(4):904–912, April 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Talarico:2023:MIA

[1157] Giancarlo G. M. Talarico, Mélissa Grégoire, Jean-Michel Weber, and Jan A. Mennigen. The mammalian insulin antagonist S961 does not exhibit insulin receptor antagonism in rainbow trout in vivo. Journal of Fish Biology, 102(4):913–923, April 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lopez-Perez:2023:EDM

[1158] Cristina López-Pérez, María Pilar Olivar, Víctor M. Tuset, Ainhoa Bernal, and Percy Alexander Hulley. Energy density of mesopelagic fishes from the Atlantic Ocean. Journal of Fish Biology, 102(4):924–935, April 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Osman:2023:DCC

[1159] Yassein A. Osman and Mohamed Samy-Kamal. Diversity and characteristics of commercial Red Sea fish species based on fish market survey: informing management to reduce the risk of overfishing. Journal of Fish Biology, 102(4):936–951, April 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kashef:2023:MBR

[1160] Neosha S Kashef, David M. Stafford, Susan M. Sogard, John C. Garza, John C. Field, and Elizabeth A. Gilbert-Horvath. Multiple-brooding rockfishes (Sebastes spp.) can utilize stored sperm from individual sires to fertilize consecutive broods. Journal of Fish Biology, 102(4):952–961, April 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fernholm:2023:NSH

[1161] Bo Fernholm and Michael Maia Mincarone. A new species of the hagfish genus Eptatretus (Myxinidae) from the Bahamas, western North Atlantic. Journal of Fish Biology, 102(4):962–967, April 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Watt:2023:RPB

[1162] Ashley M. Watt, D. Andrew R. Drake, David Lawrie, and Trevor E. Pitcher. Reproductive phenology and behaviour of endangered redside dace (Clinostomus elongatus) in urban streams. Journal of Fish Biology, 102(4):968–976, April 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Piazza:2023:ISR

[1163] Yanina G. Piazza, Ismael E. Lozano, Sabina Llamazares Vegh, Fabiana L. Lo Nostro, and Carlos M. Fuentes. Integrative study of the reproductive biology and growth of Acestrorhynchus pantaneiro Menezes, 1992 (Characiformes, Acestrorhynchidae). Journal of Fish Biology, 102(4): 977–991, April 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henderson:2023:SMM

[1164] Aaron C. Henderson, Stephan Bruns, and Shamsa Al Hameli. Smooth muscle “microsphincters” in the gastric mucosa of stingrays (Elasmobranchii, Dasyatidae). Journal of Fish Biology, 102(4):992–995, April 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2023:IIe

[1165] Anonymous. Issue information. Journal of Fish Biology, 102(5):997– 998, May 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2023:TLN

[1166] William Bernard Perry. Turn left at the next rocky atoll: small fish fare well with new satellite tags. Journal of Fish Biology, 102(5):999, May 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Desforges:2023:ERC

[1167] Jessica E. Desforges, Kim Birnie-Gauvin, Fredrik Jutfelt, Kathleen M. Gilmour, Erika J. Eliason, Terra L. Dressler, David J. McKenzie, Amanda E. Bates, Michael J. Lawrence, Nann Fangue, and Steven J. Cooke. The ecological relevance of critical thermal maxima methodology for fishes. Journal of Fish Biology, 102(5):1000–1016, May 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reis-Junior:2023:CSR

[1168] Josafá Reis-Júnior, Arnaud Bertrand, Thierry Frédou, Jonas Vasconcelos-Filho, Kátia C. Aparecido, and Paulo J. Duarte-Neto. Community-scale relationships between body shape and trophic ecology in tropical demersal marine fish of northeast Brazil. Journal of Fish Biology, 102 (5):1017–1028, May 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Naisbett-Jones:2023:MLT

[1169] Lewis C. Naisbett-Jones, Creed Branham, Shayla Birath, Savannah Paliotti, Andrew R. McMains, Frederick Joel Fodrie, James W. Morley, Jeffrey A. Buckel, and Kenneth J. Lohmann. A method for long-term retention of pop-up satellite archival tags (PSATs) on small migratory fishes. Journal of Fish Biology, 102(5):1029–1039, May 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rueda:2023:GPS

[1170] Eva C. Rueda, Carolina B. Machado, Veronica Castro, Alline Braga-Silva, Guillermo Ojeda, Facundo Vargas, Gisela Loretán, Patricia D. Freitas, Pedro M. Galetti, Jr., and Guillermo Ortí. Genetic population structure of Pseudoplatystoma corruscans (Siluriformes: Pimelodidae) and evidence of temporal variation in structure. Journal of Fish Biology, 102(5):1040–1048, May 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown-Vuillemin:2023:FER

[1171] Sarah Brown-Vuillemin, Réjean Tremblay, Denis Chabot, Pascal Sirois, and Dominique Robert. Feeding ecology of redfish (Sebastes sp.) inferred from the integrated use of fatty acid profiles as complementary dietary tracers to stomach content analysis. Journal of Fish Biology, 102(5): 1049–1066, May 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Xu:2023:EPA

[1172] Yan Xu, Zhao wei Zhong, Yan Feng, Ze yu Zhang, Lu lu Ao, Hongwei Liu, Yi lei Wang, and Yong hua Jiang. Expression pattern analysis of anti-Mullerian hormone in testis development of pearlscale angelfish (Centropyge vrolikii). Journal of Fish Biology, 102(5):1067–1078, May 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oyama:2023:BSC

[1173] Takumi Oyama, Takayuki Sonoyama, Miku Kasai, Yoichi Sakai, and Tomoki Sunobe. Bidirectional sex change and plasticity of gonadal phases in the goby Lubricogobius exiguus. Journal of Fish Biology, 102(5):1079– 1087, May 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wilson:2023:TAU

[1174] Megan Wilson, Su Sponaugle, and Kirsten Grorud-Colvert. Testing assumptions underlying cabezon (Scorpaenichthys marmoratus) otolith microstructure analysis using wild-caught juveniles and opportunistic rearing of eggs and larvae. Journal of Fish Biology, 102(5):1088–1095, May 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maamela:2023:ETD

[1175] Katja S. Maamela, Eirik R. sheim, Paul V. Debes, Andrew H. House, Jaakko Erkinaro, Petra Liljeström, Craig R. Primmer, and Kenyon B. Mobley. The effect of temperature and dietary energy content on female maturation and egg nutritional content in Atlantic salmon. Journal of Fish Biology, 102(5):1096–1108, May 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shuli:2023:MDG

[1176] Zhu Shuli, Chen Weitao, Wu Zhi, Li Yuefei, Li Jie, Li Xinhui, and Yang Jiping. Mitochondrial diversity and genetic structure of common carp (Cyprinus carpio) in Pearl River and Nandujiang River. Journal of Fish Biology, 102(5):1109–1120, May 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kajgrova:2023:CST

[1177] Lenka Kajgrová, Bořek Drozd, Ján Regenda, Vladislav Draštík, Luboš Kočvara, Tomáš Kolařík, Jiří Peterka, and Petr Blabolil. A comparative study of three fishery methods for sampling the invasive topmouth gudgeon (Pseudorasbora parva) in ponds. Journal of Fish Biology, 102 (5):1121–1128, May 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davidsen:2023:BTS

[1178] Jan Grimsrud Davidsen, Andrine Emilie Halvorsen, Sindre Håvarstein Eldøy, Eva Bonsak Thorstad, and Leif Asbjørn Vøllestad. Brown trout (Salmo trutta L. 1758) and Arctic charr [Salvelinus alpinus (L. 1758)] display different marine behaviour and feeding strategies in sympatry. Journal of Fish Biology, 102(5):1129–1140, May 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crawford:2023:NDB

[1179] Rachel Crawford, Eleanor Gee, Deborah Dupont, Brendan Hicks, and Paul Franklin. No difference between critical and sprint swimming speeds for two galaxiid species. Journal of Fish Biology, 102(5):1141–1148, May 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huo:2023:GDT

[1180] Caifei Huo, Haifeng Tian, Han Rao, Bo Zeng, Debing Li, Zhong Li, and Yan Li. Genetic diversity of Toll-like receptor 9 in swamp eels (Monopterus albus). Journal of Fish Biology, 102(5):1149–1156, May 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turan:2023:PAN

[1181] Davut Turan, Esra Bayçelebi, Müfit Özuluğ, Özcan Gaygusuz, and İsmail Aksu. Phoxinus abanticus, a new species from the Lake Abant drainage in Türkiye (Teleostei: Leuciscidae). Journal of Fish Biology, 102(5):1157– 1167, May 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamada:2023:IVB

[1182] Hiroyuki Yamada and Satoshi Wada. Interpopulation variation of behavioural and morphological traits that affect downstream displacement of the juvenile white-spotted charr Salvelinus leucomaenis. Journal of Fish Biology, 102(5):1168–1176, May 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2023:UIB

[1183] Emily R. C. Smith, Richard Heal, and Louisa E. Wood. Understanding and improving biosecurity among recreational anglers in Great Britain. Journal of Fish Biology, 102(5):1177–1190, May 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mofu:2023:TEC

[1184] Lubabalo Mofu, Tatenda Dalu, Ryan J. Wasserman, Darragh J. Woodford, and Olaf L. F. Weyl. Trophic ecology of co-occurring fishes in the Sundays River Valley irrigation ponds, assessed using stable isotope and gut content analyses. Journal of Fish Biology, 102(5):1191–1205, May 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pini-Fitzsimmons:2023:DDP

[1185] Joni Pini-Fitzsimmons, Vincent Raoult, Troy Gaston, Nathan A. Knott, and Culum Brown. Diving into the diet of provisioned smooth stingrays using stable isotope analysis. Journal of Fish Biology, 102(5):1206–1218, May 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neves:2023:COS

[1186] João Neves, Ana Veríssimo, António Múrias Santos, and Susana Garrido. Comparing otolith shape descriptors for population structure inferences in a small pelagic fish, the European sardine Sardina pilchardus (Walbaum, 1792). Journal of Fish Biology, 102(5):1219–1236, May 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aya:2023:ESG

[1187] Frolan A. Aya, John Carlo L. Unida, and Luis Maria B. Garcia. Effect of size grading on growth of yellow Pacific shortfin eel (Anguilla bicolor pacifica). Journal of Fish Biology, 102(5):1237–1244, May 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kodeeswaran:2023:NDW

[1188] Paramasivam Kodeeswaran, Anil Mohapatra, Thipramalai Thangappan Pillai Ajith Kumar, and Kuldeep Kumar Lal. A new deep-water conger eel of the genus Rhynchoconger (Anguilliformes: Congridae) from the south-west coast of India, Arabian Sea. Journal of Fish Biology, 102 (5):1245–1252, May 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Verhelst:2023:FRE

[1189] Pieterjan Verhelst and Hugo Verreycken. First record in Europe of the Asian gobiid, shimofuri (marbled) goby Tridentiger bifasciatus Steindachner, 1881. Journal of Fish Biology, 102(5):1253–1255, May 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Santos:2023:RRR

[1190] Paulo R. S. Santos, Kaliandra Klafke, Otto B. F. Gadig, and Domingos Garrone-Neto. Retrieving records of a rare and threatened shark in a mosaic of marine-protected areas of southeastern Brazil. Journal of Fish Biology, 102(5):1256–1260, May 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hancock:2023:EDT

[1191] Grace M. Hancock, Gorka Sancho, Stephan B. Munch, and Santiago Salinas. Effects of daily thermal fluctuations on the Atlantic silverside, a fish with temperature-dependent sex determination. Journal of Fish Biology, 102(5):1261–1266, May 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2023:CEC

[1192] Anonymous. Correction to “Evidence of cylindrospermopsin uptake and clearance in fish (Oreochromis niloticus) under laboratory conditions”. Journal of Fish Biology, 102(5):1267, May 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic). See [662].

Anonymous:2023:IIf

[1193] Anonymous. Issue information. Journal of Fish Biology, 102(6):1269– 1270, June 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2023:DAD

[1194] William Bernard Perry. Dining amidst the decibels is not an issue when you have company: anthropogenic noise and the topmouth Gudgeon. Journal of Fish Biology, 102(6):1271, June 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hart:2023:ELO

[1195] Paul J. B. Hart. Exploring the limits to our understanding of whether fish feel pain. Journal of Fish Biology, 102(6):1272–1280, June 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schieber:2023:AGM

[1196] Jessica J. Schieber, Daniel P. Fahy, John K. Carlson, and David W. Kerstetter. Age, growth and maturity of the yellow stingray (Urobatis jamaicensis), a biannually reproductive tropical batoid. Journal of Fish Biology, 102(6):1281–1295, June 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Le:2023:SDR

[1197] Ngoc-Son Le and Thuy-Yen Duong. Sexual dimorphism and reproductive biology of Cephalocassis borneensis (Siluriformes: Ariidae), a paternal mouth-brooding fish in the Mekong River. Journal of Fish Biology, 102 (6):1296–1310, June 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arostegui:2023:EPF

[1198] Martin C. Arostegui, Michelle R. Shero, Lawrence R. Frank, Rachel M. Berquist, and Camrin D. Braun. An enigmatic pelagic fish with internalized red muscle: a future regional endotherm or forever an ectotherm? Journal of Fish Biology, 102(6):1311–1326, June 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wynne:2023:AOB

[1199] Robert Wynne, Joshka Kaufmann, Jamie Coughlan, Karl. P. Phillips, Catherine Waters, Ross W. Finlay, Ger Rogan, Russell Poole, Philip McGinnity, and Thomas E. Reed. Autumn outmigrants in brown trout (Salmo trutta) are not a demographic dead-end. Journal of Fish Biology, 102(6):1327–1339, June 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Herlevi:2023:EAN

[1200] Heidi Herlevi, Isa Wallin Kihlberg, Katri Aarnio, Erik Bonsdorff, Ann-Britt Florin, Andreas Ljung, Karl Lundström, Johanna Mattila, and Örjan Östman. Environmental abundances of the non-native round goby Neogobius melanostomus influence feeding of native fish predators. Journal of Fish Biology, 102(6):1340–1357, June 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2023:COG

[1201] Guodong Li, Ying Xiong, Xiaming Zhong, Dade Song, Zhongjie Kang, Dongjia Li, Jianhua Tang, Yanping Wang, and Lei Wu. Changes in overwintering ground of small yellow croaker (Larimichthys polyactis) based on MaxEnt and GARP models: a case study of the southern Yellow Sea stock. Journal of Fish Biology, 102(6):1358–1372, June 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rodriguez-Madrigal:2023:REJ

[1202] José Alberto Rodríguez-Madrigal, Javier Tovar-Ávila, José Leonardo Castillo-Geniz, Carlos Javier Godínez-Padilla, J. Fernando Márquez-Farías, and David Corro-Espinosa. Re-estimation of juvenile Isurus oxyrinchus growth in the Mexican Pacific through a multimodel inference approach and verification of growth band periodicity. Journal of Fish Biology, 102(6):1373–1386, June 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sithole:2023:NSP

[1203] Yonela Sithole, Tobias Musschoot, Charlotte E. T. Huyghe, Albert Chakona, and Emmanuel J. W. M. N. Vreven. A new species of Parauchenoglanis (Auchenoglanididae: Siluriformes) from the Upper Lualaba River (Upper Congo), with further evidence of hidden species diversity within the genus. Journal of Fish Biology, 102(6):1387–1414, June 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ali:2023:EML

[1204] Bulbul Ali and Abha Mishra. Effects of monochromatic lights on the melanophores arrangement in the spotted snakehead fish Channa punctata (Bloch, 1793). Journal of Fish Biology, 102(6):1415–1424, June 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2023:HTI

[1205] Brad L. Smith, Ching-Ping Lu, Maëlle Cornic, and Jaime R. Alvarado Bremer. High-throughput identification of tuna (Thunnus spp.) larvae in the Gulf of Mexico using unlabelled-probe high-resolution melting analysis. Journal of Fish Biology, 102(6):1425–1433, June 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thore:2023:SCG

[1206] Eli S. J. Thoré and Wouter Merckx. Substrate colour guides turquoise killifish’s (Nothobranchius furzeri) choice of preferred spawning habitat. Journal of Fish Biology, 102(6):1434–1441, June 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ziegler:2023:ITP

[1207] Amanda Fern Ziegler, Bodil A. Bluhm, Paul E. Renaud, and Lis L. Jørgensen. Isotopic turnover in polar cod (Boreogadus saida) muscle determined through a controlled feeding experiment. Journal of Fish Biology, 102(6):1442–1454, June 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maguire:2023:SCR

[1208] Kylie Maguire, Helen O’Neill, Franziska Althaus, William White, and Alan Williams. Seamount coral reefs are egg case nurseries for deep-sea skates. Journal of Fish Biology, 102(6):1455–1469, June 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rojas:2023:AND

[1209] Emilie Rojas, Loïc Prosnier, Aurélie Pradeau, Nicolas Boyer, and Vincent Médoc. Anthropogenic noise does not strengthen multiple-predator effects in a freshwater invasive fish. Journal of Fish Biology, 102(6):1470– 1480, June 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dominguez-Contreras:2023:MIF

[1210] José Francisco Domínguez-Contreras, Sylvia Patricia Adelheid Jiménez-Rosenberg, Nicole Reguera-Rouzaud, Laura Sánchez-Velasco, and Noé Díaz-Viloria. Molecular identification and first morphological description of the Colorado snapper Lutjanus colorado (Perciformes: Lutjanidae) larvae. Journal of Fish Biology, 102(6):1481–1491, June 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tahri:2023:YPS

[1211] Mardja Tahri and Jacques Panfili. 13-year population survey of the critically endangered European eel in the southern Mediterranean region (Algeria). Journal of Fish Biology, 102(6):1492–1502, June 2023. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moore:2023:RHR

[1212] Alec B. M. Moore. A ray of hope? The re-appearance of Irish Sea skate decades after local extinction. Journal of Fish Biology, 102(6):1503– 1505, June 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gajic:2023:FRG

[1213] Andrej A. Gajić, Emina Karalić, Hajrudin Beširović, and James Sulikowski. The first record of gravid spiny butterfly ray (Gymnura altavela) in the northern Mediterranean Sea, with description of near-term foetuses. Journal of Fish Biology, 102(6):1506–1509, June 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

AlKatrani:2023:DTD

[1214] Layla Mustafa A. A. AlKatrani. Direct transfer to different salinities and its effect on energy and metabolic enzymes in juveniles and adults of yellowfin sea bream Acanthopagrus arabicus. Journal of Fish Biology, 102(6):1510–1516, June 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2023:IIg

[1215] Anonymous. Issue information. Journal of Fish Biology, 103(1):1–2, July 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2023:MER

[1216] William Bernard Perry. Mystery egg revealing secrets from the deep. Journal of Fish Biology, 103(1):3, July 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhou:2023:CTA

[1217] Li Zhou, Fujiang Liu, Junnan Chen, Ran Yang, Jinshou Li, Zhiyong Wang, and Mingyi Cai. Comparative transcriptome analysis reveals sex bias in expression patterns of genes related to sex steroids and immunity in the skin of spinyhead croaker Collichthys lucidus. Journal of Fish Biology, 103(1):4–12, July 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Olier:2023:CCI

[1218] Bruno Silva Olier, Otávio Mesquita De Sousa, and Eduardo Gomes Sanches. Can canthaxanthin intensify the colour of the blood swordtail Xiphophorus helleri? Journal of Fish Biology, 103(1):13–21, July 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mv:2023:CDM

[1219] Sruthi Mv and Divya Lekha. Curcumin in diet modulates fatty acid levels and mRNA expression of appetite-regulating neuropeptides in brain and enhances growth via the GH-GHR-IGF axis in tilapia, Oreochromis mossambicus. Journal of Fish Biology, 103(1):22–31, July 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Segler:2023:TDP

[1220] Philipp Segler, Klaus Heinrich Vanselow, Michael Schlachter, Mario Hasler, Frederik Kaiser, and Carsten Schulz. The thermal dependence of the protein-sparing effect in rainbow trout (Oncorhynchus mykiss, Walbaum 1792). Journal of Fish Biology, 103(1):32–43, July 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ding:2023:TRL

[1221] Weidong Ding, Liping Cao, Zheming Cao, and Xuwen Bing. Transcriptomic responses of the liver of mandarin fish (Siniperca chuatsi) under hypoxic stress. Journal of Fish Biology, 103(1):44–58, July 2023. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wang:2023:EDP

[1222] Xiangbing Wang, Yaya Li, Jiabao Hu, Youyi Zhang, Man Zhang, Guanlin Wang, Huan Jiang, Jiachu Sun, Chunyang Guo, Shanliang Xu, Yang Yang, Yajun Wang, and Xiaojun Yan. Effects of different photoperiods on growth and ovarian development and maturation of silver pomfret Pampus argenteus. Journal of Fish Biology, 103(1):59–72, July 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

White:2023:WCF

[1223] William T. White, Helen L. O’Neill, Floriaan Devloo-Delva, Kazuhiro Nakaya, and Samuel P. Iglésias. What came first, the shark or the egg?

Discovery of a new species of deepwater shark by investigation of egg case morphology. Journal of Fish Biology, 103(1):73–90, July 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Souto-Santos:2023:BGA

[1224] Igor C. A. Souto-Santos, Paulo H. F. Lucinda, and Paulo A. Buckup. Bilateral genital asymmetry in livebearer fishes of the genus Phalloceros Eigenmann, 1907, with description of a new species from coastal drainages of southern Brazil (Cyprinodontiformes: Poeciliidae). Journal of Fish Biology, 103(1):91–102, July 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crispim-Rodrigues:2023:NSM

[1225] Jefferson L. Crispim-Rodrigues, Maxwell J. Bernt, Brandon T. Waltz, Gabriel S. C. Silva, Ricardo C. Benine, Claudio Oliveira, Raphaël Covain, and Fábio F. Roxo. A new species of mailed catfish of genus Rhadinoloricaria (Siluriformes: Loricariidae: Loricariinae) from Rio Negro basin, Brazil. Journal of Fish Biology, 103(1):103–112, July 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Konhamkakkada:2023:NSS

[1226] Idreesbabu Konhamkakkada, Bineesh Kinattumkara, Rajeev Raghavan, and Sureshkumar Sivanpillai. A new species of spiny eel of the genus Notacanthus Bloch 1788 (Notacanthiformes: Notacanthidae) from the Indian Ocean. Journal of Fish Biology, 103(1):113–117, July 2023. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Andrews:2023:ILH

[1227] Adam J. Andrews, David Orton, Vedat Onar, Piero Addis, Fausto Tinti, and Michelle Alexander. Isotopic life-history signatures are retained in modern and ancient Atlantic bluefin tuna vertebrae. Journal of Fish Biology, 103(1):118–129, July 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kodeeswaran:2023:NSC

[1228] Paramasivam Kodeeswaran, David. G. Smith, Thipramalai Thangappan Pillai Ajith Kumar, and Uttam Kumar Sarkar. A new species of the conger eel genus Macrocephenchelys from deep waters of the Arabian Sea. Journal of Fish Biology, 103(1):130–135, July 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Silva-Oliveira:2023:KBN

[1229] Cárlison Silva-Oliveira, André L. Colares Canto, and Frank Raynner V. Ribeiro. Knodus borari, a new species from the lower Rio Tapajós basin (Characiformes: Characidae). Journal of Fish Biology, 103(1):136–142, July 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sanderson:2023:FFM

[1230] Sarah Sanderson, Louis Astorg, Grant E. Haines, Sandrine Beaumont-Courteau, R. Brian Langerhans, Alison M. Derry, and Andrew P. Hendry. Freshwater fishes maintain multi-trait phenotypic stability across an environmental gradient in aqueous calcium. Journal of Fish Biology, 103 (1):143–154, July 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henschel:2023:TNM

[1231] Elisabeth Henschel, Willian M. Ohara, and Wilson J. E. M. Costa. Two new miniature translucent catfish species of the rare genus Tridens (Siluriformes: Trichomycteridae) from the Madeira River basin, northern Brazil. Journal of Fish Biology, 103(1):155–171, July 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cruz:2023:CIE

[1232] Marcelo Merten Cruz, Thomas Sauvage, Anthony Chariton, and Thales Renato Ochotorena de Freitas. The challenge of implementing environmental DNA metabarcoding to detect elasmobranchs in a resourcelimited marine protected area. Journal of Fish Biology, 103(1):172–178, July 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Boardman:2023:EPE

[1233] Rose M. Boardman, Adrian C. Pinder, Adam T. Piper, Catherine Gutmann Roberts, Rosalind M. Wright, and John Robert Britton. Effects of preservation by ethanol on δ ^{13}C and δ ^{15}N of three tissues of the critically endangered European eel Anguilla anguilla. Journal of Fish Biology, 103(1):179–182, July 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Escobar-Camacho:2023:VOG

[1234] Daniel Escobar-Camacho, Wilson Andrade, Nathalia Valencia, Andrea C. Encalada, and Diana A. Pazmiño. Validating the occurrence of the giant mottled eel (Anguilla marmorata) in the Galápagos Islands.

Journal of Fish Biology, 103(1):183–188, July 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ste-Marie:2023:FRO

[1235] Eric Ste-Marie, Henrik Glenner, David J. Rees, and Nigel E. Hussey. First recorded occurrence of the parasitic barnacle (Anelasma squalicola) on a Greenland shark (Somniosus microcephalus) in the Canadian Arctic. Journal of Fish Biology, 103(1):189–193, July 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Parker:2023:LML

[1236] Ben Parker, Demetra Andreou, Iain D. Green, Katsiaryna Pabortsava, Rose M. Boardman, Adrian C. Pinder, Rosalind M. Wright, and Robert Britton. Low microplastic loads in riverine European eel (Anguilla anguilla) from southwest England during their marine–freshwater transition. Journal of Fish Biology, 103(1):194–198, July 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2023:IIh

[1237] Anonymous. Issue information. Journal of Fish Biology, 103(2):199– 200, August 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2023:EMO

[1238] William Bernard Perry. Even model organisms stress about moving house. Journal of Fish Biology, 103(2):201, August 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whitfield:2023:OGM

[1239] Alan K. Whitfield and Jean-Dominique Durand. An overview of grey mullet (Mugilidae) global occurrence and species-rich ecoregions, with indications of possible past dispersal routes within the family. Journal of Fish Biology, 103(2):202–219, August 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yu:2023:EBA

[1240] Meng-Nan Yu, Wen li Zhu, Si bu Wang, Dong-Ming Zhang, Yu-Rou Zhang, Qiu-Ju Wang, Gui-Qin Wang, Mahmoud M. Elsadek, Qi Yao, Yu-Ke Chen, and Zhi-Xin Guo. Effect of Bacillus amyloliquefaciens LSG2-8 on the intestinal barrier function of Amur minnow (Rhynchocypris lagowskii). Journal of Fish Biology, 103(2):220–234, August 2023. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wheeler:2023:NAE

[1241] Carolyn R. Wheeler, Duncan J. Irschick, John W. Mandelman, and Jodie L. Rummer. Nonlethally assessing elasmobranch ontogenetic shifts in energetics. Journal of Fish Biology, 103(2):235–246, August 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ximenes:2023:MAR

[1242] Aline Mourão Ximenes, Valéria Nogueira Machado, Emil José Hernandez-Ruz, Fábio de Lima Muniz, Tomas Hrbek, and Izeni Pires Farias. Multilocus approach reveals distinct evolutionary units of the South American apapa Pellona flavipinnis (Valenciennes, 1837) (Clupeiformes, Pristigasteridae). Journal of Fish Biology, 103(2):247–259, August 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yano:2023:HRP

[1243] Toshikazu Yano, Tsutomu Hattori, Yoji Narimatsu, Jiro Nagao, and Sho Tanaka. High reproductive potential of the commercially exploited shark Squalus suckleyi in the major fishing grounds of the western North Pacific. Journal of Fish Biology, 103(2):260–271, August 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coffill-Rivera:2023:ETS

[1244] Manuel E. Coffill-Rivera, Yvanna D. Paez Mendez, Logan Little, Patrick M. Graham, James S. Franks, Sandra Bibiana Correa, J. Wesley Neal, and Peter J. Allen. The effects of temperature and salinity on the blood chemistry and survival of juvenile Atlantic tarpon Megalops atlanticus. Journal of Fish Biology, 103(2):272–279, August 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lawrence:2023:MCI

[1245] Michael J. Lawrence, Tanya S. Prystay, Melissa Dick, Erika J. Eliason, Chris K. Elvidge, Scott G. Hinch, David A. Patterson, Andrew G. Lotto, and Steven J. Cooke. Metabolic constraints and individual variation shape the trade-off between physiological recovery and anti-predator responses in adult sockeye salmon. Journal of Fish Biology, 103(2):280– 291, August 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sharma:2023:SSC

[1246] Partigya Sharma, Shriya Purohit, Gaurav Bhatt, Sachin Kothiyal, Mahender Singh, Prakash Nautiyal, and Indrashis Bhattacharya. A study on the seasonal cyclicity of ovarian development in adult Himalayan snow trout, Schizothorax plagiostomus. Journal of Fish Biology, 103(2):292– 304, August 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Froman:2023:UUC

[1247] Niv Froman, Marie-Aude Genain, Guy M. W. Stevens, and Gareth P. Pearce. Use of underwater contactless ultrasonography to elucidate the internal anatomy and reproductive activity of manta and devil rays (family: Mobulidae). Journal of Fish Biology, 103(2):305–323, August 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anderson:2023:HUF

[1248] Antônio B. Anderson, Hudson T. Pinheiro, Ronaldo B. Francini-Filho, Carlos E. L. Ferreira, and Jean-Christophe Joyeux. Habitat use of five sympatric predatory reef fishes at a remote island in the south-western Atlantic. Journal of Fish Biology, 103(2):324–335, August 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schneider:2023:ART

[1249] Eric V. C. Schneider, Zachary C. Zuckerman, Brendan S. Talwar, Steven J. Cooke, Aaron D. Shultz, and Cory D. Suski. Aerobic response to thermal stress across ontogeny and habitats in a teleost fish. Journal of Fish Biology, 103(2):336–346, August 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huisman:2023:FID

[1250] Jeroen B. J. Huisman, Leander Höhne, Reinhold Hanel, Henry Kuipers, Peter Paul Schollema, and Leopold Nagelkerke. Factors influencing the downstream passage of European silver eels (Anguilla anguilla) through a tidal sluice. Journal of Fish Biology, 103(2):347–356, August 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alves:2023:SES

[1251] Aline Cristina Alves, Carlos Eduardo Leite Ferreira, Ricardo de Souza Rosa, Hudson Tercio Pinheiro, and Ronaldo Bastos Francini-Filho. Spawning ecology of the Saint Paul’s Gregory Stegastes sanctipauli, a damselfish endemic to the remote St Peter and St Paul’s Archipelago. Journal of Fish Biology, 103(2):357–366, August 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Philippe:2023:CHS

[1252] Charlotte Philippe, Lucia Vergauwen, Kato Huyghe, Gudrun De Boeck, and Dries Knapen. Chronic handling stress in zebrafish Danio rerio husbandry. Journal of Fish Biology, 103(2):367–377, August 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hunt:2023:QST

[1253] Georgina L. Hunt, Matthew R. D. Cobain, Rona A. R. McGill, John K. Pinnegar, Georg H. Engelhard, Benjamin D. Wigham, and Nicholas V. C. Polunin. Quantifying spatio-temporal consistency in the trophic ecology of two sympatric flatfishes. Journal of Fish Biology, 103(2):378–392, August 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kottillil:2023:ICC

[1254] Shruthi Kottillil, Trisha Gupta, Muralidharan Manoharakrishnan, Chetan Rao, and Kartik Shanker. Insights from catch composition and historical records of elasmobranchs in the Malvan fishery on the central west coast of India. Journal of Fish Biology, 103(2):393–410, August 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nieder:2023:OCT

[1255] Carolin Nieder, Jimmy Rapson, Kim Holland, Carl Meyer, John Montgomery, and Craig Radford. Operant conditioning as a tool to assess hearing abilities in sharks. Journal of Fish Biology, 103(2):411–424, August 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Azevedo:2023:MTA

[1256] Ana Azevedo, Luiz C. Navarro, Thaís Cavalheri, Hugo G. Santos, Irene Martins, and Rodrigo Ozório. Morphological traits for allometric scaling of the European sea bass Dicentrarchus labrax (Linnaeus, 1758) from Southern Portugal population. Journal of Fish Biology, 103(2):425–438, August 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jimenez-Alvarado:2023:FRP

[1257] David Jimenez-Alvarado, Eva Meyers, Airam Guerra-Marrero, Ana Espino-Ruano, Lorena Couce-Montero, Joanna Barker, Michael Sealey, Héctor Toledo, Belén Caro, Charlotte Pike, David M. P. Jacoby, Lucy Mead, and José J. Castro. First record of partial albinism in the critically endangered Angelshark (Squatina squatina) (Linnaeus, 1758). Journal of Fish Biology, 103(2):439–442, August 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wiley:2023:LHE

[1258] John Wiley and Cassandra Pardee. Life history of the endemic Hawaiian hogfish Bodianus albotaeniatus: age, growth and reproduction. Journal of Fish Biology, 103(2):443–447, August 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whitehead:2023:FRB

[1259] Darren A. Whitehead and Joel Gayford. First record of bottom-feeding behaviour in the whale shark (Rhincodon typus). Journal of Fish Biology, 103(2):448–452, August 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rodrigues:2023:ERA

[1260] Alexandre Rodrigues, Wesley Shkola, and Bianca S. Rangel. Extending records of albinism and skin disorders in American cownose rays to southeastern Brazil. Journal of Fish Biology, 103(2):453–456, August 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2023:IIi

[1261] Anonymous. Issue information. Journal of Fish Biology, 103(3):457–458, September 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2023:CSC

[1262] William Bernard Perry. The cold sting of climate change and its effect on the march of invasive fishes. Journal of Fish Biology, 103(3):459, September 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huang:2023:FAC

[1263] Kai Huang and Guoping Zhu. Fatty acid composition and energy allocation in muscle and gonad tissues indicate that the female mackerel icefish Champsocephalus gunnari is an income breeder. Journal of Fish Biology, 103(3):460–471, September 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lopez:2023:LTP

[1264] Steven G. Lopez, Andrea S. Aspbury, Sarah R. Fritts, Travis Tidwell, and Timothy H. Bonner. Long-term patterns in inland fish kills associated with cold-shock and winter stress: a regional case study from Texas. Journal of Fish Biology, 103(3):472–480, September 2023. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carnicer:2023:SSD

[1265] Cleide Carnicer, Luciano B. Lima, Tátia Leika Taguti, Fagner Junior M. Oliveira, and Dilermando Pereira Lima-Junior. Spatial and seasonal distribution of fish eggs and larvae in one free-flowing river in the Neotropical savanna. Journal of Fish Biology, 103(3):481–495, September 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mahapatra:2023:RBC

[1266] Smruti Mahapatra, Bristy Ganguly, Saswati Pani, Ashis Saha, and Mrinal Samanta. Red blood cells of Labeo rohita express Toll-like receptors, NOD-like receptors, interleukins, and interferon-I in response to Gramnegative bacterial infections and lipopolysaccharide stimulations. Journal of Fish Biology, 103(3):496–506, September 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gao:2023:BFA

[1267] Shike Gao, Xiao Zhang, Ruilin Shu, Shuo Zhang, Jikun Lu, and Guanghui Fu. Bilateral fluctuation asymmetry of otoliths of Collichthys lucidus in different functional areas of Haizhou Bay. Journal of Fish Biology, 103(3):507–515, September 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garbett:2023:HCD

[1268] Amy Garbett, Sophie L. Loca, Thomas Barreau, Manuel Biscoito, Caroline Bradley, Joe Breen, Maurice Clarke, Jim R. Ellis, Andrew M. Griffiths, Gary Hannon, Klara Jakobsdóttir, Claudia Junge, Arve Lynghammar, Matthew McCloskey, George Minos, Natasha D. Phillips, Paulo A. Prodöhl, William Roche, Samuel P. Iglésias, James Thorburn, and Patrick C. Collins. A holistic and comprehensive data approach validates the distribution of the critically endangered flapper skate (Dipturus intermedius). Journal of Fish Biology, 103(3):516–528, September 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhao:2023:SEA

[1269] Yuhe Zhao, Yang Liu, Jianfen Wu, Dan Kong, Sifan Zhao, Guopan Li, and Wei Li. Swamp eel aldehyde reductase is involved in response to nitrosative stress via regulating NO/GSH levels. Journal of Fish Biology, 103(3):529–543, September 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Briones-Hernandez:2023:MPD

[1270] Sergio A. Briones-Hernández, Ulianov Jakes-Cota, Fernando R. Elorriaga-Verplancken, Felipe Galván-Magaña, John M. Logan, John O’Sullivan, José Miguel Carvajal-Rodríguez, Esteban Elias, Joanna Alfaro-Shigueto, and Sofía Ortega-García. Movement and population dispersal of dolphinfish (Coryphaena hippurus) across the Eastern Tropical Pacific inferred from carbon and nitrogen stable isotope analyses. Journal of Fish Biology, 103(3):544–556, September 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Manda:2023:RNL

[1271] Bauchet Katemo Manda, Jos Snoeks, Eva Decru, Jonathan Brecko, and Emmanuel J. W. M. N. Vreven. Revision of Nannocharax luapulae Boulenger, 1915 (Characiformes: Distichodontidae) from the Upper Congo basin: Evidence for a species pair. Journal of Fish Biology, 103 (3):557–573, September 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leblanc:2023:IWL

[1272] John Paul Leblanc and John M. Farrell. Influence of water-level variability on fish assemblage and natural reproduction following connectivity enhancement in a Typha-dominated coastal wetland, USA. Journal of Fish Biology, 103(3):574–592, September 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huang:2023:DNM

[1273] Wen-Chien Huang, Rodulf Anthony Balisco, Florence Chan Evacitas, and Te-Yu Liao. Description of a new moray eel (Anguilliformes: Muraenidae) from the Philippines, with a preliminary assessment of genetic and morphological variations of the Uropterygius concolor species complex. Journal of Fish Biology, 103(3):593–602, September 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hashida:2023:HDE

[1274] Daisuke Hashida and Takeshi Tomiyama. Hatch date and early growth of juvenile Japanese jack mackerel Trachurus japonicus caught in the Uwa Sea. Journal of Fish Biology, 103(3):603–611, September 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nemoto:2023:UAR

[1275] Yuki Nemoto and Tomoyuki Ogawa. Underwater analysis of the reproductive behaviour and egg development of the jawfish Opistognathus iyonis. Journal of Fish Biology, 103(3):612–622, September 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shiffman:2023:PPS

[1276] David Shiffman, Charles Bangley, and Catherine Macdonald. “A prized Pacific shark”: the rise and fall (and rise again… ?) of the world’s first ecolabel certified sustainable shark fishery. Journal of Fish Biology, 103 (3):623–634, September 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Camilo:2023:ASG

[1277] Laura de Oliveira Camilo, Bruna Larissa Maganhe, Hugo Gallo Neto, Silvia Edelweiss Crusco, Carlos Eduardo Malavasi-Bruno, and Eduardo Gomes Sanches. Andrology of shortnose guitarfish Zapteryx brevirostris (Müller & Henle, 1841) (Chondrichthyes, Trygonorrhinidae). Journal of Fish Biology, 103(3):635–645, September 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2023:CDP

[1278] Geoffrey H. Smith and Debra J. Murie. Comparative diets and prey resource competition between early-juvenile common snook Centropomus undecimalis and non-native pike killifish Belonesox belizanus. Journal of Fish Biology, 103(3):646–665, September 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reis:2023:ASF

[1279] İsmail Reis, Celal Ateş, and Laith Jawad. The asymmetry in the sagitta of four mugilid species obtained from Köyceğiz Lagoon, Aegean Sea, Turkey. Journal of Fish Biology, 103(3):666–674, September 2023. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ruth:2023:PET

[1280] Andreas Ruth, Morten Bo Søndergaard Svendsen, Rasmus Nygaard, Emil Aputsiaq Flindt Christensen, Peter G. Bushnell, and John Fleng Steffensen. Physiological effects of temperature on Greenland halibut Reinhardtius hippoglossoides shows high vulnerability of Arctic stenotherms to global warming. Journal of Fish Biology, 103(3):675–683, September 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jensen:2023:LBT

[1281] Mikkel Jensen, Peter J. Nielsen, and G. Glenn Wilson. Links between the timing of life-history transitions and dietary and morphological variation during early life history in the sand goby, Pomatoschistus minutus. Journal of Fish Biology, 103(3):684–694, September 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2023:EIK

[1282] William Bernard Perry. The environmental impact of keeping a tropical aquarium in Northern Europe. Journal of Fish Biology, 103(3):695–703, September 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2023:GEU

[1283] Samantha K. Smith, Galina Aglyamova, Erik W. Oberg, Lee A. Fuiman, and Mikhail V. Matz. Gene expression underlying variation in the survival skills of red drum larvae (Sciaenops ocellatus). Journal of Fish Biology, 103(3):704–714, September 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ahmad:2023:ILL

[1284] Muhammad Talba Ahmad, Mohamed Shariff, Yong Meng Goh, Sanjoy Banerjee, and Fatimah M. Yusoff. Interaction of low-level dietary supplementation of Chlorella vulgaris Beijerinck, 1890, and feeding duration on growth hormone, growth performance and serum biochemistry of red hybrid tilapia (Oreochromis mossambicus × Oreochromis niloticus). Journal of Fish Biology, 103(3):715–726, September 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dumas:2023:CRM

[1285] Francesca Dumas, Manuela Mauro, Mirella Vazzana, Vincenzo Arizza, and Aiti Vizzini. Ciona robusta macrophage migration inhibitory factor (Mif1 and Mif2) genes are differentially regulated in the lipopolysaccharide-challenged pharynx. Journal of Fish Biology, 103(3): 727–730, September 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wargat:2023:LSH

[1286] Bryanna N. Wargat, Douglas H. Adams, and Maria Laura Habegger. Leucism in sharks: a histological examination. Journal of Fish Biology, 103(3):731–734, September 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jewell:2023:BWM

[1287] Oliver J. D. Jewell, Ben D’Antonio, Stacy Blane, Emily Gosden, Michael D. Taylor, Hannah J. Calich, Matthew W. Fraser, and Ana M. M. Sequeira. Back to the wild: movements of a juvenile tiger shark released from a public aquarium. Journal of Fish Biology, 103(3):735–740, September 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeWeirdt:2023:ODS

[1288] David De Weirdt, Charlotte Elisabeth Theodora Huyghe, and Emmanuel Jean Michel Willem Nzambemalamu Vreven. On the occurrence and distribution of Synodontis nigrita and Synodontis violaceus (Siluriformes: Mochokidae) in the Nilo-Sudan, and Upper and Lower Guinea ichthyofaunal provinces. Journal of Fish Biology, 103(3):741–746, September 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2023:IIj

[1289] Anonymous. Issue information. Journal of Fish Biology, 103(4):747–748, October 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barber:2023:CFF

[1290] Iain Barber, Andrew G. Hirst, and Carl Smith. The challenges faced by fish in a dynamic world. Journal of Fish Biology, 103(4):749–751, October 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Britton:2023:CPE

[1291] John Robert Britton. Contemporary perspectives on the ecological impacts of invasive freshwater fishes. Journal of Fish Biology, 103(4):752– 764, October 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sundin:2023:EOA

[1292] Josefin Sundin. The effects of ocean acidification on fishes — history and future outlook. Journal of Fish Biology, 103(4):765–772, October 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Candolin:2023:BRF

[1293] Ulrika Candolin and Tawfiqur Rahman. Behavioural responses of fishes to anthropogenic disturbances: Adaptive value and ecological consequences. Journal of Fish Biology, 103(4):773–783, October 2023. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Geffroy:2023:SMH

[1294] Benjamin Geffroy, Leydy Sandoval-Vargas, Myriam Boyer-Clavel, Maritza Pérez-Atehortúa, Stephane Lallement, and Iván Valdebenito Isler. A simulated marine heatwave impacts European sea bass sperm quantity, but not quality. Journal of Fish Biology, 103(4):784–789, October 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yeung:2023:BME

[1295] Emily J. Yeung, Sandra Klemet-N’Guessan, Thomas J. Hossie, and Michael G. Fox. Boldness, movement and exploration tendency in round goby (Neogobius melanostomus) in Southern Ontario. Journal of Fish Biology, 103(4):790–797, October 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cooper:2023:STP

[1296] Jack A. Cooper, John N. Griffin, René Kindlimann, and Catalina Pimiento. Are shark teeth proxies for functional traits? A framework to infer ecology from the fossil record. Journal of Fish Biology, 103(4):798– 814, October 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Santos:2023:TAI

[1297] Régis Santos, Wendell Medeiros-Leal, and Mário Pinho. Trends in abundance indices derived from commercial fisheries for priority marine stocks in the Azores. Journal of Fish Biology, 103(4):815–827, October 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alfonso:2023:WTE

[1298] Sébastien Alfonso, Camille Houdelet, Eduardo Bessa, Benjamin Geffroy, and Bastien Sadoul. Water temperature explains part of the variation in basal plasma cortisol level within and between fish species. Journal of Fish Biology, 103(4):828–838, October 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grigg:2023:TVI

[1299] A. G. Grigg, T. M. Lowi-Merri, J. A. Hutchings, and M. D. Massey. Thermal variability induces sex-specific morphometric changes in zebrafish (Danio rerio). Journal of Fish Biology, 103(4):839–850, October 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2023:IIk

[1300] Anonymous. Issue information. Journal of Fish Biology, 103(5):851–852, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2023:UDA

[1301] William Bernard Perry. The ups and downs of Atlantic halibut spawning. Journal of Fish Biology, 103(5):853, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gomez-Maldonado:2023:ITL

[1302] Sebastian Gomez-Maldonado, Adrian Calleros, Isabel Salazar-Rueda, and Morelia Camacho-Cervantes. The invasive twospot livebearer’s biology, and its current and potential global distribution. Journal of Fish Biology, 103(5):854–863, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Waller:2023:DMC

[1303] Matt J. Waller, Nuno Queiroz, Ivo da Costa, Tiago Cidade, Bruno Loureiro, Freya C. Womersley, Jorge Fontes, Pedro Afonso, Bruno C. L. Macena, Alexandra Loveridge, Nicolas E. Humphries, Emily J. Southall, and David W. Sims. Direct measurement of cruising and burst swimming speeds of the shortfin mako shark (Isurus oxyrinchus) with estimates of field metabolic rate. Journal of Fish Biology, 103(5):864–883, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mainguy:2023:OIP

[1304] Julien Mainguy, Ariel Arsenault, Lilian Tran, Mackenzie A. C. Martyniuk, Catherine Paquet, Jean-Sébastien Moore, and Michael Power. Otolith-inferred patterns of marine migration frequency in Nunavik Arctic charr. Journal of Fish Biology, 103(5):884–896, November 2023. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jothinarayanan:2023:CEL

[1305] Nivedhitha Jothinarayanan, Frank Karlsen, and Lars Eric Roseng. Comparative evaluation of loop-mediated isothermal amplification and PCR for detection of Esox lucius housekeeping genes for use in on-site environmental monitoring. Journal of Fish Biology, 103(5):897–905, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kottmann:2023:WPB

[1306] Johanna S. Kottmann, Gerd M. Berge, Katerina Kousoulaki, Tone-Kari Knutsdatter Østbye, Elisabeth Ytteborg, Bjarne Gjerde, and Ingrid Lein. Welfare and performance of ballan wrasse (Labrus bergylta) reared at two different temperatures after a preparatory feeding trial with enhanced dietary eicosapentaenoic acid. Journal of Fish Biology, 103(5): 906–923, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gendron:2023:RHH

[1307] Matthew E. Gendron and Peter M. Buston. Review of housing and husbandry of anemonefishes for use in research. Journal of Fish Biology, 103 (5):924–938, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deMeo:2023:RPR

[1308] Ilaria de Meo, Kjartan Østbye, Kimmo K. Kahilainen, and Antonio B. S. Poléo. The role of predation risk in structuring life-history traits of crucian carp (Carassius carassius) in a series of small boreal lakes. Journal of Fish Biology, 103(5):939–949, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Andrade:2023:GVL

[1309] Hector Andrade, Mikko Vihtakari, and Jorge Santos. Geographic variation in the life history of lane snapper Lutjanus synagris, with new insights from the warm edge of its distribution. Journal of Fish Biology, 103(5):950–964, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oh:2023:DNH

[1310] Han Young Oh, So Ryung Shin, Jung Jun Park, Hyeon Jin Kim, and Jung Sick Lee. Distribution of nephrons in the head kidney of three species of Sebastes (Teleostei: Scorpaenidae). Journal of Fish Biology, 103(5):965–973, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jones:2023:MDS

[1311] Nick A. R. Jones, Daphne Cortese, Amelia Munson, Helen C. Spence-Jones, Zoe Storm, Shaun S. Killen, Ruth Bethel, Amy E. Deacon, Mike M. Webster, and Libor Závorka. Maze design: size and number of choices impact fish performance in cognitive assays. Journal of Fish Biology, 103(5):974–984, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Frolova:2023:FIA

[1312] Tatyana V. Frolova, Evgeny I. Izvekov, and Galina I. Izvekova. First insights into the activity of major digestive enzymes in the intestine of the European catfish Silurus glanis and protective anti-enzymatic potential of its gut parasite Silurotaenia siluri. Journal of Fish Biology, 103(5): 985–993, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Spanik:2023:VBT

[1313] Kevin R. Spanik and Joseph C. Ballenger. Validating blueline tilefish Caulolatilus microps ages in the U.S. South Atlantic using bomb radiocarbon (F ^{14}C). Journal of Fish Biology, 103(5):994–1002, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johne:2023:MER

[1314] Alexandra S. Johne, Chris G. Carter, Simon Wotherspoon, Scott Hadley, Jane E. Symonds, Seumas P. Walker, and Julia L. Blanchard. Modeling the effects of ration on individual growth of Oncorhynchus tshawytscha under controlled conditions. Journal of Fish Biology, 103(5):1003–1014, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cortes-Hernandez:2023:NCS

[1315] Miguel Ángel Cortés-Hernández, Jeisson Alexis López-Castaño, Nadia Milani, and Carlos DoNascimiento. A new cryptic species of Imparfinis (Siluriformes: Heptapteridae) from the Orinoco River basin, revealed by an iterative approach. Journal of Fish Biology, 103(5):1015–1030, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marshall:2023:RBA

[1316] Rachel C. Marshall, Jonathan A. D. Fisher, Anthony L. Einfeldt, Paul Gatti, Dominique Robert, and Arnault Le Bris. Reproductive behavior of Atlantic halibut (Hippoglossus hippoglossus) interpreted from electronic tags. Journal of Fish Biology, 103(5):1031–1043, November 2023. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Adjei:2023:ABB

[1317] Henry Owusu Adjei, Robert Yadama Laar, Prince Ofori-Darkwah, Emmanuel Xorla Xatse, Jedida Osei Bediako, Peter Vilhelm Skov, and Kwasi Adu Obirikorang. Air-breathing behavior in Heterotis niloticus fingerlings: Response to changes in oxygen, temperature, and exercise regimes. Journal of Fish Biology, 103(5):1044–1053, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cruz-Castan:2023:ODA

[1318] Roberto Cruz-Castán, Sámar Saber, César Meiners-Mandujano, María José Gómez Vives, Sergio Curiel-Ramirez, Gabriela Galindo-Cortes, and David Macías. Oocyte dynamics for Atlantic bluefin tuna (Thunnus thynnus) in the southern Gulf of Mexico: a comparative assessment of Mediterranean spawners using two stereological methods. Journal of Fish Biology, 103(5):1054–1072, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Urbano-Bonilla:2023:NSR

[1319] Alexander Urbano-Bonilla, Alejandro Londoño-Burbano, and Tiago P. Carvalho. A new species of rheophilic armored catfish of Rineloricaria (Siluriformes: Loricariidae) from the Vaupés River, Amazonas basin, Colombia. Journal of Fish Biology, 103(5):1073–1084, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Niedrist:2023:URN

[1320] G. H. Niedrist, A. Hilpold, and P. Kranebitter. Unveiling the rise of non-native fishes in eastern alpine mountain rivers: Population trends and implications. Journal of Fish Biology, 103(5):1085–1094, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

vanderHeever:2023:MCD

[1321] G. M. van der Heever, D. Yemane, R. W. Leslie, C. D. van der Lingen, and M. J. Gibbons. Modeling and comparing the distributions and associations of two co-occurring catshark species off South Africa. Journal of Fish Biology, 103(5):1095–1105, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turan:2023:OMN

[1322] Davut Turan, Esra Bayçelebi, and Gökhan Kalaycı. Oxynoemacheilus marmaraensis, a new species from the Susurluk River, Türkiye (Teleostei: Nemacheilidae). Journal of Fish Biology, 103(5):1106–1112, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bouchard:2023:EMS

[1323] C. Bouchard and D. Nicolas. Estimating migration speed of glass eels during their colonization of a Mediterranean lagoon. Journal of Fish Biology, 103(5):1113–1121, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marinho:2023:NCS

[1324] Manoela Maria Ferreira Marinho and Fernando César Paiva Dagosta. On a new colorful species of Moenkhausia (Characiformes: Characidae) from the upper rio Madeira basin at the Chapada dos Parecis, Brazil, with comments on its conservation and putative biogeographic history. Journal of Fish Biology, 103(5):1122–1128, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gerber:2023:SIA

[1325] Ruan Gerber, Lizaan de Necker, Johan H. J. van Vuren, Yoshinori Ikenaka, Shouta M. M. Nakayama, Mayumi Ishizuka, Victor Wepener, and Nico J. Smit. A stable isotope analysis of the dietary patterns of the aquatic apex predator, the African tigerfish (Hydrocynus vittatus). Journal of Fish Biology, 103(5):1129–1143, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bonjour:2023:MTT

[1326] Sophia M. Bonjour, Keith B. Gido, Mark C. McKinstry, Charles N. Cathcart, Matthew R. Bogaard, Maria Dzul, Brian D. Healy, Zachary E. Hooley-Underwood, David L. Rogowski, and Charles B. Yackulic. Migration timing and tributary use of spawning flannelmouth sucker (Catostomus latipinnis). Journal of Fish Biology, 103(5):1144–1162, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ching:2023:UAH

[1327] Cody Ching, Jessica A. Miller, Yinphan Tsang, Kauaoa Fraiola, Hannah Clilverd, and Shaya Honarvar. Understanding amphidromy in Hawai’i: ’o’opu nākea (Awaous stamineus). Journal of Fish Biology, 103(5):1163– 1177, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ekman:2023:ENI

[1328] Drew R. Ekman, Marina G. Evich, Jonathan D. Mosley, Jon A. Doering, Kellie A. Fay, Gerald T. Ankley, and Timothy W. Collette. Expanding non-invasive approaches for fish-health monitoring: a survey of the epidermal mucous metabolomes of phylogenetically diverse freshwater fish species. Journal of Fish Biology, 103(5):1178–1189, November 2023. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cheyadmi:2023:PSR

[1329] Soumaya Cheyadmi, Housni Chadli, Bouchra El Yamlahi, Mohammed El Maadoudi, Francesco Cacciola, Ayoub Kounnoun, Hassan Nhhala, and Hicham Chairi. Physiological stress responses of European sea bass Dicentrarchus labrax related to changing operations of floating net cages in intensive farming conditions at the Moroccan M’diq Bay. Journal of Fish Biology, 103(5):1190–1198, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Verocai:2023:FFS

[1330] José E. Verocai, Fernanda Cabrera, Antoni Lombarte, and Walter Norbis. Form function of Sulcus acusticus of the Sagittal otolith in seven Sciaenidae (Acanthuriformes) species using geometric morphometrics (southwestern Atlantic). Journal of Fish Biology, 103(5):1199–1213, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Claassens:2023:FRP

[1331] Louw Claassens, Brennan Phillips, David A. Ebert, Denley Delaney, Brad Henning, Victor Nestor, Adrian Ililau, and Jonatha Giddens. First records of the Pacific sleeper shark Somniosus cf. pacificus in the western tropical Pacific. Journal of Fish Biology, 103(5):1214–1220, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garita-Alvarado:2023:CBK

[1332] Carlos A. Garita-Alvarado. Courtship behavior in the knife livebearer Alfaro cultratus (Cyprinodontiformes: Poeciliidae). Journal of Fish Biology, 103(5):1221–1225, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Freedman:2023:EIJ

[1333] Ryan M. Freedman, James M. Anderson, Chris Caldow, Brian Stirling, Patrick Rex, Emily Spurgeon, Steve McCullough, Kady Lyons, Jack May III, Connor F. White, Ryan K. Logan, Emily Meese, Echelle S. Burns, Alyssa J. Clevenstine, John O’ Sullivan, Chuck Winkler, Lizzie Duncan, Mari Cajandig, and Christopher G. Lowe. Evidence of increasing juvenile white sharks’ (Carcharodon carcharias) habitat use at the Northern Channel Islands. Journal of Fish Biology, 103(5):1226–1231, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nishino:2023:FMA

[1334] Daiki Nishino, Takayoshi Nishida, and Kohei Yoshiyama. Feeding of mussel-associated leeches Hemiclepsis kasmiana on bitterling embryos: Novel interaction between parasites in a shared host. Journal of Fish Biology, 103(5):1232–1236, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pastana:2023:NGH

[1335] Murilo N. L. Pastana, G. David Johnson, Bruce Mundy, and Alessio Datovo. A new gelatinous host for pelagic fishes: First in situ record of an association between driftfishes (Stromateiformes, Ariommatidae) and nudibranchs (Mollusca, Phylliroidae). Journal of Fish Biology, 103(5): 1237–1241, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Parmegiani:2023:FRB

[1336] Andrea Parmegiani, Jacopo Gobbato, Davide Seveso, Paolo Galli, and Simone Montano. First record of the bull shark Carcharhinus leucas (Valenciennes, 1839) from the Maldivian archipelago, central Indian Ocean. Journal of Fish Biology, 103(5):1242–1247, November 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2023:IIl

[1337] Anonymous. Issue information. Journal of Fish Biology, 103(6):1249– 1250, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2023:LSP

[1338] William Bernard Perry. Limited summer palate with no colonic irrigation required: life of a migratory shark through the metabarcoding lens. Journal of Fish Biology, 103(6):1251, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zarini:2023:ART

[1339] Sina Zarini, Federica Poli, and Sigal Balshine. Alternative reproductive tactics in goby fishes of the Caspian Sea. Journal of Fish Biology, 103(6): 1252–1263, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rohner:2023:MVH

[1340] Christoph A. Rohner, Alberto C. Naveira Garabato, and Alex R. Hearn. Movement and vertical habitat use of yellowfin tuna Thunnus albacares in a vertically compressed habitat: the Galápagos Marine Reserve. Journal of Fish Biology, 103(6):1264–1276, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Benavente:2023:UBM

[1341] Javiera N. Benavente, David Véliz, Claudio Quezada-Romegialli, and Daniel Gomez-Uchida. Uniparental and biparental markers unravel invasion pathways, population admixture, and genetic structure in naturalized rainbow trout (Oncorhynchus mykiss). Journal of Fish Biology, 103(6):1277–1288, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhang:2023:ITD

[1342] Qiyu Zhang, Xiaoqing Liu, Haoran Shi, and Yao Yang. Interaction of total dissolved gas supersaturation and suspended sediment on swimming performance of bighead carp (Hypopthalmichthys nobilis). Journal of Fish Biology, 103(6):1289–1299, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pinion:2023:LAO

[1343] Amanda K. Pinion, Ralf Britz, Kole M. Kubicek, Dustin S. Siegel, and Kevin W. Conway. The larval attachment organ of the bowfin Amia ocellicauda Richardson, 1836 (Amiiformes: Amiidae) and its phylogenetic significance. Journal of Fish Biology, 103(6):1300–1311, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Takahashi:2023:UBO

[1344] Miwa Takahashi, Corey B. Wakefield, Stephen J. Newman, Kyle B. Hillcoat, Benjamin J. Saunders, and Euan S. Harvey. Utility of body and otolith morphometry to discriminate cryptic juveniles of two sympatric red snappers (Perciformes: Lutjanidae). Journal of Fish Biology, 103(6): 1312–1320, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Berthelsen:2023:EIF

[1345] Andreas S. Berthelsen, Christian Skov, Martin Søndergaard, Martin H. Larsen, and Torben L. Lauridsen. Ecological implications of fish removal: Insights from gut-content analysis of roach (Rutilus rutilus) and European perch (Perca fluviatilis) in a eutrophic shallow lake. Journal of Fish Biology, 103(6):1321–1334, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hung:2023:SVI

[1346] Wei-Che Hung, Jhen Hsu, Yi-Jay Chang, Wen-Bin Huang, Chih-Hao Hsieh, and Ming-Tsung Chung. Spatiotemporal variations and influential factors affecting reproductive dynamics of Pacific saury in the Northwest Pacific during fishing season. Journal of Fish Biology, 103(6):1335–1346, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yao:2023:GCD

[1347] Akifumi Yao, Mayuko Nakamura, Hisanori Kohtsuka, Tomoki Sunobe, and Toru Miura. Gonadal and cellular dynamics during protogynous sex change in the harlequin sandsmelt Parapercis pulchella. Journal of Fish Biology, 103(6):1347–1356, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grant:2023:EAV

[1348] Michael I. Grant, Peter M. Kyne, Julie James, Yi Hu, Sushmita Mukherji, Yolarnie Amepou, Leontine Baje, Andrew Chin, Grant Johnson, Tegan Lee, Brandon Mahan, Christopher Wurster, William T. White, and Colin A. Simpfendorfer. Elemental analysis of vertebrae discerns diadromous movements of threatened non-marine elasmobranchs. Journal of Fish Biology, 103(6):1357–1373, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Castro-Gutierrez:2023:EMF

[1349] Jairo Castro-Gutiérrez, Sara Madera-Santana, Carlos Rodríguez-García, Ángel Rafael Domínguez-Bustos, Jésica Sarmiento-Carbajal, José Belquior Gonçalves-Neto, and Remedios Cabrera-Castro. Exploring morphometric frontiers: a comprehensive study of otolith growth patterns in brown comber Serranus hepatus (Linnaeus, 1758). Journal of Fish Biology, 103(6):1374–1381, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2023:GSR

[1350] Luojia Li, Longjun Deng, Jie Li, Tiancai Li, Pengyu Chen, Wei Luo, and Zongjun Du. Gill structure and respiratory ability of Euchiloglanis kishinouyei (Osteichthyes: Siluriformes: Sisoridae). Journal of Fish Biology, 103(6):1382–1391, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ly:2023:DGP

[1351] Tien H. Ly, Cheng-Ting Huang, Meng-Chou Lee, Vo V. Thanh, and Doan X. Diep. Determination of the growth performance and reproductive season of wild Channa lucius (Cuvier, 1831) populations of genus Channa (Channidae, Perciformes) in the Mekong Delta, Vietnam. Journal of Fish Biology, 103(6):1392–1400, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Glenn:2023:WRF

[1352] K. R. Glenn and C. M. Pennuto. Winter residency and foraging of nonnative round goby populations in Great Lakes tributary streams. Journal of Fish Biology, 103(6):1401–1408, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Olin:2023:DMC

[1353] Jill A. Olin, Hidetoshi Urakawa, Michael G. Frisk, Alisa L. Newton, Maria Manz, Michael Fogg, Colin McMullen, Lisa Crawford, and Oliver N. Shipley. DNA metabarcoding of cloacal swabs provides insight into diets of highly migratory sharks in the Mid-Atlantic Bight. Journal of Fish Biology, 103(6):1409–1418, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cote:2023:DSB

[1354] David Cote, Jordan Sutton, Sheena Roul, Hilario Murua, Fernando Gonzales, Ricardo Alpoim, and Joey Angnatok. The distribution of subarctic and boreal deep-sea demersal fish assemblages across environmental gradients of the northwest Atlantic. Journal of Fish Biology, 103(6): 1419–1429, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tremblay-Gagnon:2023:SVD

[1355] Félix Tremblay-Gagnon, Sarah Brown-Vuillemin, Katherine Skanes, Hannah Polaczek, Wojciech Walkusz, Dominique Robert, and David Deslauriers. Spatiotemporal variability in diet composition of Greenland halibut (Reinhardtius hippoglossoides) from the eastern Canadian Arctic. Journal of Fish Biology, 103(6):1430–1444, December 2023. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kar:2023:MPF

[1356] Chinmay Kar, Puthiyara Chetta Mariyambi, Rajeev Raghavan, and Sivanpillai Sureshkumar. Mitochondrial phylogeny of fusilier fishes (family Caesionidae) from the Laccadive archipelago reveals a new species and two new records from the Central Indian Ocean. Journal of Fish Biology, 103(6):1445–1451, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Strader:2023:CVH

[1357] Ryan N. Strader, Sally C. Dowd, Camryn Blawas, Richard D. Mahoney, Natalie C. Patetta, Jaelyn Leslie, and Janet A. Nye. Climate variability hypothesis is partially supported in thermal limits of juvenile Northwest Atlantic coastal fishes. Journal of Fish Biology, 103(6):1452–1462, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lehtonen:2023:EFV

[1358] Topi K. Lehtonen, David Gilljam, Lari Veneranta, Tapio Keskinen, and Mikaela Bergenius Nord. The ecology and fishery of the vendace (Coregonus albula) in the Baltic Sea. Journal of Fish Biology, 103(6):1463– 1475, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chuphal:2023:EED

[1359] Bhawna Chuphal, Priyanka Sathoria, Umesh Rai, and Brototi Roy. Exploring the effect of dihydrotestosterone on nucleotide-binding and oligomerization domain-like receptor expression in spotted snakehead Channa punctata (Bloch 1793). Journal of Fish Biology, 103(6):1476– 1487, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ouellet-Proulx:2023:PGT

[1360] Sébastien Ouellet-Proulx, Anik Daigle, André St-Hilaire, Carole-Anne Gillis, Tommi Linnansaari, Guillaume Dauphin, and Normand Émile Bergeron. A potential growth thermal index for estimating juvenile Atlantic salmon (Salmo salar) size-at-age across geographical scales. Journal of Fish Biology, 103(6):1488–1500, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Luksikova:2023:CSD

[1361] Karolína Lukšíková, Tomáš Pavlica, Marie Altmanová, Jana Štundlová, Šárka Pelikánová, Sergey A. Simanovsky, Eugene Yu. Krysanov, Marek Jankásek, Matyáš Hiřman, Martin Reichard, Petr Ráb, and Alexandr Sember. Conserved satellite DNA motif and lack of interstitial telomeric sites in highly rearranged African Nothobranchius killifish karyotypes. Journal of Fish Biology, 103(6):1501–1514, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kiyatake:2023:CGD

[1362] Itsuki Kiyatake, Thomas L. Johnson, Emy Cottrant, Yoshikazu Kitadani, Kiyoko Onda, Mikito Murata, Natalia J. Drobniewska, Timothy G. Paulet, and Kiyonori Nishida. A comparison of the growth and development of pyjama sharks (Poroderma africanum) in wild and captive populations. Journal of Fish Biology, 103(6):1515–1525, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zuercher:2023:FIB

[1363] Rachel Zuercher, David Kochan, and Alastair R. Harborne. Factors influencing the biomass of large-bodied parrotfish species in the absence of fishing on coral reefs in Florida, USA. Journal of Fish Biology, 103(6): 1526–1537, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shipley:2023:EHS

[1364] Oliver N. Shipley, Jill A. Olin, Christopher Scott, Merry Camhi, and Michael G. Frisk. Emerging human–shark conflicts in the New York Bight: a call for expansive science and management. Journal of Fish Biology, 103(6):1538–1542, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lepak:2023:RSO

[1365] Jesse M. Lepak, Adam G. Hansen, Erik T. Cristan, Demetra A. Williams, and William M. Pate. Rainbow smelt (Osmerus mordax) influence on walleye (Sander vitreus) recruitment decline: mtDNA evidence supporting the predation hypothesis. Journal of Fish Biology, 103(6):1543–1548, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Curnick:2023:NRE

[1366] David J. Curnick, Rob Deaville, Jenny R. Bortoluzzi, Luke Cameron, Jeanette E. L. Carlsson, Jens Carlsson, Haley R. Dolton, Cat A. Gordon, Phil Hosegood, Alicia Nilsson, Matthew W. Perkins, Kevin J. Purves, Simon Spiro, Marco Vecchiato, Rosie S. Williams, and Nicholas L. Payne. Northerly range expansion and first confirmed records of the smalltooth sand tiger shark, Odontaspis ferox, in the United Kingdom and Ireland. Journal of Fish Biology, 103(6):1549–1555, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brainard:2023:BRE

[1367] Miranda K. Brainard, Caitlyn P. Swango, Paityn M. Houglan, and Richard L. Londraville. Blackspotted rockskippers (Entomacrodus striatus) avoid refuges with conspecifics. Journal of Fish Biology, 103(6): 1556–1559, December 2023. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Glover:2023:NAR

[1368] Ross S. Glover and Alastair Stephen. A novel automatic release cage increases survival of Atlantic salmon (Salmo salar) smolts released at night. Journal of Fish Biology, 103(6):1560–1564, December 2023. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2024:IIa

[1369] Anonymous. Issue information. Journal of Fish Biology, 104(1):1–2, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2024:CSL

[1370] William Bernard Perry. Climbing the social ladder rewarded with sex change in Gorgeous gobies (Lythrypnus pulchellus). Journal of Fish Biology, 104(1):3, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mainguy:2024:EFM

[1371] Julien Mainguy, Martin Bélanger, Eliane Valiquette, Simon Bernatchez, Léon L’Italien, Russell B. Millar, and Rafael de Andrade Moral. Estimating fish mortality rates from catch curves: a plea for the abandonment of Ricker (1975)’s linear regression method. Journal of Fish Biology, 104(1):4–10, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Welch:2024:JSR

[1372] Rebecca J. Welch, Amber-Robyn Childs, Christian Hempel, and Nicola C. James. Juvenile sparids (Rhabdosargus holubi) consistently select structurally dense vegetated habitat in nursery seascapes. Journal of Fish Biology, 104(1):11–19, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

dePontual:2024:HOC

[1373] Hélène de Pontual, Kirsteen M. MacKenzie, Hélène Tabouret, Françoise Daverat, Kélig Mahé, Christophe Pecheyran, and Karin Hüssy. Heterogeneity of otolith chemical composition from two-dimensional mapping: Relationship with biomineralization mechanisms and implications for microchemistry analyses. Journal of Fish Biology, 104(1):20–33, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Albanesi:2024:DMD

[1374] Camila Paula Albanesi, Eugenia Méndez, María Soledad Michiels, Mariela Radonic, Andrea López, and Alejandra A. López-Mañanes. Differential modulation of digestive enzymes and energy reserves at different times after feeding in juveniles of the marine estuarine-dependent flatfish Paralichthys orbignyanus (Valenciennes, 1839). Journal of Fish Biology, 104(1):34–43, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tseng:2024:GWC

[1375] Mei-Chen Tseng, Yen-Hung Lee, Tsair-Bor Yen, and Shu-Ming Li. Genome-wide characterization of microsatellites in cobia Rachycentron canadum (Linnaeus, 1766): Survey and analysis of their abundance and diversity. Journal of Fish Biology, 104(1):44–55, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Millot:2024:AED

[1376] Rémi Millot, Lucie Vanalderweireldt, Laure Finelli Sandolo, and Eric D. H. Durieux. Age estimates derived from hard parts of swordfish Xiphias gladius from the north-western Mediterranean Sea. Journal of Fish Biology, 104(1):56–68, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ghosh:2024:TID

[1377] Shubhadeep Ghosh, Menon Muktha, Mamidi Satishkumar, Behera Pralaya Ranjan, Divipala Indira, Rohit Prathibha, Pariyappanal Ulahannan Zacharia, and Achamveetil Gopalakrishnan. Trophic interactions and the diet composition of sympatric finfish predators in the tropical demersal ecosystem in the Bay of Bengal. Journal of Fish Biology, 104(1):69–91, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eustache:2024:STA

[1378] Kim B. Eustache, Emiel van Loon, Jodie L. Rummer, Serge Planes, and Isabel Smallegange. Spatial and temporal analysis of juvenile blacktip reef shark (Carcharhinus melanopterus) demographies identifies critical habitats. Journal of Fish Biology, 104(1):92–103, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Spear:2024:IVB

[1379] Luka N. Spear, Suzanne Kohin, John A. Mohan, and R. J. David Wells. Insights into vertebral band pair deposition rate in the juvenile common thresher shark (Alopias vulpinus) in the northeastern Pacific Ocean. Journal of Fish Biology, 104(1):104–112, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sukkhee:2024:HSD

[1380] Nutchar Sukkhee, Sinlapachai Senarat, Natthawut Charoenphon, Gen Kaneko, Jes Kettratad, Kitipong Angsujinda, Piyamat Kongtueng, Tappadit Mitparian, Tassaporn Kanjanarakha, and Sumate Ampawong. Histological structure of the digestive tract, liver, and pancreas of Ambassis nalua (Hamilton, 1822) with ultrastructural details of the gastric gland. Journal of Fish Biology, 104(1):113–124, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Trehin:2024:MPA

[1381] Cécile Tréhin, Etienne Rivot, Valentin Santanbien, Rémi Patin, Stephen D. Gregory, Ludivine Lamireau, Frédéric Marchand, William R. C. Beaumont, Luke J. Scott, Robert Hillman, Anne-Laure Besnard, Pierre-Yves Boisson, Lisa Meslier, Andrew R. King, Jamie R. Stevens, and Marie Nevoux. A multi-population approach supports common patterns in marine growth and maturation decision in Atlantic salmon (Salmo salar L.) from southern Europe. Journal of Fish Biology, 104 (1):125–138, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bahamonde:2024:EVI

[1382] Paulina Andrea Bahamonde, Gustavo Chiang, Gabriela Mancilla, Tamara Contador, Caudio Quezada-Romegialli, Kelly Roland Munkittrick, and Chris Harrod. Ecological variation in invasive brown trout (Salmo trutta) within a remote coastal river catchment in northern Patagonia complicates estimates of invasion impact. Journal of Fish Biology, 104(1):139–154, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Milan:2024:HEL

[1383] Jimjohn Milan, Tyler Fidel, Watcharapong Hongjamrassilp, Benjamin W. Frable, Zachary R. Skelton, and Philip A. Hastings. Heterochrony and the evolution of the longjaw mudsucker (Gobiidae, Teleostei). Journal of Fish Biology, 104(1):155–162, January 2024. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vivas:2024:PDL

[1384] M. Vivas, J. Peñalver, J. A. Oliver, J. D. López Giraldo, and C. Mena. Population dynamics of the long-snouted seahorse (Hippocampus guttulatus Cuvier, 1829) in the Mar Menor coastal lagoon. Journal of Fish Biology, 104(1):163–170, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fam:2024:CGM

[1385] Yi Qing Fam, Jamsari Amirul Firdaus Jamaluddin, Abdullah Halim Muhammad-Rasul, Mohd Lokman Ilham-Norhakim, Nik Fadzly N. Rosely, and Sébastien Lavoué. Contrasting genetic and morphological differentiation among geographical lineages of a stenotopic miniature rasborine, Boraras maculatus, in Peninsular Malaysia. Journal of Fish Biology, 104(1):171–183, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Munoz-Arroyo:2024:BSC

[1386] Sayuri Muñoz-Arroyo, Danitzia A. Guerrero-Tortolero, Liliana Hernández-Olalde, and Eduardo F. Balart. Bidirectional sex-change behavior and physiological aspects in the Gorgeous goby Lythrypnus pulchellus (Gobiidae). Journal of Fish Biology, 104(1):184–205, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lehun:2024:FTH

[1387] Atsler Luana Lehun, Carolina Mendes Muniz, João Otávio Santos Silva, Lidiany Doreto Cavalcanti, and Ricardo Massato Takemoto. The functional traits of host fish can act as good predictors for parasite composition in a neotropical floodplain. Journal of Fish Biology, 104(1):206– 215, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nguyen:2024:EDH

[1388] Minh Chau Nguyen, Ravi Fotedar, and Hung Duc Pham. Effects of dietary hydrolysate supplementation on growth, body composition, hematological responses, and liver histology of juvenile giant trevally (Caranx ignobilis Forsskal, 1775). Journal of Fish Biology, 104(1):216–226, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kaya:2024:TEN

[1389] Cüneyt Kaya, Baran Yoğurtçuoğlu, İsmail Aksu, Esra Bayçelebi, and Davut Turan. Turcinoemacheilus ekmekciae, a new dwarf loach from upper Tigris and Euphrates (Teleostei: Nemacheilidae). Journal of Fish Biology, 104(1):227–239, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Miller:2024:CSI

[1390] Emily A. Miller, Gabriel P. Singer, Matthew L. Peterson, Molly Webb, and A. Peter Klimley. Comparative stable isotope analyses of green sturgeon (Acipenser medirostris) and white sturgeon (A. transmontanus) in the San Francisco estuary. Journal of Fish Biology, 104(1):240–251, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Knorrn:2024:RBE

[1391] Alexander H. Knorrn, Kim L. Wieben, Heino O. Fock, and Henrike Andresen. Reproductive biology of the electric lanternfish Electrona risso (Myctophidae) and the bigscale fishes Melamphaes polylepis and Scopelogadus mizolepis (Melamphaidae). Journal of Fish Biology, 104(1):252– 264, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lilly:2024:MPN

[1392] Jessie Lilly, Hannele H. Honkanen, Jessica R. Rodger, Diego del Villar, Patrick Boylan, Amy Green, Diego Pereiro, Lorna Wilkie, Richard Kennedy, Andrea Barkley, Robert Rosell, Niall Ó. Maoiléidigh, Ross O’Neill, Catherine Waters, Deirdre Cotter, David Bailey, William Roche, Ross McGill, James Barry, Samantha V. Beck, Jim Henderson, Debbie Parke, Frederick G. Whoriskey, Brian Shields, Philip Ramsden, Silas Walton, Melanie Fletcher, Ken Whelan, Colin W. Bean, Sophie Elliott, Adrian Bowman, and Colin E. Adams. Migration patterns and navigation cues of Atlantic salmon post-smolts migrating from 12 rivers through the coastal zones around the Irish Sea. Journal of Fish Biology, 104(1): 265–283, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Choi:2024:COP

[1393] Si-Won Choi, Hyo-Jae Yu, and Jin-Koo Kim. Comparative ontogeny and phylogenetic relationships of eight lizardfish species (Synodontidae) from the Northwest Pacific, with a focus on Trachinocephalus monophyly. Journal of Fish Biology, 104(1):284–303, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

ONeill:2024:SCS

[1394] Helen L. O’Neill, William T. White, John J. Pogonoski, Belinda Alvarez, Oliver Gomez, and John K. Keesing. Sharks checking in to the sponge hotel: First internal use of sponges of the genus Agelas and family Irciniidae by banded sand catsharks Atelomycterus fasciatus. Journal of Fish Biology, 104(1):304–309, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fenton:2024:CPA

[1395] M. A. Fenton, H. L. Fennell, and M. J. Kaiser. Characteristics of papers that affect citations in the Journal of Fish Biology. Journal of Fish Biology, 104(1):310–314, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vanegas-Rios:2024:LDB

[1396] James A. Vanegas-Ríos and Gustavo E. Chiaramonte. On the lectotype designation of Bryconamericus rubropictus (Berg) (Characiformes, Characidae, Stevardiinae). Journal of Fish Biology, 104(1):315–319, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Soares:2024:FRI

[1397] Marta C. Soares, Sónia C. Cardoso, Pedro Fernandes, José C. Alves, Pedro M. Anastácio, and Filipe Banha. First record of intraspecific cleaning behavior by the threatened Leuciscidae Squalius alburnoides (Steindachner, 1866) at the Guadiana River basin (Portugal). Journal of Fish Biology, 104(1):320–323, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Correia:2024:DCE

[1398] Edna Correia, José Pedro Granadeiro, Bárbara Santos, Vanessa A. Mata, Emanuel Dias, Aissa Regalla, and Teresa Catry. Diet of the critically endangered blackchin guitarfish Glaucostegus cemiculus revealed using DNA metabarcoding. Journal of Fish Biology, 104(1):324–328, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skora:2024:ESR

[1399] Michał E. Skóra, Gudni Gudbergsson, Gordon H. Copp, and J. Iwan Jones. Evidence of successful recruitment of non-native pink salmon Oncorhynchus gorbuscha in Iceland. Journal of Fish Biology, 104(1):329– 334, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ehemann:2024:EBB

[1400] N. R. Ehemann and L. A. Zambrano-Vizquel. Echinorhinus brucus (Bonnaterre, 1788) in the Caribbean Sea: a recurrent visitor, or are the artisanal fisheries exploiting deeper waters? Journal of Fish Biology, 104 (1):335–340, January 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2024:IIb

[1401] Anonymous. Issue information. Journal of Fish Biology, 104(2):341–342, February 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2024:RAD

[1402] Shan Li, Hui Wei, and Gordon H. Copp. Research advances in diversity and conservation science of freshwater fish in China. Journal of Fish Biology, 104(2):343–344, February 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cao:2024:RCS

[1403] Liang Cao, Wei-Han Shao, Wen-Jing Yi, and E Zhang. A review of conservation status of freshwater fish diversity in China. Journal of Fish Biology, 104(2):345–364, February 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Qin:2024:ITN

[1404] Jiao Qin, Fengming Liu, Bjorn Victor Schmidt, Zhuoxin Sun, Lingwei Kong, and Yunrong Yan. Interpopulation trophic niches and ontogenetic shifts of a mangrove fish predator. Journal of Fish Biology, 104(2):365– 373, February 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hu:2024:AHM

[1405] Jiaxin Hu, Jie Zhao, Xiaoyun Sui, Ren Zhu, and Dekui He. Across the highest mountain on earth: discordant phylogeographic patterns and recent dispersal of Tibetan stone loaches (Triplophysa) in the Himalayas. Journal of Fish Biology, 104(2):374–386, February 2024. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shuai:2024:ULH

[1406] Fangmin Shuai, Haiyan Li, Jie Li, Tao Jiang, Jian Yang, and Wanling Yang. Unravelling the life-history patterns and habitat preferences of the Japanese eel (Anguilla japonica) in the Pearl River, China. Journal of Fish Biology, 104(2):387–398, February 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

He:2024:AGR

[1407] Jiayang He, Zhiqiang Wu, Liangliang Huang, Yangyan Sun, Dongjie Wang, Yong Lin, Anyou He, Jie Feng, and Hao Liu. Age, growth, reproduction and status of resource development of Ptychidio jordani, a critically endangered freshwater fish in the Hongshui River, China. Journal of Fish Biology, 104(2):399–409, February 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yimer:2024:AGM

[1408] Melaku Abelneh Yimer, Liang Cao, Jiang-Zhong Shen, and E Zhang. Age, growth, maturity and mortality of the tapetail anchovy Coilia brachygnathus (Engraulidae) in Lake Honghu, China. Journal of Fish Biology, 104(2):410–421, February 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shen:2024:ICF

[1409] Wei Shen, Zhanfei Peng, and Jin Zhang. Identification and counting of fish targets using adaptive resolution imaging sonar. Journal of Fish Biology, 104(2):422–432, February 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Guo:2024:DTN

[1410] Dong-Ming Guo, Liang Cao, and E. Zhang. Descriptions of two new species of the botiid genus Leptobotia Bleeker, 1870 (Teleostei: Cypriniformes) from South China. Journal of Fish Biology, 104(2):433–449, February 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lin:2024:WIE

[1411] Li Lin, Wei-De Deng, Jin-Tao Li, and Bin Kang. Whether including exotic species alters conservation prioritization: a case study in the Min River in southeastern China. Journal of Fish Biology, 104(2):450–462, February 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Guo:2024:TCB

[1412] Xing chen Guo, Zhi peng Chu, Rong qun Song, Zhen lu Wang, Rui jiao Li, Dong mei Xiong, Mu zi Zhang, Hai bo Jiang, and Jian Shao. Toxicity of common biocides used in aquaculture to embryos and larvae of Brachymystax tsinlingensis Li. Journal of Fish Biology, 104(2):463– 472, February 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Song:2024:ESR

[1413] Yiqing Song and Songguang Xie. Effects of substrate roughening on the swimming performance of Schizothorax wangchiachii (Fang, 1936) in the Heishui River: Implications for vertical slot fishway design. Journal of Fish Biology, 104(2):473–483, February 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Luo:2024:NCS

[1414] Qi Luo, Qian Tang, Lei Deng, Qian Duan, and Renyi Zhang. A new cavefish of Sinocyclocheilus (Teleostei: Cypriniformes: Cyprinidae) from the Nanpanjiang River in Guizhou, China. Journal of Fish Biology, 104 (2):484–496, February 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2024:FRG

[1415] Xixi Li, Tong Mu, Ziwang Wang, Jianshuo Qian, and Ren Zhu. First record of the genus Schizothorax in rivers of the Central Himalayas in China. Journal of Fish Biology, 104(2):497–504, February 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2024:IIc

[1416] Anonymous. Issue information. Journal of Fish Biology, 104(3):505– 506, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2024:TGF

[1417] William Bernard Perry. They grow up so fast, especially when global warming is concerned. Journal of Fish Biology, 104(3):507, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whitfield:2024:LCC

[1418] Alan K. Whitfield. Life cycles and challenges for fish species that breed in South African estuaries. Journal of Fish Biology, 104(3):508–516, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tiddy:2024:ECA

[1419] Isabelle C. Tiddy, Kevin Schneider, and Kathryn R. Elmer. Environmental correlates of adaptive diversification in postglacial freshwater fishes.

Journal of Fish Biology, 104(3):517–535, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gomes-Goncalves:2024:ICI

[1420] Rafaela de Sousa Gomes-Gonçalves and Francisco Gerson Araújo. Interdecadal changes in ichthyofauna in a tropical bay with high anthropogenic influences: functional stability despite turnover predominance. Journal of Fish Biology, 104(3):536–547, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Seth:2024:DNS

[1421] Jaya Kishor Seth, Sanmitra Roy, Sameer Sura, Dilraj Puvala, Subhrendu Sekhar Mishra, and Anil Mohapatra. Description of a new species of the genus Awaous Valenciennes, 1837 (Gobiiformes: Oxudercidae) from the middle stretch of the Mahanadi River, Odisha, India, with comments on the Awaous species from India. Journal of Fish Biology, 104(3):548–563, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Andreeva:2024:HDL

[1422] Alla M. Andreeva, Nina Lamash, V. I. Martemyanov, A. S. Vasiliev, I. Yu. Toropygin, and D. V. Garina. High-density lipoprotein remodeling affects the osmotic properties of plasma in goldfish under critical salinity. Journal of Fish Biology, 104(3):564–575, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McAllister:2024:PEJ

[1423] Mia McAllister, Shaun Fraser, and Lea-Anne Henry. Population ecology and juvenile density hotspots of thornback ray (Raja clavata) around the Shetland Islands, Scotland. Journal of Fish Biology, 104(3):576– 589, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vijayakrishnan:2024:ACN

[1424] Balaji Vijayakrishnan and Praveenraj Jayasimhan. Amblyceps crassioris, a new sisoroid catfish from Odisha, India (Siluriformes: Amblycipitidae). Journal of Fish Biology, 104(3):590–597, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:2024:BIR

[1425] James Kennedy, Anna Heida Ólafsdóttir, Svandís Eva Aradóttir, Svanhildur Egilsdóttir, and Christophe Pampoulie. Biological information on a rare pelagic fish, black ruff Centrolophus niger, caught in Icelandic waters: Distribution, feeding, and otoliths. Journal of Fish Biology, 104 (3):598–610, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Collar:2024:SFS

[1426] David C. Collar, Trevor J. Hobbs, and Jessica S. Thompson. Scaling of fast-start performance and its thermal dependence in mummichog Fundulus heteroclitus. Journal of Fish Biology, 104(3):611–623, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yang:2024:CGC

[1427] Guokun Yang, Boya Yang, Sunan Wang, Xiaomin Liang, Chengquan Li, Yanmin Zhang, Xindang Zhang, Xulu Chang, and Xiaolin Meng. Cloning grass carp (Ctenopharyngodon idella) ccdc3 and its expression affected by nutrition state, insulin, and glucagon. Journal of Fish Biology, 104 (3):624–632, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yousaf:2024:HRI

[1428] Muhammad Naveed Yousaf, Øyvind Røn, Frederike Keitel-Gröner, Charles McGurk, and Alex Obach. Heart rate as an indicator of stress during the critical swimming speed test of farmed Atlantic salmon (Salmo salar L.). Journal of Fish Biology, 104(3):633–646, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rinaldo:2024:GWP

[1429] Adrian Rinaldo, Elvira de Eyto, Thomas Reed, Karl Øystein Gjelland, and Philip McGinnity. Global warming is projected to lead to increased freshwater growth potential and changes in pace of life in Atlantic salmon Salmo salar. Journal of Fish Biology, 104(3):647–661, March 2024. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Abdussamad:2024:UTA

[1430] Eruppakkottil M. Abdussamad, Thomas Toji, Amalraj M. R. Margaret, Kalappurakkal G. Mini, Kothanahally M. Rajesh, Pokkathappada A. Azeez, Vinothkumar Ramar, Thayichira B. Retheesh, Athikkattil M. Abbas, Ismail Shihab, Sneha M. George, Rohith Prathibha, and Achamveettil Gopalakrishnan. Untangling the taxonomic ambiguities of the spotted seerfish Scomberomorus guttatus with the description of a new species from India. Journal of Fish Biology, 104(3):662–680, March 2024. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dauphin:2024:EMY

[1431] Guillaume J. R. Dauphin, Carole-Anne Gillis, and Gérald J. Chaput. Estimating multiple years, tributary-specific, and overall Atlantic salmon smolt abundance in a large Canadian catchment using capture–mark– recapture experiments. Journal of Fish Biology, 104(3):681–697, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McCarthy:2024:MHS

[1432] James H McCarthy, Tommi Linnansaari, and R. Allen Curry. Movement and habitat shift responses of juvenile Atlantic Salmon (Salmo salar) to annually permanent stream flooding. Journal of Fish Biology, 104 (3):698–712, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pacher:2024:RMT

[1433] Korbinian Pacher, Michael Breuker, Matthew J. Hansen, Ralf H. J. M. Kurvers, Jan Häge, Felicie Dhellemmes, Paolo Domenici, John F. Steffensen, Stefan Krause, Thomas Hildebrandt, Guido Fritsch, Pascal Bach, Philippe S. Sabarros, Paul Zaslansky, Kristin Mahlow, Johannes Müller, Rogelio González Armas, Hector Villalobos Ortiz, Felipe Galván-Magaña, and Jens Krause. The rostral micro-tooth morphology of blue marlin, Makaira nigricans. Journal of Fish Biology, 104(3):713–722, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chollet-Villalpando:2024:CVS

[1434] Jorge Guillermo Chollet-Villalpando, Francisco Javier García-Rodríguez, and José De La Cruz-Agüero. Character variation in separate body regions of Gerreidae (Osteichthyes: Teleostei) fishes inferred from geometric morphometrics. Journal of Fish Biology, 104(3):723–736, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kodeeswaran:2024:NSE

[1435] Paramasivam Kodeeswaran, Ayyathurai Kathirvelpandian, Anil Mohapatra, Thipramalai Thangappan Pillai Ajith Kumar, and Uttam Kumar Sarkar. A new snake eel species of the genus Ophichthus (Anguilliformes: Ophichthidae) from the southeast coast of India, Bay of Bengal, with the taxonomic account of Ophichthus chilkensis. Journal of Fish Biology, 104(3):737–745, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baer:2024:EMS

[1436] Jan Baer, Sabrina Ziegaus, Mark Schumann, Juergen Geist, and Alexander Brinker. Escaping malnutrition by shifting habitats: a driver of three-spined stickleback invasion in Lake Constance. Journal of Fish Biology, 104(3):746–757, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2024:EEI

[1437] Zhe Li, Shanshan Zhou, Jingjing He, Jie Ying, and Kaida Xu. Environmental enrichment improves behaviors rather than the growth and physiology of rock bream Oplegnathus fasciatus. Journal of Fish Biology, 104(3):758–768, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Epur:2024:FDL

[1438] Irina V. Epur, Andrey A. Balanov, Vladimir A. Shelekhov, and Sergey V. Turanov. First description of larvae of Radulinopsis derzhavini Soldatov et Lindberg, 1930 (Scorpaeniformes: Psychrolutidae) with remarks on melanin colouration and relationships with related species. Journal of Fish Biology, 104(3):769–779, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dumith:2024:UTE

[1439] Michelle Torres Dumith and Alejandra F. G. N. Santos. Use of trophic ecology of omnivorous fish and abiotic factors as supporting tools for assessing environmental impacts in a neotropical river. Journal of Fish Biology, 104(3):780–796, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Erkinaro:2024:AFD

[1440] Jaakko Erkinaro, Panu Orell, Mikko Kytökorpi, Jan-Peter Pohjola, and Michael Power. Active feeding of downstream migrating juvenile pink salmon (Oncorhynchus gorbuscha) revealed in a large Barents Sea river using diet and stable isotope analysis. Journal of Fish Biology, 104(3): 797–806, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Todisco:2024:TDD

[1441] Vittoria Todisco, Emil Fridolfsson, Charlotte Axén, Elin Dahlgren, Maciej J. Ejsmond, Marc M. Hauber, Kjetil Hindar, Petter Tibblin, Markus Zöttl, Linda Söderberg, and Samuel Hylander. Thiamin dynamics during the adult life cycle of Atlantic salmon (Salmo salar). Journal of Fish Biology, 104(3):807–824, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schneider:2024:CDS

[1442] Eric V. C. Schneider, Brendan S. Talwar, Shaun S. Killen, Samantha Russell, Travis E. Van Leeuwen, and David M. Bailey. Colonization, diversity, and seasonality of fishes at pelagic fish aggregating devices. Journal of Fish Biology, 104(3):825–836, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Belcik:2024:RGO

[1443] John T. Belcik and Mary V. Ashley. Riverscape genetics of the orangethroat darter complex. Journal of Fish Biology, 104(3):837–850, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dodd:2024:CPW

[1444] Jamie R. Dodd, Ian G. Cowx, Domino A. Joyce, and Jonathan D. Bolland. Can’t pass or won’t pass: the importance of motivation when quantifying improved connectivity for riverine brown trout Salmo trutta. Journal of Fish Biology, 104(3):851–865, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kurita:2024:RBS

[1445] Yutaka Kurita, Kotaro Shirai, Kaoru Kubota, Hiroyuki Togashi, and Takami Morita. Relationship between stable cesium concentration and body size of Japanese flounder Paralichthys olivaceus and the effect of a size-dependent shift in diet. Journal of Fish Biology, 104(3):866–877, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rigg:2024:SPA

[1446] Alexander L. Rigg, Clara Bellotto, Ashley M. Fowler, and David J. Booth. Staining protocols affect use of otolith to estimate the demography of the damselfish sergeant major (Abudefduf vaigiensis). Journal of Fish Biology, 104(3):878–882, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Black:2024:MRC

[1447] Taylor N. Black, Aakriti Rastogi, Abby Saegert, Joseph Dib, and Rachel L. Moran. Male recognition of conspecific female chemical cues in a diverse clade of freshwater fishes. Journal of Fish Biology, 104(3):883– 886, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stewart:2024:USS

[1448] Andrew L. Stewart, Theodore W. Pietsch, Jon Moore, and Xiaotong Peng. Upside-down swimming: in situ observations of inverted orientation in Gigantactis, with a new depth record for the Ceratioidei. Journal of Fish Biology, 104(3):887–891, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Prokofiev:2024:SDO

[1449] Artem M. Prokofiev. Supplementary description and ontogenetic variability of Coryphaenoides affinis: rediscovery of the species after 145 years since its original description (Teleostei: Gadiformes: Macrouridae). Journal of Fish Biology, 104(3):892–897, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bullock:2024:JKS

[1450] Robert W. Bullock, Craig Foster, and James S. E. Lea. Just keep swimming? Observations of resting behavior in gray reef sharks Carcharhinus amblyrhynchos (Bleeker, 1856). Journal of Fish Biology, 104(3):898–900, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lechner:2024:CBT

[1451] Emily R. Lechner, Erin M. C. Stewart, Chris C. Wilson, and Graham D. Raby. CT_{max} in brook trout (Salvelinus fontinalis) embryos shows an acclimation response to developmental temperatures but is more variable than in later life stages. Journal of Fish Biology, 104(3):901–905, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Britton:2024:GHC

[1452] J. Robert Britton and Ali Serhan Tarkan. Gordon H. Copp: an obituary. Journal of Fish Biology, 104(3):906–907, March 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2024:IId

[1453] Anonymous. Issue information. Journal of Fish Biology, 104(4):909– 910, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2024:MFN

[1454] William Bernard Perry. Macabre ‘Finding Nemo’ remake, starring the reticulated damselfish (Dascyllus reticulatus). Journal of Fish Biology, 104(4):911, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kelly:2024:OCT

[1455] Lisa A. Kelly, Christopher K. Yost, and Steven J. Cooke. Opportunities and challenges with transitioning to non-lethal sampling of wild fish for microbiome research. Journal of Fish Biology, 104(4):912–919, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ogura:2024:ISM

[1456] Yuhei Ogura and Koh Hasegawa. Infection status of masu salmon Oncorhynchus masou by the parasite Salmonema cf. ephemeridarum (Nematoda, Cystidicolidae). Journal of Fish Biology, 104(4):920–928, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lin:2024:QCM

[1457] Kai Lin, Shiyu Zhang, Junjie Hu, Xingdong Lv, and Hongsong Li. Quantitative comparison of 2D and 3D monitoring dimensions in fish behavior analysis. Journal of Fish Biology, 104(4):929–938, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:2024:PRN

[1458] Morgan S. Brown, Roberto Carvalheiro, Richard S. Taylor, Wagdy Mekkawy, Timothy D. W. Luke, Lewis Rands, Damien Nieuwesteeg, Brad S. Evans, Nicholas M. Wade, Curtis E. Lind, and Pollyanna E. Hilder. Probabilistic reaction norm reveals family-related variation in the association between size, condition, and sexual maturation onset in Atlantic salmon (Salmo salar). Journal of Fish Biology, 104(4):939–949, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Waters:2024:GPD

[1459] Jonathan M. Waters, Tania M. King, and Dave Craw. Gorges partition diversity within New Zealand flathead Galaxias populations. Journal of Fish Biology, 104(4):950–956, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Landaeta:2024:MDD

[1460] Mauricio F. Landaeta, Matías Pareja, Mathias Hüne, Lisette Zenteno-Devaud, Javier Vera-Duarte, Valentina Bernal-Durán, Manuel I. Castillo, and Mario La Mesa. Morphology and diet are decoupled in nearshore notothenoids from King George Island, West Antarctica. Journal of Fish Biology, 104(4):957–968, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neuhaus:2024:NDS

[1461] Emanuel Bruno Neuhaus, Vanessa Meza-Vargas, Julio Ramírez Herrera, and Nathan K. Lujan. A new distinctively striped species of bushynose catfish (Siluriformes: Loricariidae: Ancistrus) from the Pachitea River drainage, Pasco, Peru. Journal of Fish Biology, 104(4):969–978, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sakai:2024:PEG

[1462] Yoichi Sakai, Yu Muranaka, Hiromi Nakayama, Koji Baba, Kazuya Matsushita, Hiroyuki Kuwahara, and Hisato Kuniyoshi. The potential for egg-guarding care in females of the damselfish, Dascyllus reticulatus, in the absence of uniparental male care. Journal of Fish Biology, 104(4): 979–988, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Londono-Burbano:2024:TRS

[1463] Alejandro Londoño-Burbano and Marcelo R. Britto. Taxonomic revision of Sturisoma Swainson, 1838 (Loricariidae: Loricariinae), with descriptions of four new species. Journal of Fish Biology, 104(4):989–1041, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mejia:2024:MMD

[1464] Eduardo Mejia and Roberto E. Reis. Molecular and morphometric data provide evidence of intraspecific variation in shape and pigmentation pattern in Otocinclus cocama (Siluriformes: Loricariidae) across major river drainages. Journal of Fish Biology, 104(4):1042–1053, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Boussinet:2024:JDM

[1465] Elodie Boussinet, David José Nachón, Aldo Sottolichio, Aude Lochet, Stefan Stoll, Gilles Bareille, Helene Tabouret, Christophe Pécheyran, Marie-Laure Acolas, and Françoise Daverat. Juvenile downstream migration patterns of an anadromous fish, allis shad (Alosa alosa), before and after the population collapse in the Gironde system, France. Journal of Fish Biology, 104(4):1054–1066, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vo:2024:NSE

[1466] Van Quang Vo, Hsuan-Ching Ho, Ha Viet Dao, and Thinh Cong Tran. New species of the eel genera Dysomma and Dysommina from Vietnam, South China Sea (Anguilliformes: Synaphobranchidae). Journal of Fish Biology, 104(4):1067–1078, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Simon:2024:OMM

[1467] Janek Simon. Optimizing the mass marking of fish with alizarin red S: an example with glass eels. Journal of Fish Biology, 104(4):1079–1090, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morgan:2024:HCF

[1468] Ashley N. Morgan, Susan B. Fogelson, Paul S. Wills, Tracy Mincer, Sahar Mejri, and Annie Page. Hematological changes in Florida pompano (Trachinotus carolinus) supplemented with β-glucan and Pediococcus acidilactici synbiotic. Journal of Fish Biology, 104(4):1091–1111, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martins:2024:FBY

[1469] Karla Martins, Yuri Niella, Fernanda Albuquerque, Leandro Nolé Eduardo, Paulo Oliveira, and Paulo Travassos. Feeding behavior of yellowfin tuna around two insular regions of the western Atlantic Ocean. Journal of Fish Biology, 104(4):1112–1121, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bessell:2024:PPC

[1470] Tyson J. Bessell, Rick D. Stuart-Smith, Olivia J. Johnson, Neville S. Barrett, Tim P. Lynch, Andrew J. Trotter, and Jemina Stuart-Smith. Population parameters and conservation implications for one of the world’s rarest marine fishes, the red handfish (Thymichthys politus). Journal of Fish Biology, 104(4):1122–1135, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kanjanarakha:2024:CDB

[1471] Tassaporn Kanjanarakha, Sinlapachai Senarat, Kitipong Angsujinda, Gen Kaneko, Atsuo lida, Pahol Kosiyachinda, Witchaya Tongtako, Supapong Imsonpang, and Jes Kettratad. Comparative digestive biology between the ponyfishes from the Pranburi River estuary, Thailand. Journal of Fish Biology, 104(4):1136–1151, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rodriguez-Garcia:2024:LSZ

[1472] Carlos Rodríguez-García, Isabel Abad-Rodríguez, Jesús López-Báez, and Remedios Cabrera-Castro. A life in the surf zone: Age and feeding habits of the pompano (Trachinotus ovatus, Linnaeus, 1758) on the beaches of the Gulf of Cádiz (south-west Iberian Peninsula). Journal of Fish Biology, 104(4):1152–1164, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

duToit:2024:HLM

[1473] Hendrik D. du Toit, Gavin M. Rishworth, Nadine A. Strydom, and Shaun Welman. High levels of metacercarial infestation (family: Diplostomidae) do not affect host energetics and swimming performance in the Epaulette goby (Coryogalops sordidus, Gobiidae). Journal of Fish Biology, 104(4): 1165–1179, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liu:2024:ISM

[1474] Jiehao Liu, Fahim UllahKhan, Shen Jin, Yueping Zheng, Jianan Xu, Houyong Fan, Youji Wang, and Menghong Hu. Indexing serum and mucous biochemical parameters of endangered Chinese sturgeon Acipenser sinensis with implications for health assessment. Journal of Fish Biology, 104(4):1180–1192, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

eSilva:2024:EPF

[1475] Glaucia Moises Marques e Silva, Jorge Eduardo Lins Oliveira, Naisandra B. Silva Farias, Raimundo F. A. Júnior, and Wallace S. Nascimento. Effects of the pyriproxyfen and fenthion on the gonadal morphology of Nile tilapia (Oreochromis niloticus: Perciformes, Cichlidae). Journal of Fish Biology, 104(4):1193–1201, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lubin:2024:HSD

[1476] François-Raphaël Lubin, Emilie Réalis-Doyelle, Laurent Espinat, Jean Guillard, and Allan Raffard. Heat shocks during egg incubation led to developmental, morphological, and behavioral differences in Arctic charr (Salvelinus alpinus). Journal of Fish Biology, 104(4):1202–1212, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oz:2024:EGG

[1477] Mustafa Öz, Burak Evren Inanan, Tahir Karasahin, and Suat Dikel. Effects of glutamine on growth performance, nutrient content, fatty acid profile, and blood parameters of rainbow trout (Oncorhynchus mykiss). Journal of Fish Biology, 104(4):1213–1222, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hintz:2024:WVS

[1478] William D. Hintz, Anthony P. Porreca, and James E. Garvey. Water velocity shapes fish movement behavior. Journal of Fish Biology, 104 (4):1223–1230, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sundin:2024:OPP

[1479] Josefin Sundin, Marko Freese, Lasse Marohn, Tina Blancke, and Reinhold Hanel. Occurrence of the pugnose pipefish Bryx dunckeri in the Sargasso Sea. Journal of Fish Biology, 104(4):1231–1236, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Franklin:2024:IOR

[1480] Erik C. Franklin, Madeleine T. Platt, and Pelika Andrade. Increased occurrence of the rare golden color morph of Pacific chub Kyphosus sandwicensis in a no-take marine reserve. Journal of Fish Biology, 104(4): 1237–1240, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Almeida:2024:MSB

[1481] R. Almeida, S. E. Tanner, C. S. Mateus, F. Ribeiro, and B. R. Quintella. Not so much a sea bass: Divergent European sea bass (Dicentrarchus labrax L.) freshwater incursions. Journal of Fish Biology, 104(4):1241– 1246, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:2024:NBG

[1482] James Kennedy, Ásgeir Gunnarsson, Christophe Pampoulie, and Rupert Wienerroither. Nesting behavior of greater eelpout (Lycodes esmarkii), identified through a predation event by spotted wolffish (Anarhichas minor). Journal of Fish Biology, 104(4):1247–1250, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ellis:2024:OTL

[1483] Jim R. Ellis, Paul Rumney, and Stephanie Cogan. Occurrence of tripletail Lobotes surinamensis (Bloch, 1790) in the inner Bristol Channel, UK. Journal of Fish Biology, 104(4):1251–1254, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barcelos:2024:FRG

[1484] Luís M. D. Barcelos, João B. Barreiros, and João P. Barreiros. First record of the great barracuda, Sphyraena barracuda (Perciformes, Sphyraenidae), in the Azores, north-east Atlantic. Journal of Fish Biology, 104(4):1255–1257, April 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2024:IIe

[1485] Anonymous. Issue information. Journal of Fish Biology, 104(5):1259– 1260, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2024:SLD

[1486] William Bernard Perry. Sounds like a damsel in distress! Journal of Fish Biology, 104(5):1261, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mpopetsi:2024:FDD

[1487] Pule P. Mpopetsi and Wilbert T. Kadye. Functional diversity does not explain the co-occurrence of non-native species within a flow-modified African river system. Journal of Fish Biology, 104(5):1262–1275, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blanchard:2024:PCR

[1488] Rose C. Blanchard, Thomas J. DeWitt, Sharmila Young, and Joshuah S. Perkin. Predictability and conceptual repeatability of the predatorassociated burst speed ecophenotype in western mosquitofish (Gambusia affinis). Journal of Fish Biology, 104(5):1276–1289, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Santos:2024:EMA

[1489] Lucas Santos, João V. Veloso, Theresa B. Santos, Natalia P. A. Bezerra, Paulo Oliveira, and Fabio H. V. Hazin. An equatorial mid-Atlantic Ocean archipelago as nursery area for the cookiecutter shark: Investigating foraging strategies of neonates through bite mark inferences. Journal of Fish Biology, 104(5):1290–1298, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pushchin:2024:RGC

[1490] Igor Pushchin and Naig Aleskerov. Retinal ganglion cell topography and spatial resolving power in the pajama cardinalfish Sphaeramia nematoptera (Bleeker, 1856). Journal of Fish Biology, 104(5):1299–1307, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gong:2024:FES

[1491] Yuyan Gong, Jun Zhang, Yutao Yang, Yane Jiang, Shuai Zhang, Han Tian, and Zuozhi Chen. Feeding ecology of the short-headed lanternfish Diaphus brachycephalus and Warming’s lanternfish Ceratoscopelus warmingii (Myctophidae) in the South China Sea. Journal of Fish Biology, 104(5):1308–1325, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pfennig:2024:SFI

[1492] Meredith B. Pfennig, Derek P. Crane, Nate G. Smith, and Dave L. Buckmeier. I spy: Factors influencing the observation of oxytetracycline in calcified structures of fishes viewed using standard light and fluorescence microscopy. Journal of Fish Biology, 104(5):1326–1338, May 2024. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Santos:2024:EOF

[1493] Darliane C. Santos, Lucas S. Oliveira, Fabíola K. S. Silva, Ruineris A. Cajado, and Diego M. Zacardi. Early ontogeny of the freshwater fish Rhytiodus microlepis (Characiformes, Anostomidae) from the Amazon basin. Journal of Fish Biology, 104(5):1339–1349, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liu:2024:MMB

[1494] Shikang Liu, Chengyang Cao, Yuqian Xi, Pengcheng Li, Xuefeng Chen, Lirong Xiao, and Weiwei Yao. Morphology and motor behavior of endemic fishes in the upper reaches of the Yangtze River basin. Journal of Fish Biology, 104(5):1350–1365, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brule:2024:RSM

[1495] Caroline Brûlé, Kadra Benhalima, Marie-Julie Roux, Geneviève J. Parent, Caroline Chavarria, and Caroline Senay. Reduction in size-at-maturity in unprecedentedly strong cohorts of redfish (Sebastes mentella and S. fasciatus) in the Gulf of St. Lawrence and Laurentian Channel. Journal of Fish Biology, 104(5):1366–1385, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rizzato:2024:GMB

[1496] Pedro P. Rizzato and Maria Elina Bichuette. Gross morphology of the brain and some sense organs of subterranean pencil catfishes of the genus Ituglanis Costa and Bockmann, 1993 (Siluriformes, Trichomycteridae), with a discussion on sensory compensation versus preadaptation in subterranean fishes. Journal of Fish Biology, 104(5):1386–1400, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Phillips:2024:LEC

[1497] Elizabeth W. Phillips, Davide Bottacini, Anna N. M. Schoonhoven, Ynze J. J. Kamstra, Hannah De Waele, Carlos Jimenez, Louis Hadjioannou, and Alexander Kotrschal. Limited effects of culling on the behavior of invasive lionfish (Pterois miles) in the Mediterranean. Journal of Fish Biology, 104(5):1401–1410, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Silva:2024:DAG

[1498] Jeane Campos Silva, Marcelo Soeth, Carlos Werner Hackradt, Andre Lima, and Fabiana Cézar Félix-Hackradt. Daily age, growth rate, and pelagic larval duration of commercially important snapper species in Abrolhos National Marine Park. Journal of Fish Biology, 104(5):1411– 1422, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gaffney:2024:IDO

[1499] Katie Gaffney, Erik Burgerhout, sa Espmark, and Ingrid Lein. The impact of dissolved oxygen concentration on the swimming performance and welfare of ballan wrasse (Labrus bergylta Ascanius, 1767). Journal of Fish Biology, 104(5):1423–1432, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aguilar-Santana:2024:MST

[1500] Fernando A. Aguilar-Santana, Juan Jacobo Schmitter-Soto, Gabriela Lucano-Ramírez, Omar Hernando Avila-Poveda, and Marcial Arellano-Martínez. Morphochromatic spectrum through gonad development stages of the razor surgeonfish, Prionurus laticlavius (Valenciennes, 1846) (Actinopterygii: Acanthuriformes). Journal of Fish Biology, 104 (5):1433–1444, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hata:2024:RRN

[1501] Harutaka Hata and Sébastien Lavoué. Resurrection and redescriptions of nominal species previously regarded as synonyms of Thrissina mystax (Bloch and Schneider, 1801) (Teleostei: Clupeiformes: Engraulidae). Journal of Fish Biology, 104(5):1445–1467, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yonago:2024:LBS

[1502] Takuma Yonago, Tatsuya Kawakami, and Akihide Kasai. Linkage between spatiotemporal distribution of environmental DNA and phenological activity in an amphidromous fish, ayu Plecoglossus altivelis altivelis, in a river located in its northernmost distributional area. Journal of Fish Biology, 104(5):1468–1482, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Qu:2024:IIC

[1503] Leya Qu, Liping Yang, Yiran Wang, Shaoyang Zhi, Mengjuan Zhao, Jinrui Xiong, Xiao Yan, Chaobin Qin, and Guoxing Nie. Identification of Isthmin-1 in common carp (Cyprinus carpio L.) and the effects on glucose metabolism in vivo and in vitro. Journal of Fish Biology, 104 (5):1483–1492, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tuncelli:2024:ESA

[1504] Gökhan Tunçelli and Devrim Memiş. The effect of swimming activity and feed restriction of rainbow trout (Oncorhynchus mykiss) on water quality and fish-plant growth performance in aquaponics. Journal of Fish Biology, 104(5):1493–1502, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Itsukushima:2024:FFE

[1505] Rei Itsukushima and Yuichi Kano. Fish fauna and environmental factors in river estuaries of small and medium-sized rivers on islands significantly influenced by warm currents. Journal of Fish Biology, 104(5):1503–1512, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Denton:2024:MEP

[1506] John S. S. Denton, Jeff Kneebone, Lei Yang, Arve Lynghammar, David McElroy, Shannon Corrigan, Klara Jakobsdóttir, Carolyn Miri, Mark Simpson, and Gavin J. P. Naylor. Mitogenomic evidence of population differentiation of thorny skate, Amblyraja radiata, in the North Atlantic. Journal of Fish Biology, 104(5):1513–1524, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Juza:2024:SFD

[1507] Tomáš Jůza, Milan Muška, Petr Blabolil, Luboš Kočvara, Zuzana Sajdlová, Janis Dumpis, and Ruta Medne. Spatial fish distribution in autumn in a shallow mesotrophic lake, assessed by hydroacoustic surveys, trawling, and beach seining. Journal of Fish Biology, 104(5):1525–1536, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brieske:2024:MDM

[1508] Samantha D. Brieske, Sylvia C. Mullen, and Bernard B. Rees. Method dependency of maximum oxygen uptake rate and its repeatability in the Gulf killifish, Fundulus grandis. Journal of Fish Biology, 104(5):1537– 1547, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Camins:2024:ALH

[1509] Elsa Camins, Lily M. Stanton, Miguel Correia, Sarah J. Foster, Heather J. Koldewey, and Amanada C. J. Vincent. Advances in lifehistory knowledge for 35 seahorse species from community science. Journal of Fish Biology, 104(5):1548–1565, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Xiong:2024:DOG

[1510] Chuanyu Xiong, Yuyu Chi, Bei Wang, Jianhua Yi, Yongyao Yu, Yun Li, Hua Ye, Jiyuan Yin, and Ronghua Wu. Diet with optimal glutathione supplement improves growth, nonspecific immunity, intestinal microbiota, and antioxidant ability in Micropterus salmoides. Journal of Fish Biology, 104(5):1566–1578, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hanson:2024:YDC

[1511] Kathryn A. Hanson, Brooke A. Mauland, Ananda Shastri, and Brian D. Wisenden. Yellowtail damselfish Chrysiptera parasema can associate predation risk with the acoustic call of a heterospecific damselfish following pairing with conspecific alarm cues. Journal of Fish Biology, 104(5): 1579–1586, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garcia-Rodriguez:2024:SSM

[1512] Alain García-Rodríguez, Javier Tovar-Ávila, A. Hiromi Arellano-Cuenca, David Rivas-Landa, Darío A. Chávez-Arrenquín, and Felipe Amezcua. Size at sexual maturity, seasonal variation by maturity stages, and fecundity of the spotted round ray (Urobatis maculatus) and the thorny stingray (Urotrygon rogersi) from the northern tropical eastern Pacific. Journal of Fish Biology, 104(5):1587–1602, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chupil:2024:RMS

[1513] Henrique Chupil, Andrielli Maryan Medeiros, Alisson Peter Brito Dopona, Carolina de Castro Treza, Caroline Freitas Pessi, and Eduardo Roberto Pereira. Rescue method for a stranded giant manta ray (Mobula birostris). Journal of Fish Biology, 104(5):1603–1607, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henderson:2024:MLT

[1514] Aaron C. Henderson, Stephan Bruns, and Shamsa Al Hameli. A mouthless late-term coach whipray Himantura uarnak embryo (Elasmobranchii, Dasyatidae) has implications for our understanding of matrotrophy in chondrichthyans. Journal of Fish Biology, 104(5):1608–1610, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bucair:2024:OSM

[1515] Nayara Bucair, Héllida Negrão Dias, Ana Rita Onodera Palmeira Nunes, Keyton Kylson Fonseca Coelho, Rafaela Maria Serra de Brito, João Bráullio de Luna Sales, Getulio Rincon, Ronaldo Bastos Francini-Filho, June Ferraz Dias, Israel H. A. Cintra, and Jorge Luiz Silva Nunes. Opportunistic sightings of manta rays on Brazil’s Amazon Coast. Journal of Fish Biology, 104(5):1611–1618, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mion:2024:OSR

[1516] Monica Mion, Alessandro Orio, Johan Lövgren, Francesca Vitale, and Massimiliano Cardinale. Otolith sectioning reveals higher maximum age in greater weever (Trachinus draco). Journal of Fish Biology, 104(5): 1619–1622, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leeuwen:2024:WTT

[1517] Travis E. Van Leeuwen, Donald Keefe, Mark Young, and Blair Adams. Water temperature at the time of the catch-and-release event is a better predictor of survival in Atlantic! salmon (Salmo salar) than acute water temperature changes before and after. Journal of Fish Biology, 104(5): 1623–1627, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harris:2024:FRS

[1518] Joanna L. Harris, Claire Collins, Mark Spalding, and Guy M. W. Stevens. First records of the sicklefin (Mobula tarapacana), bentfin (Mobula thurstoni), and spinetail (Mobula mobular) devil rays in the Chagos Archipelago. Journal of Fish Biology, 104(5):1628–1632, May 2024. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamada:2024:ISH

[1519] Taihei Yamada, Takahiro Nobetsu, Hirokazu Urabe, and Futoshi Nakamura. Invasion status of hatchery-origin pink salmon in an unstocked river at the Shiretoko World Natural Heritage Site in northern Japan. Journal of Fish Biology, 104(5):1633–1637, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martinengo:2024:AAE

[1520] Elena Martinengo, Primo Micarelli, Francesca Romana Reinero, Fabrizio Bertelloni, and Filippo Fratini. Antibacterial activity in egg samples from small-spotted catshark Scyliorhinus canicula and nursehound Scyliorhinus stellaris: a preliminary investigation. Journal of Fish Biology, 104 (5):1638–1644, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Finucci:2024:FOS

[1521] Brittany Finucci, Caroline Chin, Helen L. O’Neill, William T. White, and Matthew H. Pinkerton. First observation of a skate egg case nursery in the Ross Sea. Journal of Fish Biology, 104(5):1645–1650, May 2024. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).